A Complete Bibliography of ACM Transactions on Information Systems

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

04 October 2025 Version 1.92

Title word cross-reference

3 [WGL^{+}24]. k [CVS21, KZP10, LHZ^{+}25, ZTM^{+}17, ZLF^{+}17]. N [ABCZ21, CWZ^{+}20b, CWZ^{+}20a, DK04, LPLH19, PV19, SHN17, WZZ^{+}23, XHW^{+}19, ZLF^{+}23, Coh97]. O(1) [Non13].

-Gram [PV19]. -Grams [Coh97, SHN17]. -learning [GWLC06]. -nearest [KZP10]. -workspace [Non13].

1 [HRY^{+}22a, WLF^{+}25a, YSXK23a]. 100-year [Gla04, GL05]. 19 [FWG^{+}24].

2 [HRY^{+}22b, WLF^{+}25b, YSXK23b]. 2007 [Mar08b]. 2008 [Mar08b]. 2009 [Cal12]. 2012 [Cal12].

3 [ZJL^{+}24]. 3D [ABW93]. 3DIS [AM89].

5s [GFWK04].

abbreviations [YKHW06]. Abilities [CACW23, VDR25]. aboutness [WSBC01]. Abstract [MTKY92]. Abstraction [FP94]. Abstractive [BZCE22, CLG^{+}23, LYZ^{+}23]. Academic [LZL^{+}25, ZZZ^{+}25]. Accelerating [ZZL^{+}15]. Access [Gla97, MNP^{+}21b, RS97, WS90, BFF^{+}03, CDR10]. Accesses [CZ93]. accessibility [CDH^{+}09]. Accessing [Dun97]. Accounting [MK16b]. Accuracy [ES13, MMO21, MNP^{+}21a, JGP^{+}07]. Accurate [CJY23a, YLZW22]. Achieves [FJZS24]. Achieving [YXL^{+}23]. Acknowledgment [Ano95]. ACM [All89, All91, dR19]. Acoustic [MO12].

Acquisition [BWS^{+}17]. across [IH07, SMT25, TLKW16]. Act [CW94, KM97]. action [NLL^{+}24]. Actionable [YFO^{+}25]. actions [YHYO01]. Active [CZZ^{+}17, WHFG92, DEL^{+}00, HJZL09]. Activities [LAL^{+}17]. Activity [CHBS15, JMZ22, VR24]. Acyclic [ZLZ^{+}25b]. Ad [CGM05, DLA22, KSYE24, KL09]. AdaGIN [SLZ^{+}25]. Adaptation [SLD^{+}24, WWN22, WLL^{+}23b, WWL^{+}25, CBZ11]. Adapting [LMZ20, ZZW^{+}25, CDH^{+}09]. Adaption [GLZ^{+}24, ZZL^{+}22]. Adaptive [BHMW07, Bul95, CJW^{+}21, DLY^{+}25, JMZ22, LK20, LJQY19, LFW23, LZC^{+}25b, LLZ^{+}25, SLZ^{+}25, WLL^{+}23a, XMS^{+}25, ZZT^{+}21, ZDZ23, ZYY^{+}24, BDS04, CCC09, LBS08]. Adaptively [DMUN19]. adaptivity [RSG09]. AdaTaskRec [ZDZ23]. added [MLDGH04]. Additive [DLN^{+}16]. Additivity [KSVS16]. Addressing [HZF^{+}23]. Adjusted [LCL^{+}17]. Adjustment [EMS21]. Ads [YTL25]. Advancements [DLL^{+}25]. Adversarial [CZX^{+}19, CYC^{+}24, CQL^{+}24, CHZ24, GLZ^{+}24, RYC^{+}20, SLD^{+}24, WLL^{+}23b, WZG^{+}23a]. Advertisement [ZLF^{+}17]. Advertising [ZGN^{+}17, ZWL^{+}22]. Affect [Fer17, MLK^{+}18, WZL^{+}19]. affected [WCS^{+}24]. Affects [MMO21]. affiliated [BM02]. after [Gla04]. Against [ZLF^{+}17, WLW^{+}24b, WLZ^{+}24, WZG^{+}23a, ZYL^{+}25, ZWL^{+}23]. age [Mar08a]. Agent [CSH^{+}25, HLL^{+}25, WFW^{+}25, XMS^{+}25]. Agents [PYF^{+}92, WZY^{+}25]. AgentTOD [XMS^{+}25]. Aggregate [ABCZ21]. Aggregated [AC16, AC19, CSZD15]. aggregates [MS01]. Aggregation [GCC12, LYM^{+}24, XC24, TP05]. Agnostic [THL^{+}25, WZL^{+}24, XZT^{+}24]. Agree [SZ21, BM02]. Agreement [DL16]. AI [DLL^{+}25, HLL^{+}25]. Algebra [FR97, GZC89]. Algorithm [BR96, BNIL24a, CGB19, EMS21, VB98, CL03, KT05, PMPJ11]. Algorithms [AZ12, FZSG21, GR95, Kon04, Pal15, SC19, ZZL^{+}15, ZLF^{+}23, BM12, DK04, PF03, TS10]. Align [HLL^{+}17]. Alignment [PGZ^{+}25, QZW^{+}25, WWN22, XZZ^{+}22, YHS^{+}12, ZZT^{+}21]. All-MLP [GZL^{+}24]. alleviate [HCZ04]. Alleviating [QDG^{+}24]. Allows [Man97]. Alphabet [MP20]. alphabets [Non13]. Am [LCY21]. Ambiguity [KC92]. Ambiguous [YDW22]. Among [Fer17, LM90]. Analyses [MC14b]. Analysis [ASSC22, BHH^{+}16, BB93, BRS92, BGI24, CGB19, CMR91, CSZD15, DBCJ21, EMG11, FCL^{+}23, HCD^{+}23, LLG^{+}23, LZG21, LGM^{+}24, MK16a, PPGK04, PR16, RABC22, SE91, VKD21, ZSQ^{+}24, ZJW^{+}24, ZH17, dODMS20, ACS08, Car12, CJRY12, CHS99, GWR99, GL06, LM01, LWA08, MPHS03, PR09, SC09, XC00]. Analysis-Based [PR16]. analyzing [SNS^{+}05]. Anchor [LCL04]. annotate [CDH^{+}09]. Annotation [SLZ^{+}19, XCG^{+}23]. annotations [AF07]. annotator [ZWRL19]. Answer [CCC^{+}12, DXL^{+}22, DLZ^{+}22, GCR^{+}21, HWN^{+}19, LYZ^{+}23, MT21, MAL16, KSNM10, MS01, Ack98]. answer-ranking [KSNM10]. Answering [AA16, BGQT07, GCR^{+}21, JEAZ25, LLY^{+}25, LLN22, LLM22, MAL16, PHK12, TM19, YXL^{+}23, ZWMH19, ZZC^{+}25, CTM12, CKC07, KSNM10, KEW01, Lin07, MPHS03]. Anytime [MPM22]. API [XMS^{+}25]. App [CHN^{+}17, HL17, TZP22, YZW^{+}17]. Application [EMS19, GPS93, GH96, MKF91, RN96, VAJR22, WSBC01, LCH07, SSS11, PGD07]. Applications [JTL^{+}17, SS93, WLS^{+}17, WRLD22, WBB^{+}90, ZLF^{+}17, ZXZL20, DDH^{+}09, FP00, HJZL09, dOTM01]. applied [ZL04]. Applying [HCZ04]. Approach [AA16, BLH^{+}22, BK90, CLL24, CFC^{+}19, CSZ^{+}22, FB91, GPS93, HLC^{+}20, KLHK13, Kwo95, LYN^{+}21, LL23b, MAP^{+}22, MHR22, MMLP97, NXF^{+}24, NWZC12, PBS22, QZX^{+}20, SLX^{+}24, TQH15, WXLC13, WHL^{+}16, WZW^{+}22, WGL^{+}24, WZZ^{+}24a, YSZ^{+}17, YHS^{+}12, YCZ^{+}15, ZXH^{+}25, ZZL^{+}15, AC08, ASST05, CdMRB01, CC05, Fuh99, GWLC06, GWR99, GP00, HHWL01, LM01, LCL04, MS01, Owe02, PZMRN05, WMJ05, YKHW06, ZFL^{+}07]. Approaches [CCC^{+}12, Gru92, MC14b]. Appropriate [ZLF^{+}23]. Approximate [ABB^{+}15, BNIL24a, PMD01, Web13, ARSZ03]. Approximation [LJQY19]. approximations [MOT11]. Apps [AZCC21, LAL^{+}17, USN^{+}17, WLG^{+}17]. Arabic [ANW17, AA16, DM07]. Architecture [AAL^{+}20, GZL^{+}24, JXY^{+}24, KL91, KEL95, PYF^{+}92, RFB^{+}24, WZHY24, PSDB99]. Architectures [WLY^{+}14, CML00]. Archival [ANW17, ZJT16]. Archives [CCC^{+}12]. Areas [HY20]. Arithmetic [MNW98]. Arnoldi [WW10]. Array [NCZG14]. Article [USN^{+}17, XWL^{+}25]. Article-Intervened [XWL^{+}25]. Articulating [KMdRS06]. Artifact [CR92]. ARTS [LXC^{+}25]. Ask [ZHRK23, GL05]. ASM [BR96]. Aspect [FCL^{+}23, GCH^{+}19, LLZ^{+}22, LLG^{+}23, LZG21, LLC^{+}25b, SWZ^{+}22, ZWX^{+}24, WZZ^{+}24b]. Aspect-Aware [LLZ^{+}22, LLC^{+}25b, SWZ^{+}22]. Aspect-based [FCL^{+}23, LZG21]. Aspects [MSMO18, WLT^{+}22]. Assessing [CVS21, LAO^{+}24, LA08]. Assessment [BOQ^{+}25, SQZ^{+}22, PTL08]. Assessments [MHR22, STZ22, STC^{+}24, XHSD24]. Assessors [FFM17]. Assimilation [ZXZL20]. Assimilation-contrast [ZXZL20]. Assistance [CWL^{+}20]. Assistant [FSE^{+}24, LWL^{+}25, MNP^{+}21b]. Assistants [AZCC21]. Assisted [BvdHFS25, CH96, LK20, ZWMH19, SBH24, ZK20]. association [TL03, WZZ04, YWYL08]. Associations [SLHS93, ZHY14, HZSL13]. Associative [COR13, HCZ04, RBV09]. Assumptions [Coo95, FP94]. assurance [WCX24]. Asymmetric [MP20, ZWE^{+}22]. Attack [WLZ^{+}24, ZWL^{+}23]. Attacks [TQN^{+}23, WFW^{+}25, WZG^{+}23a]. Attention [HWN^{+}19, LHL23, LLZ^{+}21, LWP^{+}21, WWX^{+}23, WZY^{+}24, XZZ^{+}22, YHS^{+}21, ZHL^{+}22, ZLS^{+}23, ZCZ^{+}23, ZJW^{+}24, ZWC^{+}25, ZXZL23, APJL20, GSG^{+}25, QDZW23]. Attention-based [LLZ^{+}21, ZLS^{+}23]. Attentive [CSR^{+}20, CLM^{+}22, DXL^{+}22, GCH^{+}19, GCN^{+}19, HLW^{+}24, PCCC22, PLR^{+}22, PLS^{+}24, ZLZ25a]. Attitudes [Ols89]. Attribute [CZL^{+}25, NTV22, SYSW24]. Attribute-level [SYSW24]. Attributed [FLMD22, MLZ^{+}20, ZWF^{+}23]. Attributes [DCT^{+}17, LLW^{+}21, MLZ^{+}20, ZYW^{+}23]. Attribution [Sav12]. Auction [APJL20]. Audio [HSH93, SCL^{+}22]. augment [BM12]. Augmentation [ALS^{+}24, DSW^{+}25, LGL^{+}24, LPZ^{+}25, XLX^{+}25, YSJ25, ZZC^{+}25, ZLH^{+}25]. Augmented [JWX^{+}25, LWL^{+}25, LLN^{+}25, WGZ^{+}25, XHX^{+}22, YRC23, ZHZ^{+}19, HLW^{+}24]. Augmenting [Ack98, TGS16]. Author [Ano93, Ano96a, Ano97, Cro98, RZCG^{+}10]. author-topic [RZCG^{+}10]. authority [LS02]. authorization [BJS99]. Authors [Ano94a]. Authorship [Sav12]. Auto [CRD16, VAJR21, WLL^{+}23b, YRC23, ZZY^{+}23]. Auto-Completion [CRD16, VAJR21]. Auto-Encoder [ZZY^{+}23, WLL^{+}23b, YRC23]. Autoencoder [XHX^{+}22, ZMZ25]. Automata [SFR98]. Automated [ADW94, BOQ^{+}25, KM97, WCP^{+}25, SSCT09]. Automatic [BJL^{+}07, GS91, LMW^{+}21, PBF^{+}89, QZS^{+}24, RBV09, CdMRB01, GWR99, GIS03, LSOS06, LWA08].

Automatically [ZZH^{+}18, YKHW06]. AutoML [ZQC^{+}23]. autoregressive [CJY23a]. AutoWeb [FP00]. Auxiliary [DYL^{+}20, LDW^{+}17, YRC23]. Availability [ZLW16]. Average [SWZ^{+}24]. Aware [AC18, BWS^{+}17, CWZ^{+}20b, CDZ^{+}23, CYC^{+}24, CWZ^{+}24, CS16, CSZ^{+}22, CFE^{+}25, GXTL14, GJS^{+}25, GCH^{+}19, LJHL17, LLZ^{+}22, LXW^{+}17, LHA^{+}25b, LLC^{+}25b, MHLH23, MJH15, PLR^{+}22, SLX^{+}24, SWZ^{+}22, TWS^{+}25, TQH15, WGL^{+}24, WQL^{+}19, WMM^{+}23, YZW^{+}17, YYY^{+}24, YZC^{+}25, ZKJW22, ZGCC18, ZWE^{+}22, AZCC21, CXZ^{+}21, CHZ24, DXL^{+}22, HML^{+}23, IG08, JLC^{+}25, LXZ^{+}22, LHL23, LZZ^{+}23, LMZ23b, LZG21, SDZW24, SJZ^{+}25, WLFT18, WLT^{+}22, WWHQ25, WCZ^{+}25, WLM^{+}21, YWQ^{+}24, YLW^{+}23, YCZ^{+}15, ZWRL19, ZZD25, ZWC^{+}23, ZWX^{+}23, ZCLC23, FFM17]. Awareness [YLZW21]. Axiomatic [Gin24]. AZFin [SC09].

Backdoor [WLZ^{+}24]. Badge [WHFG92]. Bag [RDK24]. balance [PSPBY10]. balanced [YZZS24]. Balancing [YCL^{+}23]. Bandits [JMD20]. Base [ZMK^{+}20, ZKMM23]. Based [APCC15, AKFA15, AK18, AOK14, AA16, BCFG15, BC14, BA18, BLH^{+}22, CGB19, CE89, CFP95, CW94, CLL24, CWZ^{+}20a, CSH^{+}25, CDM15, DSW^{+}25, EMG11, Fuh89, GPS93, GCC12, GR96, GYC^{+}22, HLH^{+}25, IC16, JVA^{+}15, JWX^{+}25, Koi95, Kwo90, LCJ^{+}16, LLT^{+}21, LLZ^{+}22, LSG^{+}25, LYP^{+}25, LLM22, LPY94, LCL^{+}17, LGQL19, LZC^{+}24c, LCJ^{+}19, MWZ^{+}25, MAP^{+}22, MNP^{+}21b, NWZC12, PPPR13, PGW^{+}17, PR16, QLNY19, QHCY22, QFZ^{+}19, Res93, SC19, Sav12, SCL^{+}22, SZS17, SE91, SSGC89, SBH24, SF89, SFR98, SAW^{+}25, SLW^{+}20, TBC96, Tuz95, VHJC11, WWX^{+}24, WWX^{+}25, WLY^{+}14, WZW23, WGL^{+}24, WZY^{+}24, WWT^{+}25, WZHY24, Won94, XYW^{+}23a, XC98, XPL^{+}24, YW15, YLZW21, YC94, YHY^{+}17, YDXW21, YMZF22, ZLL^{+}22, ZKMM23, ZLW24, ZZL^{+}15, ZZZ^{+}16, ZLRW24, ZWE^{+}22, ZQH^{+}24b, ZSLK24, dMdSsdA^{+}08, ASSC22, ADCU08, AV02, AVA^{+}24, BZCE22, BBD11, CC01, CDH^{+}09, Car12, CZJ07, CL06, CCC09, CGD^{+}25, CSLN10]. based [CC05, Coh00, DGMG23, DK04, DDH^{+}09, EKK99, EKK00, FBN^{+}12, FCL^{+}23, FCD02, FG04, HYZ^{+}23, HLW^{+}24, HHWL01, HZS^{+}25, HHC^{+}10, IH07, JK02, KNS04, LAO^{+}24, LBS08, LBM^{+}24, LLBS18, Lin07, LHSC11, LLZ^{+}21, LZG21, LL23b, LLZ^{+}25, MPS07, MRC^{+}23, MC10b, Owe02, PCC^{+}22, PCCC22, PTKJ07, PZMRN05, QZXP06, QZW^{+}25, QHLY20, RFB^{+}24, SZ21, SZ09, TM10, TQN^{+}23, TCQ^{+}25, WLW^{+}24a, WZY^{+}25, WS90, WMJ05, WCX24, WWN^{+}25, XTF^{+}21, XLL^{+}23, XHW^{+}19, YHWZ24, YFSO24, YKHW06, ZHL^{+}22, ZZT^{+}21, ZSY^{+}21, ZHW^{+}22, ZCZW22, ZLS^{+}23, ZZY^{+}23, ZXM^{+}24, ZJK^{+}24, ZSZ^{+}24, ZHD^{+}22, ZKG08, ZRZ02, ZWL^{+}23, ZK20, dOTM01, RCR^{+}21]. Baseline [AA16]. Baselines [KSVS16]. Bases [ZWMH19]. basic [TS10]. Basis [RL94, Mel08]. Basket [LJAdR23, WC21a]. batch [HJZL09]. Bayesian [JLC^{+}25, LCJ^{+}19, SLW^{+}20, WZW^{+}22, WZZ^{+}24a, YCZ^{+}15, ZXZL23]. BCDR [ZLS^{+}23]. Be [FC17, LAO^{+}24]. Before [LLC^{+}25b]. Behavior [AC16, CLF^{+}25, HL17, LZC^{+}25a, LHH^{+}19, LCPD19, LMZ20, MYL^{+}23, MZG^{+}24, MBST17, OS96, SSY^{+}17, WZY^{+}25, XYW^{+}23b, YCG^{+}24, YCL^{+}25, ZLM^{+}23, ZLZ^{+}25b, YPYM24]. Behavioral [RSD^{+}13]. Behaviors [CCAC24, HCD^{+}23, QZS^{+}23, YCZ^{+}16]. Behaviour [BACF17, SBM^{+}17]. Belief [WH15, ZM14, LBS08]. belief-revision-based [LBS08]. Benchmark [LLN^{+}25, WCZ^{+}25]. Benchmarking [AOK14, TCQ^{+}25, WSBC01]. Beneath [KL00, PR04]. Benefit [LDX^{+}25]. Bennett [Gru92]. BERD [SYSW24]. Bernoulli [LA08]. Best [MAL16, SW90, ZOML24]. Best-Match [SW90]. Better [CCAC24, ZAK^{+}23]. between [DLA22, MTKY92]. Beyond [DXZ^{+}22, LGL^{+}24, ZTZL25]. Bi [ZXM^{+}24]. Bi-preference [ZXM^{+}24]. Bias [CDW^{+}23, FN96, LAO^{+}24, LFW23, MAP^{+}22, MMCC^{+}21, Oos23, PHK12, SLX^{+}24, WZL^{+}24]. Bias-Agnostic [WZL^{+}24]. biased [MZ08]. Biases [LCL^{+}23b, WH15]. Bidding [ZZW^{+}25]. Bidirectional [MTKY92, WLL24, ZLL^{+}20]. Big [DCT^{+}17, SSY^{+}17]. Bilateral [ZDW^{+}22]. Billion [YMZF22, ZNL^{+}17]. Billion-Edge [YMZF22]. Billion-Scale [ZNL^{+}17]. Binarized [CQZ^{+}25a]. Binary [MNZ19]. biomedical [YKHW06]. Black [MDX^{+}25, WLZ^{+}24, WFW^{+}25, WZG^{+}23a, ZWL^{+}23]. Black-Box [MDX^{+}25, WLZ^{+}24, WFW^{+}25, WZG^{+}23a, ZWL^{+}23]. Block [CWZ^{+}20b, CHS99]. Block-Aware [CWZ^{+}20b]. Blocks [FC91, LPC^{+}23]. blogosphere [CSJS10]. Bloom [AL13]. Blueprint [JS24]. Boies [Gru92]. Boolean [CGMP99]. Boosting [BMMC19, CJZ^{+}16, FPSO06, CBZ11]. BoRe [LMZ20]. Bot [PZH^{+}24]. both [WZZ04]. Bots [MTC21, ZWL^{+}22]. BotSpot [ZWL^{+}22]. bound [MOT11]. Bounded [HSS^{+}09]. Bounding [LCL^{+}23b]. Box [MDX^{+}25, WLZ^{+}24, WFW^{+}25, WZG^{+}23a, ZWL^{+}23]. Brain [YXA^{+}24]. branch [YZZS24]. Brand [LZW^{+}20]. Breaking [MK16a, SLM^{+}24, SC09]. Bridging [BNIL24b]. broadcast [LSOS06]. Broker [TGL^{+}97]. Bronze [STC^{+}24]. Brotli [AFF^{+}19]. Browse [LMW^{+}14]. Browse-to-Search [LMW^{+}14]. Browsing [CS94, SF89, SFR98, Yan15, BKGM^{+}02, MC06]. Bubbles [GWL^{+}24]. Budget [PBS22, ZAL^{+}16]. Budgets [NGMD17].

Building [DWLW09, HZG20, LWL^{+}25, MRYGM01, RP98, VL90, WW94, YLZW22]. Bundle [MHW^{+}24]. Burst [HZW02]. Bursting [GWL^{+}24]. business [KL00]. Buttons [Bie92].

C [PBF^{+}89]. C-TODOS [PBF^{+}89]. Cache [WLY^{+}14]. caching [PSPBY10, FPSO06]. CAFE [LZC^{+}25b]. Calling [XMS^{+}25]. CAME [GCB^{+}25]. Can [CHZ24, FC17, FEL22, LDX^{+}25]. cancer [SNS^{+}05]. candidate [AL13]. capacity [PSPBY10]. Capsule [LLZ^{+}22, PBS22, ZY21]. Captions [GR96]. Capture [CJY23a, RS97]. Capturing [HSH93]. card [BSST03]. Career [ZSQ^{+}24]. Carlo [LCL^{+}17]. Cascading [TQH15, YCG^{+}24]. Case [HE91, JSL^{+}24, LHA^{+}25a, LHA^{+}25b, MWA^{+}24, SWL^{+}23, SLW^{+}24, SAW^{+}25, WW94, XWL^{+}25, ZZG^{+}25, Luz12, Gul96, MC10b]. casual [LLC03]. Cat [BWS^{+}17]. Cat-Mobile [BWS^{+}17]. Categorical [QFZ^{+}19, ZHL^{+}22]. Categories [CJZ^{+}16, BBD11]. Categorization [ADW94, LPY94, YC94, CS99, GWLC06]. Category [CCC^{+}12, CZX^{+}22, LLG^{+}23, XZY21]. Category-Level [XZY21]. CATS [BZCE22]. Causal [GZW^{+}24, HZF^{+}23, LGL^{+}24, SSZ^{+}23, WLW^{+}24b, WQH^{+}25, ZJT16, ZMN^{+}24, ZWZS25]. cause [CJY^{+}23b]. Causes [LZC^{+}24b]. CBMIR [APCC15]. CD [KBD89]. CD-ROM [KBD89]. center [LQZ^{+}25]. centric [LFW23, WXZ^{+}25]. Certified [HNN^{+}25]. CETN [LSZ^{+}25]. CHA [ZHL^{+}22]. chains [AL13]. Challenges [CWM^{+}24, CCAC24, HZG20, HYM^{+}25, WWHX23]. Change [CZZ^{+}17, YGZ15]. changed [LHSC11]. Changes [MK16b]. Channel [GJS^{+}25, LPLH19, WWX^{+}24, SLZ^{+}24, WWX^{+}23]. Chao [BvdHFS25]. Characteristics [CL18, Guy18, JS24]. Characterization [TZP23, CSJS10]. Charter [Ano94b]. Chat [WLM^{+}21]. Chatbot [WAL^{+}21]. Chatbots [LLT^{+}21]. Check [LJAdR23, SLZ^{+}24, YCZ^{+}16]. Check-in [SLZ^{+}24, YCZ^{+}16]. Checking [SFR98]. Chemical [SMGM11]. Chief [Zha21]. Chimera [ATW00]. Chinese [LLN^{+}25]. Choice [RVA^{+}24]. Chronological [CLG^{+}23]. CIRS [GWL^{+}24]. Citation [LLZ^{+}22, XZH^{+}22]. City [SLQ24]. CL [YZC^{+}25]. Claim [WQH^{+}25]. Claims [CR92]. CLAIRE [TMT06]. Clarification [TTZ^{+}25, ZAK^{+}23]. Clarifying [LHA^{+}25a, YDW22, ZAK^{+}23]. Class [COR13, PYC^{+}19, YPYM24, CLPM20, GCT99]. Classical [Ruh94]. Classification [CZZ^{+}17, CZJH24, COR13, DG95, ES13, EMS19, HQZ17, IG08, LCX^{+}19, LYY^{+}24, LBC^{+}19, MMO21, MPS23, PR16, RL94, WZHY24, YHS^{+}21, YZZS24, ZWRL19, ZZD^{+}20, ZYY^{+}24, ACS08, BJL^{+}07, GIS03, LWSA10, PS05, SPS^{+}06, TY22, TMT06]. Classification-aware [IG08]. classifier [GWLC06]. Cleaning [ES13]. ClearBoard [IKG93]. Click [DRdR23, LSZ^{+}25, LXW^{+}17, MCL^{+}12, Oos23, QZS^{+}23, SLZ^{+}25, ZML^{+}19, LWSA10]. Click-Through [LSZ^{+}25, SLZ^{+}25, QZS^{+}23]. clicks [JGP^{+}07]. clip [HSS^{+}09]. Clipping [USN^{+}17]. Close [OOSC93, YWYL08]. Cluster [LGRK18, LZC^{+}24c, ADCU08, HHC^{+}10]. Cluster-Based [LZC^{+}24c, ADCU08, HHC^{+}10]. cluster-skipping [ADCU08]. Clustered [PV17]. Clustering [Can93, JDA^{+}22, LZC^{+}25a, LRZ^{+}17, TM10, CL06, KM13, MLDGH04, TG10, WNZ02, ZFL^{+}07]. Clustering-based [TM10]. Clusters [KL09]. CLV [FC91]. Co [FLMD22, GSG^{+}25, LLM22, PCCC22, TWS^{+}25, XZCZ25, GZW^{+}18]. Co-attention [GSG^{+}25]. Co-Attentive [PCCC22]. Co-Embedding [LLM22, FLMD22]. Co-Evolution [TWS^{+}25]. Co-Relation [XZCZ25]. Coalition [RP98]. Coalition-Building [RP98]. Coarse [LMW^{+}24, MZG^{+}24]. Coarse-grained [LMW^{+}24]. Coarse-to-Fine [MZG^{+}24]. Code [ZLH^{+}25]. Coding [LWL^{+}25, MNW98, MP20]. Coexisting [XZY21]. Cognition [LWC^{+}25]. Cognitive [CACW23, VDR25]. Coherence [AC16, XZD^{+}25]. Coherence-guided [XZD^{+}25]. coherent [JZ06]. Cold [CQF^{+}23, HYZ^{+}23, HLH^{+}25, HCC^{+}16, HSYX22, LLW^{+}21, SZC^{+}23, SZS17, WGL^{+}24, WLL^{+}23b, WWL^{+}25]. Cold-Start [CQF^{+}23, HLH^{+}25, HCC^{+}16, SZS17, WGL^{+}24, HYZ^{+}23, HSYX22, LLW^{+}21, WLL^{+}23b, WWL^{+}25]. Collaboration [CSB21, TWS^{+}25]. Collaboration-Aware [TWS^{+}25]. Collaborative [AK18, CCCD19, CLPM20, CJW^{+}21, CQL^{+}24, CQZ^{+}25a, DHC^{+}25, DHY^{+}19, GXTL14, HWL^{+}24, LHL23, LCW^{+}24b, LCYY25, LZW^{+}20, LXG^{+}23, LZC^{+}24c, LCNY23, LLP24, PYC^{+}19, PCC^{+}22, QJG^{+}25, RR98, SWZ^{+}24, SYX^{+}17, TM91, WHL^{+}16, WRLD22, WLL^{+}23a, WZZ^{+}24a, WLR^{+}25, XHX^{+}22, XZH^{+}22, XLL^{+}23, XHW^{+}19, YPYM24, YLYT23, YRC23, ZLL^{+}20, ZMZ^{+}24, ZSL^{+}25, ZZSS25, HKTR04, Hof04, HCZ04, IH07, WdVR08, XHYY09]. Collection [CFGM02, CFFK22, FKS19, PF03]. Collections [Gla97, KC15, ZJT16, BC13, CW02]. Collective [LGQL19, MYL^{+}23, ZZT^{+}21]. Combination [FP94, CRG02]. Combine [FFM17, MO12, MS01]. combined [WZB04]. combinFormation [KKS^{+}08]. Combining [AOU12, MC10a, WL22]. Comments [HLL^{+}17]. Commerce [DLZ^{+}22, GCR^{+}21, JSL^{+}24, KM97, ZLH^{+}22, ZLW24]. Communicating [LM90]. Communication [Mar94, CSJS10]. communicative [YHYO01]. Communities [LLM22, LZLM19, Pal15]. Community [CCC^{+}12, NXP^{+}24, PHK12]. Compact [HZS^{+}25, LZC^{+}25b, ZNL^{+}17, Zob06]. comparable [TLJ^{+}07]. Comparative [CSZD15, ZPW^{+}24]. Comparing [ANW17, Guy18, IC16, PF03, PS05]. Comparison [DLA22, HWD13, OOSC93, TFS^{+}12]. Competitive [ZLH^{+}22]. Competitively [GCB^{+}25]. Complementary [ZCZW22]. Complete [MS01, PTL08]. Completion [CRD16, ZZZ^{+}25, VAJR21]. Complex [PL94, ZHW^{+}22, CL03, PVGM11]. Complex-valued [ZHW^{+}22]. Complexity [CWL^{+}20]. Component [ASSC22, Gul96, Kwo90]. Component-based [ASSC22]. Components [Kwo90]. Composed [GJS^{+}25]. composition [KKS^{+}08]. Comprehend [DHC^{+}25]. Comprehensive [DLL^{+}25, GCF^{+}22, LWP^{+}21, LLN^{+}25]. Compressed [BCPN14, Man97, PGW^{+}17, ADCU08, SNZBY00]. Compression [BK90, BKR97, KBD89, Man97, PGW^{+}17, ZZL^{+}15, BFNP10, CW02, KT05]. Compressor [AFF^{+}19]. Computational [TY22]. computationally [GCT99]. compute [GGMW03]. Computer [All91, FN96, HE91, Jac91, Ols89, SE91, SZ93]. Computer-Human [All91]. Computerized [GH96, SM95]. Computers [FZC93]. Computing [CDF01, Ruh94, WZC^{+}17, WW10]. Concatenate [XZCZ25]. Concept [EMG11, FGM^{+}02, Owe02]. Concept-Based [EMG11, Owe02]. Concepts [BF98, CL03, FG04]. Conceptual [Gul96, ISW92, PBF^{+}89, RR98]. concerns [WH09]. Conclusions [GTW^{+}18]. Concordances [BKR97]. Conditional [LQJ^{+}24]. Conference [Cro89b].

Confidence [MMO21, Web13, WKF^{+}12]. Confident [RLC^{+}25, WXR^{+}25]. Configurations [DMUN19]. Confounder [ZWZS25]. Confounding [HZF^{+}23]. Confusion [WLR^{+}25, XWZ^{+}25]. Congestions [WZC^{+}17]. Connecting [GZW^{+}18]. Considerations [KBD89]. Consistency [Gru92, RZM24, WNS^{+}17, Wie92, ZLH^{+}25]. Consistent [CJY^{+}23b]. constant [Non13]. Constituent [MH89]. Constrained [LLM22, PBS22, ZZW^{+}25, WLL24]. Constraints [MT21, WC21b, ZLW16, PVGM11]. Constructing [ZML^{+}19]. Construction [LTC^{+}25, NCZG14, WWT^{+}25, Yan15, BC11]. Consumer [ZLW24]. Consumption [LMZ20]. Content [AK18, AOK14, BOQ^{+}25, DG95, LCPD19, MK16a, NB97, NWZC12, RS93, SSS11, WWHQ25, ZLL^{+}22, ZZG^{+}25, AF07, CDH^{+}09, EKK99, EKK00, KL06, KNS04]. Content-Based [AK18, AOK14, NWZC12, EKK99, EKK00, KNS04]. content-oriented [KL06]. Contents [CS94, EGL^{+}97]. Context [AC18, AZCC21, BB93, BLWJ15, BWS^{+}17, CML^{+}21, CDZ^{+}23, CS99, FTZZ23, GBMS99, LXZ^{+}22, MJH15, SJL^{+}25, TGS16, UY89, VAJR21, VAJR22, VR24, WGL^{+}24, WGZ^{+}25, WQL^{+}19, XC24, YCZ^{+}15, ZKJW22, FGM^{+}02, LBS08, MC06, Mel08, XC00]. Context-Aware [AC18, BWS^{+}17, CDZ^{+}23, MJH15, WGL^{+}24, WQL^{+}19, ZKJW22, AZCC21, LXZ^{+}22, YCZ^{+}15]. Context-sensitive [CS99, LBS08]. Contexts [SGJ22, LLC03]. Contextual [BTP24, BKWZ15, DL16, JMD20, LL23b, MZLK11, MYHL11, RABC22, SYX^{+}17, ZYY^{+}21, ZTZL25, ASST05]. Contextualized [DXL^{+}22, YZS^{+}24, BHMW07]. Contextualizing [KYK24]. Continuity [RRS95]. Continuous [GRSD^{+}16].

Continuous-Time [GRSD^{+}16]. Contract [WWR^{+}25]. Contrast [LSZ^{+}25, WMC^{+}23, ZXZL20]. Contrast-enhanced [LSZ^{+}25]. Contrastive [CZJH24, JZZW24, LZC^{+}25a, LZL^{+}25, LLC^{+}25a, LQZ^{+}25, MHW^{+}24, WLW^{+}24a, WZZ^{+}24b, XYCC24, XLX^{+}25, YOM24, YHDN25, YZC^{+}25, ZZZ^{+}24, ZDZW23, ZWY^{+}24, ZYL^{+}25]. contribution [WW10]. Control [Bul95, Gla97, HLNS98, BFF^{+}03, Mar08a]. Controlling [Koi95]. Conversation [CSH^{+}25, FSE^{+}24, LFL^{+}25, SLW^{+}20, TCMC21, ZLWW22, ZGF^{+}20]. Conversational [BWS^{+}17, CSB21, DZX^{+}23, DLL^{+}25, FEL22, HKS^{+}21, KMPS21, KYK24, LLW^{+}21, LLN22, LYN^{+}21, LCY21, LZW21, LPZ^{+}25, LHA^{+}25a, LHA^{+}25b, NXF^{+}24, RYC^{+}20, RCR^{+}21, SAD24, TCMC21, VKD21, VAJR21, WA22, WZC^{+}25, ZJW^{+}24, ZGL^{+}25, ZHRK23, ZSLK24]. Conversations [LCL^{+}23a, RCR^{+}21, WZW23, YLZW21]. Conversions [ZZW^{+}25]. Converting [Rad92]. Convolution [MHLH23]. Convolutional [CXZ^{+}21, DHY^{+}19, HML^{+}23, PSM24, WL22, YCG^{+}24, ZLZ25a, ZDW^{+}22]. Cooccurrence [XC98]. Cook [FEL22]. Cooking [FEL22, FSE^{+}24]. Cooking-Related [FEL22]. Cooperative [FLMM91, MLF95, Won94]. coordinate [HSS^{+}09]. Copy [AOK14, TQH15]. CORE [EGL^{+}97]. corpora [RZCG^{+}10]. Corpus [KLHK13, XC98, ZWX^{+}24, GWR99, TLJ^{+}07, YKHW06]. Corpus-Based [XC98, YKHW06]. correct [PTKJ07]. Correcting [Oos23]. correction [DM07]. Correlation [Fer17, LHSC11, PR09, PZMRN05]. Correlation-based [LHSC11, PZMRN05]. Correlations [DHY^{+}19]. Correspondence [SLM^{+}24]. CoSam [CJW^{+}21]. Cost [GXTL14, MAH24, MCF^{+}17]. Cost-Aware [GXTL14]. Cost-Effective [MCF^{+}17]. Costs [OSV19]. Count [LLZ^{+}22]. Counterfactual [CWZ^{+}24, GWL^{+}24, LGL^{+}24, LSG^{+}25, SWZ^{+}24, WLY^{+}24, YSJ25]. Countermeasure [GYW^{+}25]. Countermeasures [YYY^{+}24]. Coupling [YZH^{+}24, ZZH^{+}25]. Couplings [WC21a]. Course [WZW^{+}22, YZS^{+}24, PVGM11]. Coverage [EJ03]. Covering [ABB^{+}15]. COVID [FWG^{+}24]. COVID-19 [FWG^{+}24]. CPWS [RLC^{+}25]. crawl [PS05]. crawler [AAGY01]. Crawlers [FAZC12]. crawling [TM10]. Creating [TLJ^{+}07]. Creative [YFO^{+}25]. CrimeNet [XC05]. criminal [XC05]. Criteria [Row96]. Criterion [BvdHFS25]. Critical [EMS21, JSZL23, NCZG14, RJB89]. criticism [TL03]. Critiquing [FLMM91]. Cross [CHN^{+}17, CZT^{+}23, CJN20, EMS19, GLG^{+}23, GLZ^{+}24, GSG^{+}25, HYH15, HLH^{+}25, HWL^{+}24, JTS16, LHL23, LQJ^{+}24, LGZ^{+}24, LGY^{+}25, MXM^{+}24, MPS23, QHLY20, SLD^{+}24, SLQ24, TXC^{+}24, TL14, WLZ^{+}24, WWR^{+}25, WZW^{+}25, WYZ^{+}25, XZD^{+}25, XZZ^{+}22, XLW^{+}22, XSWX25, XYCC24, YOM24, ZZL^{+}23, ZZZ^{+}24, ZBYH23, ZYC^{+}24, ZMN^{+}24, ZZG^{+}25, ZWL^{+}23, ZPW^{+}24, ZWZS25, GNN^{+}10, PTS13, TLJ^{+}07]. Cross-City [SLQ24]. Cross-Device [HYH15]. Cross-Domain [LHL23, WWR^{+}25, WZW^{+}25, WYZ^{+}25, XSWX25, XYCC24, ZYC^{+}24, ZWZS25, GLG^{+}23, GLZ^{+}24, GSG^{+}25, HLH^{+}25, LGZ^{+}24, LGY^{+}25, MXM^{+}24, SLD^{+}24, TXC^{+}24, XZD^{+}25, YOM24, ZZL^{+}23, ZZZ^{+}24, ZMN^{+}24, PTS13]. Cross-Language [TL14, TLJ^{+}07]. Cross-Lingual [EMS19, JTS16, MPS23, XZZ^{+}22, GNN^{+}10]. Cross-Modal [HWL^{+}24, WLZ^{+}24, CJN20, XLW^{+}22, ZBYH23, ZWL^{+}23]. Cross-Model [ZPW^{+}24]. Cross-Platform [CHN^{+}17, LQJ^{+}24]. Cross-session [QHLY20]. Cross-Site [ZZG^{+}25]. Crowd [RLS^{+}24]. Crowdsourced [MHR22, XHSD24]. Crowdsourcing [BWS^{+}17, MMST17, RLS^{+}24, SC19, ZDZ23]. CRS [NXF^{+}24]. CRSAL [RYC^{+}20]. CRUD [LLN^{+}25]. CRUD-RAG [LLN^{+}25]. CSCW [Res93]. CTR [FLW^{+}25, SZC^{+}23]. Cues [LXZ^{+}22]. Cultural [GGRF23]. Cumulated [JK02, KL06]. Curation [HCD^{+}23]. Curious [BWS^{+}17]. Current [Ols89]. Curriculum [WZZ^{+}23]. Cursor [APJL20]. Customer [PR16, ZWT^{+}23]. Customizable [BZCE22]. Customizing [RS97]. Cyberchondria [WH09]. cyberspace [AC08]. Cycle [CR92]. Cyclic [NXP^{+}24].

D [NCZG14]. D-Critical [NCZG14]. DA-DAN [GLZ^{+}24]. DAIDA [JMSV92]. DAN [GLZ^{+}24]. Dangerous [WLZW21]. Danish [ANW17]. Data [AFF^{+}19, AARC20, ALS^{+}24, CST^{+}23, Coh00, DYL^{+}20, DHC^{+}25, FC17, FCL^{+}23, GC92, HCD^{+}23, JSZL23, JSL^{+}24, VST^{+}14, LGL^{+}24, MTKY92, NFZ19, QFZ^{+}19, RLC^{+}25, SLHS93, SCL^{+}22, SSZ^{+}23, SSY^{+}17, SYX^{+}17, SYSW24, TGL^{+}97, Tom89, WZC^{+}17, WS90, YLW^{+}23, YSJ25, ZWRL19, ZZC^{+}25, ZLH^{+}25, ZZZ^{+}16, BGQT07, BJS99, CDF01, CGM02, FPSO06, HZW02, KZP10, PSDB99, QZXP06, ZCD06]. Data-Driven [HCD^{+}23]. Data-to-text [JSL^{+}24]. Database [BBT92, FR97, ISW92, KBGW91, KN93, SÖ90, Tom89, VB98, BFF^{+}03, CC03, Fuh99, IG08]. Databases [CH96, KBGW91, KC95, MK94, NB97, BSST03, CC01, EKK99, EKK00, GWR99, GIS03, KZSD99, LL99, MS01, SSCT09]. Dataset [LCL^{+}23b]. Datasets [PV19, CL06]. DBface [KN93]. DBpedia [LCL^{+}17]. DBpedia-Based [LCL^{+}17]. De-Noised [MWZ^{+}25]. dead [Gla04]. Debias [CDW^{+}23]. Debiased [LZC^{+}24b, LWC^{+}25, YCL^{+}23]. Debiasing [ZXZL20]. Decaestecker [EMS21]. Decentralized [LCNY23]. Decision [ASV91, ADW94, LSG^{+}25, Won94, YGZ15, ZLW24, ZZW^{+}25, BDS04, Fuh99]. decision-theoretic [Fuh99]. Decisions [LGL^{+}24, WLWK08]. Decomposition [VFA17, YHDN25, KO98]. Deconfounding [ZWZS25]. Decoupled [HLX^{+}24, SGLS93]. Decoy [CLF^{+}25]. Deductive [KC95]. Deep [AAL^{+}20, CQL^{+}24, FZSG21, HLW^{+}24, JLC^{+}25, LMZ^{+}23a, LW20, LLZ^{+}21, LZG21, LBZ^{+}24, LCJ^{+}19, RCR^{+}18, SSY^{+}17, SLW^{+}20, SJZ^{+}25, TM19, WLZ^{+}24, XLW^{+}22, XTF^{+}21, XHW^{+}19, YW15, YZH^{+}24, ZQC^{+}23, ZWL^{+}22, dODMS20]. DeepMob [SSY^{+}17]. Defined [BBT92, Gin24]. definitional [CKC07]. Delay [GC92]. Delayed [ZZW^{+}25]. delivery [LSOS06]. Demographic [DCT^{+}17]. Denoising [DSW^{+}25, PGZ^{+}25, SWZW25, YLYT23, ZJK^{+}24, ZMZ25]. Dense [BNIL24b, LCW^{+}24a, LMZ^{+}23a, LL23a, ZOML24, ZLRW24]. Dependable [DC10]. dependencies [BSV09]. Dependency [BLWJ15, QWZ^{+}15, YW15]. dependent [CDR10, MSOH13, TFS^{+}12]. Derivation [GH96]. Deriving [LAL^{+}17]. Description [CSH^{+}25]. Descriptive [GR96, NZW^{+}14]. Design [BB93, CMR91, CR92, CW94, ERG^{+}89, GPS93, GR95, ISW92, IKG93, LL89, Mos90, OOSC93, PBF^{+}89, SZ93, WLY^{+}14, AAGY01, BSST03, FCD02, GWLC06, IM05, RR98]. Design-Evaluation [ERG^{+}89]. Designing [KN93]. Designs [HSH90]. Desktop [WDG17, CDZ08]. Detachment [YTD13]. details [ES98]. Detect [ZAL^{+}16]. Detecting [AZK12, LHG^{+}12]. Detection [AOK14, BA18, CSR^{+}20, CTH^{+}13, LYP^{+}25, LW20, MTC21, MCF^{+}17, PZH^{+}24, QWZ^{+}15, TQH15, WQH^{+}25, WNWL^{+}21, ZJT16, ZZL^{+}21, ZLH^{+}22, ZY21, ZWL^{+}22, Zob06, AC08, CCC09, CFGM02, TFS^{+}12, WKF^{+}12].

Developing [WBB^{+}90]. Development [Cro89b, Gru92, RN96, SS93, WL97, ATW00, FP00]. Device [HYH15, XYW^{+}23a, RSG09]. Devices [AK18, CMR91, WLS^{+}17, ACM^{+}02, BKGM^{+}02]. DGEKT [CYZ^{+}24]. DGeye [WLZW21]. DHyper [TCSL24]. Diagnostic [FTZ11]. Dialog [CSJ^{+}24, GCL^{+}21, HZG20, WCS^{+}24, CDH^{+}09]. dialog-based [CDH^{+}09]. Dialogue [CSH^{+}25, HZS^{+}25, LMW^{+}21, LCW^{+}24a, LMW^{+}24, LLT^{+}21, LFL^{+}25, MLZ^{+}22a, MLL^{+}24, SZZL23, SGZ^{+}24, WLL24, WGZ^{+}25, XMS^{+}25, ZHZ^{+}19, ZCZ^{+}23]. Dialogues [FTZZ23, PYDR24, VKD21, XTF^{+}21]. Dichotomy [GGRF23]. Dictionary [LPY94]. Difference [JYM13]. Different [CWL^{+}20, LM90, IH07]. differentials [TkWW99]. Difficulty [CFFK22, ZZH^{+}18]. DiffuRec [LSL24]. Diffusion [CYC^{+}24, DSW^{+}25, GRSD^{+}16, LSL24, LLZ^{+}25, QWJ^{+}24]. Digital [CDM15, EGL^{+}97, GR12, MMO21, MNP^{+}21b, SAK^{+}23, SMGM11, VR24, WBP^{+}16, AF07, Gla04, GL05, GFWK04]. Dimension [ACC18, DHY^{+}19, Koi93, RBM23]. Direct [HSH90, WKF^{+}12, MK94]. Directed [SHS^{+}23, ZHC^{+}20, ZLZ^{+}25b]. Direction [ZWE^{+}22]. Direction-Aware [ZWE^{+}22]. Directions [CCXZ25, CDW^{+}23, GZW^{+}24, ZZL^{+}23]. Directly [Man97]. directory [GKL^{+}07]. DirichletRank [WTS^{+}08]. disambiguating [YKHW06]. Disambiguation [MKZPK18, SG14]. Disclosure [CST^{+}23]. Discourse [LFL^{+}25]. Discovering [LOP^{+}13, SMT25, ZFL^{+}07]. Discovery [JTS16, OSV19, AAGY01, KKS^{+}08, LLPT10, XC05]. Discrete [YPYM24, KZP10, PRP05, QZXP06]. Discriminative [CQZ^{+}25b, MWZ^{+}25, QWZ^{+}15]. Discs [FC91]. discussions [Luz12]. Disentangled [GLG^{+}23, LLC^{+}25a, MLZ^{+}22b, WLW^{+}24b, WCP^{+}25, ZWC^{+}25]. Disentanglement [LFL^{+}25, SJL^{+}25, XZD^{+}25, ZLS^{+}23, ZWZS25]. Display [Koi95]. Displays [HM90]. Distant [SHS^{+}23]. Distantly [LYLS22]. Distillation [CZX^{+}19, FTZZ23, HLX^{+}24, HZS^{+}25, JWX^{+}25, LYLS22, ZCLC23, WMGZ24]. Distilled [KSYE24]. Distinctions [SAW^{+}25]. Distinguish [XWZ^{+}25]. Distorted [VB98]. Distributed [CCXZ25, CH96, HEA90, KEL95, KBGW91, KNS04, CML00, DDH^{+}09, KM13, LL99, MRYGM01, RSG09]. Distribution [Cum14, DSZ^{+}23, HCC^{+}17, LK20, LZWZ24, WWN22, ZMZ^{+}24, ZWC^{+}25]. Distributional [DRdR23, YWQ^{+}24]. Distributions [PSL16, SV99, XC13]. Distrust [FMSS14]. divergence [AV02]. Diversification [CRC^{+}22, DDSW24, HY20, LYS^{+}17, MSMO18, QDZW23, SDZW24]. Diversified [IYC^{+}23, WDL^{+}24]. Diversifying [CRD16, SRF^{+}24]. Diversity [QDZW23, RBCT14, SZ21, SLX^{+}24, WZL^{+}19, YLZW21, YFSO24, ZLW16]. Diversity-Novelty-Aware [SLX^{+}24]. Divide [ZYC^{+}24]. Divide-and-Transfer [ZYC^{+}24]. DNA [KT05]. Do [RLS^{+}24]. Document [ABB^{+}15, ALS^{+}24, CGB19, CFP95, CDZ^{+}23, CDM15, Cum14, CPL15, DDSW24, DMUN19, FB91, HYLC24, Kwo90, LR96, LCX^{+}19, LPC^{+}23, LYM^{+}24, LLZ^{+}21, MLL^{+}24, MPM22, MTL23, MC14b, MPS23, MK16b, Na15, NB97, RDK24, SGJ22, SF89, WWX^{+}24, WMGZ24, AOU12, Bod04, CSLN10, CFGM02, DM07, DEL^{+}00, HL10, KZSD99, KKB10, PRP05, WX10, YWYL08]. Document-at-a-time [MTL23]. Document-grounded [MLL^{+}24]. Document-Level [WWX^{+}24, HYLC24, WMGZ24]. Document-Ordered [MPM22].

Documents [AVA^{+}24, Bie92, BCPN14, Bul95, COR13, Dun97, GZC89, KLHK13, VST^{+}14, LK20, STZ22, SMGM11, TSNS19, ZK20, CTM12, FCD02, KFS10, TG10]. Does [FJZS24, Fer17, IH07, MLK^{+}18, VAJR22, WZL^{+}19, ZWS^{+}25]. Doing [LCY21]. Domain [CZT^{+}23, FLW^{+}25, GLZ^{+}24, JSL^{+}24, LLN22, LHL23, LCL^{+}23a, MLZ^{+}22a, MLK^{+}18, THL^{+}25, VL90, WWR^{+}25, WZW^{+}25, WYZ^{+}25, WWN22, WLL^{+}23b, WWL^{+}25, XSWX25, XYCC24, YSJ25, ZLS^{+}23, ZYC^{+}24, ZWX^{+}24, ZZC^{+}25, ZH17, ZWZS25, CC03, GGMW03, GLG^{+}23, GSG^{+}25, HLH^{+}25, HZG20, LMW^{+}21, LGZ^{+}24, LGY^{+}25, MXM^{+}24, MPHS03, PTS13, QXX^{+}25, SLD^{+}24, TXC^{+}24, WR98, XZD^{+}25, YLS^{+}11, YOM24, ZZL^{+}23, ZZZ^{+}24, ZMN^{+}24]. Domain-Agnostic [THL^{+}25]. Domain-Specific [VL90, CC03, YLS^{+}11]. Domains [HCC^{+}16, SBM^{+}17, IH07]. Doubly [Oos23]. Drawing [GTW^{+}18, TM91]. Drift [SKC^{+}12]. Driven [HCD^{+}23, LSG^{+}25, FP00, LYZ^{+}23, MLL^{+}24, PSPBY10]. Drug [ZWX^{+}23]. Dual [CSJ^{+}24, CJY^{+}23b, CYZ^{+}24, DSZ^{+}23, GJS^{+}25, GLZ^{+}24, HCY^{+}20, LYY^{+}24, SLD^{+}24, SLQ24, TCSL24, WYZ^{+}25, WZC^{+}25, XZZ^{+}22, XYCC24, XYZ^{+}23, YZZS24, ZGF^{+}20, ZSZ^{+}24, ZWZS25]. Dual-branch [YZZS24]. Dual-factor [ZGF^{+}20]. Dual-MRC [CJY^{+}23b]. Dual-Query [WZC^{+}25]. Dual-Target [SLQ24, WYZ^{+}25, ZWZS25]. Dual-view [XYZ^{+}23]. Duplicate [CTH^{+}13, JTL^{+}17, CFGM02, HHC^{+}10, LHSC11]. durable [GL05]. During [CHBS15, MK16a]. Dynamic [AZBC20, ASSC22, BZDW24, BR96, BFNP10, Can93, Cro06, FZH^{+}22, HLC^{+}20, IYC^{+}23, LLN22, LRZ^{+}17, LWL^{+}24, LLZ^{+}25, NXF^{+}24, PGC^{+}25, TWS^{+}25, WLFT18, WMZ^{+}21, WHL^{+}22, WCX24, XZZ^{+}22, YLZW21, ZRT91, ZCZ^{+}23, ZWC^{+}25, JBCF07, LLC03, MOT11]. Dynamics [HLX^{+}24, NSK^{+}20, WH15, RSD^{+}13].

E-Commerce [DLZ^{+}22, GCR^{+}21, JSL^{+}24, ZLH^{+}22, ZLW24]. E-Discovery [OSV19]. Early [CC03, LW20]. Earthquake [ZZL^{+}21]. Edge [YMZF22]. Editing [MRC^{+}23]. Editor [OOSC93, Zha21]. Editor-in-Chief [Zha21]. Editorial [All89, All91, All94, Cro89a, Cro91, FG93, LH94, MR92, MS93, Mar08a, Zha21]. Editors [VL90]. Education [RVA^{+}24]. EEG [MRM24]. Effect [CLF^{+}25, Dun97, FKS19, QDG^{+}24, WAL^{+}21, YTD13, YTL25, CDR10]. Effective [AVA^{+}24, CS16, FLPC24, HZS^{+}25, KC15, LMDZ20, LYLS22, LNO^{+}25, MCF^{+}17, NXF^{+}24, PPPR13, YLZW22, YSW^{+}24, MWYL01, PMPJ11]. Effectiveness [JYM13, MBST17, RZM24, SGJ22, TBC96, WWG^{+}24, IH07, MZ08, SV99, XC00]. Effects [ABCZ21, AC16, AC19, CWL^{+}20, CFFK22, KFS10, Mar94, VHJC11, WMZ^{+}21, ZXZL20, VAJR22]. Efficiency [BLMN24, HWD13]. Efficient [ALS^{+}24, AVA^{+}24, BHY^{+}25, BC13, BC14, BA18, BCPN14, BKGM^{+}02, CZX^{+}19, CZZ^{+}20, CJW^{+}21, CQZ^{+}25a, CTH^{+}13, CM10, DLY^{+}25, EKK00, FLW^{+}23, FLPC24, GZL^{+}24, HHC^{+}23, HZS^{+}25, KZSD99, KZP10, KC15, LLY^{+}25, LMW^{+}24, LMR^{+}24, LXC^{+}25, LNO^{+}25, MTL23, RDK24, Row96, SC19, TGL^{+}97, WZZ04, XYW^{+}23a, XLW^{+}22, YSW^{+}24, YHS^{+}12, ZLZ25a, ZGCC18, ZZL^{+}22, ZWL^{+}23, HZW02, IYC^{+}23, KT05, PMPJ11, PZMRN05, SSCT09]. Efficiently [PV19]. Effort [MAH24]. eFraudCom [ZLH^{+}22]. ELAKT [PLS^{+}24]. Elastic [ZWS24]. Electricity [PR16]. Electronic [KM97, Mar94, RP98, HDPM08, Mar08a]. Element [WC21b, WWR^{+}25, Cro06]. Elias [PV17]. Eliminating [SYSW24]. ELSA [CGB19]. eludes [BVKS08]. email [CHS99, EBR08, ZKG08]. Embedded [Bul95]. EmbeddedButtons [Bie92]. Embedding [AZBC20, CZX^{+}22, CGB^{+}15, DHY^{+}19, GYC^{+}22, HCY^{+}20, JEAZ25, LLM22, LLY^{+}24, LZC^{+}25b, LL23b, NLF^{+}20, TCQ^{+}25, WHL^{+}22, WNCH24, XZCZ25, XLW^{+}22, XCG^{+}23, YZS^{+}24, YLYT23, ZSY^{+}21, ZKMM23, ZWE^{+}22, FLMD22, PTS13, ZZSS25]. Embedding-Based [GYC^{+}22, LL23b, TCQ^{+}25]. Embedding-Less [ZZSS25]. Embeddings [BTP24, CCVR17, LDW^{+}17, LCYY25, MPS23, YE18, YDW22, ZY21]. Emergency [SSY^{+}17]. Emerging [CLL24]. Emotion [CJY^{+}23b, JLC^{+}25, NSK^{+}20, QWZ^{+}15, WLM^{+}21, WCS^{+}24, ZJW^{+}24]. Emotion-aware [JLC^{+}25, WLM^{+}21]. Emotion-cause [CJY^{+}23b]. Emotional [QWZ^{+}15, SLW^{+}20]. empirical [CZJ07]. Employee [WZW^{+}22]. Employer [LZW^{+}20]. Empowered [XLX^{+}25, ZXH^{+}25]. Emulators [GR12]. Enabling [WGZ^{+}25]. encapsulation [ACM^{+}02]. Encoder [QDZW23, ZZY^{+}23, WLL^{+}23b, YRC23]. Encoders [LMR^{+}24]. Encoding [HQZ17, WZC^{+}25, GL05]. Encounters [MJH15]. Encryption [KBD89]. End [CRC^{+}22, MPT21, WQH^{+}25, YLZW22]. End-to-End [MPT21, WQH^{+}25, YLZW22, CRC^{+}22]. Engagement [FSE^{+}24, LQJ^{+}24, NSVB18, TZP22]. Engine [JTS16, WLY^{+}14, YTL25, ZZL^{+}21, EJ03, LOP^{+}13, SC03, TS10]. engineer [HDPM08]. Engineering [TS10, XLW^{+}22]. Engines [DH97, RCR^{+}21, AOU12, BBMN^{+}08, FPSO06, PSPBY10]. English [ANW17]. Enhanced [AMS20, BR96, GJS^{+}25, HZS^{+}25, HML^{+}23, LLY^{+}25, LZL^{+}25, MLZ^{+}22a, MZG^{+}24, QZX^{+}20, SQZ^{+}22, WZY^{+}24, WWR^{+}25, XYZ^{+}23, YMZ^{+}22, YZC^{+}25, ZWF^{+}23, ZWX^{+}24, ZSL^{+}25, ZSH^{+}25, ZSLK24, CSJ^{+}24, GSG^{+}25, LSZ^{+}25, WLW^{+}24a, WZZ^{+}24b]. Enhancement [CQZ^{+}25a, LZZ^{+}23]. Enhances [MNP^{+}21a]. Enhancing [AZCC21, CGD^{+}25, DXL^{+}22, FSE^{+}24, LDW^{+}17, LLX^{+}22, LZLM19, LZW^{+}20, LDN^{+}25, PLS^{+}24, SSZ^{+}23, XSWX25, ZCD06, ZWrWZ23, ZPW^{+}24, FSC07]. enrich [CDH^{+}09]. Enriching [KLHK13]. enrichment [SPS^{+}06]. Ensemble [CYZ^{+}24, EMS19, MPS23, ZWL^{+}22]. Ensembles [DLN^{+}16]. Enterprises [CH96]. Entities [LMW^{+}14, ZMK^{+}20]. Entity [CFC^{+}19, JEAZ25, LCL^{+}17, LGQL19, LXG^{+}23, LL23b, XZZ^{+}22, XYCC24, YHS^{+}12, ZZT^{+}21, ZWZ^{+}24, BBD11, LCH07, PTS13]. Entity-Linking [CFC^{+}19]. Entropy [MP20, GP00]. enviroments [ATW00]. Environment [AM89, JMSV92, KMKO92, LZ96, SS93, WL97, Cro06]. Environments [CFP95, Gul96, LZWZ24, CC03, JBCF07, KM13]. Envoy [PYF^{+}92]. equalizing [TkWW99]. Equivalent [CZT^{+}23]. equivalents [PTKJ07]. Era [BOQ^{+}25]. Erase [DLY^{+}25]. Error [DM07, MPHS03]. escalation [WH09]. Estimate [LK20, Bod04]. estimates [SZ09]. Estimating [CZ93, MHR22]. Estimation [BF98, GRSD^{+}16, LYN^{+}21, LXLZ16, MTC21, Oos23, QZS^{+}23, SKC^{+}12]. Estimator [BvdHFS25]. Evaluating [ABW93, CML00, DRdR23, FKM^{+}05, GR12, HKTR04, JGP^{+}07, LCY21, MC94, SKK24, SGZ^{+}24, WCZ^{+}25, WRJV05]. Evaluation [BCPN14, CS94, ERG^{+}89, FTZ11, FFM17, Fer17, Gin24, Gru92, GR95, JMO22, JVA^{+}15, JS24, JSZL23, Kon04, LMW^{+}21, LL89, LMDZ20, LZW^{+}20, MTL23, MMST17, MMCC^{+}21, MMLP97, SZ21, STZ22, SLZ^{+}19, TBC96, TTZ^{+}25, TC91, VSS17, VB98, ZGN^{+}17, ZLF^{+}23, CJRY12, CL07, GMR09, JK02, JBCF07, KL06, LSOS06, LZW21, WCZ^{+}25]. Evaluations [FZSG21, KFS10]. Event [LYP^{+}25, LLBS18, MSMO18, TCSL24, WWX^{+}23, WWX^{+}24, WWX^{+}25, ZMZ25, ZCLC23, CCC09, LLPT10]. Event-based [LLBS18]. Events [LHG^{+}12, LXLZ16]. every [Gla04]. Evidence [MC14a, RVA^{+}24, WQH^{+}25, XHSD24, UCH03, YLW^{+}23, Gla04]. Evidence-Claim [WQH^{+}25]. Evolution [IM05, LLY^{+}24, LXLZ16, TWS^{+}25, Yan15, ZJT16]. Evolutionary [SCG^{+}19, WMZ^{+}21]. Evolving [JMSV92, WHL^{+}22]. Examining [BOQ^{+}25, KSVS16, SAK^{+}23]. Example [YC94, TFS^{+}12]. Example-Based [YC94]. Examples [KLHK13, BBD11]. Excellent [FJZS24]. Exception [SM95]. Exceptions [SM95]. Exclusion [CCVR17]. Expanding [KYK24]. Expansion [BTP24, CdMRB01, GWR99]. Expectations [MBST17]. Expected [OSV19, CL03, DWLW09]. Experience [BBT92]. Experiences [DH97]. Experimental [NLL^{+}24, WLY^{+}14]. Experiments [Kwo90, MLF95, SAK^{+}23, Car12, GMR09, IKG93]. Expert [ASV91, PXW^{+}25]. Expertise [MLK^{+}18]. Experts [GCB^{+}25, PHK12, BM02]. Explainable [AZBC20, CCZ^{+}19, LZC23, LZZ^{+}23, LYG^{+}25, LXC^{+}25, LNO^{+}25, MHLH23, SJZ^{+}25, WZW^{+}22, WQH^{+}25, XZT^{+}24, YZS^{+}24, YSJ25, ZSZ^{+}24, ZSL^{+}25, ZCS^{+}23]. Explaining [KJA20, LRSD25, Mar94]. Explanation [Gul96, LGL^{+}24, LMW^{+}23, SWZ^{+}22, WLY^{+}24, ZZD25]. Explanations [LRA23, LMW^{+}23, TBCW20]. Explicit [EMG11, FMSS14, MSMO18, SZM^{+}23]. Exploitation [LLC^{+}25b]. Exploiting [CL18, FFM17, GGMW03, GNN^{+}10, MKF91, QHLY20, QHCY22, TLS^{+}22, TL14, WX10, WLT^{+}22, ZZL^{+}21, FPSO06, TLJ^{+}07]. Exploration [JMD20, Lin07]. Exploratory [CS94, HYH15, HL17, LMW^{+}14, MLK^{+}18, RPE^{+}18]. Explore [YSW^{+}24]. explorer [XC05]. Exploring [CCC^{+}12, EBR08, LCW^{+}24a, LMZ23b, PHK12, SPKW11, WZW^{+}19, ZZG^{+}25]. Exposure [MAP^{+}22, XYW^{+}23b]. Expressive [KM97, SW94]. Extended [YWYL08, KL06]. Extending [DEL^{+}00, GSR96]. Extensible [AM89, KL91]. Extension [BR96, LLY^{+}24, GWR99]. External [CZX^{+}19, LDN^{+}25, LHA^{+}25a, NCZG14, NCHW15, TSNS19, ZLZ25a]. Extra [ACC18]. Extra-Topical [ACC18]. Extraction [CJY^{+}23b, CSJS10, HYLC24, JZ06, LYLS22, MKZPK18, PR16, RL94, SHS^{+}23, SZM^{+}23, WWX^{+}23, WWX^{+}24, WWX^{+}25, WMGZ24, WWR^{+}25, MC10a, WX10]. Extractive [CLG^{+}23, LRA23, RCR^{+}18]. Extremely [AVA^{+}24]. Extrinsic [SJL^{+}25]. Eye [Jac91].

Facet [WCZ^{+}25, LWA08]. Facet-aware [WCZ^{+}25]. Faceted [MT21, NFZ19]. Facility [MK94]. Fact [WWN^{+}25]. Factoid [JEAZ25, CTM12, Lin07]. Factor [LGL^{+}24, NFZ19, SJL^{+}25, WYZ^{+}25, ZGF^{+}20]. Factor-level [LGL^{+}24]. Factorization [CZZ^{+}20, FMSS14, LJHL17, LZG^{+}18, LPLH19, SZS17, WRLD22, ZLW^{+}23]. Factors [ACC18, FZSG21, HJZ^{+}23]. Factual [MDX^{+}25]. FairGap [CWZ^{+}24]. Fairness [CWZ^{+}24, LAO^{+}24, LRSD25, LFW23, SKK24, SWZ^{+}24, WMZ^{+}23, WLY^{+}24, WMM^{+}23, WCX24, XYX^{+}25, YWQ^{+}24]. Fairness-assurance [WCX24]. Fairness-Aware [CWZ^{+}24, WMM^{+}23, YWQ^{+}24]. Fairness-centric [LFW23]. Faithfully [XZT^{+}24]. Fake [AZK12, LPYM23, LW20, WQH^{+}25]. False [MMCC^{+}21]. False-positive [MMCC^{+}21]. family [DWLW09]. Fano [PV17]. Fast [ABB^{+}15, AL13, BC11, CJY23a, DLA15, DLN^{+}16, LZC^{+}25b, Man97, MZ96, NTV22, SNZBY00, WZB04, ZTM^{+}17, CFGM02, HZW02, ZHC^{+}13]. FASTER [WCX24]. Favorite [GCL^{+}21]. FDKT [LBZ^{+}24]. Feature [CWZ^{+}20a, CXZ^{+}21, CLM^{+}22, HZF^{+}23, HWN^{+}19, SCG^{+}19, XLW^{+}22, YCL^{+}23, ZSL^{+}25, ACS08]. Feature-aware [CXZ^{+}21]. Feature-Based [CWZ^{+}20a]. Feature-Enhanced [ZSL^{+}25]. Feature-Level [CLM^{+}22]. Features [LPY94, MAL16, ZZT^{+}21, dODMS20, GBMS99, MSOH13]. Federated [DSW^{+}25, FWG^{+}24, GYW^{+}25, HNN^{+}25, IYC^{+}23, LPYM23, LWH^{+}23, LGY^{+}25, TXC^{+}24, YPYM24, YPF^{+}25, YYY^{+}24, YYQ^{+}25, ZLW^{+}23]. Feedback [DLA15, Dun97, LMZ^{+}23a, LZC^{+}24b, LPLH19, MYHL11, MHZ16, MCL^{+}12, NSVB18, Oos23, RK19, RRS95, VHJC11, WC21a, WWHQ25, XYX^{+}25, YXA^{+}24, CRG02, CZJ07, JGP^{+}07, WRJV05, ZCD06]. Feeling [MRM24]. Feeling-of-Knowing [MRM24]. Felicity [KM97]. Few [CLL24, CZJH24, DPSH23, FZXdR24, LLG^{+}23, LLZ^{+}21, ZK20, GMR09]. Few-Shot [CLL24, CZJH24, FZXdR24, LLG^{+}23, LLZ^{+}21]. Fidelity [HWD13]. Field [Ack98, PR16, QFZ^{+}19]. figure [GWLC06]. figure-of-merit [GWLC06]. figures [BM12]. File [CZ93, Man97]. Files [LL89, LR96, MZ96, ZRT91, ADCU08, BB10, HDPM08]. Filter [GWL^{+}24, LBM^{+}24, AL13, CL07]. Filter-based [LBM^{+}24]. Filtering [AARC20, CCCD19, CLPM20, CQL^{+}24, CQZ^{+}25a, CDM15, DLY^{+}25, DHY^{+}19, GXTL14, LCX^{+}19, LCYY25, LLZ^{+}21, LXG^{+}23, LZC^{+}24c, MNZ19, MMLP97, NTV22, PYC^{+}19, QJG^{+}25, SWZ^{+}24, WHL^{+}16, WRLD22, WLL^{+}23a, XLL^{+}23, XHW^{+}19, YPYM24, YLYT23, ZLL^{+}20, ZMZ^{+}24, ZZSS25, ZDW^{+}22, DC10, HKTR04, Hof04, HCZ04, IH07, TKD09, WdVR08, ZKG08]. Filters [QJG^{+}25, Row96]. financial [SC09].

Finding

[Mar94, PXW^{+}25, SBM^{+}17, VB98, UCH03]. Fine [CSR^{+}20, CL18, CL19, CCZ^{+}23, LGZ^{+}24, MZG^{+}24, XPL^{+}24]. Fine-Grained [CSR^{+}20, CL18, CCZ^{+}23, XPL^{+}24, CL19]. Fine-tune [LGZ^{+}24]. Finite [CCVR17]. First [GCF^{+}22, GCB^{+}25, MMH^{+}24]. First-Stage [GCF^{+}22, GCB^{+}25, MMH^{+}24]. Físchlár [LSOS06]. Físchlár-News [LSOS06]. Fish [ABW93]. Fit [FC17, QZX^{+}20, IH07, MDX^{+}25]. FIT-RAG [MDX^{+}25]. FITE [PTKJ07]. Fitting [WWN^{+}25]. Flexible [DC92, HM90, Tom89, XMS^{+}25, ZTM^{+}17, BJS99, SNZBY00]. Flow [LLM22]. Flow-Based [LLM22]. FNED [LW20]. Focused [FAZC12, ACM^{+}02]. Follow [CLG^{+}23]. Following [ZXH^{+}25]. Footprint [WRLD22]. Footprints [GZW^{+}18]. Force [SHS^{+}23]. Force-Directed [SHS^{+}23]. Forecasting [YZH^{+}24]. Forest [IC16]. Forgetting [HLC^{+}20]. Form [AARC20, HEA90, BKGM^{+}02]. Formal [Gru92, AF07, GFWK04]. formalisms [GBMS99]. formalization [SM99]. Formation [RFB^{+}24]. Formative [ERG^{+}89]. Formulae [SMGM11]. Formulation [CFFK22, MC14a]. Forums [NZW^{+}14, ACS08]. Forward [DLA15, LMR^{+}24]. Foundational [JXY^{+}24]. Fractal [BF98, Koi95]. Fractal-Based [Koi95]. Frames [OG94]. Framework [ASMZ25, BHH^{+}16, BA18, CML^{+}21, CJY^{+}23b, CGB^{+}15, CFE^{+}25, DXZ^{+}22, DLY^{+}25, DGMG23, DZX^{+}23, DC92, FLW^{+}25, GCL^{+}21, GR12, HHC^{+}23, HLNS98, LAO^{+}24, LCJ^{+}16, LYLS22, LLX^{+}22, LL23a, LPYM23, LXC^{+}25, LCNY23, MTKY92, MZG^{+}24, MHZ16, MCL^{+}12, PYF^{+}92, SKK24, SWZW25, SZY^{+}24, SZZL23, SYSW24, SLZ^{+}24, VL90, WLL^{+}23a, WZL^{+}24, WYZ^{+}25, WMM^{+}23, XPSC24, XZT^{+}24, YDXW21, ZWRL19, ZCZ^{+}23, ZZD25, ZGCC18, ZDZ23, CL06, CCC09, DWLW09, DDH^{+}09, EKK99, KL00, MR10, SSCT09, XC05]. Fraud [ZLH^{+}22, ZWL^{+}22]. Free [YYQ^{+}25, MC10a]. Frequency [PR16, PTKJ07, HO07, SV99]. Frequency-based [PTKJ07]. Frequent [BCFG15, CGB19]. Frequently [HY20]. FSR [YYQ^{+}25]. full [Leh06, MRYGM01]. full-text [Leh06, MRYGM01]. function [CRG02, CBZ11]. functional [WSBC01, BSV09]. Functions [BGI24, Coh97, FCDH91, Fuh89, IC16]. Funnelling [EMS19, MPS23]. Further [GTW^{+}18]. Fusing [LYN^{+}21, YFO^{+}25]. Fusion [APCC15, AK18, BGI24, LWL^{+}24, WZW23, ZWZ^{+}22]. Future [CDW^{+}23, GZW^{+}24, Ols89, ZZL^{+}23, ZLL^{+}22]. Fuzzy [LBZ^{+}24, BC13].

Gain [DGMG23, JK02, KL06]. Gain-based [DGMG23, JK02]. Gains [ZQH^{+}24a]. Game [MHR22]. Games [ZWT^{+}23]. Gap [AMS20, WTS^{+}08]. garbling [DM07]. Garden [Ack98]. Gated [XZZ^{+}22]. Gating [MWZ^{+}25, WZW^{+}25]. Gaussian [LBC^{+}19]. GCN [MHLH23]. GDESA [QDZW23]. General [AFF^{+}19, CGB^{+}15, DXZ^{+}22, Gul96, HJZ^{+}23, JTL^{+}17, LXC^{+}25, MTKY92, YE18, ZZL^{+}15, ZWT^{+}23, CW02]. General-Purpose [AFF^{+}19, YE18, CW02]. generalization [CL03]. Generalized [LZC24a, MPS23, SZY^{+}24, XPSC24]. Generate [ZSN^{+}22]. Generated [DGMG23, MAA^{+}25]. Generating [CWZ^{+}24, LCL^{+}23a, LHA^{+}25a, CLG^{+}23]. Generation [BC14, BOQ^{+}25, DLZ^{+}22, GCR^{+}21, JSL^{+}24, LJZ^{+}25, LLN^{+}25, MLL^{+}24, QZS^{+}24, RCR^{+}21, SZZL23, SLW^{+}20, SWZ^{+}22, WLL24, WZC^{+}25, WCS^{+}24, ZGF^{+}20, ZGC^{+}22, ZCZ^{+}23, ZZD25, ZCS^{+}23, AL13, CC05, RBV09]. Generative [CSJ^{+}24, LMW^{+}21, LJZ^{+}25, TZG^{+}24, TQN^{+}23, WZL^{+}24, ZYD^{+}24].

Generator [LYZ^{+}23]. Generic [LLX^{+}22, LPYM23, SYSW24]. genomic [Wil05]. Genre [YHYO01]. Geographic [HY20]. Geographical [LZG^{+}18]. Geolocation [CL18, CL19]. Geometric [JWX^{+}25]. Geometric-Augmented [JWX^{+}25]. geometrical [CHS99]. geometry [HZSL13]. GeoMF [LZG^{+}18]. Geoparsing [MK16a, MKZPK18]. Georeferencing [VST^{+}14]. Geosemantics [MK16a]. Geospatial [TGS16]. Geotagging [MKZPK18]. Gerard [Ano96b]. Get [Jac91]. Getting [CR92]. Global [CH96, WNS^{+}17, ZLWW22, CRNZ^{+}03]. GMRES [WW10]. GNN [LLP24, TQN^{+}23, WZHY24]. GNN-Based [WZHY24, TQN^{+}23]. GNN-transformers [LLP24]. GNNs [WLW^{+}24a, ZZZ^{+}25]. goal [DZX^{+}23]. Gold [STC^{+}24]. good [GMR09]. Goods [WLZW21]. Google [WW10]. Gould [Gru92]. GPS [WZC^{+}17]. GPTs [BOQ^{+}25]. Gradient [YZZS24, ZWX^{+}23]. Gradient-balanced [YZZS24]. Grained [CSR^{+}20, CL18, CCZ^{+}23, XPL^{+}24, CL19, DDSW24, LMW^{+}24]. Gram [PV19]. Grammar [BBT92]. Grammar-Defined [BBT92]. Grammars [BK90, RS93]. Grams [Coh97, SHN17]. granularity [SLZ^{+}24, YLS^{+}11]. Graph [APCC15, AM25, BLH^{+}22, CQF^{+}23, CSR^{+}20, CWZ^{+}24, CQZ^{+}25a, CYZ^{+}24, DXZ^{+}22, FLW^{+}23, FWG^{+}24, GLL^{+}23, HRY^{+}22a, HRY^{+}22b, HML^{+}23, JEAZ25, JWX^{+}25, LLN22, LHL23, LSG^{+}25, LCYY25, LZL^{+}25, LMMO23, LXG^{+}23, LZC^{+}24c, LLZ^{+}25, MHLH23, MAP^{+}22, PCC^{+}22, PCCC22, PGC^{+}25, PBS22, PZD^{+}22, PSM24, PL94, QGW^{+}24, QJG^{+}25, QHLY20, RVA^{+}24, SLZ^{+}25, SWZW25, SHS^{+}23, SZY^{+}24, TWS^{+}25, TLS^{+}22, TXC^{+}24, WWX^{+}23, WWX^{+}25, WZW^{+}19, WL22, WLL^{+}23a, WZZ^{+}24b, WZY^{+}24, WWR^{+}25, WZHY24, WWN^{+}25, XZCZ25, XZZ^{+}22, XZH^{+}22, XLL^{+}23, XXDW24, XLX^{+}25, YCG^{+}24, YHS^{+}21, YMZ^{+}22, YZS^{+}24, YLYT23, YOM24, YHS^{+}12, YZL^{+}25, ZJL^{+}24, ZSY^{+}21, ZLH^{+}22, ZCZW22, ZWZ^{+}22, ZLY^{+}22, ZZY^{+}23, ZJW^{+}24, ZWC^{+}25, ZWS^{+}25, ZLZ25a, ZZSS25, ZMZ25, ZDW^{+}22, ZXZL23, ZWC^{+}23, ZYL^{+}25, HHC^{+}10, LWSA10, MC10b, PH07, ZHD^{+}22]. Graph-Based [APCC15, BLH^{+}22, JWX^{+}25, LSG^{+}25, MAP^{+}22, WWX^{+}25, ZSY^{+}21, ZCZW22, ZZY^{+}23]. Graph-Enhanced [WWR^{+}25]. Graph-guided [ZLY^{+}22]. Graph-Regularized [CSR^{+}20]. graph-walk-based [MC10b]. GraphHINGE [JDZ^{+}22]. Graphical [VL90, CDF01]. Graphs [TCSL24, YMZF22, ZLZ^{+}25b]. GRAS [PMPJ11]. Greedy [LCL^{+}17, QDZW23]. Greene [Gru92]. Grounded [FTZZ23, ZKJW22, HZS^{+}25, MLL^{+}24]. Grounding [MK16a]. Group [ASV91, CSB21, GYC^{+}22, LMZ23b, OOSC93, SKK24, TLS^{+}22, XPL^{+}24, TkWW99]. Group-Based [XPL^{+}24]. Groups [LM90, CSJS10]. Groupwork [OOSC93]. Grudin [Wie92]. GSLAE [WYZ^{+}25]. Guest [LH94, MS93]. Guided [CYC^{+}24, DLY^{+}25, DSW^{+}25, LCX^{+}19, MLZ^{+}22b, PZD^{+}22, WLM^{+}21, XZD^{+}25, XCG^{+}23, YSW^{+}24, ZLY^{+}22, ZZH^{+}25]. GVoS [JTL^{+}17].

H3GNN [YHWZ24]. HAL [YWYL08]. Hallucination [HYM^{+}25]. Hand [SZ93]. handheld [BKGM^{+}02]. Handling [Mye90, PV19, SM95, Owe02]. Happy [Mar94]. Harmonious [WGZ^{+}25]. Harp [LL99]. Harvesting [SMN^{+}16]. Hash [FCDH91]. Hashing [CJN20, Coh97, HHC^{+}23, WLZ^{+}24, ZZL^{+}22, ZWL^{+}23, ZHC^{+}13]. HDM [GPS93]. Healthcare [YLW^{+}23, ZTZL25]. heavily [LHSC11]. HEK [YZC^{+}25].

HEK-CL [YZC^{+}25]. Help [CE89, CHZ24, VAJR22, Bod04]. Heterogeneity [JLC^{+}25, SAK^{+}23]. Heterogeneity-aware [JLC^{+}25]. Heterogeneous [CQF^{+}23, CLPM20, EMS19, FZH^{+}22, FZXdR24, GLL^{+}23, HCC^{+}16, JDZ^{+}22, LTC^{+}25, LLY^{+}24, MPS23, PYC^{+}19, SWZW25, SSY^{+}17, TLKW16, WZZ^{+}23, WWN22, WWL^{+}25, YHS^{+}21, YPYM24, ZJL^{+}24, ZWZ^{+}22, ZXM^{+}24, ATW00, CGMP99, Leh06]. HGAT [YHS^{+}21]. HGDNet [MWZ^{+}25]. hidden [GIS03, IG08, JZ06]. hidden-web [IG08, GIS03]. Hierarchical [CSR^{+}20, CZX^{+}22, CS94, GYC^{+}22, LYZ^{+}23, LWH^{+}23, MWZ^{+}25, QGW^{+}24, WWX^{+}23, WZW^{+}25, XC24, YHWZ24, YZC^{+}25, ZZD^{+}20, ZWX^{+}24, ZY21, ZWL^{+}22, ZCLX22, BFF^{+}03, GGMW03]. Hierarchies [BRS92, LM90]. Hierarchy [Yan15, ZHL^{+}22]. Hierarchy-based [ZHL^{+}22]. High [BS12, DC92, FC91, LJHL17, OS96, RL94, RLC^{+}25, WZW^{+}19, YMZF22, ZK20, HZSL13]. High-Level [DC92, OS96]. High-Order [LJHL17, WZW^{+}19, HZSL13]. High-performance [BS12]. High-Precision [RL94]. High-Quality [RLC^{+}25, YMZF22]. High-recall [ZK20]. Highly [AVA^{+}24, KM13, MWYL01]. Hilbert [CCVR17]. Historical [CJY23a, TP05, ZYY^{+}24, FPSO06]. History [CSH^{+}25, LLT^{+}21]. HMM [CCC09]. HMM-based [CCC09]. Hoc [DLA22, KSYE24, CGM05, KL09]. Holistic [MNP^{+}21b]. Home [Ols89, UCH03]. Hop [ZQH^{+}24b]. HORAE [HWT^{+}25]. Human [All91, BHH^{+}16, Jac91, NFZ19, SSY^{+}17, SZ93, WGZ^{+}25, WWHQ25, Wil98, YXL^{+}23, ZZZ^{+}16, GKL^{+}07]. Human-Computer [Jac91, SZ93]. Human-Machine [WGZ^{+}25]. HYB [BC11]. Hybrid [AK18, BGI24, CML^{+}21, HJZ^{+}23, YHWZ24]. Hyper [HCY^{+}20, HO07, SM99].

Hyper-Network [HCY^{+}20]. hyper-object [SM99]. hyper-parameters [HO07]. Hyperbolic [WLT^{+}22, ZCS^{+}23]. Hyperdocuments [SFR98]. Hyperedge [GYC^{+}22]. Hyperform [WL97]. Hypergraph [AGT24, CXZ^{+}21, LTC^{+}25, TCSL24, WS90, ZSQ^{+}24, ZXM^{+}24, YHWZ24]. Hypergraph-based [WS90]. Hypergraphs [WLL^{+}23b]. hyperlinks [KL10, LS02]. Hypermedia [Bul95, HLNS98, LZ96, MC94, SLHS93, UY89, WL97, ATW00, BHMW07, dOTM01]. HyperSoRec [WLT^{+}22]. Hyperspherical [FLMD22]. Hypertext [CE89, GPS93, KL91, Rad92, STSM95, SF89, Tom89, WR98]. Hypertext-Based [CE89]. Hypertexts [BRS92].

ICF [CLM^{+}22]. ID [CGD^{+}25, WXZ^{+}25]. ID-based [CGD^{+}25]. ID-centric [WXZ^{+}25]. identification [AC08, PTKJ07, SSCT09]. Identify [PHK12]. Identifying [BRS92, MJH15, SMGM11]. Identity [ZZG^{+}25, AC08]. identity-level [AC08]. IDF [SHN17, WLWK08]. IF [CDR21]. ILMART [LNO^{+}25]. Image [CLZ^{+}14, GJS^{+}25, GR95, NWZC12, SLZ^{+}19, SLM^{+}24, VB98, WZL^{+}19, CZJ07, EKK99, EKK00, HJZL09, KKS^{+}08, TMT06, XLW^{+}22, ZCD06, ZRZ02]. Image-Text [SLM^{+}24]. Images [CCZ^{+}19, GR96, dODMS20]. Impact [BACF17, CST^{+}23, MTC21, RABC22, RZM24, XHSD24, EJ03, SV99, TkWW99]. Impacts [JYM13]. imperfect [Owe02]. Implementation [CW94, KEL95, TGL^{+}97, WW94]. Implementing [DC92]. Implicit [CMZ^{+}23, DYL^{+}20, LZLM19, LZC^{+}24b, NSVB18, SZM^{+}23, FKM^{+}05, JGP^{+}07, WRJV05]. Importance [LYN^{+}21]. Imprecise [Mor90]. Improve [HL17, HCC^{+}16, FKM^{+}05, LLC03]. Improved [BBMN^{+}08, CCVR17, CPL15, FMSS14, CDH^{+}09]. Improvement [GS91]. Improves [VAJR21]. Improving [CRG02, DYL^{+}20, ES13, HCC^{+}17, LYY^{+}24, MZLK11, MRC^{+}23, MMH^{+}24, MC10b, RBCT14, XC00, PPGK04]. in-memory [TS10]. In-Situ [YTL25]. Including [XZCZ25]. Incomplete [KC95, LTC^{+}25]. Inconsistencies [Coo95]. Incorporating [ASST05, MZLK11, MWA^{+}24, MBST17, VDR25, WZC^{+}17]. Incremental [ADCU08, Can93, SM99, PSPBY10, TM10]. indefinite [WMZ10]. Independence [DLA22]. Independent [LSG^{+}25, TFS^{+}12, UCH03]. Index [Ano93, Ano96a, Ano97, BC14, Cro98, DLA15, TM20, AOU12, BC11, MRYGM01]. Index-Based [BC14]. Indexes [LMR^{+}24, MPM22, PV17, BS12, FBN^{+}12, WZB04]. Indexing [FB91, HCC14, MZ96, STSM95, SMGM11, WXLC13, ZNL^{+}17, CM10, EKK00, HSS^{+}09, KO98, QZXP06, TKD09, TMT06]. Indices [PR16]. Individual [FWG^{+}24]. Individual-Level [FWG^{+}24]. Induced [LCL^{+}23b, NCHW15, KL10]. Inductive [CQF^{+}23, FP94, ZYY^{+}21]. Industry [HE91]. Infection [FWG^{+}24]. Infer [TLKW16]. Inference [Cum14, GZW^{+}24, HGBY16, KJA20, LHH^{+}19, LYZ^{+}23, TC91, WQH^{+}25, WY95, WWN^{+}25, ZYW^{+}23, ZZZ^{+}16, TL03]. Inferior [RK19]. Inferring [GZW^{+}18, LRZ^{+}17]. Influence [GRSD^{+}16, JMZ22, LLBS18, NGMD17, VFA17, WLFT18]. Influences [CACW23]. Influencing [ACC18]. Influential [FZSG21]. Inform [TBCW20]. Information [AM89, AMS20, APCC15, All89, Ano94a, APJL20, ACC18, BVKS08, BK90, BLMN24, CE89, Can93, CCC^{+}12, CJY23a, CWL^{+}20, CHBS15, Coo95, CV98, Cro89b, CPL15, DGMG23, EMG11, FTZ11, FZH^{+}22, FZXdR24, FLPC24, FMSS14, FCDH91, FEL22, FAZC12, FP94, FR97, Gin24, JMSV92, JVA^{+}15, JDZ^{+}22, JSHS96, KMPS21, Koi95, KC95, KC92, Kwo90, Kwo95, LCJ^{+}16, LPC^{+}23, LJZ^{+}25, LZC24a, LZM^{+}19, LWH^{+}23, LMMO23, LPZ^{+}25, LLZ^{+}25, LZ96, MZLK11, MWA^{+}24, MMST17, MKF91, MDX^{+}25, MAH24, MH89, MRM24, Mor90, MMLP97, MBJK90, Na15, OS96, OG94, PBJ^{+}16, PR04, PZH^{+}24, PSL16, QHLY20, QHCY22, RABC22, RRS95, RL94, RN96, RP98, Row96, Ruh94, RS97, SL95, SM95, SZM^{+}23, TY22, TLS^{+}22, THT12, TKD09, TL14, VKD21, VdRK18, VSS17, VR24, WW94, WZC^{+}25, WY95, XPSC24, YSW^{+}24, YMZ^{+}22, YTD13, ZHY14, ZK20, dR19]. information [ASST05, AV02, BCD^{+}08, BBMN^{+}08, BDS04, CML00, CRNZ^{+}03, CdMRB01, Car12, CGMP99, Coh00, CGM02, FG04, GBMS99, HT99, HO07, HZSL13, KMdRS06, KKS^{+}08, KNS04, KO98, KL09, LBS08, LA08, MR10, MC10a, Mel08, MC10b, Owe02, PMPJ11, PPGK04, PH07, PGD07, PSDB99, SS08, TLJ^{+}07, Wil05, WSBC01, XC00, YLS^{+}11, ZL04]. Information-guided [YSW^{+}24]. Information-Seeking [VKD21]. information-theoretic [CdMRB01, MR10]. Informative [HCC^{+}16, LCL^{+}23a]. Informed [XXDW24]. Ingredients [FEL22]. Inhibition [AC19]. Initiative [VKD21, KKS^{+}08]. Inner [BNIL24a, BNIL24b]. Innovation [WBP^{+}16]. Input [CMR91, KMKO92, Mye90]. Insights [GTW^{+}18, ZTZL25]. Instability [LMZ20]. Install [ZWL^{+}22]. Instance [CFE^{+}25, NLL^{+}24]. Instruction [LHZ^{+}25, ZXH^{+}25]. Integrated [BA18, WDL^{+}24]. Integrating [ASV91, CHS99, CDZ^{+}23, CSB21, FC17, HLL^{+}25, JS24, MHZ16]. Integration [Bul95, FR97, HE91, IKG93, Coh00, GBMS99]. Intelligent [HZG20, LLY^{+}24, MMLP97, QZS^{+}24, SQZ^{+}22, SSGC89, GBMS99, Owe02]. Intent [BZDW24, GJS^{+}25, QZW^{+}25, RPE^{+}18, SLW^{+}24, SYX^{+}17, ZCLX22, LWSA10]. Intent-Oriented [BZDW24]. Intention [LZZ^{+}23, WMZ^{+}21]. Intentions [CDR21, XZY21]. Inter [HE91, SAW^{+}25, KKB10]. Inter-Case [SAW^{+}25]. inter-document [KKB10]. Inter-Organization [HE91]. inter-passage [KKB10]. Interaction [All91, CDZ^{+}23, CLF^{+}25, CCZ^{+}23, GYW^{+}25, HLX^{+}24, Jac91, JVA^{+}15, JDZ^{+}22, MLK^{+}18, MTC21, SLZ^{+}25, SMT25, SZ93, WGZ^{+}25, ZWZ^{+}22, ZWX^{+}23, ZCS^{+}23, CC03, LGM^{+}24]. Interaction-aware [ZWX^{+}23]. Interactions [HYH15, SZM^{+}23, ZLWW22, ZAK^{+}23, GKL^{+}07]. Interactive [GWL^{+}24, HLL^{+}25, HM90, LMW^{+}14, RPE^{+}18, SLX^{+}24, WC21a, WBB^{+}90, XTF^{+}21, ZJW^{+}24, MLDGH04, ZFL^{+}07, CL06]. interchange [GBMS99]. Interdependence [HE91]. Interest [AC18, AK18, BZDW24, HWT^{+}25, LJHL17, MPS07, MZG^{+}24, RABC22, YCZ^{+}16, ZZZ^{+}24, CRC^{+}22, MMH^{+}24, XC24]. Interest-based [MPS07]. Interesting [ZHD^{+}22]. Interests [AGT24, GZW^{+}18, LRZ^{+}17]. Interface [Gru92, HSH90, HM90, MR92, PCD92, PSDB99, SS93, TM91, Wie92]. Interfaces [CS94, DC92, KN93, SE91]. Interleaved [HWD13, CJRY12]. Interleaving [CSZD15]. Intermediary [SSGC89]. Internet [CDR21, LLC03]. interoperability [PR04]. Interpersonal [IKG93]. Interpretable [LRA23, LLZ^{+}22, LBZ^{+}24, WNCH24]. Interpreting [WLWK08]. intersection [CM10]. Intervals [Web13]. Intervened [XWL^{+}25]. Interview [SQZ^{+}22]. Interviews [QZS^{+}24]. Intrinsic [RBCT14, SJL^{+}25]. Introduction [BYFM06, BCD^{+}08, Cro89a, HKS^{+}21, HRY^{+}22a, HRY^{+}22b, Kli94, Kon04, SL95, WLS^{+}17, WLF^{+}25a, WLF^{+}25b, Wil05, YSXK23a, YSXK23b]. Invariant [LZWZ24, YCL^{+}25]. invariants [PH07]. Inverted [KNCLO17, MZ96, TM20, ADCU08, BB10, CM10]. Investigating [SGJ22, TBCW20]. Investigation [RJB89, RS93, CZJ07]. Investment [CHZ24]. Invisible [WLZ^{+}24]. IR [Fuh99, GP00, JK02, JS24, TCMC21]. Isometric [CCVR17]. Issue [BKWZ15, Cro89a, HKS^{+}21, HRY^{+}22a, HZF^{+}23, Kli94, PBJ^{+}16, SL95, WLS^{+}17, WLF^{+}25a, WLF^{+}25b, YSXK23a, BYFM06, LdJKR12]. Issues [LL89, BSST03, MPHS03]. Item [CWZ^{+}20b, CYC^{+}24, DK04, DSZ^{+}23, LRSD25, LWP^{+}21, SYSW24, WMZ^{+}21, WLT^{+}22, WQL^{+}19, WWG^{+}24, XZY21, XHW^{+}19, YCS^{+}14, ZLL^{+}20, ZCS^{+}23, ZCLX22]. Item- [SYSW24]. Item-based [DK04, XHW^{+}19]. Items [CLL24, CWZ^{+}20a, HCC^{+}17, LLW^{+}21, LFW23, SMT25]. Itemsets [BCFG15, CGB19]. Iterative [XZZ^{+}22, ZWMH19]. iVIBRATE [CL06].

January [Mar07]. Job [QZX^{+}20, QZS^{+}24, SQZ^{+}22, SJZ^{+}25]. Join [BF98]. joins [Coh00]. Joint [BLWJ15, CCCD19, LZG^{+}18, LLBS18, NLF^{+}20, SQZ^{+}22, WZC^{+}25, XLW^{+}22, YLZW22, YCZ^{+}16, PTS13]. Joint-Encoding [WZC^{+}25]. Jointly [CHN^{+}17, MLZ^{+}20, OSV19, YSZ^{+}17, ZLL^{+}22]. Judge [ZZH^{+}18]. Judgment [RZM24, XWZ^{+}25, ZDZW23]. Judgments [GTW^{+}18, MAA^{+}25, RLS^{+}24, SWL^{+}23, Wil98]. Judicious [ZNL^{+}17]. June [Mar08b]. Jurisprudential [XWL^{+}25]. Justification [WAL^{+}21].

Kabiria [CFP95]. Keeping [AARC20, BCD^{+}08]. Key [LPC^{+}23]. keyblock [ZRZ02]. keyblock-based [ZRZ02]. keyphrase [WX10]. keys [HZW02]. Keyword [ZGCC18]. Keywords [YDW22]. KG4RecEval [ZWS^{+}25]. Kinematics [CV98]. KNET [CGB^{+}15]. Know [SGJ22]. Knowing [MRM24]. Knowledge [AMS20, ACM^{+}02, BWS^{+}17, CFP95, CZX^{+}19, CSJ^{+}24, CACW23, CGB^{+}15, CSZ^{+}22, CCZ^{+}23, CYZ^{+}24, CQZ^{+}25b, DXL^{+}22, FTZZ23, FSE^{+}24, HLW^{+}24, HLH^{+}25, HZS^{+}25, HQZ17, HLC^{+}20, JEAZ25, LHL23, LZZ^{+}23, LBZ^{+}24, LDN^{+}25, LHA^{+}25a, LWC^{+}25, MHLH23, MLL^{+}24, MZG^{+}24, MLZ^{+}22b, MBJK90, PLS^{+}24, QGW^{+}24, Ruh94, SSGC89, SSY^{+}17, TSNS19, TCSL24, Tuz95, WZW^{+}19, WWT^{+}25, Wil98, Won94, WWN22, XZCZ25, XWZ^{+}25, YZS^{+}24, YZC^{+}25, ZJL^{+}24, ZWMH19, ZMK^{+}20, ZKMM23, ZWF^{+}23, ZWS^{+}25, ZZH^{+}25, ZWC^{+}23, ZCLC23, ZH17, ZSLK24, DDH^{+}09, WX10, XC05, YHYO01]. Knowledge-Aware [MHLH23, YZC^{+}25, DXL^{+}22]. Knowledge-Based [CFP95, SSGC89, Tuz95, WWT^{+}25, Won94]. Knowledge-Enhanced [AMS20, MZG^{+}24, ZWF^{+}23, ZSLK24, CSJ^{+}24]. Knowledge-Enhancing [FSE^{+}24]. Knowledge-grounded [HZS^{+}25]. Knowledge-Guided [MLZ^{+}22b]. Korean [ANW17]. KR [MHLH23]. KR-GCN [MHLH23].

Label [ES13, WL22, ZZD^{+}20, YZZS24]. Labeling [AZK12, RLC^{+}25]. Labels [WXR^{+}25]. Lake [AARC20]. LambdaGap [ASMZ25]. Language [AAL^{+}20, ANW17, CSJ^{+}24, CGD^{+}25, CJY23a, CPL15, DHC^{+}25, FLW^{+}25, GR96, HLL^{+}25, HYM^{+}25, JSHS96, KJA20, LMZ^{+}23a, LWL^{+}25, LYP^{+}25, LDX^{+}25, LLC^{+}25b, LHZ^{+}25, LLN^{+}25, MWC^{+}25, MAA^{+}25, MK16b, PGZ^{+}25, QZW^{+}25, QXX^{+}25, THL^{+}25, TL14, Tuz95, WCP^{+}25, WWHQ25, WZY^{+}25, WZC^{+}25, XSWX25, YFO^{+}25, ZHW^{+}22, ZGC^{+}22, ZWS24, ZXH^{+}25, ZLRW24, ZPW^{+}24, Coh00, ES98, FBN^{+}12, FG04, KL09, KL10, Owe02, TLJ^{+}07, TFS^{+}12, XHYY09, ZL04]. language-dependent [TFS^{+}12]. language-independent [TFS^{+}12]. Languages [PYDR24, SW94, ACS08, KMdRS06]. Large [CJRY12, CGD^{+}25, CS94, CTH^{+}13, DHC^{+}25, FLW^{+}25, Gla97, HLL^{+}25, HYM^{+}25, KC15, LMDZ20, LDX^{+}25, LZC^{+}25b, LLC^{+}25b, LHZ^{+}25, LLN^{+}25, MWC^{+}25, MAA^{+}25, MP20, NLF^{+}20, PGZ^{+}25, QZW^{+}25, SC19, THL^{+}25, VKD21, VFA17, VB98, WXLC13, WCP^{+}25, WWHQ25, WZY^{+}25, WZC^{+}25, XSWX25, YFO^{+}25, ZJL^{+}24, ZTM^{+}17, ZXH^{+}25, BC13, BJL^{+}07, CW02, CL06, EKK00, SSCT09, YKHW06]. Large-Alphabet [MP20]. Large-Scale [CTH^{+}13, LMDZ20, NLF^{+}20, SC19, VFA17, WXLC13, ZJL^{+}24, CJRY12, LZC^{+}25b, VKD21]. Last [ZK20]. Late [LGM^{+}24, GL05]. Late-interaction [LGM^{+}24]. Latency [BACF17, TM20]. Latent [BLWJ15, CGB19, Hof04, LHH^{+}19, LZWZ24, LLBS18, QWZ^{+}15, SMT25, WXLC13, KO98, PR09]. Latinne [EMS21]. lattice [CSLN10]. lattice-based [CSLN10]. Law [PSL16, XWL^{+}25, ZDW^{+}22]. Layers [ZSY^{+}21]. LBSNs [LMZ23b]. LCARS [YCS^{+}14]. LDA [ZHD^{+}22]. LDA-based [ZHD^{+}22]. Leadership [CSB21]. Leakage [JSZL23, XYW^{+}23b]. Lean [Ruh94]. Learn [HSYX22]. Learned [MTL23, PCD92]. Learning [AMS20, AAL^{+}20, AYWM21, AGT24, ADW94, CZZ^{+}17, CLL24, CSR^{+}20, CWM^{+}24, CZJH24, CQZ^{+}25a, CGB^{+}15, CSZ^{+}22, CYZ^{+}24, DXZ^{+}22, DZX^{+}23, DMUN19, DSZ^{+}23, EMS19, FZSG21, FZH^{+}22, FZXdR24, FC17, FTZZ23, FSE^{+}24, FWG^{+}24, FB91, FP94, GCL^{+}21, GYC^{+}22, GLG^{+}23, GCB^{+}25, GSG^{+}25, HLW^{+}24, HLL^{+}17, HCC^{+}16, HWL^{+}24, HLC^{+}20, HSYX22, IC16, ISW92, JMZ22, JDZ^{+}22, JZZW24, KSYE24, LAO^{+}24, LZC^{+}25a, LWSA10, LZC23, LZWZ24, LYY^{+}24, LZL^{+}25, LLC^{+}25a, LWL^{+}24, LCNY23, LLZ^{+}25, LGY^{+}25, LNO^{+}25, LWC^{+}25, MHW^{+}24, MXM^{+}24, MC94, MLZ^{+}20, MZG^{+}24, MCL^{+}12, MPS23, MLZ^{+}22b, NXP^{+}24, NXF^{+}24, NZW^{+}14, NLF^{+}20, Oos23, PCC^{+}22, PGC^{+}25, PS05, QLNY19, QZW^{+}25, QZS^{+}23, QGW^{+}24, RSD^{+}13, RYC^{+}20, RZCG^{+}10, SC19, SHS^{+}23, SQZ^{+}22, SLX^{+}24, SSZ^{+}23, SSY^{+}17, SCG^{+}19, SZM^{+}23, SJZ^{+}25, TSNS19, TLKW16, TZG^{+}24, TWS^{+}25, TXC^{+}24, UA22, WLW^{+}24a, WNS^{+}17, WC21a, WZZ^{+}24b, WGL^{+}24, WWR^{+}25, WWHQ25, WQL^{+}19, WXR^{+}25, WPS^{+}25, XLW^{+}22, XCG^{+}23, XLL^{+}23, XYCC24, XLX^{+}25, XYZ^{+}23]. Learning [YW15, YLZW21, YLZW22, YSW^{+}24, YCL^{+}25, YE18, YDXW21, YZZS24, YHDN25, YZC^{+}25, ZZZ^{+}24, ZZT^{+}21, ZSQ^{+}24, ZWRL19, ZYY^{+}21, ZLL^{+}22, ZCZW22, ZDZW23, ZWF^{+}23, ZBYH23, ZXM^{+}24, ZJK^{+}24, ZSZ^{+}24, ZSH^{+}25, ZGL^{+}25, ZZH^{+}18, ZHC^{+}20, ZYL^{+}25, ZCLX22, ZHRK23, ZYY^{+}24, AAGY01, CL03, CS99, HJZL09, LCH07, LLG^{+}23, MSOH13, SZC^{+}23, SC03, GWLC06]. Learning-Based [SC19, YDXW21, HLW^{+}24, XLL^{+}23, ZZT^{+}21]. Learning-enhanced [WLW^{+}24a]. Learning-to-Rank [IC16, LNO^{+}25]. legacy [LL99]. Legal [LHA^{+}25a, LHA^{+}25b, MWA^{+}24, SWL^{+}23, SLW^{+}24, SAW^{+}25, XWZ^{+}25, XWL^{+}25, YMZ^{+}22, ZDZW23]. LegalGNN [YMZ^{+}22]. Length [Na15, QDG^{+}24]. Less [PSM24, ZZSS25]. Lessons [PCD92, LWA08]. Level [CLM^{+}22, DC92, FWG^{+}24, OS96, WWX^{+}24, WGL^{+}24, XZY21, AC08, HYLC24, LGL^{+}24, LCW^{+}24b, SYSW24, WMGZ24, ZLS^{+}23, ZYW^{+}23]. Levels [ZZH^{+}18]. Leveraging [BACF17, CCZ^{+}19, FAZC12, ZSN^{+}22]. Lexical [KC92, LL23a, PPGK04]. libraries [GFWK04, LL99, SNS^{+}05]. Library [EGL^{+}97]. lies [PR04]. Life [Mar94, CCC09]. Lifestyle [ZZZ^{+}16]. Lifestyle-Based [ZZZ^{+}16]. Lifetime [ZWT^{+}23]. Light [PGC^{+}25]. LightFR [ZLW^{+}23]. Lightweight [LMR^{+}24, LCYY25, TCQ^{+}25, YLZW22, ZLW^{+}23, ZLZ25a, ZCLC23, BFNP10]. Like [JXY^{+}24]. likelihood [KT05, WNCH24]. Limit [ZZSS25]. Limited [ZAL^{+}16]. linear [Non13]. linear-time [Non13]. Lingual [EMS19, JTS16, MPS23, XZZ^{+}22, GNN^{+}10]. Linguistic [GJS^{+}25, CHS99]. Linguistic-Visual [GJS^{+}25]. Link [CP13, SG14, YMZF22, ZMK^{+}20, CRNZ^{+}03, LM01]. Link-Based [YMZF22]. link-structure [LM01]. Linkage [ZZG^{+}25]. Linking [CFC^{+}19, LCL^{+}17, LGQL19]. Links [ZHC^{+}20, KL10]. Lists [ABCZ21, SKK24, SKC16]. Listwise [IC16, TZG^{+}24, WHL^{+}16]. LkeRec [YLZW22]. LLM [GJS^{+}25, HZS^{+}25]. LLM-Enhanced [GJS^{+}25]. LLMCDSR [XSWX25]. Load [MAH24, PSPBY10, PSPBY10]. Load-driven [PSPBY10]. Local [CRNZ^{+}03, CWZ^{+}20a, LPC^{+}23, LJQY19, Row96, SZS17, WNS^{+}17, ZLWW22, XC00]. Locality [PLS^{+}24, dMdSsdA^{+}08]. Locality-Based [dMdSsdA^{+}08]. Localization [CTH^{+}13, HWL^{+}24]. Localized [WLL^{+}23a]. Locating [ZK20]. Location [CS16, LCW^{+}24b, LZG^{+}18, MKZPK18, MPT21, WLFT18, WHFG92, YE18]. Location-Aware [CS16, WLFT18]. Log [JTS16]. Logic [Tuz95, YWL^{+}25]. Logical [BSST03, HYLC24, ISL95, LYG^{+}25]. Logical/Physical [ISL95]. Logs [LHG^{+}12, BJL^{+}07, GNN^{+}10, LOP^{+}13, WNZ02]. Long [AVA^{+}24, GCN^{+}19, LPC^{+}23, SJZ^{+}25, XPL^{+}24, XYX^{+}25, YZZS24, ZLF^{+}17, TZP22]. Long- [XPL^{+}24]. Long-tailed [YZZS24]. Long-Term [XYX^{+}25, SJZ^{+}25]. Look [Jac91]. Loops [XYX^{+}25]. Loose [WC21b]. Loss [WWG^{+}24, YZZS24, ZPW^{+}24]. Lossless [GYW^{+}25]. Low [LJQY19, SZZL23, ZZZ^{+}25]. Low-Pass [ZZZ^{+}25]. Low-rank [LJQY19]. Low-Resource [SZZL23]. LTP [XYX^{+}25]. LTP-MMF [XYX^{+}25].

M [WGL^{+}24]. M3GAT [ZJW^{+}24]. Machine [LAO^{+}24, LPY94, TL14, WGZ^{+}25, WLM^{+}21]. Machine-Readable [LPY94]. Machines [LPLH19]. Magnitudes [MMST17]. Mail [RP98]. Maintaining [ZLH^{+}25]. Make [CR92]. Making [EGL^{+}97, OG94, Won94, BDS04, WLWK08]. MAN [HLW^{+}24]. Management [AM89, HEA90, LAL^{+}17, WLZW21, ZHZ^{+}19, BBMN^{+}08, BVKS08, BJS99, DEL^{+}00, Mar08a, MC10b]. managing [HDPM08]. Manipulating [YYY^{+}24]. Manipulation [HSH90, PL94, ZQH^{+}24a]. Manner [SSZ^{+}23]. manuscript [Mar08a]. Many [LGZ^{+}24, RLS^{+}24]. Many-to-one [LGZ^{+}24]. Mapping [YC94]. Market [LCJ^{+}16, SJZ^{+}25, SC09, WMJ05]. Market-aware [SJZ^{+}25]. market-based [WMJ05]. Marketing [YFO^{+}25]. Markets [GGRF23]. Markov [JZ06, YGZ15, ZZW^{+}25]. Marks [LPYM23]. Massive [PGW^{+}17, PV19, CC03]. Match [SW90, VB98]. Matching [AARC20, DXZ^{+}22, GJS^{+}25, KJA20, LJZ^{+}25, LL23a, LGM^{+}24, MKF91, MJH15, MH89, MPT21, SLM^{+}24, XPL^{+}24, XXDW24, XWL^{+}25, ZLF^{+}17, CKC07, LCH07]. matrices [GCT99]. Matrix [CZZ^{+}20, FMSS14, LZG^{+}18, LJQY19, SZS17, ZLW^{+}23, ZZZ^{+}25, KO98]. Matter [ZWS^{+}25]. Matters [LLT^{+}21, SLQ24]. Max [XYX^{+}25]. Max-Min [XYX^{+}25]. maximal [GWLC06]. Maximization [GRSD^{+}16, VFA17, WLFT18]. maximizing [CL03]. Maximum [BNIL24a, BNIL24b, CZZ^{+}17, GP00, KT05]. May [Mar07, Mar08b]. MCAP [ZZZ^{+}25]. MCN4Rec [LCW^{+}24b]. MCRPL [LGZ^{+}24]. Me [YHY^{+}17]. Meal [YHY^{+}17]. Mean [FJZS24]. Meaningful [GCR^{+}21, ZFL^{+}07]. means [BGQT07]. measure [WMZ10]. Measurement [BA18, GR12, MBST17, MZ08]. Measures [FFM17, Fer17, Gin24, RJB89, RZM24, SZ21, FKM^{+}05, KL06]. Measuring [CLF^{+}25, LXLZ16, TL03, AV02]. mechanical [HDPM08]. Mechanism [CST^{+}23, HCY^{+}20, ZYW^{+}23, ZMN^{+}24, BJS99]. Mechanisms [LGM^{+}24]. Media [VST^{+}14, LW20, LBC^{+}19, MK16a, MKZPK18, PBJ^{+}16, WZC^{+}17, ZLL^{+}22, ZY21, CDR10, GC92]. mediated [LWA08]. Medical [AZK12, Luz12, WH09]. Medication [BLH^{+}22, LQZ^{+}25]. Medicine [ZWF^{+}23]. Medium [ACC18, Mar94]. Meet [ZAK^{+}23]. meetings [Luz12]. Meets [BOQ^{+}25, PXW^{+}25, ZLW16]. MEGA [CL03]. MEGCF [LXG^{+}23]. Memoriam [Ano96b]. Memory [Ack98, AC19, HLW^{+}24, NCZG14, NCHW15, ZHZ^{+}19, ZQH^{+}24b, EBR08, TS10]. Memory-Augmented [ZHZ^{+}19, HLW^{+}24]. Mental [TBCW20]. merge [SC03]. MergeDTS [LMDZ20]. merging [SZ09]. merit [GWLC06]. merits [Mar08a]. Message [KM97, Zha21]. Meta [BHH^{+}16, CSZ^{+}22, KMPS21, LYY^{+}24, LLG^{+}23, LZW21, LWL^{+}24, NXF^{+}24, SZC^{+}23, SS08, WGL^{+}24, WCZ^{+}25, WPS^{+}25, ZMN^{+}24, ZZH^{+}25, NXF^{+}24]. Meta-CRS [NXF^{+}24]. Meta-evaluation [LZW21, WCZ^{+}25]. Meta-Framework [BHH^{+}16]. Meta-Information [KMPS21]. Meta-Learning [CSZ^{+}22, LYY^{+}24, LWL^{+}24, NXF^{+}24, WPS^{+}25, LLG^{+}23, SZC^{+}23]. Meta-Level [WGL^{+}24]. Meta-relation [ZZH^{+}25]. Meta-representations [ZMN^{+}24]. metadata [RBV09]. Metagraphs [FZH^{+}22]. Metaphorical [SGZ^{+}24]. Metasearch [DH97, MWYL01]. Method [CYZ^{+}24, EMS19, HYZ^{+}23, Koi95, LMDZ20, LGQL19, SZ93, YC94, ZZD^{+}20, ZLW24, FSC07, MWYL01, PRP05, PH07, QZXP06, SC03]. Methodologies [SLZ^{+}19]. methodology [FCD02, SNS^{+}05]. Methods [APCC15, AK18, CSZD15, DLA15, GCC12, HT99, HWD13, MC94, SBH24, WWHX23, ZZL^{+}23, CS99, SS08, TFS^{+}12, ZL04, dMdSsdA^{+}08]. Metric [CCVR17, ES98, LAO^{+}24, LMW^{+}21, PLR^{+}22, WCZ^{+}25, ARSZ03]. Metric-Aware [PLR^{+}22]. Metrics [BRS92, LZW21, MMCC^{+}21, Pal15]. MF [WYZ^{+}25]. MF-GSLAE [WYZ^{+}25]. MFoM [GWLC06]. Micro [LWL^{+}24, QDG^{+}24, YHDN25]. Micro-Video [LWL^{+}24, QDG^{+}24]. Microblog [ZNL^{+}17, ZHC^{+}20]. Microblogging [MCF^{+}17]. Microtask [WNWL^{+}21]. MiDTD [LYLS22]. Min [XYX^{+}25]. Minimal [FCDH91]. Minimizing [OSV19]. Minimum [Yan15]. Mining [AARC20, HL17, HZSL13, HHC^{+}10, KLHK13, WLS^{+}17, WC21b, YHS^{+}12, ZH17, LCH07, LCL04, WZZ04]. Misidentified [Coo95]. Misinformation [ZAL^{+}16]. Mitigating [LFW23, MAP^{+}22]. Mixed [KKS^{+}08, VKD21]. Mixed-initiative [KKS^{+}08]. Mixture [GCB^{+}25, PYDR24]. Mixture-of-Experts [GCB^{+}25]. Mixture-of-Languages [PYDR24]. MKCRec [ZZH^{+}25]. MLI [ZYW^{+}23]. MLP [GZL^{+}24, JXY^{+}24]. MLP-Like [JXY^{+}24]. MMALFM [CCZ^{+}19]. MMF [XYX^{+}25]. Mneme [Mos90]. Mobile [AZCC21, AK18, BWS^{+}17, DCT^{+}17, Guy18, HYH15, HL17, LAL^{+}17, TZP22, TZP23, USN^{+}17, WLS^{+}17, WLG^{+}17, WDG17, YSZ^{+}17, YZW^{+}17, ZML^{+}19, ZWL^{+}22].

Mobility [SSY^{+}17, ZSQ^{+}24]. Modal [HWL^{+}24, WLZ^{+}24, CJN20, LXZ^{+}22, XLW^{+}22, ZBYH23, ZJW^{+}24, ZZL^{+}22, ZWL^{+}23]. Modality [GCL^{+}21, LLC^{+}25a, YHDN25]. Modality-Disentangled [LLC^{+}25a]. mode [HJZL09]. Model [AZBC20, CZZ^{+}17, CLL24, CSJ^{+}24, CZJH24, CJY23a, CH96, CDM15, CPL15, FJZS24, FCL^{+}23, FP00, GPS93, GJS^{+}25, GCB^{+}25, HZS^{+}25, JEAZ25, KMKO92, KSYE24, LAO^{+}24, LCX^{+}19, LSL24, LWL^{+}25, LLM22, LXW^{+}17, LLZ^{+}21, LQZ^{+}25, LHZ^{+}25, MO12, MMLP97, Mye90, NB97, PGZ^{+}25, PL94, QWJ^{+}24, RN96, SLM^{+}24, SFR98, SWZ^{+}22, SLQ24, TBC96, THL^{+}25, Tom89, TC91, WGL^{+}24, WZY^{+}25, WZC^{+}25, WS90, XZH^{+}22, XZT^{+}24, YGZ15, YZZS24, ZWMH19, ZGF^{+}20, ZMK^{+}20, ZWX^{+}24, ZXH^{+}25, ZZZ^{+}16, ZHD^{+}22, ZML^{+}19, ZWL^{+}22, ZCLX22, ZPW^{+}24, AF07, BFF^{+}03, GFWK04, LLPT10, PZMRN05, RR98, SS08, TKD09, XHYY09, YLS^{+}11, dOTM01]. Model-agnostic [XZT^{+}24]. Model-Assisted [CH96]. Model-Based [TBC96, WGL^{+}24, LAO^{+}24, WZY^{+}25]. Model-driven [FP00]. Modeling [BZDW24, BHH^{+}16, BLWJ15, BKR97, CHN^{+}17, CML^{+}21, Coo95, CCZ^{+}23, DLZ^{+}22, DDSW24, DCT^{+}17, DHY^{+}19, GCH^{+}19, Gul96, GCN^{+}19, HRY^{+}22a, HRY^{+}22b, ISL95, LDW^{+}17, LZG^{+}18, LLBS18, LZLM19, LWH^{+}23, QXX^{+}25, RPE^{+}18, SLHS93, SG14, WXLC13, WHL^{+}22, WA22, WY95, XZY21, XC13, YGZ15, YSZ^{+}17, YE18, YCZ^{+}16, ZLWW22, ZHY14, ZXZL23, ZH17, ZZG^{+}25, ZYY^{+}24, ZSLK24, BBD11, PGD07, RSD^{+}13]. Modelling [ZXZL20]. Models [CWZ^{+}20b, CWZ^{+}20a, CQL^{+}24, CGD^{+}25, Cum14, DLA22, DRdR23, DL16, DHC^{+}25, FTZ11, FLW^{+}25, Gin24, GCF^{+}22, HLL^{+}25, HYM^{+}25, JDZ^{+}22, LMZ^{+}23a, LYP^{+}25, LZW^{+}20, LDX^{+}25, LGM^{+}24, LZC^{+}25b, LLC^{+}25b, LNO^{+}25, LLN^{+}25, MWC^{+}25, MMH^{+}24, MAA^{+}25, MNP^{+}21b, PARS22, QZW^{+}25, QWZ^{+}15, RBM23, SQZ^{+}22, TZG^{+}24, THL^{+}25, TBCW20, WRLD22, WCP^{+}25, WWHQ25, WLF^{+}25a, WLF^{+}25b, WZG^{+}23b, WZG^{+}23a, XSWX25, YFO^{+}25, ZHW^{+}22, ZWS24, ZOML24, ZLRW24, ZWrWZ23, AV02, Bod04, CKC07, GP00, Hof04, JZ06, KSNM10, KL09, KL10, LA08, MSOH13, RZCG^{+}10, WdVR08, WRJV05, ZL04]. Modern [ZQH^{+}24a]. modes [IH07]. modular [TMT06]. Moment [HWL^{+}24, WWT^{+}25]. Monitoring [CTH^{+}13]. Monitors [GH96]. Monte [LCL^{+}17]. Morphological [CMR91, CGB^{+}15]. Motion [DG95]. Mouse [APJL20]. Movements [Jac91]. Movie [KEL95, ZSN^{+}22]. MovieLens [FJZS24]. MR [SGLS93]. MRC [CJY^{+}23b]. MUADDIB [RSG09]. Multi [AGT24, BHY^{+}25, CRC^{+}22, CJY23a, CCZ^{+}23, DLZ^{+}22, DZX^{+}23, DDSW24, FTZZ23, FLW^{+}25, GCL^{+}21, GC92, GLG^{+}23, HYZ^{+}23, HWT^{+}25, HJZ^{+}23, JDA^{+}22, LZC^{+}25a, LLT^{+}21, LXZ^{+}22, LCW^{+}24b, LMZ23b, LYN^{+}21, LGQL19, LXC^{+}25, LQZ^{+}25, LPLH19, LLP24, MHW^{+}24, MAP^{+}22, MZG^{+}24, PBS22, PZD^{+}22, QZW^{+}25, QFZ^{+}19, QXX^{+}25, SCG^{+}19, SZM^{+}23, SLZ^{+}24, WWX^{+}23, WWX^{+}24, WZW23, WZZ^{+}24b, WZY^{+}24, WFW^{+}25, WYZ^{+}25, WMM^{+}23, XTF^{+}21, XCG^{+}23, XPSC24, YLZW21, YCG^{+}24, YCL^{+}25, YPYM24, YLW^{+}23, YZZS24, YRC23, ZZZ^{+}24, ZWRL19, ZZD^{+}20, ZLL^{+}22, ZWZ^{+}22, ZWF^{+}23, ZLS^{+}23, ZYW^{+}23, ZCZ^{+}23, ZJW^{+}24, ZWX^{+}24, ZDW^{+}22, ZWrWZ23, ZZL^{+}22, ZLZ^{+}25b, ZQH^{+}24b, HL10]. Multi-Agent [WFW^{+}25]. Multi-annotator [ZWRL19]. Multi-aspect [WZZ^{+}24b]. Multi-auxiliary [YRC23]. Multi-Behavior [LZC^{+}25a, MZG^{+}24, YCG^{+}24, YCL^{+}25, ZLZ^{+}25b, YPYM24].

Multi-center [LQZ^{+}25]. Multi-Channel [LPLH19, WWX^{+}24, SLZ^{+}24, WWX^{+}23]. multi-document [HL10]. Multi-Domain [FLW^{+}25, ZWX^{+}24, QXX^{+}25]. Multi-evidence [YLW^{+}23]. Multi-Factor [WYZ^{+}25]. Multi-factors [HJZ^{+}23]. Multi-Field [QFZ^{+}19]. Multi-goal [DZX^{+}23]. Multi-grained [DDSW24]. Multi-granularity [SLZ^{+}24]. Multi-Graph [ZWZ^{+}22]. Multi-Hop [ZQH^{+}24b]. Multi-Interest [HWT^{+}25, MZG^{+}24, CRC^{+}22]. Multi-Label [ZZD^{+}20, YZZS24]. Multi-level [LCW^{+}24b, ZLS^{+}23, ZYW^{+}23]. Multi-media [GC92]. Multi-modal [LXZ^{+}22, ZJW^{+}24, ZZL^{+}22]. Multi-Modality [GCL^{+}21]. Multi-Objective [PBS22, SCG^{+}19, WMM^{+}23, JDA^{+}22]. Multi-party [WZW23]. Multi-pattern [LMZ23b]. Multi-perspective [DLZ^{+}22]. Multi-Relational [CCZ^{+}23, PZD^{+}22, ZDW^{+}22]. Multi-Response [YLZW21]. Multi-Sided [MAP^{+}22]. Multi-Stage [LYN^{+}21]. Multi-Stakeholder [WMM^{+}23]. Multi-strategy-based [HYZ^{+}23]. Multi-target [GLG^{+}23]. Multi-Task [LXC^{+}25, ZWF^{+}23, BHY^{+}25, DZX^{+}23, SZM^{+}23, XCG^{+}23, XPSC24, ZLL^{+}22, ZJW^{+}24]. Multi-Turn [GCL^{+}21, LLT^{+}21, CJY23a, FTZZ23]. Multi-types [XTF^{+}21]. Multi-View [AGT24, WZY^{+}24, ZWrWZ23, ZQH^{+}24b, LLP24, MHW^{+}24, QZW^{+}25, ZZZ^{+}24]. Multi-View-Based [LGQL19]. MultiCBR [MHW^{+}24]. Multidimensional [GH96, ASST05, QZXP06]. multidisciplinary [Luz12]. Multidocument [MLDGH04, YW15]. Multifocal [WWX^{+}25]. Multilevel [MMLP97]. Multilingual [BCFG15, CGB19, JTS16, KMKO92, LZG21, PYDR24, ZJL^{+}24, ZOML24, KSNM10].

Multimedia [CJN20, LTC^{+}25, RRS95, ZZL^{+}22, DDH^{+}09, MR10, ZHC^{+}13]. Multimodal [CSJ^{+}24, JDA^{+}22, LLC^{+}25a, LXG^{+}23, LWL^{+}24, MYHL11, TQH15]. Multiple [Car12, CH96, DPSH23, FC17, FFM17, LZWZ24, LPY94, MC14a, MK94, NGMD17, SMN^{+}16, WLT^{+}22, WMC^{+}23, Wil98, XZY21, ZLL^{+}22, ACS08, CRG02, GWR99, MSOH13]. Multiplex [YSW^{+}24]. Multiuser [DC92]. Multivariate [YZH^{+}24, LA08]. Music [CS16, JLC^{+}25, MO12, RVA^{+}24, WZY^{+}24, PH07, SPKW11, SSCT09]. MVideoRec [YHDN25]. MvStHgL [AGT24]. MWI [BCFG15]. MWI-Sum [BCFG15]. MyrrorBot [MNP^{+}21b].

N [SC19]. Name [All89]. Named [LCH07, XYCC24, ZWZ^{+}24, PTS13]. Names [SMGM11]. Narrative [ZSN^{+}22]. Natural [AAL^{+}20, GR96, KJA20, ZGC^{+}22, ES98, FBN^{+}12, Owe02]. Natural-Language [GR96, ES98]. Navigation [AKFA15, BBMN^{+}08, MS01]. Near [CTH^{+}13, JTL^{+}17, NGMD17, HHC^{+}10, LHSC11]. Near-Duplicate [CTH^{+}13, JTL^{+}17, HHC^{+}10, LHSC11]. Near-Optimal [NGMD17]. nearest [KZP10]. Need [LMW^{+}23, RLS^{+}24, SGJ22, YTD13, LWA08]. Needs [FEL22, VR24, KMdRS06]. Negative [CMZ^{+}23, LZL^{+}25, Mar94, WC21a, WC21b, WLL^{+}23b, WZZ04]. Negotiation [CW94, LWA08]. neighbor [KZP10]. Neighborhood [JDZ^{+}22, PZD^{+}22, WX10]. Neighboring [LPZ^{+}25]. Neighbors [SMT25]. NER [ZYC^{+}24]. Nested [PL94]. Nested-Graph [PL94]. Net [SF89]. Nets [OS96]. Network [AM25, CQF^{+}23, CXZ^{+}21, DXL^{+}22, GGR22, GLZ^{+}24, HCY^{+}20, HML^{+}23, JDZ^{+}22, JLC^{+}25, Kwo95, LLZ^{+}22, LZZ^{+}23, LBM^{+}24, LCW^{+}24b, LSZ^{+}25, LW20, LWP^{+}21, LLY^{+}24, MHLH23, MAL16, QZX^{+}20, SLZ^{+}25, SLW^{+}20, TCSL24, WWX^{+}23, WWX^{+}25, WZY^{+}24, WZW^{+}25, WWN^{+}25, XHX^{+}22, YCG^{+}24, YSZ^{+}17, YMZ^{+}22, YZH^{+}24, YHWZ24, ZHW^{+}22, ZJW^{+}24, ZWC^{+}25, ZLZ25a, ZY21, ZDW^{+}22, ZWE^{+}22, FCD02, LLPT10, XC05]. Network-based [ZHW^{+}22, FCD02]. Network=based [TC91]. networked [Fuh99, PSDB99]. Networks [DCT^{+}17, FLMD22, FZH^{+}22, FZXdR24, FC17, GRSD^{+}16, GLL^{+}23, HGBY16, HE91, HLW^{+}24, HQZ17, LHL23, LLBS18, MWZ^{+}25, MTC21, MLZ^{+}20, PGC^{+}25, PZD^{+}22, PSM24, QHLY20, QFZ^{+}19, RCR^{+}18, SZY^{+}24, SMN^{+}16, TLKW16, TLS^{+}22, UY89, VFA17, WL22, WHL^{+}22, XZZ^{+}22, XXDW24, YSZ^{+}17, YHS^{+}21, ZAL^{+}16, ZTM^{+}17, ZLH^{+}22, ZYW^{+}23, ZXM^{+}24, ZGCC18, ZDW^{+}22, ZXZL23, CGM05, KNS04, RBV09]. Neural [AM25, BLMN24, CQF^{+}23, CLL24, CCCD19, CZZ^{+}20, CXZ^{+}21, DXL^{+}22, DHY^{+}19, FLPC24, GLL^{+}23, HQZ17, KJA20, LMR^{+}24, LLX^{+}22, LCW^{+}24b, LZC24a, LYN^{+}21, LMMO23, MPT21, PZD^{+}22, QZX^{+}20, QHLY20, QFZ^{+}19, RFB^{+}24, RVA^{+}24, RCR^{+}18, SZY^{+}24, SLW^{+}20, TCSL24, TLS^{+}22, VdRK18, WWX^{+}25, WL22, WZHY24, WZG^{+}23b, WZG^{+}23a, XZH^{+}22, XPSC24, YSZ^{+}17, YMZ^{+}22, YWL^{+}25, YLYT23, YHWZ24, ZHW^{+}22, ZLH^{+}22, ZSL^{+}25]. Neuronal [ZPW^{+}24]. News [HLL^{+}17, HJZ^{+}23, LW20, LXLZ16, LMZ20, MYL^{+}23, MK16a, WQH^{+}25, WWHX23, XPL^{+}24, ZLF^{+}17, LSOS06, SC09, LSOS06]. Next [AGT24, CSZ^{+}22, HLH^{+}25, LHH^{+}19, LJAdR23, LCW^{+}24b, LCNY23, SLZ^{+}24, XC24, XZY21, ZHL^{+}22, ZLL^{+}20, ZCLX22]. Next-Item [ZLL^{+}20, ZCLX22]. NIR [XPSC24]. NIR-Prompt [XPSC24]. Node [LZWZ24, WHL^{+}22]. Nodes [MLZ^{+}20, NB97]. Noise [CQL^{+}24, CFE^{+}25, ZWRL19].

Noise-aware [ZWRL19]. Noise-Oriented [CFE^{+}25]. Noised [MWZ^{+}25]. Noisy [GTW^{+}18, SLM^{+}24, ZYL^{+}25]. Non [CMZ^{+}23, CJY23a, DLN^{+}16, GLZ^{+}24, GSG^{+}25, LGZ^{+}24, BM02, QZXP06]. non-affiliated [BM02]. Non-autoregressive [CJY23a]. Non-Oblivious [DLN^{+}16]. non-ordered [QZXP06]. Non-overlapping [GLZ^{+}24, GSG^{+}25, LGZ^{+}24]. Non-sampling [CMZ^{+}23]. Nonlinguistic [SCL^{+}22]. Nonmatching [Dun97]. nonordered [KZP10]. Nonparametric [YCZ^{+}15]. Nonuniform [HCY^{+}20]. nonverbal [Luz12]. Normalization [Na15, HO07]. normalized [KT05]. Note [Gru92]. Novel [LMW^{+}21, MS01, PRP05, PH07, SSCT09]. Novelty [SLX^{+}24]. NSP [WC21b]. number [KFS10]. Numeral [MP20]. Nutrient [YHY^{+}17]. Nutrient-Based [YHY^{+}17].

Object [AM89, GSR96, JSHS96, KBGW91, MT21, MH89, Mos90, Sci89, SÖ90, SM99]. Object-Oriented [AM89, GSR96, JSHS96, KBGW91, SÖ90]. Objective [IC16, PBS22, SCG^{+}19, UA22, WMM^{+}23, JDA^{+}22]. Objects [PL94, Gla04, GL05]. Oblivious [DLN^{+}16]. obtained [PTKJ07]. Occurrences [BKR97]. OCPC [ZZW^{+}25]. OCR [DM07, TBC96]. October [Cal12]. OED [BBT92]. Offensive [YTL25]. Office [CFP95, GZC89, PBF^{+}89]. Offline [AYWM21, CWM^{+}24, GZW^{+}18, JMO22, JSZL23, LCY21, RZM24, TSNS19, TTZ^{+}25, WGL^{+}24, ZLF^{+}23]. On-Device [XYW^{+}23a]. Once [NGMD17]. One [CLPM20, PYC^{+}19, THL^{+}25, YPYM24, IH07, LGZ^{+}24, WTS^{+}08]. One-Class [PYC^{+}19, YPYM24, CLPM20]. one-size [IH07]. Online [AYWM21, CP13, CL07, GGRF23, GZW^{+}18, LMDZ20, LZC^{+}25b, MCF^{+}17, MCL^{+}12, MNP^{+}21b, TSNS19, TTZ^{+}25, WZW23, ZLWW22, ZAL^{+}16, ZXZL20, ZZH^{+}18, ZWT^{+}23, ZZL^{+}22, ZQH^{+}24a, CC03]. Ontological [MSR04]. Ontology [YLW^{+}23]. Ontology-aware [YLW^{+}23]. Open [HZG20, HYM^{+}25, LMW^{+}21, LLN22, LCL^{+}23a, MLZ^{+}22a, PYF^{+}92, WWX^{+}23, ZZC^{+}25, BHMW07, MPHS03]. Open-Domain [LLN22, LCL^{+}23a, MLZ^{+}22a, HZG20, LMW^{+}21, MPHS03]. Operator [XZCZ25]. Operators [LYG^{+}25, EJ03, GCT99]. Opinion [FAZC12, GCC12, ZSZ^{+}24, ACS08]. Opinion-based [ZSZ^{+}24]. Opinions [NSK^{+}20]. Opportunities [CWM^{+}24, DLL^{+}25, MJH15]. Optical [FC91]. Optimal [FC91, NGMD17, QJG^{+}25]. Optimality [Row96]. Optimization [SZY^{+}24, WMM^{+}23]. Optimizing [JMD20]. Optimum [Fuh89]. Oracle [NWZC12]. Order [CLG^{+}23, FCDH91, LJHL17, WZW^{+}19, HZSL13]. Ordered [MPM22, QZXP06]. Ordering [STC^{+}24]. Organization [HE91]. Organizational [Ack98, BB93, KL00]. Organizations [CZ93, OG94]. organized [JDA^{+}22]. Organizing [HDPM08]. Orientation [UY89, TL03]. Oriented [ASMZ25, AM89, BZDW24, CFE^{+}25, DXZ^{+}22, GSR96, HGBY16, HSH90, HWN^{+}19, JSHS96, KL91, KBGW91, QZS^{+}24, SÖ90, WHL^{+}16, XMS^{+}25, ZHZ^{+}19, ZYD^{+}24, KL06, SGZ^{+}24, WLL24]. Our [FJZS24, BVKS08]. Out-of-Distribution [ZWC^{+}25]. Out-of-sequence [LDN^{+}25]. out-of-vocabulary [WKF^{+}12]. Outcome [MLK^{+}18]. Outdated [DLY^{+}25]. Outdoor [dODMS20]. OutdoorSent [dODMS20]. Output [KMKO92]. Oval [MLF95]. Overcome [RVA^{+}24]. overlapping [GLZ^{+}24, GSG^{+}25, LGZ^{+}24]. Overload [RVA^{+}24]. Overview [BKWZ15, PBJ^{+}16, THT12]. Ownership [YPF^{+}25].

Paced [CZJH24]. Package [ZWX^{+}23].

Packages [Ruh94]. Page

[JDA^{+}22, LGRK18, BKGM^{+}02, UCH03]. PageRank [BSV09, KL10, WTS^{+}08, WW10]. Pages [ANW17, STC^{+}24]. Pair [CJY^{+}23b, BS12]. Pairwise [YMZF22, ZMK^{+}20]. Palmtop [FZC93]. Paper [LCPD19, ZZD^{+}20, ZZH^{+}25]. Paper-Reviewer [ZZD^{+}20]. PARADE [LYM^{+}24]. Paradigm [LGZ^{+}24, WCZ^{+}25, XMS^{+}25, ZYC^{+}24, CDH^{+}09]. paradox [KM13, JMO22]. Parallelism [MKF91]. Parallelization [PGW^{+}17]. Parallelized [YHS^{+}12]. Parameter [YYQ^{+}25]. parameters [Bod04, HO07]. Parity [YXL^{+}23]. Parser [MH89]. Parsimonious [YE18]. parsing [CHS99]. Part [HRY^{+}22a, HRY^{+}22b, WLF^{+}25a, WLF^{+}25b, YSXK23a, YSXK23b]. Partial [SGJ22]. Partially [LM90]. Participant [LLBS18]. Partitioned [CZ93, LL89, LR96]. Partitioning [ZRT91, QZXP06]. Parts [RK19]. party [WZW23]. Pass [ZZZ^{+}25]. Passage [LYM^{+}24, LYN^{+}21, SDZW24, TM19, KZSD99, KKB10]. Passage-aware [SDZW24]. Passages [MNP^{+}21a, JZ06]. Patent [MC14a, ZYY^{+}24]. Path [LL23b, ZWC^{+}23]. Pathway [UA22]. Pathways [YLW^{+}23]. patient [Luz12]. Pattern [LHH^{+}19, MKF91, VB98, WC21b, CKC07, LMZ23b]. Patterns [CHBS15, RN96, ZTZL25, ZYY^{+}24, IM05]. Peer [CGM02, LLZ^{+}22, THT12, CGM05, KNS04]. Peer-to-Peer [THT12, CGM02, CGM05, KNS04]. Peers [XHSD24]. people [Tee08]. PEPT [PXW^{+}25]. Perception [LMW^{+}23, WLZW21]. Perceptions [CCAC24]. Perceptual [AC19]. perCLTV [ZWT^{+}23]. Perfect [FCDH91]. Performance [ASSC22, ABW93, BMMC19, Cum14, DGMG23, FJZS24, FKS19, GH96, LL89, MAA^{+}25, MPHS03, NWZC12, RJB89, SKC^{+}12, SKC16, TM20, TCQ^{+}25, VAJR22, ZQH^{+}24a, BS12, CML00, FPSO06, KFS10, PF03]. Periodic [AGT24]. Permission [WLG^{+}17]. Perseus [MC94]. persistence [PPGK04]. Persistent [Mos90]. Person [LXZ^{+}22, NLL^{+}24, QZX^{+}20]. Person-action [NLL^{+}24]. Person-Job [QZX^{+}20]. Persona [HZS^{+}25]. Persona-based [HZS^{+}25]. Personal [AZCC21, GZW^{+}18, HJZ^{+}23, YDW22, BCD^{+}08, BBMN^{+}08, BVKS08, HDPM08, MKR04, MC10b, SS08]. Personalised [LCPD19]. Personality [CSH^{+}25, WCS^{+}24]. Personality-affected [WCS^{+}24]. Personalization [MWC^{+}25, ZYY^{+}21, ZHD^{+}22]. Personalized [ABCZ21, AC18, BZDW24, BTP24, DLZ^{+}22, GCN^{+}19, JLC^{+}25, LJHL17, LZC23, LZLM19, LXG^{+}23, LDN^{+}25, MNP^{+}21b, PLR^{+}22, PXW^{+}25, SZZL23, TLS^{+}22, WZW^{+}22, WDL^{+}24, WWHX23, XPL^{+}24, YHY^{+}17, YDXW21, YDW22, YHWZ24, ZZD25, ZWT^{+}23, ZLZ^{+}25b, MPS07, XHYY09, LLT^{+}21]. Personalizing [BLH^{+}22]. personally [ZFL^{+}07]. Perspective [LRSD25, ZZY^{+}23, DLZ^{+}22, HDPM08, PVGM11]. Perspectives [Kli94, NFZ19]. Perturbation [YLYT23]. Perturbations [CHZ24]. pervasive [ACM^{+}02]. Petri [OS96, SF89]. Phases [ZLW24]. Philosophy [DC10]. Phone [Res93]. Phone-Based [Res93]. phrase [WZB04]. Physical [ISL95, WNS^{+}17, YTD13, BSST03]. PIC [GCT99]. PicASHOW [LS02]. Pictorial [MTKY92, LS02]. Piggyback [CFC^{+}19]. Pipelines [LAO^{+}24]. Pitfalls [LMZ^{+}23a]. Placement [FC91]. places [ZFL^{+}07].

Placing [FGM^{+}02]. Planning [WLL24]. Platform [CHN^{+}17, LQJ^{+}24, RVA^{+}24]. Player [HZS^{+}25]. Plus [SS93]. PocketLens [MKR04]. POI [AGT24, CSZ^{+}22, HLH^{+}25, JMZ22, LHH^{+}19, LCNY23, QLNY19, QWJ^{+}24, SLQ24, SLZ^{+}24, ZHL^{+}22, ZLW16]. Poincaré [GLL^{+}23]. Point [AC18, LJHL17, MMH^{+}24, RABC22, SBH24, XC24, YCZ^{+}16, ZCS^{+}23]. Point-of-Interest [LJHL17, RABC22, YCZ^{+}16, MMH^{+}24, XC24]. Points [AK18]. Points-of-Interest [AK18]. Pointwise [IC16]. POIs [HY20]. Poisoning [TQN^{+}23]. Policy [DRdR23, MLL^{+}24, ZGL^{+}25, ZWX^{+}23]. Policy-driven [MLL^{+}24]. Pólya [CPL15]. PolyCF [QJG^{+}25]. Polynomial [Fuh89]. PONE [LMW^{+}21]. Pooled [STZ22, STC^{+}24]. popular [BM02]. popularities [AOU12]. Popularity [GGRF23, LFW23, MMCC^{+}21, MCF^{+}17, SLX^{+}24, ZMZ^{+}24]. portals [FSC07]. Position [HML^{+}23, Oos23, KFS10]. Position-Enhanced [HML^{+}23]. Positional [QHCY22, ZLZ25a]. Positive [LPLH19, WC21a, WLL^{+}23b, CQZ^{+}25a, MMCC^{+}21, WZZ04]. Positive/Negative [WC21a]. Post [CZL^{+}25]. Post-Training [CZL^{+}25]. Posterior [EMS21, WKF^{+}12]. Potential [LXLZ16, PHK12]. Power [LPC^{+}23, PSL16, RLS^{+}24, SW94, ZDW^{+}22]. Power-law [ZDW^{+}22]. Practical [Non13, WZG^{+}23a, CC05]. Practices [ZOML24]. PRADA [WZG^{+}23a]. praise [TL03]. Pre [AARC20, HYZ^{+}23, HWT^{+}25, LPC^{+}23, LYP^{+}25, LMMO23, MMH^{+}24, PXW^{+}25, QXX^{+}25, SWZW25, SAW^{+}25, WLL^{+}23a, WZZ^{+}23, WYZ^{+}25, WLF^{+}25a, WLF^{+}25b, WXZ^{+}25, YZL^{+}25]. Pre-Filtering [AARC20]. Pre-ranking [LPC^{+}23]. Pre-Trained [LYP^{+}25, WLF^{+}25a, WLF^{+}25b, QXX^{+}25]. Pre-Training [PXW^{+}25, WYZ^{+}25, HYZ^{+}23, HWT^{+}25, LMMO23, MMH^{+}24, SWZW25, SAW^{+}25, WLL^{+}23a, WZZ^{+}23, WXZ^{+}25, YZL^{+}25]. Precision [ASMZ25, PGD07, RJB89, RL94, MZ08]. Precision-Oriented [ASMZ25]. Predicate [CGMP99]. Predict [DHC^{+}25, NSVB18]. Predicting [LCJ^{+}16, SKC^{+}12, SAD24, VR24, ZLL^{+}22]. Prediction [CP13, Cum14, DGMG23, FHW^{+}19, FWG^{+}24, FLW^{+}25, HLX^{+}24, LLZ^{+}22, LQJ^{+}24, LSZ^{+}25, LLZ^{+}25, MWZ^{+}25, MAA^{+}25, MCF^{+}17, MAL16, MPT21, NWZC12, QFZ^{+}19, SLZ^{+}25, SZC^{+}23, SKC16, SMN^{+}16, SYX^{+}17, TCSL24, TZP22, TZP23, TM20, WLZW21, XWZ^{+}25, YZW^{+}17, ZM14, ZMK^{+}20, ZDZW23, ZKMM23, ZWT^{+}23, ZZG^{+}25, RSD^{+}13, SC09, WdVR08]. Predictions [ZTZL25]. Predictive [YFO^{+}25]. Preference [ABCZ21, CJZ^{+}16, DLZ^{+}22, DSZ^{+}23, GCL^{+}21, GCN^{+}19, LHL23, LYG^{+}25, LZLM19, MRC^{+}23, SZ21, WZW^{+}19, WLW^{+}24b, Won94, XZD^{+}25, YCL^{+}25, ZGL^{+}25, ZZG^{+}25, BBMN^{+}08, ZXM^{+}24]. Preference-aware [LHL23]. Preference-Based [Won94, SZ21]. Preferences [CZT^{+}23, CVS21, DLY^{+}25, DL16, Gin24, IYC^{+}23, LAL^{+}17, SZ21, SZC^{+}23]. Prefetching [WDG17, FPSO06]. Preparing [Leh06]. Prescription [YLW^{+}23]. Presentation [LGRK18, BDS04]. Preservation [GR12]. preserve [CGM02]. Preserving [FCDH91, FWG^{+}24, NXP^{+}24, WLR^{+}25, WWN22, YPYM24, LGY^{+}25, TXC^{+}24, ZLW^{+}23, ZBYH23]. Pretrain [LGZ^{+}24, LQZ^{+}25]. Pretrained [CSJ^{+}24, ZY21, ZLRW24]. preview [PSDB99]. Price [APJL20, CJZ^{+}16]. Price-per-attention [APJL20]. Principle [Fuh89, DWLW09]. Principles [GC92, HYM^{+}25, Lin07]. Prioritise [STZ22]. Prioritization [ZKJW22]. Priors [HCC^{+}16]. Privacy [CSR^{+}20, FWG^{+}24, GYW^{+}25, LGY^{+}25, TXC^{+}24, WLR^{+}25, YPYM24, ZLW^{+}23, ZBYH23]. Privacy-Preserving [FWG^{+}24, WLR^{+}25, YPYM24, LGY^{+}25, TXC^{+}24, ZLW^{+}23, ZBYH23]. Private [KBGW91]. Privilege [OSV19]. Proactive [DLL^{+}25, WLL24, ZBYH23]. Probabilistic [AV02, BB10, Coo95, DL16, FB91, FP94, FR97, KSNM10, Kwo90, Kwo95, MHZ16, MO12, PARS22, WLZW21, WY95, ZHY14, ZZZ^{+}16, ZHD^{+}22, GCT99, GP00, YWYL08]. Probability [CV98, EMS21, FC91, Fuh89, DWLW09]. Probe [WZC^{+}17]. Problem [FLMM91, GGR22, HCZ04, WTS^{+}08, WW10]. Problems [ZWZ^{+}24, ZZH^{+}18]. Procedural [CCAC24]. Process [BHH^{+}16, KL91, OOSC93, RN96, TZG^{+}24, YGZ15, ZZW^{+}25, ZCS^{+}23, KL00, RR98]. Process-Oriented [KL91]. Processes [HCD^{+}23, LBC^{+}19, SBH24, SM95, KL00]. Processing [AAL^{+}20, Can93, KM97, SÖ90, ZLF^{+}17, BS12]. Product [AZBC20, BNIL24a, BNIL24b, GCN^{+}19, LDN^{+}25, PLR^{+}22, QFZ^{+}19, ZJT16, ZCZW22, ZHRK23]. Product-Based [QFZ^{+}19]. Professionals [Ols89]. Proficiency [HLC^{+}20]. Profile [AKFA15, FC17]. Profile-Based [AKFA15]. profiles [CCC09, JD07]. Profiling [LLM22, SCL^{+}22, MSR04, SNS^{+}05]. Program [JVA^{+}15]. Programmatic [RLC^{+}25]. Programming [BR96]. Progress [DBCJ21]. Progressive [HLX^{+}24, WMGZ24]. Project [EGL^{+}97, MC94]. promotes [KKS^{+}08]. Promotion [CYC^{+}24]. Prompt [BHY^{+}25, GSG^{+}25, LZC23, LYP^{+}25, LGZ^{+}24, LXC^{+}25, LQZ^{+}25, MWC^{+}25, WWR^{+}25, YOM24, ZZD25, XPSC24]. Prompt-aware [ZZD25]. Prompt-Based [LYP^{+}25]. Prompt-tuning [YOM24]. Prompting [DHC^{+}25, YZL^{+}25]. Prompts [ZLZ25a]. Propagating [MK16a]. Propagation [CQL^{+}24, SZZL23, WL22, WBP^{+}16, ZM14]. Properties [SFR98, DEL^{+}00]. Prospective [SRF^{+}24]. Prospects [Ols89]. Protocol [CW94]. prototype [RR98]. prototypical [CSJS10]. Provider [XYX^{+}25]. PROXHY [KL91]. Proximal [NB97]. Proximity [AARC20, CL19, FLW^{+}23, GCC12, ARSZ03]. Proximity-Based [GCC12]. proxy [HHWL01]. proxy-based [HHWL01]. PRUM [PGD07]. pruned [BS12]. Pruning [LGM^{+}24, AOU12, BB10, MOT11, dMdSsdA^{+}08]. Pseudo [CQZ^{+}25a, DLA15, LMZ^{+}23a, MHZ16, WWT^{+}25]. Pseudo-positive [CQZ^{+}25a]. Pseudo-Relevance [DLA15]. Pseudo-Supervision [WWT^{+}25]. psychiatric [YWYL08]. Psychometric [AAL^{+}20]. PTF [YYQ^{+}25]. PTF-FSR [YYQ^{+}25]. Public [NSK^{+}20, ZGN^{+}17, LL99]. publications [BM12]. pure [HZSL13]. Purpose [AFF^{+}19, WLG^{+}17, YE18, CW02].

QProber [GIS03]. QUADEN [SMT25]. Quadrants [SMT25]. Qualitative [BDS04, MC14b]. Quality [BOQ^{+}25, GGRF23, HCC^{+}17, RLC^{+}25, WWHQ25, YMZF22]. Quality-aware [WWHQ25]. Quantitative [MC14b, ZGN^{+}17]. Quantum [ZHW^{+}22]. Queries [AVA^{+}24, CFC^{+}19, Guy18, KYK24, SÖ90, WPS^{+}25, ZLF^{+}17, ZZL^{+}21, ZGCC18, BGQT07, BJL^{+}07, BS12, CGMP99, CDF01, JD07, LCL04, Owe02]. Query [ABB^{+}15, BBD11, BTP24, BMMC19, BLWJ15, CRD16, CC01, CH96, CFFK22, Cum14, DGMG23, GS91, JTS16, LYN^{+}21, MTL23, MC14a, MAA^{+}25, MK94, MLP22, NB97, PPPR13, SW94, SPS^{+}06, SKC^{+}12, SKC16, SÖ90, TM20, UCH03, VAJR21, WZC^{+}25, WNZ02, YGZ15, ZZL^{+}22, ZWL^{+}23, AOU12, BJL^{+}07, Bod04, CdMRB01, CL03, DM07, EJ03, FPSO06, FG04, GNN^{+}10, GWR99, GCT99, JGP^{+}07, KMdRS06, LWSA10, LL99, LOP^{+}13, MSOH13, PSDB99, TFS^{+}12, TKD09, XC13, YWYL08]. Query-Based [PPPR13, CC01, ZWL^{+}23]. query-by-example [TFS^{+}12]. query-dependent [MSOH13]. Query-Drift [SKC^{+}12]. Query-independent [UCH03]. querying [MS01, WZB04]. Question [CCC^{+}12, GCR^{+}21, HWN^{+}19, LLY^{+}25, LLN22, LYZ^{+}23, LLM22, MAL16, NLF^{+}20, PHK12, QZS^{+}24, TM19, YXL^{+}23, ZWMH19, ZLY^{+}22, ZGC^{+}22, ZZC^{+}25, ZK20, CTM12, CKC07, KSNM10, KEW01, Lin07, MPHS03]. Question-Answer [CCC^{+}12]. Question-Answering [GCR^{+}21, LLY^{+}25, TM19, CTM12, KSNM10]. Question-based [ZK20]. Question-driven [LYZ^{+}23]. Question-Oriented [HWN^{+}19]. Question-Topic [NLF^{+}20]. Questions [AA16, HYM^{+}25, JEAZ25, LCL^{+}23a, LHA^{+}25a, NZW^{+}14, Pal15, ZAK^{+}23]. Quotation [WZW23]. Quote [TWLX18]. QuoteRec [TWLX18].

Radically [MLF95]. RAG [LLN^{+}25, MDX^{+}25]. Random [IC16, LCL^{+}17]. Random-Forest-Based [IC16]. Randomise [STZ22]. Randomized [LCL^{+}23b]. Randomly [FC91]. randomness [AV02]. Rank [AYWM21, DXZ^{+}22, DMUN19, IC16, KSYE24, LNO^{+}25, MZ08, Oos23, PYC^{+}19, SCG^{+}19, WPS^{+}25, BM02, DWLW09, LJQY19]. Rank-biased [MZ08]. Ranked [HCC14, SKK24]. Ranker [AVA^{+}24, LMDZ20, LHZ^{+}25]. Ranking [ASMZ25, ALS^{+}24, CBZ11, CDZ^{+}23, DXZ^{+}22, DLN^{+}16, FHW^{+}19, Fuh89, LR96, LRA23, LMR^{+}24, MPM22, MT21, RDK24, RBM23, SMGM11, WHL^{+}16, WDL^{+}24, WZZ^{+}24a, WZG^{+}23b, WZG^{+}23a, XTF^{+}21, ZKJW22, ZLY^{+}22, ZYD^{+}24, CRG02, CC00, CDZ08, DWLW09, EJ03, KZSD99, KSNM10, LPC^{+}23, PZMRN05, VDR25, ZLM^{+}23]. Ranking-Oriented [WHL^{+}16, ZYD^{+}24]. rankings [WMZ10]. Rapidly [WBB^{+}90]. Rate [ANW17, LSZ^{+}25, QZS^{+}23, SLZ^{+}25]. Rating [MWZ^{+}25, YZW^{+}17, ZXZL20, WdVR08]. Ratings [ABCZ21, CHN^{+}17, CDR10]. Raw [GZW^{+}18]. Re [AVA^{+}24, SBM^{+}17, TY22, VDR25, ZLM^{+}23]. Re-classification [TY22]. Re-Finding [SBM^{+}17]. Re-Ranker [AVA^{+}24]. Re-ranking [VDR25, ZLM^{+}23]. Readable [LPY94]. Reader [LMZ20]. Reading [BHH^{+}16, DPSH23]. Real [SYX^{+}17, YCZ^{+}16, ZLW16, ZZL^{+}21, AL13, HSS^{+}09]. Real-Time [SYX^{+}17, YCZ^{+}16, ZZL^{+}21, AL13, HSS^{+}09]. Real-World [ZLW16]. Reality [FZC93, LJAdR23, SGLS93]. realizations [DC10]. Really [ZWS^{+}25]. Reappraisal [BB93]. Reasoner [ZWZ^{+}24]. Reasoning [HYLC24, LLN22, LYG^{+}25, LZC^{+}24b, LBZ^{+}24, MHLH23, YWL^{+}25, ZSL^{+}25, ZWC^{+}23]. Reassessment [EMS21]. Rebalancing [CQZ^{+}25b]. Rec [WGL^{+}24]. Recall [RJB89, Web13, PGD07, Tee08, ZK20]. Recency [MCL^{+}12]. Recognition [XYCC24, ZJW^{+}24, ZWZ^{+}24, PTS13]. Recognize [MYL^{+}23, Tee08]. Recommend [NZW^{+}14]. Recommendation [AZ12, ABCZ21, AC18, AZCC21, AGT24, BHY^{+}25, BKWZ15, BLH^{+}22, CQF^{+}23, CCXZ25, CHN^{+}17, CJZ^{+}16, CZX^{+}19, CWZ^{+}20b, CZZ^{+}20, CJW^{+}21, CXZ^{+}21, CRC^{+}22, CMZ^{+}23, CZT^{+}23, CWZ^{+}24, CGD^{+}25, CS16, CCZ^{+}19, CLM^{+}22, CHZ24, CSZ^{+}22, DLY^{+}25, DSZ^{+}23, DHC^{+}25, FZSG21, FLW^{+}23, FMSS14, GCH^{+}19, GYW^{+}25, GYC^{+}22, GLG^{+}23, GLL^{+}23, GLZ^{+}24, GSG^{+}25, HYZ^{+}23, HKS^{+}21, HRY^{+}22a, HRY^{+}22b, HZF^{+}23, HLH^{+}25, HCC^{+}16, HWT^{+}25, HJZ^{+}23, HNN^{+}25, JMO22, JMZ22, JWX^{+}25, JXY^{+}24, JLC^{+}25, LZC^{+}25a, LJHL17, LHH^{+}19, LLW^{+}21, LLX^{+}22, LJAdR23, LZC23, LZZ^{+}23, LSL24, LCW^{+}24b, LRSD25, LSG^{+}25, LZG^{+}18, LMZ23b, LLBS18, LZL^{+}25, LZLM19, LPYM23, LZC^{+}24b, LLC^{+}25a, LTC^{+}25, LJQY19, LMMO23, LXG^{+}23, LFW23, LWL^{+}24, LGZ^{+}24, LXC^{+}25, LPZ^{+}25, LQZ^{+}25, LZC^{+}25b, LLC^{+}25b, LCNY23, LPLH19, LCJ^{+}19, LMW^{+}23, LGY^{+}25, LHZ^{+}25, LMZ20, MHLH23, MHW^{+}24, MXM^{+}24, MWC^{+}25, MYHL11, MYL^{+}23, MZG^{+}24, NXP^{+}24, PCC^{+}22, PCCC22, PVGM11, PBS22, PSM24, PGZ^{+}25, QLNY19, QZW^{+}25, QZS^{+}24, QWJ^{+}24, QGW^{+}24, QHLY20]. Recommendation [QHCY22, QXX^{+}25, QDG^{+}24, RABC22, SC19, SWZW25, SZS17, SRF^{+}24, SLX^{+}24, SSZ^{+}23, SLD^{+}24, SJL^{+}25, SWZ^{+}22, SYSW24, SLQ24, SLZ^{+}24, SJZ^{+}25, TWLX18, THL^{+}25, TWS^{+}25, TLS^{+}22, TXC^{+}24, WLW^{+}24a, WC21a, WMZ^{+}21, WLT^{+}22, WZZ^{+}23, WZW23, WMC^{+}23, WLW^{+}24b, WLY^{+}24, WZZ^{+}24b, WGL^{+}24, WZY^{+}24, WCP^{+}25, WZW^{+}25, WWHQ25, WYZ^{+}25, WLR^{+}25, WLF^{+}25a, WLF^{+}25b, WQL^{+}19, WLL^{+}23b, WMM^{+}23, WWHX23, WWG^{+}24, WWL^{+}25, WXZ^{+}25, XYW^{+}23a, XZH^{+}22, XC24, XSWX25, XZY21, XPL^{+}24, XZT^{+}24, XLX^{+}25, XYX^{+}25, XHW^{+}19, XYZ^{+}23, YLZW22, YCG^{+}24, YSW^{+}24, YCL^{+}25, YMZ^{+}22, YCL^{+}23, YZS^{+}24, YWQ^{+}24, YPYM24, YPF^{+}25, YZW^{+}17, YLW^{+}23, YWL^{+}25, YRC23, YOM24, YCZ^{+}16, YSXK23a, YSXK23b, YHWZ24, YFSO24, YSJ25, YZL^{+}25, YCZ^{+}15, YZC^{+}25, ZHL^{+}22, ZZL^{+}23, ZZZ^{+}24, ZLL^{+}20, ZLWW22, ZLW16, ZZD^{+}20, ZSY^{+}21, ZCZW22, ZWZ^{+}22, ZWF^{+}23, ZLW^{+}23, ZXM^{+}24, ZWY^{+}24, ZJK^{+}24, ZMN^{+}24, ZSZ^{+}24, ZWC^{+}25, ZSL^{+}25, ZLZ25a, ZZH^{+}25, ZXH^{+}25, ZGL^{+}25, ZMZ25, ZDW^{+}22, ZDZ23, ZXZL23, ZLF^{+}23, ZWC^{+}23, ZYL^{+}25, ZWX^{+}23]. Recommendation [ZWE^{+}22, ZWrWZ23, ZCLX22, ZWZS25, ZLZ^{+}25b, ZQH^{+}24b, ZSLK24, DK04, IH07]. Recommendations [GZL^{+}24, GXTL14, HL17, HCC^{+}17, HML^{+}23, HLL^{+}25, LCPD19, LHL23, LYG^{+}25, LLP24, WZW^{+}22, WFW^{+}25, WAL^{+}21, XZD^{+}25, YFO^{+}25, YHDN25, ZZZ^{+}25]. Recommended [HY20]. Recommender [AM25, CCCD19, CDW^{+}23, CST^{+}23, CYC^{+}24, CWM^{+}24, CZL^{+}25, CSB21, DBCJ21, DZX^{+}23, DSW^{+}25, DYL^{+}20, GWL^{+}24, GZW^{+}24, HSYX22, HLL^{+}25, IYC^{+}23, JSZL23, JZZW24, LDX^{+}25, LZG21, LCL^{+}23b, MZLK11, MAP^{+}22, MMCC^{+}21, MLZ^{+}22b, NXF^{+}24, RVA^{+}24, RYC^{+}20, SAD24, SMT25, TQN^{+}23, TCQ^{+}25, WZW^{+}19, WMZ^{+}23, WZZ^{+}24a, WZL^{+}24, WCX24, XHX^{+}22, XYW^{+}23b, YHY^{+}17, YCS^{+}14, YYY^{+}24, YYQ^{+}25, ZZY^{+}23, ZWS^{+}25, ZSH^{+}25, ZZSS25, ZQC^{+}23, ZQH^{+}24a, ASST05, HKTR04, Kon04, MSR04, MKR04, RSG09, WMJ05]. Recommenders [MRC^{+}23, WXR^{+}25]. Recommending [CWZ^{+}20a]. Reconciling [GGRF23]. Recovery [DG95]. RecRanker [LHZ^{+}25]. ReCRec [LZC^{+}24b]. Recruitment [QZX^{+}20]. Recurrent [TCSL24]. Recursive [AZK12, Coh97]. redesign [KL00]. Reduce [AMS20, CHZ24, PSPBY10]. Reducing [GKL^{+}07]. Reduction [MDX^{+}25, GMR09]. Redundancy [CFE^{+}25, PSM24, Lin07, SSS11]. Redundancy-Aware [CFE^{+}25]. redundancy-based [Lin07]. Redundant [Row96]. Referees [Ano95]. Reference [SKC16]. refinding [BCD^{+}08, EBR08]. Refinement [LSG^{+}25]. Refining [MCL^{+}12]. Reflection [XHX^{+}22]. Reflection-Augmented [XHX^{+}22]. Reformulation [BLWJ15, GS91, XC13]. reformulations [JGP^{+}07]. ReFRS [IYC^{+}23]. Region [ARSZ03]. Regression [DLN^{+}16, LZW^{+}20]. Regularization [ZSY^{+}21]. Regularized [CSR^{+}20, WXLC13]. Regulation [WBP^{+}16]. Reinforced [FTZZ23, MWC^{+}25, PZD^{+}22]. Reinforcement [CWM^{+}24, FLW^{+}23, SLX^{+}24, SJZ^{+}25, WGL^{+}24, YSW^{+}24, YDXW21, ZZT^{+}21]. Related [FEL22, JTL^{+}17]. Relation [AZBC20, HYLC24, LYLS22, LL23b, SHS^{+}23, SQZ^{+}22, WMGZ24, XZCZ25, XWZ^{+}25, ZYY^{+}21, ZZH^{+}25]. Relation-Enhanced [SQZ^{+}22]. Relational [CCZ^{+}23, FHW^{+}19, FR97, LYP^{+}25, PZD^{+}22, BJS99, ZDW^{+}22]. Relations [RCR^{+}18, WMZ^{+}21, ZCZW22, ES98, MC10a]. Relationship [ZJT16]. Relative [DGMG23]. Relevance [Bod04, DXZ^{+}22, DLA15, Dun97, LMZ^{+}23a, LGL^{+}24, LLZ^{+}21, MMST17, MC14a, MYHL11, MAA^{+}25, MHZ16, MHR22, RK19, RBCT14, RLS^{+}24, STZ22, SKK24, SWL^{+}23, Wil98, XWL^{+}25, YXA^{+}24, EJ03, PTL08, WdVR08, WLWK08, ZCD06]. Relevant [LK20, LCL^{+}23a, ZK20, JZ06, KFS10]. Reliability [MHR22]. Reliable [LLY^{+}25]. Relieving [SLX^{+}24]. remote [HHWL01]. Removing [PSM24]. Repeatable [JBCF07]. Repetition [SE91]. Replicates [VSS17]. reporter [ZKG08]. reporter-based [ZKG08]. reporters [ZKG08]. Repository [CTH^{+}13, YHYO01]. Representation [CSR^{+}20, CZX^{+}22, CDZ^{+}23, CACW23, FZH^{+}22, GYC^{+}22, HSH90, HYLC24, LZWZ24, LYM^{+}24, LL23a, LWC^{+}25, MLZ^{+}22b, PL94, QLNY19, RS93, SQZ^{+}22, TSNS19, WYZ^{+}25, WQL^{+}19, YLZW21, YLZW22, ZSQ^{+}24, ZZD^{+}20, ZSH^{+}25, ZCLC23, ZHRK23, Coh00, KL00]. Representations [AMS20, CQZ^{+}25a, GLG^{+}23, LWP^{+}21, MTL23, MLZ^{+}20, RDK24, Ruh94, TL14, VR24, WLT^{+}22, XTF^{+}21, XYZ^{+}23, ZHC^{+}20, ZMN^{+}24].

Representative [SZS17]. Representative-Based [SZS17]. Representing [HSH93, MBJK90]. Reproducibility [DBCJ21]. reputation [ZKG08]. Requests [ACC18]. Reranking [LCPD19, LYM^{+}24, LZM^{+}19, TM19, HHC^{+}10, KKB10, KL10, MC10b]. Research [CCXZ25, Cro89b, DBCJ21, MAH24]. Residual [LK20, SZC^{+}23, YCG^{+}24]. Resource [IYC^{+}23, MC14b, SZZL23, AAGY01]. Resource-efficient [IYC^{+}23]. Resources [ZJL^{+}24]. Respond [GCL^{+}21, YLZW21]. Response [BACF17, FTZZ23, LCW^{+}24a, LMW^{+}24, LLT^{+}21, MLL^{+}24, QFZ^{+}19, RCR^{+}21, Wie92, XTF^{+}21, YLZW21]. Responses [CQZ^{+}25b]. Responsible [WBP^{+}16]. Responsiveness [OSV19]. Result [DDSW24, LYS^{+}17, LZM^{+}19, SDZW24, WDG17, ZLM^{+}23, SZ09]. Results [LGRK18, MC94, MCL^{+}12, NTV22, QDZW23, WDL^{+}24, ZZL^{+}21, BM12, EJ03, FPSO06, HHC^{+}10, KKB10, SC03, Tee08]. RESUS [SZC^{+}23]. Rethinking [FCL^{+}23, ZTZL25, ZQH^{+}24a]. Retrieval [AMS20, ABB^{+}15, AVA^{+}24, BGI24, BLMN24, CE89, CCC^{+}12, CFP95, CJN20, CR95, Coo95, CDM15, CV98, CPL15, DLA15, DLA22, DMUN19, EMG11, FTZ11, FKS19, FLPC24, FCDH91, Fuh89, FP94, FR97, GJS^{+}25, GC92, GCC12, Gin24, GR95, GR96, GCF^{+}22, GCB^{+}25, HCC14, JYM13, KSYE24, KBD89, KC92, Kwo90, Kwo95, LCW^{+}24a, LLT^{+}21, LPC^{+}23, LWL^{+}25, LJZ^{+}25, LZC24a, LYN^{+}21, LHA^{+}25a, LHA^{+}25b, LL23b, LZ96, LLN^{+}25, MWA^{+}24, MMST17, MKF91, MAH24, MMH^{+}24, MH89, MO12, MZ96, Na15, PARS22, PSL16, RJB89, RRS95, Row96, SZ21, SGJ22, SWL^{+}23, SLW^{+}24, SW90, SL95, SAW^{+}25, TBC96, TZG^{+}24, THT12, TL14, TC91, VHJC11, VdRK18, VSS17, VAJR22, WLZ^{+}24, WWT^{+}25, WY95, XLW^{+}22, XTF^{+}21, XPSC24, XWL^{+}25, YLZW21, YC94, YMZF22, ZNL^{+}17, ZBYH23, ZOML24, ZHY14, ZLRW24, ZYD^{+}24, ZZL^{+}22, ZWL^{+}23, ZK20, ADCU08, AV02, BYFM06, CML00, CRG02]. retrieval [Car12, CSLN10, CC00, Cro06, DM07, EKK99, FG04, GMR09, HO07, HJZL09, HZSL13, HCZ04, KL06, KNS04, KO98, KL09, LBS08, LHSC11, LA08, MR10, MLDGH04, Mel08, MZ08, PMPJ11, PMD01, PRP05, PH07, PGD07, PTL08, SV99, TLJ^{+}07, TKD09, Wil05, WSBC01, XC00, YLS^{+}11, YWYL08, ZL04, ZCD06, ZRZ02, Cro89b]. Retrieval-Augmented [LWL^{+}25, LLN^{+}25]. Retrieval-Based [LLT^{+}21, YLZW21, XTF^{+}21]. Retrieve [QZS^{+}23]. Retrieved [FC91, KFS10]. Retrievers [LMZ^{+}23a]. Retrospective [SRF^{+}24]. Reusable [XZT^{+}24]. reuse [Tee08]. RevGNN [LZL^{+}25]. Review [BvdHFS25, GCH^{+}19, GCF^{+}22, LLZ^{+}22, LZG21, LLC^{+}25b, MWZ^{+}25, MAH24, MMO21, OSV19, WZZ^{+}24b, WGZ^{+}25, ZGC^{+}22, ZLW24, ZSH^{+}25, Mar08a]. Review-Aware [GCH^{+}19, LZG21]. Review-Based [MWZ^{+}25]. Review-Enhanced [ZSH^{+}25, WZZ^{+}24b]. Reviewer [LZL^{+}25, ZZD^{+}20, Mar08a]. Reviewers [Cal12, dR19, Mar06, Mar07, Mar08b, ACM03]. Reviewing [LK20]. Reviews [CCZ^{+}19, LK20, SBH24, ZJT16, ZXZL23, ZK20]. Revised [XWZ^{+}25]. revision [LBS08]. Revisited [MNW98, FGM^{+}02]. Revisiting [CMZ^{+}23, LFL^{+}25, RDK24, ZZY^{+}23, ZLF^{+}23]. Rewiring [LCYY25]. Rewriting [LYN^{+}21, CGMP99]. rHDP [ZWX^{+}24]. Rich [Ruh94]. Risk [SCG^{+}19, WLZW21, WA22, CHZ24]. Risk-aware [CHZ24]. Risk-Sensitive [SCG^{+}19]. Risks [CHZ24]. RLPS [YDXW21]. Road [ZGCC18]. Robust [ALS^{+}24, CQL^{+}24, CJN20, HHC^{+}23, Oos23, PPPR13, SLM^{+}24, SZ09, WZG^{+}23b, WXR^{+}25, YLYT23, ZMZ^{+}24, EKK99, GWLC06]. Robustness [DRdR23, FCL^{+}23]. ROGER [ZYD^{+}24]. Role [FLMM91, HZS^{+}25, Koi93, Ruh94, WWX^{+}24, WHL^{+}22]. Roles [GSR96]. ROM [KBD89]. Routing [FLW^{+}23, Pal15, PYDR24, PSPBY10]. RPRS [AVA^{+}24]. RST [AA16]. RST-Based [AA16]. Rules [ADW94, CDR21, YWL^{+}25, PTKJ07, WZZ04]. Rumor [LBM^{+}24]. Rumors [MK16a]. Rumour [LBC^{+}19].

Saerens [EMS21]. Safe [JMD20]. SAGE [SNS^{+}05]. Salient [CJY23a]. SALSA [LM01]. Salton [Ano96b]. Sample [ALS^{+}24, CQZ^{+}25a, SZ09]. sample-based [SZ09]. Sampled [WWG^{+}24]. Sampler [CJW^{+}21]. Sampling [CZZ^{+}20, CMZ^{+}23, LK20, LZL^{+}25, LLZ^{+}25, ZKMM23, CC01]. Sampling-Based [ZKMM23, LLZ^{+}25]. Satisfaction [SAD24, WZL^{+}19]. scalability [KM13]. Scalable [FZH^{+}22, LZG^{+}18, MWYL01]. Scale [CTH^{+}13, LMDZ20, NGMD17, NLF^{+}20, SC19, VFA17, WXLC13, ZJL^{+}24, ZNL^{+}17, ZHC^{+}20, CJRY12, LZC^{+}25b, VKD21, YKHW06]. Scaling [KEW01, YMZF22]. Scenario [CR92]. Scenarios [SJL^{+}25, GFWK04]. Schema [AARC20, RR98]. Scheme [APJL20, LM90, Man97, WWX^{+}25, CW02]. schemes [PS05]. Scholar [WCZ^{+}25]. Scholars [Ruh94]. Science [Kli94, NFZ19]. scientific [BM12]. Score [Cum14, MTL23, SZ09]. Score-at-a-time [MTL23]. scraps [BVKS08]. Screenshots [USN^{+}17]. Script [ZSN^{+}22]. Scrutability [XZT^{+}24]. SEA [CDH^{+}09]. Seamlessly [LLW^{+}21]. Search [AZBC20, ASSC22, AC16, AC19, BACF17, BZDW24, BMMC19, BKWZ15, BNIL24a, BNIL24b, CLF^{+}25, CWL^{+}20, CCAC24, CSZD15, CHBS15, CCVR17, CLZ^{+}14, DDSW24, DH97, FEL22, GS91, GCN^{+}19, Guy18, HYH15, HKS^{+}21, HHC^{+}23, JTS16, KMPS21, KC15, LGRK18, LXZ^{+}22, LYS^{+}17, LCY21, LHG^{+}12, LCL^{+}17, LZM^{+}19, LZW21, LDN^{+}25, LMW^{+}14, MT21, MLK^{+}18, MAH24, MMH^{+}24, MRM24, MBST17, MCL^{+}12, MK16b, NTV22, NWZC12, NFZ19, NLL^{+}24, PLR^{+}22, QDZW23, RBCT14, RCR^{+}21, RPE^{+}18, STZ22, SLZ^{+}19, SSZ^{+}23, SSGC89, SDZW24, TTZ^{+}25, TBCW20, USN^{+}17, VKD21, VAJR21, VDR25, WLY^{+}14, WLS^{+}17, WA22, WDL^{+}24, WLF^{+}25a, WLF^{+}25b, WZHY24, WH15, WDG17, WZL^{+}19, YGZ15, YDXW21, YDW22, YSXK23a, YSXK23b, YTL25, YCZ^{+}15, ZTM^{+}17, ZZL^{+}21, ZLM^{+}23, ZLH^{+}25, ZZD25, ZHD^{+}22, ZML^{+}19, ZHRK23, ZAK^{+}23, AOU12, ARSZ03, ACM^{+}02, AL13, BBD11, BC13, BBMN^{+}08, BM12, CZJ07, CJRY12, CDZ08, CGM05, EJ03, FPSO06]. search [FGM^{+}02, FKM^{+}05, HSS^{+}09, HHC^{+}10, IH07, JBCF07, JGP^{+}07, KM13, KKB10, Leh06, LS02, LWA08, LOP^{+}13, MPS07, PSPBY10, SC03, Tee08, TS10, WH09, XHYY09, ZHC^{+}13, dMdSsdA^{+}08]. Searcher [TBCW20, UA22, WRJV05]. Searchers [CACW23]. searches [GKL^{+}07]. Searching [Man97, ZNL^{+}17, KZP10, LdJKR12, LLC03, SNZBY00, MC06]. SECON [ZLH^{+}25]. Secondary [MNZ19]. Secret [LPYM23]. Section [BLMN24, HRY^{+}22b, YSXK23b]. Seed [LCX^{+}19]. Seed-Guided [LCX^{+}19]. Seeking [TY22, VKD21, WZC^{+}25]. Segment [CDH^{+}09]. Segment-enrich-annotate [CDH^{+}09]. Segmentation [JDA^{+}22]. segments [CC05]. Selecting [DH97, LPC^{+}23]. Selection [AZCC21, CFE^{+}25, DXL^{+}22, FTZZ23, HWN^{+}19, LCW^{+}24a, LMW^{+}24, LLT^{+}21, MLL^{+}24, MC14b, PHK12, PZD^{+}22, SCG^{+}19, ZLW24, ACS08, CC03, Fuh99, HT99, IG08, PF03]. Selective [KC15, LGRK18]. Selectivity [BF98]. Self [BF98, CZJH24, JDA^{+}22, JWX^{+}25, JZZW24, KSYE24, LXC^{+}25, MZ96, QDZW23, WMGZ24, ZZC^{+}25, CGM05, FBN^{+}12, PR09]. Self-attention [QDZW23]. self-correlation [PR09]. Self-Distillation [JWX^{+}25, WMGZ24]. Self-Distilled [KSYE24]. self-indexes [FBN^{+}12]. Self-Indexing [MZ96]. Self-organized [JDA^{+}22]. Self-Paced [CZJH24]. Self-Prompt [LXC^{+}25]. Self-Spatial [BF98]. Self-supervised [JZZW24]. self-supervising [CGM05]. Semantic [AMS20, CGB19, DHC^{+}25, EMG11, GCF^{+}22, GSG^{+}25, HHC^{+}23, HWL^{+}24, LPY94, LL23a, LLC^{+}25b, LGY^{+}25, LXLZ16, MPT21, SLHS93, TG10, TM19, WXLC13, WZW23, XCG^{+}23, ZLH^{+}25, ZHC^{+}20, dODMS20, FCD02, Hof04, KO98, MR10, PR09, TL03, YLS^{+}11]. Semantic-enhanced [GSG^{+}25]. semantic-multimedia [MR10]. Semantics [SF89, WR98]. Semi [LYY^{+}24, MP20, YHS^{+}21]. Semi-Static [MP20]. Semi-Supervised [LYY^{+}24, YHS^{+}21]. Semiconductor [HE91]. semidiscrete [KO98]. Semisupervised [HJZL09, SC03]. Sense [OG94]. Sensed [CLZ^{+}14]. Sensing [PGW^{+}17]. Sensitive [GC92, SCG^{+}19, CS99, LBS08]. Sensitivity [MMO21]. Sensors [FC17]. Sentence [RCR^{+}18, XXDW24]. Sentiment [ACS08, BHH^{+}16, FCL^{+}23, HQZ17, LLG^{+}23, LZG21, ZJW^{+}24, ZH17, dODMS20]. Sentimental [FAZC12]. Separate [HYLC24]. September [Cal12]. Sequence [MXM^{+}24, ZSH^{+}25, ZCLC23, ZCS^{+}23, KT05, LDN^{+}25]. Sequence-aware [ZCLC23]. Sequences [HCC14, Zob06]. Sequencing [CR95]. Sequential [CRC^{+}22, CSZ^{+}22, FZSG21, GZL^{+}24, GLL^{+}23, HLX^{+}24, HWT^{+}25, HSYX22, HML^{+}23, JXY^{+}24, LSL24, LXC^{+}25, LDN^{+}25, LLP24, MRC^{+}23, PLR^{+}22, RN96, SRF^{+}24, SYSW24, SLQ24, TZG^{+}24, WC21a, WMZ^{+}21, WC21b, WMC^{+}23, WCP^{+}25, WZW^{+}25, WLR^{+}25, WXR^{+}25, XSWX25, XYZ^{+}23, YSW^{+}24, YYQ^{+}25, ZZZ^{+}24, ZKJW22, ZCZW22, ZWY^{+}24, ZWC^{+}25, ZLZ25a, ZTZL25, ZWrWZ23, ZSLK24]. Sequential-Knowledge-Aware [CSZ^{+}22]. Series [YZH^{+}24]. SERP [RCR^{+}21, SZ21]. SERP-based [RCR^{+}21]. server [HT99]. service [DDH^{+}09, FSC07]. service-based [DDH^{+}09]. ServiceFinder [FSC07]. Services [MNP^{+}21b, JBCF07]. Session [CML^{+}21, PCC^{+}22, PCCC22, QZW^{+}25, QHCY22, RBCT14, SGJ22, WLW^{+}24a, WZY^{+}24, WCX24, XYW^{+}23a, YGZ15, YHWZ24, YFSO24, ZXM^{+}24, ZQH^{+}24b, QHLY20]. Session-Based [QHCY22, WZY^{+}24, XYW^{+}23a, ZQH^{+}24b, PCC^{+}22, PCCC22, QZW^{+}25, QHLY20, WLW^{+}24a, WCX24, YHWZ24, YFSO24, ZXM^{+}24]. Sessions [SE91]. Set [PZMRN05, CM10, GMR09]. Set-based [PZMRN05]. SetRank [WZZ^{+}24a]. setting [HO07]. Setwise [WZZ^{+}24a]. Shallow [TM19]. Shards [FKS19]. Shared [IKG93, KBGW91, LM90, ZH17]. Sharing [LPYM23, ZWX^{+}24, BCD^{+}08, SS08]. Sharing-Enhanced [ZWX^{+}24]. Shell [VFA17]. Shift [DRdR23, ZMZ^{+}24]. Shifts [HE91, LZWZ24, WLW^{+}24b]. Short [GCN^{+}19, LDW^{+}17, LRZ^{+}17, LYS^{+}17, XPL^{+}24, YHS^{+}21, ZLF^{+}17]. Short-Term [GCN^{+}19, XPL^{+}24]. Shortening [BBT92]. Shortest [CC00]. Shortest-substring [CC00]. Shot [CLL24, CZJH24, FZXdR24, LLG^{+}23, LLZ^{+}21]. Shots [BA18]. Should [LAO^{+}24, STZ22]. Showing [XHSD24]. Sibyl [CTM12]. Side [FMSS14, LMMO23, SWZ^{+}24]. Sided [MAP^{+}22]. SIGIR [Cro89b]. Signal [CDM15]. Signals [SYSW24, SLZ^{+}24, WMC^{+}23, YXA^{+}24]. Signature [BA18, CZ93, LL89, LR96, ZRT91, CHS99]. Signature-Based [BA18]. signatures [PPGK04, Zob06]. Signed [CXZ^{+}21]. SIMD [ZZL^{+}15]. SIMD-Based [ZZL^{+}15]. Similar [HHC^{+}23]. similarities [KKB10]. Similarity [BA18, CWZ^{+}20b, CWZ^{+}20a, GR95, MO12, MK16b, YMZF22, ZLL^{+}20, ZTM^{+}17, AC08, ARSZ03, Coh00, EKK99, GGMW03, MC10b, SPKW11, WMZ10]. Simple [LYLS22, OOSC93, PCD92]. Simplify [ZZSS25]. Simpson [JMO22]. Simulating [WA22]. Simulation [CST^{+}23, SGLS93, VHJC11, WZY^{+}25, ZLM^{+}23]. Simulation-Based [VHJC11]. simulations [WRJV05]. Simulators [SGZ^{+}24]. singer [SSCT09]. Single [Kwo90, ZWZ^{+}24, WX10]. Single-stream [ZWZ^{+}24]. Site [AKFA15, ZZG^{+}25, IM05]. Sites [AZK12, ZLL^{+}22]. Situ [YTL25]. Size [MT21, IH07]. Sizes [FKS19]. Sketches [MNZ19]. skewed [SV99]. Skill [QZS^{+}24, RFB^{+}24, SJZ^{+}25]. Skill-based [RFB^{+}24]. Skill-Oriented [QZS^{+}24]. skipping [ADCU08]. SLED [ZJK^{+}24]. Sliding [LYZ^{+}23]. Small [MC14b]. Small-Document [MC14b]. SMAPH [CFC^{+}19]. smart [BSST03]. SMLP4Rec [GZL^{+}24]. smoothing [ZL04]. Smoothness [ZWrWZ23]. Snippet [BC14]. Social [CP13, CLZ^{+}14, DCT^{+}17, FC17, FMSS14, HGBY16, Kli94, VST^{+}14, LXZ^{+}22, LLX^{+}22, LSG^{+}25, LYP^{+}25, LLBS18, LW20, LBC^{+}19, MZLK11, Mar94, MJH15, MTC21, MK16a, MKZPK18, MAL16, NGMD17, NXP^{+}24, NZW^{+}14, PBJ^{+}16, PZH^{+}24, QWZ^{+}15, RN96, SMN^{+}16, TLKW16, WZC^{+}17, WLFT18, WLT^{+}22, WFW^{+}25, YSZ^{+}17, YTD13, ZAL^{+}16, ZLL^{+}22, ZWZ^{+}22, ZYW^{+}23, ZZZ^{+}16, ZHC^{+}20, ZY21, ZDW^{+}22, CDR10, KL00, LLPT10]. Social-aware [WLT^{+}22]. Social-Sensed [CLZ^{+}14]. societies [GFWK04]. Society [WW94]. Soft [CKC07, TQH15, WXR^{+}25, ZWY^{+}24]. Softmax [WWG^{+}24]. Software [Koi93, MR92, Tuz95, ATW00]. Solutions [NGMD17, TY22]. Solving [FLMM91, ZWZ^{+}24, WTS^{+}08]. Some [Coo95, Gin24]. Sorted [NTV22]. Sorting [NCHW15, Non13]. Sound [GTW^{+}18, PTL08]. Soundness [HWD13]. Source [VAJR22]. Sources [VST^{+}14, LZM^{+}19, MC14a]. Space [CMR91, IKG93, MHZ16, QZXP06, SPKW11]. space-partitioning-based [QZXP06]. Spaces [VdRK18, ARSZ03, GFWK04, KZP10, QZXP06]. spam [CL07, ZKG08]. Sparse [BNIL24a, BNIL24b, FLPC24, MTL23, SMT25, WZC^{+}17, ZHC^{+}13]. Sparsely [WPS^{+}25]. sparsity [HCZ04]. Spatial [AGT24, BF98, GR95, Koi93, YCS^{+}14, ZGCC18, ZDZ23, ES98, EKK99]. Spatial-Temporal [AGT24]. Spatio [WLW^{+}24a, XC24, PMD01, TP05, LLPT10]. Spatio-temporal [WLW^{+}24a, XC24, PMD01, TP05, LLPT10]. Spatiotemporal [JMZ22, MK16a, QLNY19]. SPContrastNet [CZJH24]. Special [BKWZ15, BLMN24, Cro89a, HKS^{+}21, HRY^{+}22a, HRY^{+}22b, Kli94, LdJKR12, PBJ^{+}16, SL95, WLS^{+}17, WLF^{+}25a, WLF^{+}25b, YSXK23a, YSXK23b, BYFM06]. Specialization [Sci89]. Specific [Sav12, VL90, CC03, DEL^{+}00, YLS^{+}11]. Specification [HM90, JSHS96, OS96]. Specifications [Tuz95]. Spectral [QJG^{+}25]. Spectrum [NGMD17]. Speech [CW94, KM97, LdJKR12]. Speed [AC19]. SPEX [LLX^{+}22]. Spidering [FAZC12]. Split [CHZ24]. Spoken [CJY23a, Guy18, VAJR21, CSLN10, CTM12, TFS^{+}12, WKF^{+}12]. Spreadsheet [ISL95]. SSD [WLY^{+}14]. SSD-Based [WLY^{+}14]. SSR [ZWZ^{+}24]. Stability [AZ12]. Stack [SZZL23]. Stack-Propagation [SZZL23]. Stage [GCF^{+}22, LYN^{+}21, Na15, GCB^{+}25, MMH^{+}24]. Stakeholder [WMM^{+}23]. Stance [LBM^{+}24, LBC^{+}19, ZY21].

Standards [Gru92]. STARec [JMZ22]. Start [CQF^{+}23, HLH^{+}25, HCC^{+}16, SZS17, WGL^{+}24, HYZ^{+}23, HSYX22, LLW^{+}21, WLL^{+}23b, WWL^{+}25]. statechart [dOTM01]. statechart-based [dOTM01]. Static [AOU12, MP20, XZCZ25, ZCZ^{+}23, BB10, GL06]. Statistical [CSLN10, RLS^{+}24, TL14, Car12]. statistics [CFGM02]. status [TkWW99]. Stemming [PPPR13, XC98, PMPJ11]. Stepped [VFA17]. STEvent [LLPT10]. Stickers [GCL^{+}21]. Stochastic [ABB^{+}15, LM01]. Stock [CHZ24, FHW^{+}19, LCJ^{+}16, SC09]. Stop [LK20]. Stopping [BvdHFS25, SBH24]. Storage [GC92]. Store [Mos90]. Storing [KBD89]. Storm [JTL^{+}17]. Story [NLL^{+}24]. Strategy [HJZ^{+}23, LSG^{+}25, WW94, WCX24, HYZ^{+}23]. Stream [CTH^{+}13, ZLF^{+}17, ZWZ^{+}24]. Streaming [BNIL24a]. Streams [GFWK04, LRZ^{+}17, LYS^{+}17, WLFT18]. Strength [ZZZ^{+}16]. string [HZW02]. stripper [MMP^{+}07]. Structural [BRS92, JYM13, MWA^{+}24, PZH^{+}24, WHL^{+}22, KL10]. Structure [COR13, LZM^{+}19, LHA^{+}25b, MH89, NB97, QGW^{+}24, SF89, WWX^{+}24, YLYT23, ZJK^{+}24, GGMW03, HZW02, LM01, Luz12]. Structure-Aware [LHA^{+}25b]. Structure-Based [WWX^{+}24]. Structured [BB93, GZC89, JDZ^{+}22, WR98, BDS04, Cro06, LL99, PGD07]. Structures [CZX^{+}22, TGL^{+}97, TM19, BHMW07, GFWK04]. Structuring [HSH93]. Students [HLC^{+}20]. studies [MC10b, WH09]. Study [Ack98, CV98, FKS19, JSZL23, JSL^{+}24, MRM24, MLP22, NLL^{+}24, VHJC11, WW94, WLY^{+}14, ZLF^{+}23, ZZG^{+}25, ZL04]. Studying [CST^{+}23, KM13]. Styles [WAL^{+}21]. Stylized [JSL^{+}24]. stylometric [AC08]. Subgraph [WZZ^{+}23]. Subspace [PR16]. Substitutable [ZCZW22]. substrate [SM99]. substring [CC00]. Substrings [NCZG14]. Substructure [YW15]. Substructure-Based [YW15]. Successes [LMZ^{+}23a]. Suffix [NCZG14, MMP^{+}07, Non13]. Suffixes [NCHW15]. Suggesting [AK18]. Suggestion [MLP22]. suggestions [GNN^{+}10]. SUIT [PCD92]. Sum [DPSH23, RK19, BCFG15]. Summaries [WCZ^{+}25, BGQT07]. Summarisation [AKFA15]. Summarization [BZCE22, CGB19, LYZ^{+}23, MSMO18, RCR^{+}18, YW15, BKGM^{+}02, HL10, MLDGH04, WX10]. Summarizer [BCFG15]. Summarizing [BM12]. Summary [CLG^{+}23, MC06, ZZL^{+}21]. SuperBook [ERG^{+}89]. Supervised [LYLS22, LYY^{+}24, SHS^{+}23, WPS^{+}25, CL07, JZZW24, YHS^{+}21]. supervising [CGM05]. Supervision [KYK24, LZC24a, RLC^{+}25, WWT^{+}25]. Support [ASV91, KM97, MK16a, TkWW99, TMT06]. Supporting [Bie92, Bul95, HYH15, KBGW91, RSG09]. surface [KL00]. Surrogate [TQN^{+}23]. Surrogate-based [TQN^{+}23]. surrogates [KKS^{+}08]. Survey [AM25, CCXZ25, CLPM20, CDW^{+}23, DLL^{+}25, GZW^{+}24, HYM^{+}25, JZZW24, LJZ^{+}25, LDX^{+}25, MLZ^{+}22a, WMZ^{+}23, ZZL^{+}23, ZLRW24, ZQC^{+}23]. SVM [HJZL09]. SWIM [VFA17]. Swipe [NSVB18]. Symmetric [ZYL^{+}25]. Synchronization [RRS95]. Synopsis [TM20]. Syntactic [HQZ17, MH89, TM19, XXDW24]. Syntactic-Informed [XXDW24]. Syntactic/Semantic [TM19]. Synthesizing [MC14a]. System [CE89, CFP95, CDW^{+}23, CLF^{+}25, CH96, CR95, DMUN19, DSW^{+}25, GWL^{+}24, IYC^{+}23, JSZL23, JTL^{+}17, KEL95, KBGW91, LLY^{+}25, LZG21, LCL^{+}23b, LCJ^{+}19, MLZ^{+}22a, MMLP97, NXF^{+}24, OS96, PBF^{+}89, RJB89, SSGC89, SMT25, WZZ^{+}24a, WGZ^{+}25, WHFG92, WL97, XYW^{+}23b, XZT^{+}24, YHY^{+}17, YCS^{+}14, YYQ^{+}25, ZLH^{+}22, ZWT^{+}23, CGMP99, CTM12, CC03, FP00, GL06, GIS03, HSS^{+}09, IH07, KFS10, KNS04, LBS08, LSOS06, LL99, MKR04, MPHS03, RR98, RSG09, SC09, TkWW99, TMT06, ZKG08]. System-Induced [LCL^{+}23b]. Systematic [MLZ^{+}22a, MAH24, RABC22]. Systems [ASV91, AM25, CCXZ25, CP13, CCCD19, CST^{+}23, CYC^{+}24, CSJ^{+}24, CWM^{+}24, CZL^{+}25, CSB21, DBCJ21, DZX^{+}23, DYL^{+}20, FN96, FR97, GZW^{+}24, GSR96, GH96, HZG20, JMO22, JMSV92, JZZW24, JSHS96, KM97, KBD89, LMW^{+}21, LDX^{+}25, LCY21, LCL^{+}23b, LLY^{+}24, LZ96, MZLK11, MAP^{+}22, MJH15, MMCC^{+}21, MP20, Mor90, MLZ^{+}22b, MBJK90, RVA^{+}24, RYC^{+}20, RN96, SC19, SAD24, SÖ90, SGZ^{+}24, THL^{+}25, TQN^{+}23, Tom89, TCQ^{+}25, TM19, WW94, WZW^{+}19, WMZ^{+}23, WZL^{+}24, WCS^{+}24, Won94, WCX24, XHX^{+}22, YYY^{+}24, ZGN^{+}17, ZHZ^{+}19, ZXZL20, ZZY^{+}23, ZWS^{+}25, ZSH^{+}25, ZZSS25, ZZH^{+}18, ZQC^{+}23, ZQH^{+}24a, ASST05, BJS99, BFF^{+}03, Car12, DEL^{+}00, HKTR04, IH07, Kon04, Mar08a, MSR04, PVGM11, PR04, PSDB99, SS08, WMJ05, All89, dR19]. systems-based [Car12].

Table [WWN^{+}25]. Table-based [WWN^{+}25]. Tables [CS94, BM12]. Tablets [NSVB18]. Tactics [SSGC89]. TAE [SS93]. Tag [YRC23, ZHC^{+}20]. Tagging [CLL24, CP13, NLF^{+}20, ZLY^{+}22, SSS11]. Tags [CLL24, MO12, NZW^{+}14, CDR10]. Tail [HCC^{+}17]. tailed [YZZS24]. Tailorable [MLF95]. Talent [QZX^{+}20, YZS^{+}24]. Tank [ABW93]. Target [AZCC21, ACC18, SLQ24, WLL24, WYZ^{+}25, WLM^{+}21, ZMN^{+}24, ZWZS25, GLG^{+}23].

Target-constrained [WLL24]. Target-guided [WLM^{+}21]. Target-oriented [WLL24]. Target-Unknown [ZMN^{+}24]. Targeted [ZGN^{+}17, CZJ07]. TASC [TQH15]. Task [ABW93, CR92, CWL^{+}20, JVA^{+}15, JS24, LXC^{+}25, SGZ^{+}24, XMS^{+}25, ZHZ^{+}19, ZKJW22, ZWF^{+}23, ZDZ23, BHY^{+}25, CDR10, DZX^{+}23, SZM^{+}23, XCG^{+}23, XPSC24, ZLL^{+}22, ZJW^{+}24]. Task-Artifact [CR92]. Task-Based [JVA^{+}15]. task-dependent [CDR10]. Task-Oriented [XMS^{+}25, ZHZ^{+}19, SGZ^{+}24]. Tasks [CCAC24, HYH15, SC19, TY22, TZP23, LOP^{+}13]. Taxonomies [ZZL^{+}23]. Taxonomy [CC05, HYM^{+}25, SLW^{+}24, ZYY^{+}24, YHYO01]. TCGC [TWS^{+}25]. TCKT [HLH^{+}25]. Teach [YSW^{+}24]. Team [RFB^{+}24, Luz12]. Techniques [Jac91, RRS95, SW90, CDZ08, HCZ04, JK02]. Technological [OG94]. Technologies [HRY^{+}22a, HRY^{+}22b]. Technology [BvdHFS25, LK20, MR92, OG94, RP98, SBH24, ZJT16, ZK20]. Technology-Assisted [BvdHFS25, LK20, SBH24, ZK20]. Tell [NSVB18]. Telos [MBJK90]. Templar [Tuz95]. Temporal [AGT24, CL19, FHW^{+}19, HWT^{+}25, JD07, LLY^{+}24, MSMO18, PGC^{+}25, RBM23, TCSL24, TWS^{+}25, Tuz95, WMZ^{+}21, XZCZ25, ZCS^{+}23, PMD01, TP05, WLW^{+}24a, XC24, YWYL08, LLPT10]. Temporality [JYM13]. Tendency [SMN^{+}16]. Tensor [LCJ^{+}16, LJHL17, LCJ^{+}19]. Tensor-Based [LCJ^{+}16, LCJ^{+}19]. Term [BLWJ15, DLA22, GCN^{+}19, LYN^{+}21, XPL^{+}24, XYX^{+}25, ZHY14, AOU12, BS12, CRG02, CZJ07, HO07, PR09, SJZ^{+}25, TFS^{+}12, WKF^{+}12, WLWK08]. Term-Independence [DLA22]. term-ranking [CRG02]. Terms [Kwo90, SKK24, BBD11]. Test [CW94]. testing [Car12, HHWL01]. Text [ADW94, BOQ^{+}25, CZJH24, ES13, EMS19, HHC^{+}23, HWN^{+}19, KMKO92, KJA20, KBD89, LRA23, LYY^{+}24, LYS^{+}17, LPY94, MLZ^{+}22a, Man97, MZ96, MPS23, RS93, RL94, RBM23, STSM95, SLM^{+}24, SWZ^{+}22, TBC96, XLW^{+}22, YC94, YHS^{+}21, YZZS24, ZLF^{+}17, ZGC^{+}22, ZLRW24, BC13, BFNP10, BS12, CC01, CZJ07, CC05, CS99, FBN^{+}12, GWLC06, IG08, JSL^{+}24, KKS^{+}08, Leh06, LCL04, MC10a, MRYGM01, RZCG^{+}10, SC09, SNZBY00, TS10]. text-based [CZJ07]. Text-image [XLW^{+}22]. Textbook [Rad92]. Texts [CHN^{+}17, LDW^{+}17, LRZ^{+}17]. Textstream [PGW^{+}17]. Textual [APCC15, KC15, MAL16, SC09]. Texture [VB98]. TF [WLWK08]. TF-IDF [WLWK08]. TFGIN [WWN^{+}25]. Their [LAL^{+}17, TY22, WLS^{+}17, WRLD22]. Them [ZAL^{+}16]. themes [HL10]. Theoretic [PBS22, CdMRB01, Fuh99, MR10]. Theoretical [MC14b, WW10]. Theories [TCMC21]. Theory [BK90, ISL95, JVA^{+}15, KM97, Kwo90, MHR22, PZH^{+}24, ZRZ02]. these [FEL22]. Things [CDR21]. Thorough [TCQ^{+}25]. Thoroughly [QXX^{+}25]. Three [CS94, LWP^{+}21]. Three-tier [LWP^{+}21]. threshold [CCC09]. tier [LWP^{+}21]. Ties [SHS^{+}23, TLKW16]. Tight [WWN^{+}25]. Tight-Fitting [WWN^{+}25]. Time [GRSD^{+}16, HML^{+}23, LJHL17, LMZ23b, LXW^{+}17, MK16b, SYX^{+}17, YZH^{+}24, ZLW16, ZGCC18, ZWC^{+}23, ZML^{+}19, AL13, HSS^{+}09, MTL23, Non13, YCZ^{+}16, ZZL^{+}21]. Time-Aware [LJHL17, LXW^{+}17, ZGCC18, HML^{+}23, LMZ23b, ZWC^{+}23]. Timeline [CLG^{+}23]. TISoN [HGBY16]. TME [XCG^{+}23]. TODOS [PBF^{+}89]. TOIS [All91, Ano94b, Cal12, Mar06, Mar07, Mar08b, ACM03]. Token [LGM^{+}24, MDX^{+}25, WWX^{+}24].

Token-Event-Role [WWX^{+}24]. too

[GL05]. Tool [CWL^{+}20, CACW23, MLF95, PBF^{+}89, WBB^{+}90]. Toolkit [SGLS93, PCD92]. Tools [Res93, BVKS08]. Top [ABCZ21, CWZ^{+}20b, CWZ^{+}20a, CVS21, LPLH19, LHZ^{+}25, SC19, WZZ^{+}23, XHW^{+}19, ZTM^{+}17, ZLF^{+}17, ZLF^{+}23, DK04]. Top- [CVS21, LHZ^{+}25, SC19, ZTM^{+}17, ZLF^{+}17, ABCZ21, CWZ^{+}20b, CWZ^{+}20a, LPLH19, WZZ^{+}23, XHW^{+}19, ZLF^{+}23, DK04]. Topic [BZCE22, BLWJ15, CLL24, CFFK22, JTS16, LDW^{+}17, LCX^{+}19, LZZ^{+}23, LZW^{+}20, MHZ16, MCF^{+}17, NLF^{+}20, SQZ^{+}22, SG14, WWX^{+}25, WXLC13, WZW23, XZH^{+}22, ZWX^{+}24, ZHD^{+}22, GMR09, HL10, RZCG^{+}10]. Topic-aware [LZZ^{+}23]. Topic-based [BZCE22]. Topic-graph [ZHD^{+}22]. Topical [ACC18, AAGY01]. Topics [HLL^{+}17, LLBS18, LHG^{+}12, ZZH^{+}18, ZHD^{+}22, BM02, GMR09]. TopPRF [MHZ16]. Touch [HYH15]. Tour [GXTL14]. Trace [SFR98]. Trace-Based [SFR98]. Tracing [CCZ^{+}23, CYZ^{+}24, CQZ^{+}25b, HLW^{+}24, LBZ^{+}24, LWC^{+}25, PLS^{+}24]. Tracking [HLC^{+}20, LHG^{+}12]. tractable [GCT99]. trading [CGM02]. Traditional [SZ21]. Traffic [WZC^{+}17]. Trained [LYP^{+}25, WLF^{+}25a, WLF^{+}25b, QXX^{+}25]. Training [CQL^{+}24, CZL^{+}25, CHZ24, ES13, FCL^{+}23, PXW^{+}25, WZW^{+}22, WYZ^{+}25, WNCH24, XZZ^{+}22, XPSC24, YZS^{+}24, ZWMH19, ZOML24, HYZ^{+}23, HWT^{+}25, LMMO23, MMH^{+}24, SWZW25, SAW^{+}25, WLL^{+}23a, WZZ^{+}23, WXZ^{+}25, YZL^{+}25]. Trajectories [TGS16, YSZ^{+}17]. Trajectory [ZSQ^{+}24, ZZZ^{+}16]. Transactions [All89, dR19]. Transcripts [SE91]. Transfer [EMS19, HLH^{+}25, NXP^{+}24, PTS13, PYC^{+}19, TLKW16, ZYC^{+}24]. Transferable [YZL^{+}25]. Transferring [ZMN^{+}24]. transform [PRP05].

Transformation [CZT^{+}23, TQH15, PTKJ07]. Transformation-Aware [TQH15]. Transformer [AVA^{+}24, CCZ^{+}23, MRC^{+}23, WZZ^{+}23, XC24, ZQH^{+}24b, ZSLK24]. Transformer-Based [ZSLK24, AVA^{+}24, MRC^{+}23]. Transformers [MLP22, RDK24, SRF^{+}24, LLP24]. Transient [WS90]. Transition [MYL^{+}23]. transitive [LCL04]. translating [CGMP99]. Translation [MTKY92, QLNY19, TL14, YHS^{+}12, LCH07, LCL04, PTKJ07]. Translation-Based [QLNY19]. translations [LCH07]. Transmission [YYQ^{+}25]. Transmission-Free [YYQ^{+}25]. Transparent [ZZD^{+}20]. Transportable [SS93]. Transportation [ZGN^{+}17]. Travel [GXTL14, LGL^{+}24]. Traveling [ZLW16]. Treaps [KNCLO17]. Treatment [YLW^{+}23]. TRECVID [AOK14]. Tree [HLH^{+}25, XCG^{+}23]. Tree-Based [HLH^{+}25]. Tree-guided [XCG^{+}23]. treelike [MS01]. Trees [DLN^{+}16, TGS16, CBZ11]. Trending [MCF^{+}17]. Trials [Res93]. tries [HZW02]. TriMLP [JXY^{+}24]. Trip [ZLW16]. Triple [MXM^{+}24, ZSZ^{+}24]. TROLL [JSHS96]. Troubling [DBCJ21]. Truncated [PARS22]. Trust [AZK12, FMSS14, HGBY16, MK16a, PBJ^{+}16, ZM14]. Trust-Oriented [HGBY16]. Trusting [ZKG08]. Trustworthy [Gla04, GL05, YSXK23a, YSXK23b]. Truthfulness [XHSD24]. TSSuBERT [DPSH23]. tunable [BS12]. tune [LGZ^{+}24]. Tuning [BHY^{+}25, LQZ^{+}25, LHZ^{+}25, PSPBY10, YOM24]. Turn [GCL^{+}21, LLT^{+}21, ZCZ^{+}23, CJY23a, FTZZ23]. Tutoring [LLY^{+}24]. Tweet [CL18, FC17, AL13]. Tweets [CL19, DPSH23, JYM13, MPT21]. Twitter [NSK^{+}20, YCZ^{+}15]. Two [Na15]. Two-Stage [Na15]. Type [LM90, ZMZ25].

Typed [Guy18]. Typed-in [Guy18]. Types [CWL^{+}20, XTF^{+}21].

UAN [HSH90]. Ubiquitous [HSH93].

Ukkonen [BR96]. Unbiased [AYWM21, Oos23]. Uncertainty [GJS^{+}25, LXLZ16, Mor90, ZLW16]. Uncertainty-Aware [GJS^{+}25]. underlying [Lin07]. Understanding [BACF17, CLF^{+}25, CJY23a, CCAC24, FEL22, HYH15, LLC^{+}25b, MRM24, NFZ19, RBCT14, SWL^{+}23, SAD24, UA22, WLG^{+}17, YFSO24, ZJT16, ZXZL20, ZPW^{+}24, ZQH^{+}24a, ZAK^{+}23]. Unexplored [CJZ^{+}16]. Unidraw [VL90]. Unified [BK90, DZX^{+}23, FLW^{+}25, LTC^{+}25, WdVR08, ZSQ^{+}24, ZZD25]. Uniform [YZL^{+}25]. Unifying [GCL^{+}21, LLW^{+}21, SZY^{+}24, WNS^{+}17]. Universal [USN^{+}17, WZL^{+}24, ZSH^{+}25]. Unknown [ZMN^{+}24]. unlabeled [BJL^{+}07, ZCD06]. Unlearning [CZL^{+}25, DLY^{+}25, HNN^{+}25]. Unreliable [SYSW24]. unseen [LCH07]. Unstructured [MLZ^{+}22a]. Unsupervised [AMS20, APCC15, CJN20, GLZ^{+}24, PZH^{+}24, SWZ^{+}22, VdRK18, WWT^{+}25, YE18]. Upper [MOT11]. Upper-bound [MOT11]. Urban [WZC^{+}17, WLZW21, dODMS20]. Urn [CPL15]. usability [HHWL01]. Usage [VHJC11, FPSO06]. Usage-Based [VHJC11]. Use [CW94, CWL^{+}20, Jac91, LM90, WLG^{+}17, ARSZ03, LLC03]. Useful [BRS92]. User [ABCZ21, ACC18, BACF17, CQF^{+}23, CJZ^{+}16, CZT^{+}23, CLF^{+}25, CL18, CHBS15, CDR21, DC92, DCT^{+}17, FSE^{+}24, GCL^{+}21, HSH90, HRY^{+}22a, HRY^{+}22b, HCC^{+}16, HM90, IYC^{+}23, JS24, LSOS06, LCPD19, LGL^{+}24, LQJ^{+}24, LRSD25, LRZ^{+}17, LAL^{+}17, LWP^{+}21, LWH^{+}23, LMW^{+}23, MR92, MLK^{+}18, MYHL11, MBST17, MAL16, MCL^{+}12, MNP^{+}21b, NSVB18, Pal15, PCD92, QZS^{+}23, QFZ^{+}19, SZ21, SWZ^{+}24, SCL^{+}22, SZC^{+}23, SE91, SAD24, SAK^{+}23, SGZ^{+}24, SS93, WZW^{+}19, WMZ^{+}21, WHL^{+}22, WLT^{+}22, WLW^{+}24b, WYZ^{+}25, WZY^{+}25, Wie92, WQL^{+}19, WZL^{+}19, WWL^{+}25, XYW^{+}23b, XHYY09, XYZ^{+}23, YCL^{+}25, YE18, YCL^{+}23, YCZ^{+}16, ZYW^{+}23, ZLM^{+}23, ZXZL23, ZWE^{+}22, ZCS^{+}23, ZZG^{+}25, ZAK^{+}23, CZJ07, CC03, DEL^{+}00, KFS10, MSR04, PGD07, WNZ02]. User-Item [WQL^{+}19, ZCS^{+}23]. User-Oriented [HSH90]. User-Side [SWZ^{+}24]. user-specific [DEL^{+}00]. User/Item [LRSD25]. Users [HCC^{+}17, LLW^{+}21, LLM22, LPY94, LFW23, LMW^{+}23, SZC^{+}23, SMT25, YTL25, YCZ^{+}15, ZHC^{+}20, ZAK^{+}23]. Using [AZK12, APCC15, AM25, APJL20, BF98, BMMC19, BvdHFS25, CE89, CGB^{+}15, DH97, EMG11, FKS19, HL10, JDA^{+}22, VST^{+}14, LMR^{+}24, LCPD19, LNO^{+}25, MWZ^{+}25, MAA^{+}25, NSVB18, NCZG14, RS93, RVA^{+}24, Row96, SKC16, TSNS19, TM20, Tuz95, VSS17, XC98, YLW^{+}23, ZZD^{+}20, dODMS20, ASST05, ADCU08, BJL^{+}07, BM02, BKGM^{+}02, CML00, CCC09, Coh00, JZ06, KL10, LPLH19, PRP05, PH07, RBV09, SC09, SZ09, WNZ02, WRJV05, XC13, ZCD06, Zob06]. Utility [DXZ^{+}22, ZPW^{+}24, CSJS10]. Utility-Oriented [DXZ^{+}22]. Utilizing [KKB10, WZC^{+}25].

Vague [ZGL^{+}25]. validation [CJRY12]. Valuable [MJH15]. Value [ZWT^{+}23, MLDGH04]. valued [ZHW^{+}22]. Variability [RZM24]. Variants [DGMG23, XC98]. Variational [CWZ^{+}20a, CHZ24, FLMD22, RFB^{+}24, WZW^{+}22, WQH^{+}25, YRC23, ZZY^{+}23, ZMZ25]. Variations [BMMC19]. variety [CML00]. Various [SJL^{+}25]. Varying [CACW23]. Vector [VdRK18, PZMRN05, TMT06]. Vectors [BNIL24a]. Venue [CZX^{+}22, CL18, LMZ23b, XCG^{+}23].

Veracity [MK16a, PBJ^{+}16]. Verification [LBM^{+}24, SFR98, WBP^{+}16, WWN^{+}25, YPF^{+}25, ZWMH19]. Versatile [SKK24]. Version [HLNS98, YZW^{+}17]. Version-Aware [YZW^{+}17]. versus [CRNZ^{+}03, MC06, TSNS19, WW10]. Vertical [SBM^{+}17, WC21b]. very [BJL^{+}07]. Via [MP20, WQH^{+}25, AK18, BLWJ15, CQF^{+}23, CST^{+}23, CWZ^{+}24, CQL^{+}24, CHZ24, DLZ^{+}22, DHY^{+}19, FZH^{+}22, FWG^{+}24, FLW^{+}25, GS91, HZSL13, HYLC24, LJHL17, LHH^{+}19, LXZ^{+}22, LWL^{+}25, LZG^{+}18, LLM22, LLG^{+}23, LPYM23, LCL^{+}17, LZG21, LWP^{+}21, LWL^{+}24, LBZ^{+}24, LLC^{+}25b, LLP24, NLF^{+}20, PZH^{+}24, SZC^{+}23, SMN^{+}16, SLD^{+}24, TZG^{+}24, WZC^{+}25, WWL^{+}25, XZCZ25, XWZ^{+}25, YLZW21, YLYT23, ZLL^{+}20, ZM14, ZLH^{+}22, ZLY^{+}22, ZWZ^{+}24, ZWX^{+}23, ZWZS25, ZHRK23, ZSLK24]. Video [AOK14, BA18, Bul95, CTH^{+}13, CR95, DG95, JTL^{+}17, LWL^{+}24, LCJ^{+}19, MYHL11, QDG^{+}24, SL95, TM91, TQH15, VHJC11, WWT^{+}25, YHDN25, BFF^{+}03, HSS^{+}09, HHC^{+}10, SSS11, Zob06]. Video-length [QDG^{+}24]. Video-Related [JTL^{+}17]. VideoDraw [TM91]. Videos [LXZ^{+}22, NLL^{+}24, LHSC11]. View [AGT24, LGQL19, WZY^{+}24, ZWrWZ23, ZQH^{+}24b, LLP24, MHW^{+}24, QZW^{+}25, XYZ^{+}23, ZZZ^{+}24]. Viewpoint [JVA^{+}15]. Viewport [ZML^{+}19]. Views [Koi95, LM90, ZYL^{+}25, AOU12]. Virtual [ABW93, FG93, FZC93, SGLS93, WNS^{+}17]. Visibility [GGR22]. Visited [HY20]. Visual [APCC15, BA18, GJS^{+}25, LZM^{+}19, LMW^{+}14, LZ96, WGZ^{+}25, YXL^{+}23, KNS04]. Visual-Context [WGZ^{+}25]. Visualization [Koi93, CL06]. visualization-based [CL06]. Visually [CYC^{+}24, YYY^{+}24]. Visually-Aware [CYC^{+}24]. Vital [XWL^{+}25]. VM [WC21b]. VM-NSP [WC21b]. Vocabulary [Sav12, ZMK^{+}20, WKF^{+}12]. Voice [Guy18]. Volume [dR19]. Volunteerism [SMN^{+}16]. vs [AA16, CMZ^{+}23, DM07, GGRF23]. Vulnerability [CLF^{+}25, GYW^{+}25, XLL^{+}23].

Walk [LCL^{+}17, MC10b]. Wardrobe [PBS22]. Warm [SZC^{+}23]. Warm-up [SZC^{+}23]. wavelet [PRP05]. Way [WDL^{+}24]. Weak [KSVS16, LZC24a, RLC^{+}25]. Wearable [FC17]. web [HHWL01, IG08, SPS^{+}06, dMdSsdA^{+}08, ACS08, AZK12, AKFA15, ANW17, AOU12, BACF17, BZDW24, BJL^{+}07, BLWJ15, BKGM^{+}02, CRNZ^{+}03, CC05, CFC^{+}19, EJ03, FPSO06, FKM^{+}05, FP00, GKL^{+}07, GIS03, Guy18, HYH15, HHC^{+}10, IM05, JDA^{+}22, JGP^{+}07, KEW01, LLY^{+}25, LS02, LHG^{+}12, LCL04, MRYGM01, PPGK04, RSD^{+}13, RBCT14, STZ22, STC^{+}24, SPKW11, TM10, VAJR21, VDR25, WH09, WH15, ZHD^{+}22]. Web-based [CC05]. Web-Enhanced [LLY^{+}25]. web-query [SPS^{+}06]. WebGLM [LLY^{+}25]. WebQuilt [HHWL01]. Weight [LR96]. Weight-Partitioned [LR96]. Weighted [BCFG15, PARS22, VFA17, DWLW09]. Weighting [MNP^{+}21a]. weights [WLWK08]. Wellness [FC17]. Where [YCZ^{+}15]. Who [YCZ^{+}15]. Whole [RK19, RBCT14, SZ93]. Whole-Session [RBCT14]. Why-Questions [AA16]. Wide [PPGK04]. Wiecha [Gru92]. Wiki [ZJL^{+}24]. Wikipedia [VST^{+}14, SG14]. Wildfires [WBP^{+}16]. Window [VAJR22]. Windowing [KMKO92]. Without [OOSC93, CZZ^{+}20, KYK24, KL10, LHA^{+}25a, ZZG^{+}25]. witness [Gla04]. Word [BTP24, BKR97, CGB^{+}15, FBN^{+}12, HCC14, LDW^{+}17, SHN17, XC98, YDW22, Coh00, HZSL13, SNZBY00]. Word-based [FBN^{+}12, Coh00]. Words [RDK24]. Work [KM97, MLF95, Mar94, Ols89, Ruh94].

Workers [RLS^{+}24, SC19]. Working

[AC19]. workloads [CML00]. Workspace [IKG93, Non13]. World [GZW^{+}18, ZLW16, PPGK04]. Worlds [ABW93, FG93, WNS^{+}17]. Writeprints [AC08]. Writing [TWLX18].

X [COR13]. X-Class [COR13]. XIRQL [FG04]. XLORE [ZJL^{+}24]. XML [BYFM06, BGQT07, BCPN14, CDF01, COR13, FCD02, FG04, KMdRS06, KL06, Leh06, PH07, PTL08, TG10]. XMovie [KEL95]. XPath [BCPN14, GL06]. XXS [BCPN14].

YASS [MMP^{+}07]. year [Gla04, GL05]. Years [DPSH23]. YouTube [SSS11]. Yum [YHY^{+}17]. Yum-Me [YHY^{+}17].

zero [WTS^{+}08]. zero-one [WTS^{+}08].

References

Azmi:2016:AAW

[AA16] Aqil M. Azmi and Nouf A. Alshenaifi. Answering Arabic why-questions: Baseline vs. RST-based approach. ACM Transactions on Information Systems, 35(1):6:1–6:??, October 2016. CODEN ATISET. ISSN 1046-8188.

Aggarwal:2001:DLC

[AAGY01] Charu C. Aggarwal, Fatima Al-Garawi, and Philip S. Yu. On the design of a learning crawler for topical resource discovery. ACM Transactions on Information Systems, 19 (3):286–309, July 2001. CO-DEN ATISET. ISSN 1046-

Ahmad:2020:DLA

[AAL^{+}20] Faizan Ahmad, Ahmed Abbasi, Jingjing Li, David G. Dobolyi, Richard G. Netemeyer, Gari D. Clifford, and Hsinchun Chen. A deep learning architecture for psychometric natural language processing. ACM Transactions on Information Systems, 38(1):6:1– 6:29, February 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3365211.

Alserafi:2020:KDL

[AARC20] Ayman Alserafi, Alberto Abelló, Oscar Romero, and Toon Calders. Keeping the data lake in form: Proximity mining for pre-filtering schema matching. ACM Transactions on Information Systems, 38(3):26:1–26:30, June 2020. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/doi/abs/10.
1145/3388870.

Anagnostopoulos:2015:SQC

[ABB^{+}15] Aris Anagnostopoulos, Luca Becchetti, Ilaria Bordino, Stefano Leonardi, Ida Mele, and Piotr Sankowski. Stochastic query covering for fast approximate document retrieval. ACM Transactions on Information Systems, 33(3):11:1– 11:??, February 2015. CO-DEN ATISET. ISSN 1046-8188.

Adomavicius:2021:EPA

[ABCZ21] Gediminas Adomavicius, Jesse Bockstedt, Shawn Curley, and Jingjing Zhang. Effects of personalized and aggregate top-N recommendation lists on user preference ratings. ACM Transactions on Information Systems, 39(2):13:1– 13:38, March 2021. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3430028.

Arthur:1993:ETP

[ABW93] Kevin W. Arthur, Kellogg S. Booth, and Colin Ware. Evaluating 3D task performance for fish tank virtual worlds. ACM Transactions on Information Systems, 11(3):239– 265, July 1993. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:

80. Special Issue on Virtual Worlds.

Abbasi:2008:WSA

[AC08] Ahmed Abbasi and Hsinchun Chen. Writeprints: a stylometric approach to identitylevel identification and similarity detection in cyberspace. ACM Transactions on Information Systems, 26(2):7:1– 7:??, March 2008. CODEN ATISET. ISSN 1046-8188.

Arguello:2016:EAS

[AC16] Jaime Arguello and Rob Capra. The effects of aggregated search coherence on search behavior. ACM Transactions on Information Systems, 35(1):2:1–2:??, October 2016. CODEN ATISET. ISSN 1046-8188.

Aliannejadi:2018:PCA

[AC18] Mohammad Aliannejadi and Fabio Crestani. Personalized context-aware point of interest recommendation. ACM Transactions on Information Systems, 36(4):45:1– 45:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Arguello:2019:EWM

[AC19] Jaime Arguello and Bogeum Choi. The effects of working memory, perceptual speed, and inhibition in aggregated search. ACM Transactions on Information Systems, 37 (3):36:1–36:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3322128.

Arguello:2018:FIU

[ACC18] Jaime Arguello, Bogeum Choi, and Robert Capra. Factors influencing users’ information requests: Medium, target, and extra-topical dimension. ACM Transactions on Information Systems, 36 (4):41:1–41:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Ackerman:1998:AOM

[Ack98] Mark S. Ackerman. Augmenting organizational memory: a field study of Answer Garden. ACM Transactions on Information Systems, 16(3): 203–224, July 1998. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/pubs/citations/journals/
tois/1998-16-3/p203-ackerman/

.

Aridor:2002:KEF

[ACM^{+}02] Yariv Aridor, David Carmel, Yoelle S. Maarek, Aya Soffer, and Ronny Lempel. Knowledge encapsulation for focused search from pervasive devices. ACM Transactions on Information Systems, 20(1):25– 46, January 2002. CODEN ATISET. ISSN 1046-8188.

Staff:2003:TR

[ACM03] ACM Transactions on Information Systems Staff. TOIS reviewers. ACM Transactions on Information Systems, 21(4):492–493, October 2003. CODEN ATISET. ISSN 1046-8188.

Abbasi:2008:SAM

[ACS08] Ahmed Abbasi, Hsinchun Chen, and Arab Salem. Sentiment analysis in multiple languages: Feature selection for opinion classification in Web forums. ACM Transactions on Information Systems, 26(3):12:1–12:??, June 2008. CODEN ATISET. ISSN 1046-

Altingovde:2008:ICB

[ADCU08] Ismail Sengor Altingovde, Engin Demir, Fazli Can, and Özgür Ulusoy. Incremental cluster-based retrieval using compressed cluster-skipping inverted files. ACM Transactions on Information Systems, 26(3):15:1–15:??, June 2008. CODEN ATISET. ISSN 1046-8188.

Apte:1994:ALD

[ADW94] Chidanand Apte, Fred Damerau, and Sholom M. Weiss. Automated learning of decision rules for text categorization. ACM Transactions on Information Systems, 12(3): 233–251, July 1994. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Text Categorization.

Agosti:2007:FMA

[AF07] Maristella Agosti and Nicola Ferro. A formal model of annotations of digital content. ACM Transactions on Information Systems, 26(1):3:1– 3:??, November 2007. CO-DEN ATISET. ISSN 1046-8188.

Alakuijala:2019:BGP

[AFF^{+}19] Jyrki Alakuijala, Andrea Farruggia, Paolo Ferragina, Eugene Kliuchnikov, Robert Obryk, Zoltan Szabadka, and Lode Vandevenne. Brotli: a general-purpose data compressor. ACM Transactions on Information Systems, 37 (1):4:1–4:??, January 2019. CODEN ATISET. ISSN 1046-8188.

An:2024:MMV

[AGT24] Jingmin An, Ming Gao, and Jiafu Tang. MvStHgL: Multi-view hypergraph learning with spatial-temporal periodic interests for next POI recommendation. ACM Transactions on Information Systems, 42(6):152:1–152:??, November 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3664651.

Arampatzis:2018:SPI

[AK18] Avi Arampatzis and Georgios Kalamatianos. Suggesting points-of-interest via contentbased, collaborative, and hybrid fusion methods in mobile devices. ACM Transactions on Information Systems, 36(3):23:1–23:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Alhindi:2015:PBS

[AKFA15] Azhar Alhindi, Udo Kruschwitz, Chris Fox, and M-Dyaa Albakour. Profile-based summarisation for Web site navigation. ACM Transactions on Information Systems, 33(1):4:1–4:??, March 2015. CODEN ATISET. ISSN 1046-

Asadi:2013:FCG

[AL13] Nima Asadi and Jimmy Lin.

Fast candidate generation for real-time tweet search with Bloom filter chains. ACM Transactions on Information Systems, 31(3):13:1–13:??, July 2013. CODEN ATISET. ISSN 1046-8188.

Allen:1989:ENN

[All89] R. B. Allen. Editorial: a new name — ACM Transactions on Information Systems. ACM Transactions on Information Systems, 7 (1):1–2, January 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Allen:1991:ECH

[All91] Robert B. Allen. Editorial:

Computer-human interaction and ACM TOIS. ACM Transactions on Information Systems, 9(2):97–98, April 1991. CODEN ATISET. ISSN 1046-

8188. URL http://www.

acm.org:80. Special Issue on Computer-Human Interaction.

Allen:1994:E

[All94] Robert B. Allen. Editorial.

ACM Transactions on Information Systems, 12(1):1, January 1994. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Anand:2024:DAS

[ALS^{+}24] Abhijit Anand, Jurek Leonhardt, Jaspreet Singh, Koustav Rudra, and Avishek Anand. Data augmentation for sample efficient and robust document ranking. ACM Transactions on Information Systems, 42(5):119:1–119:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3634911.

Afsarmanesh:1989:EOO

[AM89] Hamideh Afsarmanesh and Dennis McLeod. The 3DIS: An extensible, object-oriented information management environment. ACM Transactions on Information Systems, 7(4):339–377, October 1989. CODEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Anand:2025:SRS

[AM25] Vineeta Anand and Ashish Kumar Maurya. A survey on recommender systems using graph neural network. ACM Transactions on Information Systems, 43(1):9:1– 9:??, January 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3694784.

Agosti:2020:LUK

[AMS20] Maristella Agosti, Stefano Marchesin, and Gianmaria Silvello. Learning unsupervised knowledge-enhanced representations to reduce the semantic gap in information retrieval. ACM Transactions on Information Systems, 38 (4):38:1–38:48, October 2020. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3417996.

Anonymous:1993:AI

[Ano93] Anonymous. 1993 author index. ACM Transactions on Information Systems, 11(4): 425–426, October 1993. CO-DEN ATISET. ISSN 1046-8188.

Anonymous:1994:IA

[Ano94a] Anonymous. Information for authors. ACM Transactions on Information Systems, 12 (3):333–337, July 1994.

Anonymous:1994:TC

[Ano94b] Anonymous. TOIS charter.

ACM Transactions on Information Systems, 12(1):3, January 1994. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Anonymous:1995:AR

[Ano95]  Anonymous.  Acknowledg-

ment to referees. ACM Transactions on Information Systems, 13(1):112–113, January 1995. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Anonymous:1996:AI

[Ano96a] Anonymous. 1996 author index. ACM Transactions on Information Systems, 14(4): 471–472, October 1996. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80/tois/abstracts/cheung.
html.

Anonymous:1996:MGS

[Ano96b] Anonymous. In memoriam: Gerard Salton. ACM Transactions on Information Systems, 14(1):1, January 1996. CODEN ATISET. ISSN 1046-8188.

Anonymous:1997:AI

[Ano97] Anonymous. 1997 author index. ACM Transactions on Information Systems, 15(4): 436–437, October 1997. CO-DEN ATISET. ISSN 1046-8188.

Alkwai:2017:CAR

[ANW17] Lulwah M. Alkwai, Michael L. Nelson, and Michele C. Weigle. Comparing the archival rate of Arabic, English, Danish, and Korean language Web pages. ACM Transactions on Information Systems, 36(1): 1:1–1:??, August 2017. CO-DEN ATISET. ISSN 1046-8188.

Awad:2014:CBV

[AOK14] George Awad, Paul Over, and Wessel Kraaij. Contentbased video copy detection benchmarking at TRECVID. ACM Transactions on Information Systems, 32(3):14:1– 14:??, June 2014. CODEN ATISET. ISSN 1046-8188.

Altingovde:2012:SIP

[AOU12] Ismail S. Altingovde, Rifat Ozcan, and Özgür Ulusoy. Static index pruning in Web search engines: Combining term and document popularities with query views. ACM Transactions on Information Systems, 30(1):2:1– 2:??, February 2012. CODEN ATISET. ISSN 1046-8188.

Ah-Pine:2015:UVT

[APCC15] Julien Ah-Pine, Gabriela Csurka, and Stéphane Clinchant. Unsupervised visual and textual information fusion in CBMIR using graphbased methods. ACM Transactions on Information Systems, 33(2):9:1–9:??, February 2015. CODEN ATISET. ISSN 1046-8188.

Arapakis:2020:PPA

[APJL20] Ioannis Arapakis, Antonio Penta, Hideo Joho, and Luis A. Leiva. A price-per-attention auction scheme using mouse cursor information. ACM Transactions on Information Systems, 38(2):13:1– 13:30, January 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3374210.

Amato:2003:RPM

[ARSZ03] Giuseppe Amato, Fausto Rabitti, Pasquale Savino, and Pavel Zezula. Region proximity in metric spaces and its use for approximate similarity search. ACM Transactions on Information Systems, 21 (2):192–227, 2003. CODEN ATISET. ISSN 1046-8188.

Adalia:2025:LFP

[ASMZ25] Ramon Adàlia, Gemma Sanjuan, Tomàs Margalef, and Ismael Zamora. The Lambda-Gap framework for precisionoriented ranking. ACM Transactions on Information Systems, 43(4):97:1–97:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Albahem:2022:CBA

[ASSC22] Ameer Albahem, Damiano Spina, Falk Scholer, and Lawrence Cavedon. Componentbased analysis of dynamic search performance. ACM Transactions on Information Systems, 40(3):61:1– 61:47, July 2022. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3483237.

Adomavicius:2005:ICI

[ASST05] Gediminas Adomavicius, Ramesh Sankaranarayanan, Shahana Sen, and Alexander Tuzhilin. Incorporating contextual information in recommender systems using a multidimensional approach. ACM Transactions on Information Systems, 23(1):103–145, January 2005. CODEN ATISET. ISSN 1046-8188.

Aiken:1991:IES

[ASV91] Milam W. Aiken, Olivia R. Liu Sheng, and Douglas R. Vogel. Integrating expert systems with group decision support systems. ACM Transactions on Information Systems, 9(1):75–95, January 1991. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Anderson:2000:CHH

[ATW00] Kenneth M. Anderson, Richard N.

Taylor, and E. James Whitehead. Chimera: hypermedia for heterogeneous software development enviroments. ACM Transactions on Information Systems, 18 (3):211–245, 2000. URL

http://www.acm.org/pubs/
citations/journals/tois/
2000-18-3/p211-anderson/.

Amati:2002:PMI

[AV02] Gianni Amati and Cornelis Joost Van Rijsbergen. Probabilistic models of information retrieval based on measuring the divergence from randomness. ACM Transactions on Information Systems, 20(4):357–389, 2002. CODEN ATISET. ISSN 1046-8188.

Askari:2024:REL

[AVA^{+}24] Arian Askari, Suzan Verberne, Amin Abolghasemi, Wessel Kraaij, and Gabriella Pasi. Retrieval for extremely long queries and documents with RPRS: a highly efficient and effective transformer-based re-ranker. ACM Transactions on Information Systems, 42(5):115:1– 115:??, September 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3631938.

Ai:2021:ULR

[AYWM21] Qingyao Ai, Tao Yang, Huazheng Wang, and Jiaxin Mao. Unbiased learning to rank: Online or offline? ACM Transactions on Information Systems, 39(2): 21:1–21:29, March 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3439861.

Adomavicius:2012:SRA

[AZ12] Gediminas Adomavicius and Jingjing Zhang. Stability of recommendation algorithms. ACM Transactions on Information Systems, 30(4):23:1– 23:??, November 2012. CO-DEN ATISET. ISSN 1046-8188.

Ai:2020:EPS

[AZBC20] Qingyao Ai, Yongfeng Zhang, Keping Bi, and W. Bruce Croft. Explainable product search with a dynamic relation embedding model. ACM Transactions on Information Systems, 38(1):4:1– 4:??, February 2020. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.

acm.org/ft_gateway.cfm?
id=3361738.

Aliannejadi:2021:CAT

[AZCC21] Mohammad Aliannejadi, Hamed Zamani, Fabio Crestani, and W. Bruce Croft. Contextaware target apps selection and recommendation for enhancing personal mobile assistants. ACM Transactions on Information Systems, 39(3): 29:1–29:30, July 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3447678.

Abbasi:2012:DFM

[AZK12] Ahmed Abbasi, Fatemeh ‘Mariam’ Zahedi, and Siddharth Kaza. Detecting fake medical Web sites using recursive trust labeling. ACM Transactions on Information Systems, 30(4): 22:1–22:??, November 2012. CODEN ATISET. ISSN 1046-8188.

Bekhet:2018:ISB

[BA18] Saddam Bekhet and Amr Ahmed. An integrated signature-based framework for efficient visual similarity detection and measurement in video shots. ACM Transactions on Information Systems, 36(4):37:1–37:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Bai:2017:ULI

[BACF17] Xiao Bai, Ioannis Arapakis, B. Barla Cambazoglu, and Ana Freire. Understanding and leveraging the impact of response latency on user behaviour in Web search. ACM Transactions on Information Systems, 36(2):21:1– 21:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Bansler:1993:RSA

[BB93] Jørgen P. Bansler and Keld Bødker. A reappraisal of structured analysis: Design in an organizational context. ACM Transactions on Information Systems, 11(2):165– 193, April 1993. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80.

Blanco:2010:PSP

[BB10] Roi Blanco and Alvaro Barreiro. Probabilistic static pruning of inverted files. ACM Transactions on Information Systems, 28(1):1:1–1:??, January 2010. CODEN ATISET. ISSN 1046-8188.

Balog:2011:QME

[BBD11] Krisztian Balog, Marc Bron, and Maarten De Rijke. Query modeling for entity search based on terms, categories, and examples. ACM Transactions on Information Systems, 29(4):22:1–22:??, December 2011. CODEN ATISET. ISSN 1046-8188.

Bergman:2008:ISE

[BBMN^{+}08] Ofer Bergman, Ruth Beyth-Marom, Rafi Nachmias, Noa Gradovitch, and Steve Whittaker. Improved search engines and navigation preference in personal information management. ACM Transactions on Information Systems, 26(4):20:1–20:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Blake:1992:SOE

[BBT92] G. Elizabeth Blake, Tim Bray, and Frank Wm. Tompa. Shortening the OED: Experience with a grammar-defined database. ACM Transactions on Information Systems, 10 (3):213–232, July 1992. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Bast:2011:FCH

[BC11] Hannah Bast and Marjan Celikik. Fast construction of the HYB index. ACM Transactions on Information Systems, 29(3):16:1–16:??, July 2011. CODEN ATISET. ISSN 1046-8188.

Bast:2013:EFS

[BC13] Hannah Bast and Marjan Celikik. Efficient fuzzy search in large text collections. ACM Transactions on Information Systems, 31(2):10:1–10:??, May 2013. CODEN ATISET. ISSN 1046-8188.

Bast:2014:EIB

[BC14] Hannah Bast and Marjan Celikik. Efficient index-based snippet generation. ACM Transactions on Information Systems, 32(2):6:1–6:??, April 2014. CODEN ATISET. ISSN 1046-8188.

Barreau:2008:IKR

[BCD^{+}08] Deborah Barreau, Robert Capra, Susan Dumais, William Jones, and Manuel Pérez-Quiñones. Introduction to keeping, refinding and sharing personal information. ACM Transactions on Information Systems, 26(4):18:1–18:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Baralis:2015:MSM

[BCFG15] Elena Baralis, Luca Cagliero, Alessandro Fiori, and Paolo Garza. MWI-Sum: a multilingual summarizer based on frequent weighted itemsets. ACM Transactions on Information Systems, 34(1):5:1– 5:??, October 2015. CODEN ATISET. ISSN 1046-8188.

Brisaboa:2014:XEX

[BCPN14] Nieves R. Brisaboa, Ana Cerdeira-Pena, and Gonzalo Navarro. XXS: Efficient XPath evaluation on compressed XML documents. ACM Transactions on Information Systems, 32(3):13:1– 13:??, June 2014. CODEN ATISET. ISSN 1046-8188.

Brafman:2004:QDM

[BDS04] Ronen I. Brafman, Carmel Domshlak, and Solomon E. Shimony. Qualitative decision making in adaptive presentation of structured information. ACM Transactions on Information Systems, 22(4): 503–539, October 2004. CO-DEN ATISET. ISSN 1046-8188.

Belussi:1998:SSJ

[BF98] Alberto Belussi and Christos Faloutsos. Self-spatial join selectivity estimation using fractal concepts. ACM Transactions on Information Systems, 16(2):161–201, April 1998. CODEN ATISET. ISSN 1046-8188.

Bertino:2003:HAC

[BFF^{+}03] Elisa Bertino, Jianping Fan, Elena Ferrari, Mohand-Said Hacid, Ahmed K. Elmagarmid, and Xingquan Zhu. A hierarchical access control model for video database systems. ACM Transactions on Information Systems, 21 (2):155–191, 2003. CODEN ATISET. ISSN 1046-8188.

Brisaboa:2010:DLT

[BFNP10] Nieves Brisaboa, Antonio Fariña, Gonzalo Navarro, and José Paramá. Dynamic lightweight text compression. ACM Transactions on Information Systems, 28(3):10:1– 10:??, June 2010. CODEN ATISET. ISSN 1046-8188.

Bruch:2024:AFF

[BGI24] Sebastian Bruch, Siyu Gai, and Amir Ingber. An analysis of fusion functions for hybrid retrieval. ACM Transactions on Information Systems, 42 (1):20:1–20:??, January 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3596512.

Baralis:2007:AXQ

[BGQT07] Elena Baralis, Paolo Garza, Elisa Quintarelli, and Letizia Tanca. Answering XML queries by means of data summaries. ACM Transactions on Information Systems, 25(3): 10:1–10:??, July 2007. CO-DEN ATISET. ISSN 1046-8188.

Baly:2016:MFM

[BHH^{+}16] Ramy Baly, Roula Hobeica, Hazem Hajj, Wassim El-Hajj, Khaled Bashir Shaban, and Ahmad Al-Sallab. A metaframework for modeling the human reading process in sentiment analysis. ACM Transactions on Information Systems, 35(1):7:1–7:??, October 2016. CODEN ATISET. ISSN 1046-8188.

Bailey:2007:AHT

[BHMW07] Christopher Bailey, Wendy Hall, David E. Millard, and Mark J. Weal. Adaptive hypermedia through contextualized open hypermedia structures. ACM Transactions on Information Systems, 25 (4):16:1–16:??, October 2007. CODEN ATISET. ISSN 1046-8188.

Bai:2025:EMT

[BHY^{+}25] Ting Bai, Le Huang, Yue Yu, Cheng Yang, Cheng Hou, Zhe Zhao, and Chuan Shi. Efficient multi-task prompt tuning for recommendation. ACM Transactions on Information Systems, 43(4):104:1– 104:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Bier:1992:ESB

[Bie92] Eric A. Bier. EmbeddedButtons: Supporting buttons in documents. ACM Transactions on Information Systems, 10(4):381–407, October 1992. CODEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Beitzel:2007:ACW

[BJL^{+}07] Steven M. Beitzel, Eric C.

Jensen, David D. Lewis, Abdur Chowdhury, and Ophir Frieder. Automatic classification of Web queries using very large unlabeled query logs. ACM Transactions on Information Systems, 25(2): 9:1–9:??, April 2007. CODEN ATISET. ISSN 1046-8188.

Bertino:1999:FAM

[BJS99] Elisa Bertino, Sushil Jajodia, and Pierangela Samarati. A flexible authorization mechanism for relational data management systems. ACM Transactions on Information Systems, 17(2):101–140, April 1999. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org/pubs/citations/
journals/tois/1999-17-2/
p101-bertino/.

Bookstein:1990:CIT

[BK90] Abraham Bookstein and Shmuel T. Klein. Compression, information theory, and grammars: a unified approach. ACM Transactions on Information Systems, 8(1): 27–49, January 1990. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Buyukkokten:2002:EWB

[BKGM^{+}02] Orkut Buyukkokten, Oliver Kaljuvee, Hector Garcia-Molina, Andreas Paepcke, and Terry Winograd. Efficient Web browsing on handheld devices using page and form summarization. ACM Transactions on Information Systems, 20(1):82–115, January 2002. CODEN ATISET. ISSN 1046-8188.

Bookstein:1997:MWO

[BKR97] A. Bookstein, S. T. Klein, and T. Raita. Modeling word occurrences for the compression of concordances. ACM Transactions on Information Systems, 15(3):254– 290, July 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/bookstein.
html.

Bennett:2015:OSI

[BKWZ15] Paul N. Bennett, Diane Kelly, Ryen W. White, and Yi Zhang. Overview of the special issue on contextual search and recommendation. ACM Transactions on Information Systems, 33(1):1:1– 1:??, March 2015. CODEN ATISET. ISSN 1046-8188.

Bhoi:2022:PMR

[BLH^{+}22] Suman Bhoi, Mong Li Lee, Wynne Hsu, Hao Sen Andrew Fang, and Ngiap Chuan Tan. Personalizing medication recommendation with a graphbased approach. ACM Transactions on Information Systems, 40(3):55:1–55:23, July 2022. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3488668.

Bruch:2024:SSE

[BLMN24] Sebastian Bruch, Claudio Lucchese, Maria Maistro, and Franco Maria Nardini. Special section on efficiency in neural information retrieval. ACM Transactions on Information Systems, 42(5):113:1– 113:??, September 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3641203.

Bing:2015:WQR

[BLWJ15] Lidong Bing, Wai Lam, Tak-Lam Wong, and Shoaib Jameel. Web query reformulation via joint modeling of latent topic dependency and term context. ACM Transactions on Information Systems, 33(2):6:1–6:??, February 2015. CODEN ATISET. ISSN 1046-8188.

Bharat:2002:WEA

[BM02] Krishna Bharat and George A. Mihaila. When experts agree: using non-affiliated experts to rank popular topics. ACM Transactions on Information Systems, 20(1):47–58, January 2002. CODEN ATISET. ISSN 1046-8188.

Bhatia:2012:SFT

[BM12] Sumit Bhatia and Prasenjit Mitra. Summarizing figures, tables, and algorithms in scientific publications to augment search results. ACM Transactions on Information Systems, 30(1):3:1–3:??, February 2012. CODEN ATISET. ISSN 1046-8188.

Benham:2019:BSP

[BMMC19] Rodger Benham, Joel Mackenzie, Alistair Moffat, and J. Shane Culpepper. Boosting search performance using query variations. ACM Transactions on Information Systems, 37(4):41:1– 41:??, December 2019. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3345001.

Bruch:2024:AAM

[BNIL24a] Sebastian Bruch, Franco Maria Nardini, Amir Ingber, and Edo Liberty. An approximate algorithm for maximum inner product search over streaming sparse vectors. ACM Transactions on Information Systems, 42(2):42:1–42:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3609797.

Bruch:2024:BDS

[BNIL24b] Sebastian Bruch, Franco Maria Nardini, Amir Ingber, and Edo Liberty. Bridging dense and sparse maximum inner product search. ACM Transactions on Information Systems, 42(6):151:1–151:??, November 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3665324.

Bodoff:2004:RMH

[Bod04]  David Bodoff.  Relevance

models to help estimate document and query parameters. ACM Transactions on Information Systems, 22(3):357– 380, July 2004. CODEN ATISET. ISSN 1046-8188.

Bevilacqua:2025:WAA

[BOQ^{+}25] Marialena Bevilacqua, Kezia Oketch, Ruiyang Qin, Will Stamey, Xinyuan Zhang, Yi Gan, Kai Yang, and Ahmed Abbasi. When automated assessment meets automated content generation:

Examining text quality in the era of GPTs. ACM Transactions on Information Systems, 43(2):43:1–43:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3702639.

Berghel:1996:EUE

[BR96] Hal Berghel and David Roach. An extension of Ukkonen’s enhanced dynamic programming ASM algorithm. ACM Transactions on Information Systems, 14(1):94–106, January 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/berghel.
html.

Botafogo:1992:SAH

[BRS92] Rodrigo A. Botafogo, Ehud Rivlin, and Ben Shneiderman. Structural analysis of hypertexts: Identifying hierarchies and useful metrics. ACM Transactions on Information Systems, 10(2):142–180, April 1992. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Broschart:2012:HPP

[BS12] Andreas Broschart and Ralf Schenkel. High-performance processing of text queries with tunable pruned term and term pair indexes. ACM Transactions on Information Systems, 30(1):5:1–5:??, February 2012. CODEN ATISET. ISSN 1046-8188.

Bolchini:2003:LPD

[BSST03] Cristiana Bolchini, Fabio Salice, Fabio A. Schreiber, and Letizia Tanca. Logical and physical design issues for smart card databases. ACM Transactions on Information Systems, 21(3):254–285, July 2003. CODEN ATISET. ISSN 1046-8188.

Boldi:2009:PFD

[BSV09] Paolo Boldi, Massimo Santini, and Sebastiano Vigna. PageRank: Functional dependencies. ACM Transactions on Information Systems, 27(4):19:1–19:??, November 2009. CODEN ATISET. ISSN 1046-8188.

Bassani:2024:PQE

[BTP24] Elias Bassani, Nicola Tonellotto, and Gabriella Pasi. Personalized query expansion with contextual word embeddings. ACM Transactions on Information Systems, 42(2): 61:1–61:??, March 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3624988.

Bulterman:1995:EVH

[Bul95] Dick C. A. Bulterman. Embedded video in hypermedia documents: Supporting integration and adaptive control. ACM Transactions on Information Systems, 13(4): 440–470, October 1995. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Video Information Retrieval.

Bron:2025:UCE

[BvdHFS25] Michiel Bron, Peter G. M. van der Heijden, Ad Feelders, and Arno Siebes. Using Chao’s estimator as a stopping criterion for technologyassisted review. ACM Transactions on Information Systems, 43(3):81:1–81:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Bernstein:2008:ISH

[BVKS08] Michael Bernstein, Max Van Kleek, David Karger, and M. C. Schraefel. Information scraps: How and why information eludes our personal information management tools. ACM Transactions on Information Systems, 26(4):24:1– 24:??, September 2008. CO-DEN ATISET. ISSN 1046-8188.

Bradesko:2017:CCM

[BWS^{+}17] Luka Bradesko, Michael Witbrock, Janez Starc, Zala Herga, Marko Grobelnik, and Dunja Mladenić. Curious catmobile, context-aware conversational crowdsourcing knowledge acquisition. ACM Transactions on Information Systems, 35(4):33:1–33:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Baeza-Yates:2006:ISI

[BYFM06] Ricardo Baeza-Yates, Norbert Fuhr, and Yoelle Maarek. Introduction to the special issue on XML retrieval. ACM Transactions on Information Systems, 24(4):405–406, October 2006. CODEN ATISET. ISSN 1046-8188.

Bahrainian:2022:CCA

[BZCE22] Seyed Ali Bahrainian, George Zerveas, Fabio Crestani, and Carsten Eickhoff. CATS: Customizable abstractive topicbased summarization. ACM Transactions on Information Systems, 40(1):5:1–5:24, January 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3464299.

Bai:2024:IOD

[BZDW24] Yutong Bai, Yujia Zhou, Zhicheng Dou, and Ji-Rong Wen. Intent-oriented dynamic interest modeling for personalized Web search. ACM Transactions on Information Systems, 42(4):96:1– 96:??, July 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3639817.

Choi:2023:IKR

[CACW23] Bogeum Choi, Jaime Arguello, Robert Capra, and Austin R. Ward. The influences of a knowledge representation tool on searchers with varying cognitive abilities. ACM Transactions on Information Systems, 41 (1):18:1–18:??, January 2023.

CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3527661.

Callan:2012:TRO

[Cal12] Jamie Callan. TOIS reviewers: October 2009 to September 2012. ACM Transactions on Information Systems, 30(4):27:1–27:??, November 2012. CODEN ATISET. ISSN 1046-8188.

Can:1993:ICD

[Can93] Fazli Can. Incremental clustering for dynamic information processing. ACM Transactions on Information Systems, 11(2):143–164, April 1993. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

Carterette:2012:MTS

[Car12] Benjamin A. Carterette. Multiple testing in statistical analysis of systems-based information retrieval experiments. ACM Transactions on Information Systems, 30(1): 4:1–4:??, February 2012. CO-DEN ATISET. ISSN 1046-8188.

Chen:2011:RFA

[CBZ11] Keke Chen, Jing Bai, and Zhaohui Zheng. Ranking function adaptation with boosting trees. ACM Transactions on Information Systems, 29(4):18:1–18:??, December 2011. CODEN ATISET. ISSN 1046-8188.

Clarke:2000:SSR

[CC00] Charles L. A. Clarke and Gordon V. Cormack. Shortestsubstring retrieval and ranking. ACM Transactions on Information Systems, 18(1): 44–78, January 2000. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.
acm.org/pubs/citations/
journals/tois/2000-18-1/
p44-clarke/.

Callan:2001:QBS

[CC01] Jamie Callan and Margaret Connell. Query-based sampling of text databases. ACM Transactions on Information Systems, 19(2):97–130, April 2001. CODEN ATISET. ISSN 1046-8188.

Conrad:2003:EUS

[CC03] Jack G. Conrad and Joanne R. S. Claussen. Early user–system interaction for database selection in massive domain-specific online environments. ACM Transactions on Information Systems, 21 (1):94–131, 2003. CODEN ATISET. ISSN 1046-8188.

Chuang:2005:TGT

[CC05] Shui-Lung Chuang and Lee-Feng Chien. Taxonomy generation for text segments: a practical Web-based approach. ACM Transactions on Information Systems, 23(4): 363–396, October 2005. CO-DEN ATISET. ISSN 1046-

Choi:2024:BUP

[CCAC24] Bogeum Choi, Sarah Casteel, Jaime Arguello, and Robert Capra. Better understanding procedural search tasks: Perceptions, behaviors, and challenges. ACM Transactions on Information Systems, 42(3): 65:1–65:??, May 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3630004.

Chen:2009:ATF

[CCC09] Chien Chin Chen, Meng Chang Chen, and Ming-Syan Chen. An adaptive threshold framework for event detection using HMM-based life profiles. ACM Transactions on Information Systems, 27(2):9:1– 9:??, February 2009. CODEN ATISET. ISSN 1046-8188.

Cao:2012:AEC

[CCC^{+}12] Xin Cao, Gao Cong, Bin Cui, Christian S. Jensen, and Quan Yuan. Approaches to exploring category information for question retrieval in community question-answer archives. ACM Transactions on Information Systems, 30(2):7:1– 7:??, May 2012. CODEN ATISET. ISSN 1046-8188.

Chen:2019:JNC

[CCCD19] Wanyu Chen, Fei Cai, Honghui Chen, and Maarten De Rijke. Joint neural collaborative filtering for recommender systems. ACM Transactions on Information Systems, 37(4):

39:1–39:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3343117.

Connor:2017:HEI

[CCVR17] Richard Connor, Franco Alberto Cardillo, Lucia Vadicamo, and Fausto Rabitti. Hilbert exclusion: Improved metric search through finite isometric embeddings. ACM Transactions on Information Systems, 35(3):17:1–17:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Cai:2025:DRS

[CCXZ25] Qiqi Cai, Jian Cao, Guandong Xu, and Nengjun Zhu. Distributed recommendation systems: Survey and research directions. ACM Transactions on Information Systems, 43 (1):10:1–10:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3694783.

Cheng:2019:MER

[CCZ^{+}19] Zhiyong Cheng, Xiaojun Chang, Lei Zhu, Rose C. Kanjirathinkal, and Mohan Kankanhalli. MMALFM: Explainable recommendation by leveraging reviews and images. ACM Transactions on Information Systems, 37 (2):16:1–16:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3291060.

Cui:2023:FGI

[CCZ^{+}23] Jiajun Cui, Zeyuan Chen, Aimin Zhou, Jianyong Wang, and Wei Zhang. Finegrained interaction modeling with multi-relational transformer for knowledge tracing. ACM Transactions on Information Systems, 41(4):104:1– 104:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3580595.

Comai:2001:CGQ

[CDF01] Sara Comai, Ernesto Damiani, and Piero Fraternali. Computing graphical queries over XML data. ACM Transactions on Information Systems, 19(4):371–430, October 2001. CODEN ATISET. ISSN 1046-8188.

Candan:2009:SSE

[CDH^{+}09] K. Selçuk Candan, Mehmet E. Dönderler, Terri Hedgpeth, Jong Wook Kim, Qing Li, and Maria Luisa Sapino. SEA: Segment-enrich-annotate paradigm for adapting dialog-based content for improved accessibility. ACM Transactions on Information Systems, 27(3): 15:1–15:??, May 2009. CO-DEN ATISET. ISSN 1046-8188.

Costa:2015:DRM

[CDM15] Alberto Costa, Emanuele Di Buccio, and Massimo Melucci. A document retrieval model based on digital signal filtering. ACM Transactions on Information Systems, 34(1):6:1– 6:??, October 2015. CODEN ATISET. ISSN 1046-8188.

Carpineto:2001:ITA

[CdMRB01] Claudio Carpineto, Renato de Mori, Giovanni Romano, and Brigitte Bigi. An information-theoretic approach to automatic query expansion. ACM Transactions on Information Systems, 19(1):1–27, 2001. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org/
pubs/articles/journals/
tois/2001-19-1/p1-carpineto/
p1-carpineto.pdf;   http:
//www.acm.org/pubs/citations/
journals/tois/2001-19-1/
p1-carpineto/.

Clements:2010:TDE

[CDR10] Maarten Clements, Arjen P.

De Vries, and Marcel J. T. Reinders. The task-dependent effect of tags and ratings on social media access. ACM Transactions on Information Systems, 28(4):21:1–21:??, November 2010. CODEN ATISET. ISSN 1046-8188.

Corno:2021:UII

[CDR21] Fulvio Corno, Luigi De Russis, and Alberto Monge Roffarello. From users’ intentions to IF–THEN rules in the Internet of Things. ACM Transactions on Information Systems, 39(4):53:1– 53:33, October 2021. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3447264.

Chen:2023:BDR

[CDW^{+}23] Jiawei Chen, Hande Dong, Xiang Wang, Fuli Feng, Meng Wang, and Xiangnan He. Bias and debias in recommender system: a survey and future directions. ACM Transactions on Information Systems, 41(3):67:1–67:??, July 2023. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3564284.

Cohen:2008:RTD

[CDZ08] Sara Cohen, Carmel Domshlak, and Naama Zwerdling. On ranking techniques for desktop search. ACM Transactions on Information Systems, 26(2):11:1–11:??, March 2008. CODEN ATISET. ISSN 1046-8188.

Chen:2023:IRI

[CDZ^{+}23] Haonan Chen, Zhicheng Dou, Qiannan Zhu, Xiaochen Zuo, and Ji-Rong Wen. Integrating representation and interaction for context-aware document ranking. ACM Transactions on Information Systems, 41 (1):21:1–21:??, January 2023. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3529955.

Campagnoni:1989:IRU

[CE89] F. R. Campagnoni and Kate Ehrlich. Information retrieval using a hypertext-based help system. ACM Transactions on Information Systems, 7(3): 271–291, July 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Cornolti:2019:SPA

[CFC^{+}19] Marco Cornolti, Paolo Ferragina, Massimiliano Ciaramita, Stefan Rüd, and Hinrich Schütze. SMAPH: a piggyback approach for entity-linking in Web queries. ACM Transactions on Information Systems, 37(1):13:1– 13:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Cunha:2025:NOR

[CFE^{+}25] Washington Cunha, Alejandro Moreo Fernández, Andrea Esuli, Fabrizio Sebastiani, Leonardo Rocha, and Marcos André Gonçalves. A noiseoriented and redundancyaware instance selection framework. ACM Transactions on Information Systems, 43(2): 45:1–45:??, March 2025. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3705000.

Culpepper:2022:TDC

[CFFK22] J. Shane Culpepper, Guglielmo Faggioli, Nicola Ferro, and Oren Kurland. Topic difficulty: Collection and query formulation effects. ACM Transactions on Information Systems, 40(1):19:1–19:36, January 2022. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3470563.

Chowdhury:2002:CSF

[CFGM02] Abdur Chowdhury, Ophir Frieder, David Grossman, and Mary Catherine Mc-Cabe. Collection statistics for fast duplicate document detection. ACM Transactions on Information Systems, 20 (2):171–191, 2002. CODEN ATISET. ISSN 1046-8188.

Celentano:1995:KBD

[CFP95] Augusto Celentano, Maria Grazia Fugini, and Silvano Pozzi. Knowledge-based document retrieval in office environments: The Kabiria system. ACM Transactions on Information Systems, 13(3):237– 268, July 1995. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80.

Cui:2015:KGF

[CGB^{+}15] Qing Cui, Bin Gao, Jiang Bian, Siyu Qiu, Hanjun Dai, and Tie-Yan Liu. KNET: a general framework for learning word embedding using morphological knowledge. ACM Transactions on Information Systems, 34(1):4:1– 4:??, October 2015. CODEN ATISET. ISSN 1046-8188.

Cagliero:2019:EMD

[CGB19] Luca Cagliero, Paolo Garza, and Elena Baralis. ELSA: a multilingual document summarization algorithm based on frequent itemsets and latent semantic analysis. ACM Transactions on Information Systems, 37(2):21:1– 21:??, March 2019. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.
acm.org/ft_gateway.cfm?
id=3298987.

Chen:2025:EIB

[CGD^{+}25] Lei Chen, Chen Gao, Xiaoyi Du, Hengliang Luo, Depeng Jin, Yong Li, and Meng Wang. Enhancing ID-based recommendation with large language models. ACM Transactions on Information Systems, 43(5):112:1–112:30, September 2025. CODEN ATISET. ISSN 1046-8188.

Cooper:2002:PPD

[CGM02] Brian F. Cooper and Hector Garcia-Molina. Peer-to-peer data trading to preserve information. ACM Transactions on Information Systems, 20 (2):133–170, 2002. CODEN ATISET. ISSN 1046-8188.

Cooper:2005:AHS

[CGM05] Brian F. Cooper and Hector Garcia-Molina. Ad hoc, self-supervising peer-to-peer search networks. ACM Transactions on Information Systems, 23(2):169–200, April 2005. CODEN ATISET. ISSN 1046-8188.

Chang:1999:PRT

[CGMP99] Chen-Chuan K. Chang, Héctor Garcia-Molina, and Andreas Paepcke. Predicate rewriting for translating Boolean queries in a heterogeneous information system. ACM Transactions on Information Systems, 17(1):1–39, January 1999. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org/pubs/citations/
journals/tois/1999-17-1/
p1-chang/.

Cheung:1996:MAG

[CH96] Waiman Cheung and Cheng Hsu. The model-assisted global query system for multiple databases in distributed enterprises. ACM Transactions on Information Systems, 14(4):421–470, October 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/cheung.
html.

Cole:2015:UAP

[CHBS15] Michael J. Cole, Chathra Hendahewa, Nicholas J. Belkin, and Chirag Shah. User activity patterns during information search. ACM Transactions on Information Systems, 33(1):1:1–1:??, March 2015. CODEN ATISET. ISSN 1046-8188.

Cao:2017:CPA

[CHN^{+}17] Da Cao, Xiangnan He, Liqiang Nie, Xiaochi Wei, Xia Hu, Shunxiang Wu, and Tat-Seng Chua. Crossplatform app recommendation by jointly modeling ratings and texts. ACM Transactions on Information Systems, 35(4):37:1–37:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Chen:1999:IGL

[CHS99] Hao Chen, Jianying Hu, and Richard W. Sproat. Integrating geometrical and linguistic analysis for email signature block parsing. ACM Transactions on Information Systems, 17(4):343–366, October 1999. CODEN ATISET. ISSN 1046-8188. URL

http://www.acm.org/pubs/
articles/journals/tois/
1999-17-4/p343-chen/p343-
chen.pdf;   http://www.
acm.org/pubs/citations/
journals/tois/1999-17-4/
p343-chen/.

Cheng:2024:CPH

[CHZ24] Jiezhu Cheng, Kaizhu Huang, and Zibin Zheng. Can perturbations help reduce investment risks? Risk-aware stock recommendation via split variational adversarial training. ACM Transactions on Information Systems, 42 (4):101:1–101:??, July 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3643131.

Cheng:2020:RUC

[CJN20] Miaomiao Cheng, Liping Jing, and Michael K. Ng. Robust unsupervised crossmodal hashing for multimedia retrieval. ACM Transactions on Information Systems, 38(3):30:1–30:25, June 2020. CODEN ATISET. ISSN 1046-8188. URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3389547.

Chapelle:2012:LSV

[CJRY12] Olivier Chapelle, Thorsten Joachims, Filip Radlinski, and Yisong Yue. Large-scale validation and analysis of interleaved search evaluation. ACM Transactions on Information Systems, 30(1):6:1– 6:??, February 2012. CODEN ATISET. ISSN 1046-8188.

Chen:2021:CEC

[CJW^{+}21] Jiawei Chen, Chengquan Jiang, Can Wang, Sheng Zhou, Yan Feng, Chun Chen, Martin Ester, and Xiangnan He. CoSam: an efficient collaborative adaptive sampler for recommendation. ACM Transactions on Information Systems, 39(3):34:1– 34:24, July 2021. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3450289.

Cheng:2023:CSH

[CJY23a] Lizhi Cheng, Weijia Jia, and Wenmian Yang. Capture salient historical information: a fast and accurate non-autoregressive model for multi-turn spoken language understanding. ACM Transactions on Information Systems, 41(2):41:1–41:??, April 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3545800.

Cheng:2023:CDM

[CJY^{+}23b] Zifeng Cheng, Zhiwei Jiang, Yafeng Yin, Cong Wang, Shiping Ge, and Qing Gu. A consistent dual-MRC framework for emotion-cause pair extraction. ACM Transactions on Information Systems, 41(4):105:1–105:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3558548.

Chen:2016:BRU

[CJZ^{+}16] Jia Chen, Qin Jin, Shiwan Zhao, Shenghua Bao, Li Zhang, Zhong Su, and Yong Yu. Boosting recommendation in unexplored categories by user price preference. ACM Transactions on Information Systems, 35(2): 12:1–12:??, December 2016. CODEN ATISET. ISSN 1046-

Cui:2007:SPM

[CKC07] Hang Cui, Min-Yen Kan, and Tat-Seng Chua. Soft pattern matching models for definitional question answering. ACM Transactions on Information Systems, 25(2):8:1– 8:??, April 2007. CODEN ATISET. ISSN 1046-8188.

Chang:2003:MME

[CL03] Edward Chang and Beitao Li. MEGA—the maximizing expected generalization algorithm for learning complex query concepts. ACM Transactions on Information Systems, 21(4):347–382, October 2003. CODEN ATISET. ISSN 1046-8188.

Chen:2006:IIV

[CL06] Keke Chen and Ling Liu. iVI-BRATE: Interactive visualizationbased framework for clustering large datasets. ACM Transactions on Information Systems, 24(2):245–294, April 2006. CODEN ATISET. ISSN 1046-8188.

Cormack:2007:OSS

[CL07] Gordon V. Cormack and Thomas R. Lynam. Online supervised spam filter evaluation. ACM Transactions on Information Systems, 25(3): 11:1–11:??, July 2007. CO-DEN ATISET. ISSN 1046-8188.

Chong:2018:EUV

[CL18] Wen-Haw Chong and Ee-Peng Lim. Exploiting user and venue characteristics for finegrained tweet geolocation. ACM Transactions on Information Systems, 36(3):26:1– 26:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Chong:2019:FGG

[CL19] Wen-Haw Chong and Ee-Peng Lim. Fine-grained geolocation of tweets in temporal proximity. ACM Transactions on Information Systems, 37(2):17:1–17:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3291059.

Chen:2025:DES

[CLF^{+}25] Nuo Chen, Jiqun Liu, Hanpei Fang, Yuankai Luo, Tetsuya Sakai, and Xiao-Ming Wu. Decoy effect in search interaction: Understanding user behavior and measuring system vulnerability. ACM Transactions on Information Systems, 43(2):53:1–53:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3708884.

Chen:2023:FTG

[CLG^{+}23] Xiuying Chen, Mingzhe Li, Shen Gao, Zhangming Chan, Dongyan Zhao, Xin Gao, Xiangliang Zhang, and Rui Yan. Follow the timeline! Generating an abstractive and extractive timeline summary in chronological order. ACM Transactions on Information Systems, 41(1):9:1– 9:??, January 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3517221.

Che:2024:TIE

[CLL24] Shangkun Che, Hongyan Liu, and Shen Liu. Tagging items with emerging tags: a neural topic model based few-shot learning approach. ACM Transactions on Information Systems, 42(4):102:1– 102:??, July 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3641859.

Cheng:2022:FLA

[CLM^{+}22] Zhiyong Cheng, Fan Liu, Shenghan Mei, Yangyang Guo, Lei Zhu, and Liqiang Nie. Feature-level attentive ICF for recommendation. ACM Transactions on Information Systems, 40(4):75:1– 75:24, October 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3490477.

Chen:2020:SHO

[CLPM20] Xiancong Chen, Lin Li, Weike Pan, and Zhong Ming. A survey on heterogeneous oneclass collaborative filtering. ACM Transactions on Information Systems, 38(4):35:1– 35:54, October 2020. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3402521.

Cui:2014:SSI

[CLZ^{+}14] Peng Cui, Shao-Wei Liu, Wen-Wu Zhu, Huan-Bo Luan, Tat-Seng Chua, and Shi-Qiang Yang. Social-sensed image search. ACM Transactions on Information Systems, 32(2):8:1–8:??, April 2014. CODEN ATISET. ISSN 1046-8188.

Culpepper:2010:ESI

[CM10] J. Shane Culpepper and Alistair Moffat. Efficient set intersection for inverted indexing. ACM Transactions on Information Systems, 29(1):1:1– 1:??, December 2010. CO-DEN ATISET. ISSN 1046-8188.

Cahoon:2000:EPD

[CML00] Brendon Cahoon, Kathryn S.

McKinley, and Zhihong Lu. Evaluating the performance of distributed architectures for information retrieval using a variety of workloads. ACM Transactions on Information Systems, 18(1):1–43, January 2000. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org/pubs/citations/
journals/tois/2000-18-1/
p1-cahoon/.

Chen:2021:HFS

[CML^{+}21] Jia Chen, Jiaxin Mao, Yiqun Liu, Ziyi Ye, Weizhi Ma, Chao Wang, Min Zhang, and Shaoping Ma. A hybrid framework for session context modeling. ACM Transactions on Information Systems, 39(3):30:1–30:35, July 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3448127.

Card:1991:MAD

[CMR91] Stuart K. Card, Jock D. Mackinlay, and George G. Robertson. A morphological analysis of the design space of input devices. ACM Transactions on Information Systems, 9(2):99–122, April 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.

acm.org:80. Special Issue on Computer-Human Interaction.

Chen:2023:RNS

[CMZ^{+}23] Chong Chen, Weizhi Ma, Min Zhang, Chenyang Wang, Yiqun Liu, and Shaoping Ma. Revisiting negative sampling vs. non-sampling in implicit recommendation. ACM Transactions on Information Systems, 41(1):12:1– 12:??, January 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3522672.

Cohen:1997:RHF

[Coh97] Jonathan D. Cohen. Recursive hashing functions for ngrams. ACM Transactions on Information Systems, 15(3): 291–320, July 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:

80/tois/abstracts/cohen.
html.

Cohen:2000:DIU

[Coh00] William W. Cohen. Data integration using similarity joins and a word-based information representation language. ACM Transactions on Information Systems, 18 (3):288–321, 2000. URL

http://www.acm.org/pubs/
citations/journals/tois/
2000-18-3/p288-cohen/.

Cooper:1995:SIM

[Coo95] William S. Cooper. Some inconsistencies and misidentified modeling assumptions in probabilistic information retrieval. ACM Transactions on Information Systems, 13(1): 100–111, January 1995. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Costa:2013:XCA

[COR13] Gianni Costa, Riccardo Ortale, and Ettore Ritacco. Xclass: Associative classification of XML documents by structure. ACM Transactions on Information Systems, 31(1):3:1–3:??, January 2013. CODEN ATISET. ISSN 1046-8188.

Chelmis:2013:SLP

[CP13] Charalampos Chelmis and Viktor K. Prasanna. Social link prediction in online social tagging systems.

ACM Transactions on Information Systems, 31(4):20:1– 20:??, November 2013. CO-DEN ATISET. ISSN 1046-8188.

Cummins:2015:PUD

[CPL15] Ronan Cummins, Jiaul H. Paik, and Yuanhua Lv. A Pólya urn document language model for improved information retrieval. ACM Transactions on Information Systems, 33(4):21:1–21:??, May 2015. CODEN ATISET. ISSN 1046-8188.

Cai:2023:UCS

[CQF^{+}23] Desheng Cai, Shengsheng Qian, Quan Fang, Jun Hu, and Changsheng Xu. User cold-start recommendation via inductive heterogeneous graph neural network. ACM Transactions on Information Systems, 41(3):64:1– 64:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3560487.

Chen:2024:TRD

[CQL^{+}24] Hai Chen, Fulan Qian, Chang Liu, Yanping Zhang, Hang Su, and Shu Zhao. Training robust deep collaborative filtering models via adversarial noise propagation. ACM Transactions on Information Systems, 42(1):9:1– 9:??, January 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3589000.

Chen:2025:LBR

[CQZ^{+}25a] Yankai Chen, Yue Que, Xinni Zhang, Chen Ma, and Irwin King. Learning binarized representations with pseudopositive sample enhancement for efficient graph collaborative filtering. ACM Transactions on Information Systems, 43(5):131:1–131:28, September 2025. CODEN ATISET. ISSN 1046-8188.

Cui:2025:RDR

[CQZ^{+}25b] Jiajun Cui, Hong Qian, Chanjin Zheng, Lu Wang, Mo Yu, and Wei Zhang. Rebalancing discriminative responses for knowledge tracing. ACM Transactions on Information Systems, 43(3):75:1–75:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Carroll:1992:GAT

[CR92] John M. Carroll and Mary Beth Rosson. Getting around the task-artifact cycle: How to make claims and design by scenario. ACM Transactions on Information Systems, 10 (2):181–212, April 1992. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Chua:1995:VRS

[CR95] Tat-Seng Chua and Li-Qun Ruan. A video retrieval and sequencing system. ACM Transactions on Information Systems, 13(4):373–407, October 1995. CODEN ATISET.

ISSN 1046-8188. URL http:/

                                        /www.acm.org:80. Special Is-

sue on Video Information Retrieval.

Chen:2022:MID

[CRC^{+}22] Wanyu Chen, Pengjie Ren, Fei Cai, Fei Sun, and Maarten De Rijke. Multi-interest diversification for end-to-end sequential recommendation. ACM Transactions on Information Systems, 40(1):20:1– 20:30, January 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3475768.

Cai:2016:DQA

[CRD16] Fei Cai, Ridho Reinanda, and Maarten De Rijke. Diversifying query auto-completion. ACM Transactions on Information Systems, 34(4):25:1– 25:??, September 2016. CO-DEN ATISET. ISSN 1046-8188.

Carpineto:2002:IRF

[CRG02] Claudio Carpineto, Giovanni Romano, and Vittorio Giannini. Improving retrieval feedback with multiple termranking function combination. ACM Transactions on Information Systems, 20(3):259– 290, 2002. CODEN ATISET. ISSN 1046-8188.

Calado:2003:LVG

[CRNZ^{+}03] Pável Calado, Berthier Ribeiro-Neto, Nivio Ziviani, Edleno Moura, and Ilmério Silva. Local versus global link information in the Web. ACM Transactions on Information Systems, 21(1):42–63, 2003. CO-DEN ATISET. ISSN 1046-8188.

Croft:1989:EIS

[Cro89a] W. B. Croft. Editorial: Introduction to the special issue. ACM Transactions on Information Systems, 7(3): 181–182, July 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Croft:1989:SCR

[Cro89b] W. Bruce Croft, editor. SI-GIR Conference on Research and Development in Information Retrieval, volume 7(3) of ACM Transactions on Information Systems. ACM Press, New York, NY 10036, USA, July 1989. CODEN ATISET. ISSN 1046-8188.

Croft:1991:E

[Cro91]  W. Bruce Croft.  Edito-

rial. ACM Transactions on Information Systems, 9(3): 185–186, July 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Croft:1998:AI

[Cro98] W. Bruce Croft. Author index. ACM Transactions on Information Systems, 16(4):

413–414, October 1998. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80/pubs/citations/journals/
                                        tois/1998-16-4/p413-croft/

.

Crouch:2006:DER

[Cro06] Carolyn J. Crouch. Dynamic element retrieval in a structured environment. ACM Transactions on Information Systems, 24(4):437–454, October 2006. CODEN ATISET. ISSN 1046-8188.

Chimera:1994:EET

[CS94] Richard Chimera and Ben Shneiderman. An exploratory evaluation of three interfaces for browsing large hierarchical tables of contents. ACM Transactions on Information Systems, 12(4):383–406, October 1994. CODEN ATISET. ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Cohen:1999:CSL

[CS99] William W. Cohen and Yoram Singer. Context-sensitive learning methods for text categorization. ACM Transactions on Information Systems, 17(2):141–173, April 1999. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org/pubs/citations/
                                        journals/tois/1999-17-2/
                                        p141-cohen/.

Cheng:2016:ELA

[CS16] Zhiyong Cheng and Jialie Shen. On effective locationaware music recommendation. ACM Transactions on Information Systems, 34(2):13:1– 13:??, April 2016. CODEN ATISET. ISSN 1046-8188.

Contreras:2021:ICL

[CSB21] David Contreras, Maria Salamó, and Ludovico Boratto. Integrating collaboration and leadership in conversational group recommender systems. ACM Transactions on Information Systems, 39 (4):41:1–41:32, October 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3462759.

Chu:2025:PDA

[CSH^{+}25] Yuxing Chu, Mingxu Sun, Ke Huang, Haokun Geng, Yanyu Zhang, Lili Zhang, and Menghua Zhang. Personality dialogue agent based on personality description and conversation history. ACM Transactions on Information Systems, 43(4):98:1–98:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Chen:2024:MDS

[CSJ^{+}24] Xiaolin Chen, Xuemeng Song, Liqiang Jing, Shuo Li, Linmei Hu, and Liqiang Nie. Multimodal dialog systems with dual knowledgeenhanced generative pretrained language model. ACM Transactions on Information Systems, 42(2):53:1– 53:??, March 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3606368.

Choudhury:2010:ECU

[CSJS10] Munmun De Choudhury, Hari Sundaram, Ajita John, and Doree Duncan Seligmann. Extraction, characterization and utility of prototypical communication groups in the blogosphere. ACM Transactions on Information Systems, 29 (1):6:1–6:??, December 2010. CODEN ATISET. ISSN 1046-8188.

Chia:2010:SLB

[CSLN10] Tee Kiah Chia, Khe Chai Sim, Haizhou Li, and Hwee Tou Ng. Statistical lattice-based spoken document retrieval. ACM Transactions on Information Systems, 28(1):2:1– 2:??, January 2010. CODEN ATISET. ISSN 1046-8188.

Chen:2020:FGP

[CSR^{+}20] Xiaolin Chen, Xuemeng Song, Ruiyang Ren, Lei Zhu, Zhiyong Cheng, and Liqiang Nie. Fine-grained privacy detection with graph-regularized hierarchical attentive representation learning. ACM Transactions on Information Systems, 38(4):37:1–37:26, October 2020. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3406109.

Chen:2023:SID

[CST^{+}23] Ziqian Chen, Fei Sun, Yifan Tang, Haokun Chen, Jinyang Gao, and Bolin Ding. Studying the impact of data disclosure mechanism in recommender systems via simulation. ACM Transactions on Information Systems, 41(3): 60:1–60:??, July 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3569452.

Cui:2022:SKA

[CSZ^{+}22] Yue Cui, Hao Sun, Yan Zhao, Hongzhi Yin, and Kai Zheng. Sequential-knowledge-aware next POI recommendation: a meta-learning approach. ACM Transactions on Information Systems, 40(2): 23:1–23:22, April 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3460198.

Chuklin:2015:CAI

[CSZD15] Aleksandr Chuklin, Anne Schuth, Ke Zhou, and Maarten De Rijke. A comparative analysis of interleaving methods for aggregated search. ACM Transactions on Information Systems, 33(2): 5:1–5:??, February 2015. CO-DEN ATISET. ISSN 1046-8188.

Chiu:2013:EVS

[CTH^{+}13] Chih-Yi Chiu, Tsung-Han Tsai, Guei-Wun Han, Cheng-Yu Hsieh, and Sheng-Yang Li. Efficient video stream monitoring for near-duplicate detection and localization in a large-scale repository. ACM Transactions on Information Systems, 31(4):22:1– 22:??, November 2013. CO-DEN ATISET. ISSN 1046-8188.

Comas:2012:SFQ

[CTM12] Pere R. Comas, Jordi Turmo, and Lluís Màrquez. Sibyl, a factoid question-answering system for spoken documents. ACM Transactions on Information Systems, 30(3):19:1– 19:??, August 2012. CODEN ATISET. ISSN 1046-8188.

Cummins:2014:DSD

[Cum14] Ronan Cummins. Document score distribution models for query performance inference and prediction. ACM Transactions on Information Systems, 32(1):2:1–2:??, January 2014. CODEN ATISET. ISSN 1046-8188.

Crestani:1998:SPK

[CV98] F. Crestani and C. J. Van Rijsbergen. A study of probability kinematics in information retrieval. ACM Transactions on Information Systems, 16(3):225– 255, July 1998. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tois/1998-16-3/p225-crestani/

.

Clarke:2021:ATP

[CVS21] Charles L. A. Clarke, Alexandra Vtyurina, and Mark D.

Smucker. Assessing top-k preferences. ACM Transactions on Information Systems, 39(3):33:1–33:21, July 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3451161.

Chang:1994:SAB

[CW94] Man Kit Chang and Carson C. Woo. A speech act based negotiation protocol: Design, implementation, and test use. ACM Transactions on Information Systems, 12(4):360–382, October 1994. CODEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Cannane:2002:GPC

[CW02] Adam Cannane and Hugh E.

Williams. A general-purpose compression scheme for large collections. ACM Transactions on Information Systems, 20(3):329–355, 2002. CODEN ATISET. ISSN 1046-8188.

Choi:2020:ETC

[CWL^{+}20] Bogeum Choi, Austin Ward, Yuan Li, Jaime Arguello, and Robert Capra. The effects of task complexity on the use of different types of information in a search assistance tool. ACM Transactions on Information Systems, 38(1):9:1– 9:28, February 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3371707.

Chen:2024:OCO

[CWM^{+}24] Xiaocong Chen, Siyu Wang, Julian McAuley, Dietmar Jannach, and Lina Yao. On the opportunities and challenges of offline reinforcement learning for recommender systems. ACM Transactions on Information Systems, 42(6):150:1– 150:??, November 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3661996.

Chen:2020:LVF

[CWZ^{+}20a] Yifan Chen, Yang Wang, Xiang Zhao, Hongzhi Yin, Ilya Markov, and Maarten De Rijke. Local variational featurebased similarity models for recommending top-N new items. ACM Transactions on Information Systems, 38(2): 12:1–12:33, March 2020. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3372154.

Chen:2020:BAI

[CWZ^{+}20b] Yifan Chen, Yang Wang, Xiang Zhao, Jie Zou, and Maarten De Rijke. Blockaware item similarity models for top-N recommendation. ACM Transactions on Information Systems, 38(4):42:1– 42:26, October 2020. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3411754.

Chen:2024:FFA

[CWZ^{+}24] Wei Chen, Yiqing Wu, Zhao Zhang, Fuzhen Zhuang, Zhongshi He, Ruobing Xie, and Feng Xia. FairGap: Fairness-aware recommendation via generating counterfactual graph. ACM Transactions on Information Systems, 42(4):94:1–94:??, July 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3638352.

Chen:2021:NFA

[CXZ^{+}21] Xu Chen, Kun Xiong, Yongfeng Zhang, Long Xia, Dawei Yin, and Jimmy Xiangji Huang. Neural feature-aware recommendation with signed hypergraph convolutional network. ACM Transactions on Information Systems, 39(1): 8:1–8:22, January 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3423322.

Chen:2024:AIP

[CYC^{+}24] Lijian Chen, Wei Yuan, Tong Chen, Guanhua Ye, Nguyen Quoc Viet Hung, and Hongzhi Yin. Adversarial item promotion on visually-aware recommender systems by guided diffusion. ACM Transactions on Information Systems, 42(6): 156:1–156:??, November 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3666088.

Cui:2024:DDG

[CYZ^{+}24] Chaoran Cui, Yumo Yao, Chunyun Zhang, Hebo Ma, Yuling Ma, Zhaochun Ren, Chen Zhang, and James Ko. DGEKT: a dual graph ensemble learning method for knowledge tracing. ACM Transactions on Information Systems, 42(3):78:1– 78:??, May 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3638350.

Ciaccia:1993:EAP

[CZ93] Paulo Ciaccia and Pavel Zezula. Estimating accesses in partitioned signature file organizations. ACM Transactions on Information Systems, 11(2):133–142, April 1993. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

Chai:2007:EIU

[CZJ07] Joyce Y. Chai, Chen Zhang, and Rong Jin. An empirical investigation of user term feedback in text-based targeted image search. ACM Transactions on Information Systems, 25(1):3:1–3:??, February 2007. CODEN ATISET. ISSN 1046-8188.

Chen:2024:SSP

[CZJH24] Junfan Chen, Richong Zhang, Xiaohan Jiang, and Chunming Hu. SPContrastNet: a self-paced contrastive learning model for few-shot text classification. ACM Transactions on Information Systems, 42 (5):130:1–130:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3652600.

Chen:2025:PTA

[CZL^{+}25] Chaochao Chen, Yizhao Zhang, Yuyuan Li, Jun Wang, Lianyong Qi, Xiaolong Xu, Xiaolin Zheng, and Jianwei Yin. Post-training attribute unlearning in recommender systems. ACM Transactions on Information Systems, 43 (1):25:1–25:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3701987.

Chen:2023:TET

[CZT^{+}23] Xu Chen, Ya Zhang, Ivor W.

Tsang, Yuangang Pan, and Jingchao Su. Toward equivalent transformation of user preferences in cross domain recommendation. ACM Transactions on Information Systems, 41(1):14:1– 14:??, January 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3522762.

Chen:2019:ADE

[CZX^{+}19] Xu Chen, Yongfeng Zhang, Hongteng Xu, Zheng Qin, and Hongyuan Zha. Adversarial distillation for efficient recommendation with external knowledge. ACM Transactions on Information Systems, 37(1):12:1–12:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Chen:2022:EHS

[CZX^{+}22] Meng Chen, Lei Zhu, Ronghui Xu, Yang Liu, Xiaohui Yu, and Yilong Yin. Embedding hierarchical structures for venue category representation. ACM Transactions on Information Systems, 40(3): 57:1–57:29, July 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3478285.

Cai:2017:ALC

[CZZ^{+}17] Wenbin Cai, Yexun Zhang, Ya Zhang, Siyuan Zhou, Wenquan Wang, Zhuoxiang Chen, and Chris Ding. Active learning for classification with maximum model change. ACM Transactions on Information Systems, 36(2):15:1– 15:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Chen:2020:ENM

[CZZ^{+}20] Chong Chen, Min Zhang, Yongfeng Zhang, Yiqun Liu, and Shaoping Ma. Efficient neural matrix factorization without sampling for recommendation. ACM Transactions on Information Systems, 38(2):14:1–14:28, January 2020. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/doi/abs/10.
1145/3373807.

Dacrema:2021:TAR

[DBCJ21] Maurizio Ferrari Dacrema, Simone Boglio, Paolo Cremonesi, and Dietmar Jannach. A troubling analysis of reproducibility and progress in recommender systems research. ACM Transactions on Information Systems, 39(2): 20:1–20:49, March 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3434185.

Dewan:1992:HLF

[DC92] Prasun Dewan and Rajiv Choudhary. A high-level and flexible framework for implementing multiuser user interfaces. ACM Transactions on Information Systems, 10(4): 345–380, October 1992. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

DellAmico:2010:DFP

[DC10] Matteo Dell’Amico and Licia Capra. Dependable filtering: Philosophy and realizations. ACM Transactions on Information Systems, 29 (1):5:1–5:??, December 2010. CODEN ATISET. ISSN 1046-8188.

Dong:2017:UMD

[DCT^{+}17] Yuxiao Dong, Nitesh V.

Chawla, Jie Tang, Yang Yang, and Yang Yang. User modeling on demographic attributes in big mobile social networks. ACM Transactions on Information Systems, 35(4):35:1– 35:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Dupplaw:2009:DSB

[DDH^{+}09] David Dupplaw, Srinandan Dasmahapatra, Bo Hu, Paul Lewis, and Nigel Shadbolt. A distributed, service-based framework for knowledge applications with multimedia. ACM Transactions on Information Systems, 27(4):22:1– 22:??, November 2009. CO-DEN ATISET. ISSN 1046-8188.

Deng:2024:MGD

[DDSW24] Zhirui Deng, Zhicheng Dou, Zhan Su, and Ji-Rong Wen. Multi-grained document modeling for search result diversification. ACM Transactions on Information Systems, 42 (5):126:1–126:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3652852.

Dourish:2000:EDM

[DEL^{+}00] Paul Dourish, W. Keith Edwards, Anthony LaMarca, John Lamping, Karin Petersen, Michael Salisbury, Douglas B. Terry, and James Thornton. Extending document management systems with user-specific active properties. ACM Transactions on Information Systems, 18(2): 140–170, April 2000. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.
acm.org/pubs/citations/
journals/tois/2000-18-2/
p140-dourish/.

Dimitrova:1995:MRV

[DG95] Nevenka Dimitrova and Forouzan Golshani. Motion recovery for video content classification. ACM Transactions on Information Systems, 13(4):408– 439, October 1995. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:

80. Special Issue on Video Information Retrieval.

Datta:2023:RIG

[DGMG23] Suchana Datta, Debasis Ganguly, Mandar Mitra, and Derek Greene. A relative information gain-based query performance prediction framework with generated query variants. ACM Transactions on Information Systems, 41(2):38:1–38:??, April 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3545112.

Dreilinger:1997:ESS

[DH97] Daniel Dreilinger and Adele E.

Howe. Experiences with selecting search engines using metasearch. ACM Transactions on Information Systems, 15(3):195– 222, July 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:

                                        80/tois/abstracts/dreilinger.
                                        html.

Dong:2025:CTP

[DHC^{+}25] Zhiang Dong, Liya Hu, Jingyuan Chen, Zhihua Wang, and Fei Wu. Comprehend then predict: Prompting large language models for recommendation with semantic and collaborative data. ACM Transactions on Information Systems, 43(5):115:1–115:26, September 2025. CODEN ATISET. ISSN 1046-8188.

Du:2019:MED

[DHY^{+}19] Xiaoyu Du, Xiangnan He, Fajie Yuan, Jinhui Tang, Zhiguang Qin, and Tat-Seng Chua. Modeling embedding dimension correlations via convolutional neural collaborative filtering. ACM Transactions on Information Systems, 37(4): 47:1–47:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3357154.

Deshpande:2004:IBT

[DK04] Mukund Deshpande and George Karypis. Item-based top-N recommendation algorithms. ACM Transactions on Information Systems, 22(1): 143–177, January 2004. CO-DEN ATISET. ISSN 1046-8188.

Do:2016:PMC

[DL16] Loc Do and Hady W. Lauw.

Probabilistic models for contextual agreement in preferences. ACM Transactions on Information Systems, 34(4): 21:1–21:??, September 2016. CODEN ATISET. ISSN 1046-8188.

Dang:2015:FFI

[DLA15] Edward Kai Fung Dang, Robert Wing Pong Luk, and James Allan. Fast forward index methods for pseudorelevance feedback retrieval. ACM Transactions on Information Systems, 33(4):19:1– 19:??, May 2015. CODEN ATISET. ISSN 1046-8188.

Dang:2022:CBT

[DLA22] Edward Kai Fung Dang, Robert Wing Pong Luk, and James Allan. A comparison between term-independence retrieval models for ad hoc retrieval. ACM Transactions on Information Systems, 40(3): 62:1–62:37, July 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3483612.

Deng:2025:PCA

[DLL^{+}25] Yang Deng, Lizi Liao, Wenqiang Lei, Grace Hui Yang, Wai Lam, and Tat-Seng Chua. Proactive conversational AI: a comprehensive survey of advancements and opportunities. ACM Transactions on Information Systems, 43(3):

67:1–67:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

Dato:2016:FRA

[DLN^{+}16] Domenico Dato, Claudio Lucchese, Franco Maria Nardini, Salvatore Orlando, Raffaele Perego, Nicola Tonellotto, and Rossano Venturini. Fast ranking with additive ensembles of oblivious and non-oblivious regression trees. ACM Transactions on Information Systems, 35(2):15:1– 15:??, December 2016. CO-DEN ATISET. ISSN 1046-8188.

Dang:2025:EAR

[DLY^{+}25] Yizhou Dang, Yuting Liu, Enneng Yang, Guibing Guo, Linying Jiang, Jianzhe Zhao, and Xingwei Wang. Efficient and adaptive recommendation unlearning: a guided filtering framework to erase outdated preferences. ACM Transactions on Information Systems, 43(2):49:1– 49:??, March 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3706633.

Deng:2022:TPA

[DLZ^{+}22] Yang Deng, Yaliang Li, Wenxuan Zhang, Bolin Ding, and Wai Lam. Toward personalized answer generation in e-commerce via multiperspective preference modeling. ACM Transactions on Information Systems, 40 (4):87:1–87:28, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3507782.

Darwish:2007:ECV

[DM07] Kareem Darwish and Walid Magdy. Error correction vs. query garbling for Arabic OCR document retrieval. ACM Transactions on Information Systems, 26(1):5:1– 5:??, November 2007. CO-DEN ATISET. ISSN 1046-8188.

deMoura:2008:LBP

[dMdSsdA^{+}08] Edleno Silva de Moura, Celia Francisca dos Santos, Bruno Dos santos de Araujo, Altigran Soares da Silva, Pavel Calado, and Mario A. Nascimento. Locality-based pruning methods for web search. ACM Transactions on Information Systems, 26(2): 9:1–9:??, March 2008. CO-DEN ATISET. ISSN 1046-8188.

Deveaud:2019:LAR

[DMUN19] Romain Deveaud, Josiane Mothe, Md Zia Ullah, and Jian-Yun Nie. Learning to adaptively rank document retrieval system configurations. ACM Transactions on Information Systems, 37(1):3:1– 3:??, January 2019. CODEN ATISET. ISSN 1046-8188.

deOliveira:2020:OSA

[dODMS20] Wyverson Bonasoli de Oliveira, Leyza Baldo Dorini, Rodrigo Minetto, and Thiago H. Silva. OutdoorSent: Sentiment analysis of urban outdoor images by using semantic and deep features. ACM Transactions on Information Systems, 38(3):23:1– 23:28, June 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3385186.

deOliveira:2001:SBM

[dOTM01] Maria Cristina Ferreira de Oliveira, Marcelo Augusto Santos Turine, and Paulo Cesar Masiero. A statechartbased model for hypermedia applications. ACM Transactions on Information Systems, 19(1):28–52, 2001. CODEN ATISET. ISSN 1046-8188. URL

http://www.acm.org/pubs/
articles/journals/tois/
2001-19-1/p28-de_oliveira/
p28-de_oliveira.pdf;   http:
//www.acm.org/pubs/citations/
journals/tois/2001-19-1/
p28-de_oliveira/.

Dusart:2023:THS

[DPSH23] Alexis Dusart, Karen Pinel-Sauvagnat, and Gilles Hubert. TSSuBERT: How to sum up multiple years of reading in a few tweets. ACM Transactions on Information Systems, 41(4):109:1–109:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3581786.

deRijke:2019:RAT

[dR19] Maarten de Rijke. Reviewers for ACM Transactions on Information Systems volume 37. ACM Transactions on Information Systems, 37(4): 49:1–49:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3365367.

Deffayet:2023:ERC

[DRdR23] Romain Deffayet, Jean-Michel Renders, and Maarten de Rijke. Evaluating the robustness of click models to policy distributional shift. ACM Transactions on Information Systems, 41(4):84:1– 84:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3569086.

Di:2025:FRS

[DSW^{+}25] Yicheng Di, Hongjian Shi, Xiaoming Wang, Ruhui Ma, and Yuan Liu. Federated recommender system based on diffusion augmentation and guided denoising. ACM Transactions on Information Systems, 43 (2):31:1–31:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3688570.

Dong:2023:DPD

[DSZ^{+}23] Xue Dong, Xuemeng Song, Na Zheng, Yinwei Wei, and Zhongzhou Zhao. Dual preference distribution learning for item recommendation. ACM Transactions on Information Systems, 41(3):56:1– 56:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3565798.

Dunlop:1997:EAN

[Dun97] Mark D. Dunlop. The effect of accessing nonmatching documents on relevance feedback. ACM Transactions on Information Systems, 15(2): 137–153, April 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/dunlop.
html.

Dang:2009:BFP

[DWLW09] Edward Kai Fung Dang, Ho Chung Wu, Robert Wing Pong Luk, and Kam Fai Wong. Building a framework for the probability ranking principle by a family of expected weighted rank. ACM Transactions on Information Systems, 27(4):20:1– 20:??, November 2009. CO-DEN ATISET. ISSN 1046-8188.

Deng:2022:CKA

[DXL^{+}22] Yang Deng, Yuexiang Xie, Yaliang Li, Min Yang, Wai Lam, and Ying Shen. Contextualized knowledge-aware attentive neural network: Enhancing answer selection with knowledge. ACM Transactions on Information Systems, 40(1):2:1–2:33, January 2022.

CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3457533.

Dai:2022:BRR

[DXZ^{+}22] Xinyi Dai, Yunjia Xi, Weinan Zhang, Qing Liu, Ruiming Tang, Xiuqiang He, Jiawei Hou, Jun Wang, and Yong Yu. Beyond relevance ranking: a general graph matching framework for utilityoriented learning to rank. ACM Transactions on Information Systems, 40(2):25:1– 25:29, April 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3464303.

Ding:2020:IIR

[DYL^{+}20] Jingtao Ding, Guanghui Yu, Yong Li, Xiangnan He, and Depeng Jin. Improving implicit recommender systems with auxiliary data. ACM Transactions on Information Systems, 38(1):11:1–11:27, February 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3372338.

Deng:2023:UMT

[DZX^{+}23] Yang Deng, Wenxuan Zhang, Weiwen Xu, Wenqiang Lei, Tat-Seng Chua, and Wai Lam. A unified multi-task learning framework for multi-goal conversational recommender systems. ACM Transactions on Information Systems, 41(3): 77:1–77:??, July 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3570640.

Elsweiler:2008:EME

[EBR08] David Elsweiler, Mark Baillie, and Ian Ruthven. Exploring memory in email refinding. ACM Transactions on Information Systems, 26(4): 21:1–21:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Entlich:1997:MDL

[EGL^{+}97] Richard Entlich, Lorrin Garson, Michael Lesk, Lorraine Normore, Jan Olsen, and Stuart Weibel. Making a digital library: The contents of the CORE project. ACM Transactions on Information Systems, 15(2):103– 123, April 1997. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/entlich.
                                        html.

Eastman:2003:CRR

[EJ03] Caroline M. Eastman and Bernard J. Jansen. Coverage, relevance, and ranking: The impact of query operators on Web search engine results. ACM Transactions on Information Systems, 21(4): 383–411, October 2003. CO-DEN ATISET. ISSN 1046-8188.

El-Kwae:1999:RFC

[EKK99] Essam A. El-Kwae and Mansur R. Kabuka. A robust framework for content-based retrieval by spatial similarity in image databases. ACM Transactions on Information Systems, 17(2):174–198, April 1999. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org/pubs/citations/
journals/tois/1999-17-2/
p174-el-kwae/.

El-Kwae:2000:ECB

[EKK00] Essam A. El-Kwae and Mansur R. Kabuka. Efficient content-based indexing of large image databases. ACM Transactions on Information Systems, 18(2):171– 210, April 2000. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org/
pubs/citations/journals/
tois/2000-18-2/p171-el-
kwae/.

Egozi:2011:CBI

[EMG11] Ofer Egozi, Shaul Markovitch, and Evgeniy Gabrilovich. Concept-based information retrieval using explicit semantic analysis. ACM Transactions on Information Systems, 29(2):8:1–8:??, April 2011. CODEN ATISET. ISSN 1046-8188.

Esuli:2019:FNE

[EMS19] Andrea Esuli, Alejandro Moreo, and Fabrizio Sebastiani. Funnelling: a new ensemble method for heterogeneous transfer learning and its application to cross-lingual text classification. ACM Transactions on Information Systems, 37(3):37:1–37:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3326065.

Esuli:2021:CRS

[EMS21] Andrea Esuli, Alessio Molinari, and Fabrizio Sebastiani. A critical reassessment of the Saerens–Latinne– Decaestecker algorithm for posterior probability adjustment. ACM Transactions on Information Systems, 39(2): 19:1–19:34, March 2021. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3433164.

Egan:1989:FDE

[ERG^{+}89] Dennis E. Egan, Joel R. Remde, Louis M. Gomez, Thomas K. Landauer, Jennifer Eberhardt, and Carol C. Lochbaum. Formative designevaluation of SuperBook. ACM Transactions on Information Systems, 7(1):30– 57, January 1989. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80.

Egenhofer:1998:MDN

[ES98] Max J. Egenhofer and A. Rashid B. M. Shariff. Metric details for natural-language spatial relations. ACM Transactions on Information Systems, 16(4):295–321, October 1998. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/pubs/citations/journals/
tois/1998-16-4/p295-egenhofer/

.

Esuli:2013:ITC

[ES13] Andrea Esuli and Fabrizio Sebastiani. Improving text classification accuracy by training label cleaning. ACM Transactions on Information Systems, 31(4):19:1–19:??, November 2013. CODEN ATISET. ISSN 1046-8188.

Fu:2012:SSL

[FAZC12] Tianjun Fu, Ahmed Abbasi, Daniel Zeng, and Hsinchun Chen. Sentimental spidering: Leveraging opinion information in focused crawlers. ACM Transactions on Information Systems, 30(4):24:1– 24:??, November 2012. CO-DEN ATISET. ISSN 1046-8188.

Fuhr:1991:PLA

[FB91] Norbert Fuhr and Chris Buckley. A probabilistic learning approach for document indexing. ACM Transactions on Information Systems, 9(3): 223–248, July 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Farina:2012:WBS

[FBN^{+}12] Antonio Fariña, Nieves R.

Brisaboa, Gonzalo Navarro, Francisco Claude, Ángeles S.

Places, and Eduardo Rodríguez. Word-based self-indexes for natural language text. ACM Transactions on Information Systems, 30(1):1:1–1:??, February 2012. CODEN ATISET. ISSN 1046-8188.

Ford:1991:OPH

[FC91] Daniel Alexander Ford and Stavros Christodoulakis. Optimal placement of high probability randomly retrieved blocks on CLV optical discs. ACM Transactions on Information Systems, 9(1):1– 30, January 1991. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80.

Farseev:2017:TCF

[FC17] Aleksandr Farseev and Tat-Seng Chua. Tweet can be fit: Integrating data from wearable sensors and multiple social networks for wellness profile learning. ACM Transactions on Information Systems, 35(4):42:1–42:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Feng:2002:SNB

[FCD02] Ling Feng, Elizabeth Chang, and Tharam Dillon. A semantic network-based design methodology for XML documents. ACM Transactions on Information Systems, 20 (4):390–421, 2002. CODEN ATISET. ISSN 1046-8188.

Fox:1991:OPM

[FCDH91] Edward A. Fox, Qi Fan Chen, Amjad M. Daoud, and Lenwood S. Heath. Order preserving minimal perfect hash functions and information retrieval. ACM Transactions on Information Systems, 9(3): 281–308, July 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Fei:2023:RAB

[FCL^{+}23] Hao Fei, Tat-Seng Chua, Chenliang Li, Donghong Ji, Meishan Zhang, and Yafeng Ren. On the robustness of aspect-based sentiment analysis: Rethinking model, data, and training. ACM Transactions on Information Systems, 41(2):50:1–50:??, April 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3564281.

Frummet:2022:WCC

[FEL22] Alexander Frummet, David Elsweiler, and Bernd Ludwig. “What Can I Cook with these Ingredients?” — understanding cooking-related information needs in conversational search. ACM Transactions on Information Systems, 40 (4):81:1–81:32, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3498330.

Ferro:2017:WDA

[Fer17] Nicola Ferro. What does affect the correlation among evaluation measures? ACM Transactions on Information Systems, 36(2):19:1–19:??, September 2017. CODEN ATISET. ISSN 1046-8188.

Ferrante:2017:AEE

[FFM17] Marco Ferrante, Nicola Ferro, and Maria Maistro. AWARE: Exploiting evaluation measures to combine multiple assessors. ACM Transactions on Information Systems, 36(2): 20:1–20:??, September 2017. CODEN ATISET. ISSN 1046-8188.

Feiner:1993:EVW

[FG93] Steven K. Feiner and Simon J. Gibbs. Editorial: Virtual worlds. ACM Transactions on Information Systems, 11 (3):195–196, July 1993. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.

org:80. Special Issue on Virtual Worlds.

Fuhr:2004:XXQ

[FG04] Norbert Fuhr and Kai Großjohann. XIRQL: An XML query language based on information retrieval concepts. ACM Transactions on Information Systems, 22(2):313–356, April 2004. CODEN ATISET. ISSN 1046-8188.

Finkelstein:2002:PSC

[FGM^{+}02] Lev Finkelstein, Evgeniy Gabrilovich, Yossi Matias, Ehud Rivlin, Zach Solan, Gadi Wolfman, and Eytan Ruppin. Placing search in context: The concept revisited. ACM Transactions on Information Systems, 20(1): 116–131, January 2002. CO-DEN ATISET. ISSN 1046-8188.

Feng:2019:TRR

[FHW^{+}19] Fuli Feng, Xiangnan He, Xiang Wang, Cheng Luo, Yiqun Liu, and Tat-Seng Chua. Temporal relational ranking for stock prediction. ACM Transactions on Information Systems, 37(2):27:1– 27:??, March 2019. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.
acm.org/ft_gateway.cfm?
id=3309547.

Fan:2024:OMA

[FJZS24] Yu-Chen Fan, Yitong Ji, Jie Zhang, and Aixin Sun. Our model achieves excellent performance on MovieLens: What does it mean? ACM Transactions on Information Systems, 42(6):159:1–159:??, November 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3675163.

Fox:2005:EIM

[FKM^{+}05] Steve Fox, Kuldeep Karnawat, Mark Mydland, Susan Dumais, and Thomas White. Evaluating implicit measures to improve Web search. ACM Transactions on Information Systems, 23(2):147–168, April 2005. CODEN ATISET. ISSN 1046-8188.

Ferro:2019:UCS

[FKS19] Nicola Ferro, Yubin Kim, and Mark Sanderson. Using collection shards to study retrieval performance effect sizes. ACM Transactions on Information Systems, 37(3):30:1–30:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3310364.

Fang:2022:HVC

[FLMD22] Jinyuan Fang, Shangsong Liang, Zaiqiao Meng, and Maarten De Rijke. Hyperspherical variational coembedding for attributed networks. ACM Transactions on Information Systems, 40(3): 58:1–58:36, July 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3478284.

Fischer:1991:RCC

[FLMM91] Gerhard Fischer, Andreas C. Lemke, Thomas Mastaglio, and Anders I. Morch. The role of critiquing in cooperative problem solving. ACM Transactions on Information Systems, 9(2):123–151, April 1991. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Computer-Human Interaction.

Formal:2024:TEE

[FLPC24] Thibault Formal, Carlos Lassance, Benjamin Piwowarski, and Stéphane Clinchant. Towards effective and efficient sparse neural information retrieval. ACM Transactions on Information Systems, 42 (5):116:1–116:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3634912.

Feng:2023:RRP

[FLW^{+}23] Chao Feng, Defu Lian, Xiting Wang, Zheng Liu, Xing Xie, and Enhong Chen. Reinforcement routing on proximity graph for efficient recommendation. ACM Transactions on Information Systems, 41(1):8:1–8:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3512767.

Fu:2025:UFM

[FLW^{+}25] Zichuan Fu, Xiangyang Li, Chuhan Wu, Yichao Wang, Kuicai Dong, Xiangyu Zhao, Mengchen Zhao, Huifeng Guo, and Ruiming Tang. A unified framework for multidomain CTR prediction via large language models. ACM Transactions on Information Systems, 43(5):117:1–117:33, September 2025. CODEN ATISET. ISSN 1046-8188.

Forsati:2014:MFE

[FMSS14] Rana Forsati, Mehrdad Mahdavi, Mehrnoush Shamsfard, and Mohamed Sarwat. Matrix factorization with explicit trust and distrust side information for improved social recommendation. ACM Transactions on Information Systems, 32(4):17:1– 17:??, October 2014. CODEN ATISET. ISSN 1046-8188.

Friedman:1996:BCS

[FN96] Batya Friedman and Helen Nissenbaum. Bias in computer systems. ACM Transactions on Information Systems, 14(3):330– 347, July 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/friedman.
html.

Fuhr:1994:PIR

[FP94] Norbert Fuhr and Ulrich Pfeifer. Probabilistic information retrieval as a combination of abstraction, inductive learning, and probabilistic assumptions. ACM Transactions on Information Systems, 12(1):92–115, January 1994. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Fraternali:2000:MDD

[FP00] Piero Fraternali and Paolo Paolini. Model-driven development of Web applications: the AutoWeb system. ACM Transactions on Information Systems, 18(4):323– 382, 2000. URL http://www.

acm.org/pubs/citations/
journals/tois/2000-18-4/
p323-fraternali/.

Fagni:2006:BPW

[FPSO06] Tiziano Fagni, Raffaele Perego, Fabrizio Silvestri, and Salvatore Orlando. Boosting the performance of Web search engines: Caching and prefetching query results by exploiting historical usage data. ACM Transactions on Information Systems, 24(1): 51–78, January 2006. CODEN ATISET. ISSN 1046-8188.

Fuhr:1997:PRA

[FR97] Norbert Fuhr and Thomas Rölleke. A probabilistic relational algebra for the integration of information retrieval and database systems. ACM Transactions on Information Systems, 15(1):32– 66, January 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/fuhr.
html.

Fang:2007:SMT

[FSC07] Xiao Fang, Olivia R. Liu Sheng, and Michael Chau. ServiceFinder: a method towards enhancing service portals. ACM Transactions on Information Systems, 25 (4):17:1–17:??, October 2007. CODEN ATISET. ISSN 1046-

Frummet:2024:CCE

[FSE^{+}24] Alexander Frummet, Alessandro Speggiorin, David Elsweiler, Anton Leuski, and Jeff Dalton. Cooking with conversation: Enhancing user engagement and learning with a knowledge-enhancing assistant. ACM Transactions on Information Systems, 42 (5):122:1–122:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3649500.

Fang:2011:DEI

[FTZ11] Hui Fang, Tao Tao, and Chengxiang Zhai. Diagnostic evaluation of information retrieval models. ACM Transactions on Information Systems, 29(2):7:1–7:??, April 2011. CODEN ATISET. ISSN 1046-8188.

Feng:2023:LMT

[FTZZ23] Jiazhan Feng, Chongyang Tao, Xueliang Zhao, and Dongyan Zhao. Learning multi-turn response selection in grounded dialogues with reinforced knowledge and context distillation. ACM Transactions on Information Systems, 41(4):115:1–115:??, October 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3584701.

Fuhr:1989:OPR

[Fuh89]  Norbert Fuhr.  Optimum

polynomial retrieval functions based on the probability ranking principle. ACM Transactions on Information Systems, 7(3):183–204, July 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Fuhr:1999:DTA

[Fuh99] Norbert Fuhr. A decisiontheoretic approach to database selection in networked IR. ACM Transactions on Information Systems, 17(3):229, July 1999. CODEN ATISET. ISSN 1046-8188. URL

http://www.acm.org/pubs/
citations/journals/tois/
1999-17-3/p229-fuhr/.

Fu:2024:PPI

[FWG^{+}24] Wenjie Fu, Huandong Wang, Chen Gao, Guanghua Liu, Yong Li, and Tao Jiang. Privacy-preserving individuallevel COVID-19 infection prediction via federated graph learning. ACM Transactions on Information Systems, 42(3):82:1–82:??, May 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3633202.

Fitzmaurice:1993:VRP

[FZC93] George W. Fitzmaurice, Shumin Zhai, and Mark H. Chignell. Virtual reality for palmtop computers. ACM Transactions on Information Systems, 11(3):197–218, July 1993. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Virtual Worlds.

Fang:2022:SRL

[FZH^{+}22] Yang Fang, Xiang Zhao, Peixin Huang, Weidong Xiao, and Maarten de Rijke. Scalable representation learning for dynamic heterogeneous information networks via metagraphs. ACM Transactions on Information Systems, 40 (4):64:1–64:27, October 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3485189.

Fang:2021:DLS

[FZSG21] Hui Fang, Danning Zhang, Yiheng Shu, and Guibing Guo. Deep learning for sequential recommendation: Algorithms, influential factors, and evaluations. ACM Transactions on Information Systems, 39(1):10:1–10:42, January 2021. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3426723.

Fang:2024:FSL

[FZXdR24] Yang Fang, Xiang Zhao, Weidong Xiao, and Maarten de Rijke. Few-shot learning for heterogeneous information networks. ACM Transactions on Information Systems, 42 (4):107:1–107:??, July 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3649311.

Goh:1999:CIN

[GBMS99] Cheng Hian Goh, Stéphane Bressan, Stuart Madnick, and Michael Siegel. Context interchange: new features and formalisms for the intelligent integration of information. ACM Transactions on Information Systems, 17(3): 270, July 1999. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org/
pubs/citations/journals/
tois/1999-17-3/p270-goh/.

Gemmell:1992:PDS

[GC92] Jim Gemmell and Stavros Christodoulakis. Principles of delay sensitive multi-media data storage and retrieval. ACM Transactions on Information Systems, 10(1):51– 90, January 1992. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Guo:2025:CCL

[GCB^{+}25] Jiafeng Guo, Yinqiong Cai, Keping Bi, Yixing Fan, Wei Chen, Ruqing Zhang, and Xueqi Cheng. CAME: Competitively learning a mixtureof-experts model for firststage retrieval. ACM Transactions on Information Systems, 43(2):35:1–35:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3678880.

Gerani:2012:AMP

[GCC12] Shima Gerani, Mark Carman, and Fabio Crestani. Aggregation methods for proximitybased opinion retrieval. ACM Transactions on Information Systems, 30(4):26:1–26:??, November 2012. CODEN ATISET. ISSN 1046-8188.

Guo:2022:SMF

[GCF^{+}22] Jiafeng Guo, Yinqiong Cai, Yixing Fan, Fei Sun, Ruqing Zhang, and Xueqi Cheng. Semantic models for the firststage retrieval: a comprehensive review. ACM Transactions on Information Systems, 40(4):66:1–66:42, October 2022. CODEN ATISET. ISSN 1046-8188. URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3486250.

Guan:2019:AAM

[GCH^{+}19] Xinyu Guan, Zhiyong Cheng, Xiangnan He, Yongfeng Zhang, Zhibo Zhu, Qinke Peng, and Tat-Seng Chua. Attentive aspect modeling for review-aware recommendation. ACM Transactions on Information Systems, 37 (3):28:1–28:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3309546.

Gao:2021:LRY

[GCL^{+}21] Shen Gao, Xiuying Chen, Li Liu, Dongyan Zhao, and Rui Yan. Learning to respond with your favorite stickers: a framework of unifying multi-modality and user preference in multi-turn dialog. ACM Transactions on Information Systems, 39(2): 12:1–12:32, March 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3429980.

Guo:2019:ALS

[GCN^{+}19] Yangyang Guo, Zhiyong Cheng, Liqiang Nie, Yinglong Wang, Jun Ma, and Mohan Kankanhalli. Attentive long short-term preference modeling for personalized product search. ACM Transactions on Information Systems, 37(2):19:1–19:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3295822.

Gao:2021:MAG

[GCR^{+}21] Shen Gao, Xiuying Chen, Zhaochun Ren, Dongyan Zhao, and Rui Yan. Meaningful answer generation of e-commerce questionanswering. ACM Transactions on Information Systems, 39(2):18:1–18:26, March 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3432689.

Greiff:1999:PMC

[GCT99] Warren R. Greiff, W. Bruce Croft, and Howard Turtle. PIC matrices: a computationally tractable class of probabilistic query operators. ACM Transactions on Information Systems, 17(4):367– 405, October 1999. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tois/1999-17-4/
                                        p367-greiff/.

Goncalves:2004:SSS

[GFWK04] Marcos André Gonçalves, Edward A. Fox, Layne T. Watson, and Neill A. Kipp. Streams, structures, spaces, scenarios, societies (5s): a formal model for digital libraries. ACM Transactions on Information Systems, 22(2):270– 312, April 2004. CODEN ATISET. ISSN 1046-8188.

Ganesan:2003:EHD

[GGMW03] Prasanna Ganesan, Hector Garcia-Molina, and Jennifer Widom. Exploiting hierarchical domain structure to compute similarity. ACM Transactions on Information Systems, 21(1):64–93, 2003. CO-DEN ATISET. ISSN 1046-8188.

Gatmiry:2022:NVP

[GGR22] Khashayar Gatmiry and Manuel Gomez-Rodriguez. The network visibility problem. ACM Transactions on Information Systems, 40(2): 22:1–22:42, April 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3460475.

Gaeta:2023:RQV

[GGRF23] Rossano Gaeta, Michele Garetto, Giancarlo Ruffo, and Alessandro Flammini. Reconciling the quality vs popularity dichotomy in online cultural markets. ACM Transactions on Information Systems, 41(1):25:1– 25:??, January 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3530790.

Grant:1996:CPM

[GH96] Rebecca A. Grant and Chris A.

Higgins. Computerized performance monitors as multidimensional systems: Derivation and application. ACM Transactions on Information Systems, 14(2):212– 235, April 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/grant.
html.

Giner:2024:IRE

[Gin24] Fernando Giner. Information retrieval evaluation measures defined on some axiomatic models of preferences. ACM Transactions on Information Systems, 42(3):69:1– 69:??, May 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3632171.

Gravano:2003:QSA

[GIS03] Luis Gravano, Panagiotis G.

Ipeirotis, and Mehran Sahami. QProber: a system for automatic classification of hidden-Web databases. ACM Transactions on Information Systems, 21(1):1–41, 2003. CODEN ATISET. ISSN 1046-8188.

Ge:2025:LEC

[GJS^{+}25] Hongfei Ge, Yuanchun Jiang, Jianshan Sun, Kun Yuan, and Yezheng Liu. LLM-enhanced composed image retrieval: an intent uncertainty-aware linguistic-visual dual channel matching model. ACM Transactions on Information Systems, 43(2):37:1–37:??, March 2025. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3699715.

Gerstel:2007:RHI

[GKL^{+}07] Ori Gerstel, Shay Kutten, Eduardo Sany Laber, Rachel Matichin, David Peleg, Artur Alves Pessoa, and Criston Souza. Reducing human interactions in Web directory searches. ACM Transactions on Information Systems, 25 (4):20:1–20:??, October 2007. CODEN ATISET. ISSN 1046-8188.

Gladney:2005:TYD

[GL05] H. M. Gladney and R. A. Lorie. Trustworthy 100-year digital objects: durable encoding for when it’s too late to ask. ACM Transactions on Information Systems, 23(3):

299–324, July 2005. CODEN ATISET. ISSN 1046-8188.

Geneves:2006:SSA

[GL06] Pierre Genevès and Nabil Layaïda. A system for the static analysis of XPath. ACM Transactions on Information Systems, 24(4):475– 502, October 2006. CODEN ATISET. ISSN 1046-8188.

Gladney:1997:ACL

[Gla97]  H. M. Gladney.  Access

control for large collections. ACM Transactions on Information Systems, 15(2):154– 194, April 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/gladney.
html.

Gladney:2004:TYD

[Gla04] Henry M. Gladney. Trustworthy 100-year digital objects: Evidence after every witness is dead. ACM Transactions on Information Systems, 22 (3):406–436, July 2004. CO-DEN ATISET. ISSN 1046-8188.

Guo:2023:DRL

[GLG^{+}23] Xiaobo Guo, Shaoshuai Li, Naicheng Guo, Jiangxia Cao, Xiaolei Liu, Qiongxu Ma, Runsheng Gan, and Yunan Zhao. Disentangled representations learning for multitarget cross-domain recommendation. ACM Transactions on Information Systems, 41(4):85:1–85:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3572835.

Guo:2023:PHG

[GLL^{+}23] Naicheng Guo, Xiaolei Liu, Shaoshuai Li, Qiongxu Ma, Kaixin Gao, Bing Han, Lin Zheng, Sheng Guo, and Xiaobo Guo. Poincaré heterogeneous graph neural networks for sequential recommendation. ACM Transactions on Information Systems, 41(3):63:1–63:??, July 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3568395.

Guo:2024:DDA

[GLZ^{+}24] Lei Guo, Hao Liu, Lei Zhu, Weili Guan, and Zhiyong Cheng. DA-DAN: a dual adversarial domain adaption network for unsupervised nonoverlapping cross-domain recommendation. ACM Transactions on Information Systems, 42(2):48:1–48:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3617825.

Guiver:2009:FGT

[GMR09] John Guiver, Stefano Mizzaro, and Stephen Robertson. A few good topics: Experiments in topic set reduction for retrieval evaluation. ACM Transactions on Information Systems, 27(4):21:1– 21:??, November 2009. CO-DEN ATISET. ISSN 1046-8188.

Gao:2010:EQL

[GNN^{+}10] Wei Gao, Cheng Niu, Jian-Yun Nie, Ming Zhou, Kam-Fai Wong, and Hsiao-Wuen Hon. Exploiting query logs for cross-lingual query suggestions. ACM Transactions on Information Systems, 28(2): 6:1–6:??, May 2010. CODEN ATISET. ISSN 1046-8188.

Greiff:2000:MEA

[GP00] Warren R. Greiff and Jay M.

Ponte. The maximum entropy approach and probabilistic IR models. ACM Transactions on Information Systems, 18(3):246–287, 2000. URL

http://www.acm.org/pubs/
citations/journals/tois/
2000-18-3/p246-greiff/.

Garzotto:1993:HMB

[GPS93] Franca Garzotto, Paolo Paolini, and Daniel Schwabe. HDM — A model based approach to hypertext application design. ACM Transactions on Information Systems, 11(1): 1–26, January 1993. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80.

Gudivada:1995:DEA

[GR95] Venkat N. Gudivada and Vijay V. Raghavan. Design and evaluation of algorithms for image retrieval by spatial similarity. ACM Transactions on Information Systems, 13(2):

115–144, April 1995. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Guglielmo:1996:NLR

[GR96] Eugene J. Guglielmo and Neil C. Rowe. Naturallanguage retrieval of images based on descriptive captions. ACM Transactions on Information Systems, 14(3):237– 267, July 1996. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/guglielmo.
                                        html.

Guttenbrunner:2012:MFE

[GR12] Mark Guttenbrunner and Andreas Rauber. A measurement framework for evaluating emulators for digital preservation. ACM Transactions on Information Systems, 30(2): 14:1–14:??, May 2012. CO-DEN ATISET. ISSN 1046-8188.

Gomez-Rodriguez:2016:IEM

[GRSD^{+}16] Manuel Gomez-Rodriguez, Le Song, Nan Du, Hongyuan Zha, and Bernhard Schölkopf. Influence estimation and maximization in continuous-time diffusion networks. ACM Transactions on Information Systems, 34(2):9:1–9:??, April 2016. CODEN ATISET. ISSN 1046-8188.

Grudin:1992:CSF

[Gru92] Jonathan Grudin. Consistency, standards, and formal approaches to interface development and evaluation: a note on Wiecha, Bennett, Boies, Gould, and Greene. ACM Transactions on Information Systems, 10(1):103– 111, January 1992. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Gauch:1991:SIA

[GS91] Susan Gauch and John B.

Smith. Search improvement via automatic query reformulation. ACM Transactions on Information Systems, 9(3): 249–280, July 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Guo:2025:SEC

[GSG^{+}25] Lei Guo, Chenlong Song, Feng Guo, Xiaohui Han, Xiaojun Chang, and Lei Zhu. Semantic-enhanced coattention prompt learning for non-overlapping cross-domain recommendation. ACM Transactions on Information Systems, 43(5):137:1–137:27, September 2025. CODEN ATISET. ISSN 1046-8188.

Gottlob:1996:EOO

[GSR96] Georg Gottlob, Michael Schrefl, and Brigitte Rock. Extending object-oriented systems with roles. ACM Transactions on Information Systems, 14(3): 268–296, July 1996. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/gottlob.
                                        html.

Goldberg:2018:FID

[GTW^{+}18] David Goldberg, Andrew Trotman, Xiao Wang, Wei Min, and Zongru Wan. Further insights on drawing sound conclusions from noisy judgments. ACM Transactions on Information Systems, 36(4):36:1–36:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Gulla:1996:GEC

[Gul96] Jon Atle Gulla. A general explanation component for conceptual modeling in CASE environments. ACM Transactions on Information Systems, 14(3):297– 329, July 1996. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/gulla.
                                        html.

Guy:2018:CVS

[Guy18] Ido Guy. The characteristics of voice search: Comparing spoken with typed-in mobile Web search queries. ACM Transactions on Information Systems, 36(3):30:1– 30:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Gao:2024:CBF

[GWL^{+}24] Chongming Gao, Shiqi Wang, Shijun Li, Jiawei Chen, Xiangnan He, Wenqiang Lei, Biao Li, Yuan Zhang, and Peng Jiang. CIRS: Bursting filter bubbles by counterfactual interactive recommender system. ACM Transactions on Information Systems, 42 (1):14:1–14:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3594871.

Gao:2006:MFM

[GWLC06] Sheng Gao, Wen Wu, Chin-Hui Lee, and Tat-Seng Chua. A maximal figure-of-merit (MFoM)-learning approach to robust classifier design for text categorization. ACM Transactions on Information Systems, 24(2):190–218, April 2006. CODEN ATISET. ISSN 1046-8188.

Gauch:1999:CAA

[GWR99] Susan Gauch, Jianying Wang, and Satya Mahesh Rachakonda. A corpus analysis approach for automatic query expansion and its extension to multiple databases. ACM Transactions on Information Systems, 17(3):250, July 1999. CODEN ATISET. ISSN 1046-

8188. URL http://www.
acm.org/pubs/citations/
journals/tois/1999-17-3/
p250-gauch/.

Ge:2014:CAC

[GXTL14] Yong Ge, Hui Xiong, Alexander Tuzhilin, and Qi Liu. Cost-aware collaborative filtering for travel tour recommendations. ACM Transactions on Information Systems, 32(1):4:1–4:??, January 2014. CODEN ATISET. ISSN 1046-8188.

Guo:2022:HHE

[GYC^{+}22] Lei Guo, Hongzhi Yin, Tong Chen, Xiangliang Zhang, and Kai Zheng. Hierarchical hyperedge embeddingbased representation learning for group recommendation. ACM Transactions on Information Systems, 40(1): 3:1–3:27, January 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3457949.

Gui:2025:IPV

[GYW^{+}25] Xiaoqiang Gui, Guoxian Yu, Jun Wang, Shuguang Han, Qingzhong Li, Yongqing Zheng, and Wei Wang. Interaction privacy vulnerability in federated recommendation and lossless countermeasure. ACM Transactions on Information Systems, 43(5):134:1– 134:31, September 2025. CO-DEN ATISET. ISSN 1046-8188.

Guting:1989:ASO

[GZC89] Ralf Hartmut Guting, Roberto Zicari, and David M. Choy. An algebra for structured office documents. ACM Transactions on Information Systems, 7(2):123–157, April 1989. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Gao:2024:SEA

[GZL^{+}24] Jingtong Gao, Xiangyu Zhao, Muyang Li, Minghao Zhao, Runze Wu, Ruocheng Guo, Yiding Liu, and Dawei Yin. SMLP4Rec: an efficient All-MLP architecture for sequential recommendations. ACM Transactions on Information Systems, 42(3):86:1– 86:??, May 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3637871.

Guo:2018:CIP

[GZW^{+}18] Long Guo, Dongxiang Zhang, Yuan Wang, Huayu Wu, Bin Cui, and Kian-Lee Tan. CO 2: Inferring personal interests from raw footprints by connecting the offline world with the online world. ACM Transactions on Information Systems, 36(3):31:1–31:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Gao:2024:CIR

[GZW^{+}24] Chen Gao, Yu Zheng, Wenjie Wang, Fuli Feng, Xiangnan He, and Yong Li. Causal inference in recommender systems: a survey and future directions. ACM Transactions on Information Systems, 42(4): 88:1–88:??, July 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3639048.

Huston:2014:IWS

[HCC14] Samuel Huston, J. Shane Culpepper, and W. Bruce Croft. Indexing word sequences for ranked retrieval. ACM Transactions on Information Systems, 32(1):3:1– 3:??, January 2014. CODEN ATISET. ISSN 1046-8188.

Hu:2016:LIP

[HCC^{+}16] Liang Hu, Longbing Cao, Jian Cao, Zhiping Gu, Guandong Xu, and Dingyu Yang. Learning informative priors from heterogeneous domains to improve recommendation in cold-start user domains. ACM Transactions on Information Systems, 35(2):13:1– 13:??, December 2016. CO-DEN ATISET. ISSN 1046-8188.

Hu:2017:IQR

[HCC^{+}17] Liang Hu, Longbing Cao, Jian Cao, Zhiping Gu, Guandong Xu, and Jie Wang. Improving the quality of recommendations for users and items in the tail of distribution. ACM Transactions on Information Systems, 35(3):25:1– 25:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Han:2023:DDA

[HCD^{+}23] Lei Han, Tianwa Chen, Gianluca Demartini, Marta Indulska, and Shazia Sadiq. A data-driven analysis of behaviors in data curation processes. ACM Transactions on Information Systems, 41(3): 72:1–72:??, July 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3567419.

Huang:2020:NHN

[HCY^{+}20] Jie Huang, Chuan Chen, Fanghua Ye, Weibo Hu, and Zibin Zheng. Nonuniform hyper-network embedding with dual mechanism. ACM Transactions on Information Systems, 38(3):28:1– 28:18, June 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3388924.

Huang:2004:AAR

[HCZ04] Zan Huang, Hsinchun Chen, and Daniel Zeng. Applying associative retrieval techniques to alleviate the sparsity problem in collaborative filtering. ACM Transactions on Information Systems, 22(1):116– 142, January 2004. CODEN ATISET. ISSN 1046-8188.

Hicks:2008:OMP

[HDPM08] B. J. Hicks, A. Dong, R. Palmer, and H. C. Mcalpine. Organizing and managing personal electronic files: a mechanical engineer’s perspective. ACM Transactions on Information Systems, 26(4):23:1–23:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Hart:1991:ION

[HE91] Paul Hart and Deborah Estrin. Inter-organization networks, computer integration, and shifts in interdependence:

The case of the semiconductor industry. ACM Transactions on Information Systems, 9(4): 370–398, October 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Hammainen:1990:DFM

[HEA90] Heikki Hammainen, Eero Eloranta, and Jari Alasuvanto. Distributed form management. ACM Transactions on Information Systems, 8(1): 50–76, January 1990. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Hamdi:2016:TTI

[HGBY16] Sana Hamdi, Alda Lopes Gancarski, Amel Bouzeghoub, and Sadok Ben Yahia. TI-SoN: Trust inference in trustoriented social networks. ACM Transactions on Information Systems, 34(3):17:1– 17:??, May 2016. CODEN ATISET. ISSN 1046-8188.

Huang:2010:MND

[HHC^{+}10] Zi Huang, Bo Hu, Hong Cheng, Heng Tao Shen, Hongyan Liu, and Xiaofang Zhou. Mining nearduplicate graph for clusterbased reranking of Web video search results. ACM Transactions on Information Systems, 28(4):22:1–22:??, November 2010. CODEN ATISET. ISSN 1046-8188.

He:2023:ERS

[HHC^{+}23] Liyang He, Zhenya Huang, Enhong Chen, Qi Liu, Shiwei Tong, Hao Wang, Defu Lian, and Shijin Wang. An efficient and robust semantic hashing framework for similar text search. ACM Transactions on Information Systems, 41 (4):91:1–91:??, October 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3570725.

Hong:2001:WPB

[HHWL01] Jason I. Hong, Jeffrey Heer, Sarah Waterson, and James A. Landay. WebQuilt: a proxybased approach to remote web usability testing. ACM Transactions on Information Systems, 19(3):263–285, July 2001. CODEN ATISET. ISSN 1046-8188.

Huang:2023:PGH

[HJZ^{+}23] Zhenya Huang, Binbin Jin, Hongke Zhao, Qi Liu, Defu Lian, Bao Tengfei, and Enhong Chen. Personal or general? A hybrid strategy with multi-factors for news recommendation. ACM Transactions on Information Systems, 41(2):44:1–44:??, April 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3555373.

Hoi:2009:SSB

[HJZL09] Steven C. H. Hoi, Rong Jin, Jianke Zhu, and Michael R. Lyu. Semisupervised SVM batch mode active learning with applications to image retrieval. ACM Transactions on Information Systems, 27(3): 16:1–16:??, May 2009. CO-DEN ATISET. ISSN 1046-8188.

Hauff:2021:CSR

[HKS^{+}21] Claudia Hauff, Julia Kiseleva, Mark Sanderson, Hamed Zamani, and Yongfeng Zhang. Conversational search and recommendation: Introduction to the special issue. ACM Transactions on Information Systems, 39(4):38:1– 38:6, October 2021. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3465272.

Herlocker:2004:ECF

[HKTR04] Jonathan L. Herlocker, Joseph A. Konstan, Loren G. Terveen, and John T. Riedl. Evaluating collaborative filtering recommender systems. ACM Transactions on Information Systems, 22(1):5–53, January 2004. CODEN ATISET. ISSN 1046-8188.

Harabagiu:2010:UTT

[HL10] Sanda Harabagiu and Finley Lacatusu. Using topic themes for multi-document summarization. ACM Transactions on Information Systems, 28(3):13:1–13:??, June 2010. CODEN ATISET. ISSN 1046-8188.

He:2017:MEB

[HL17] Jiangning He and Hongyan Liu. Mining exploratory behavior to improve mobile app recommendations. ACM Transactions on Information Systems, 35(4):32:1– 32:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Huang:2020:LFD

[HLC^{+}20] Zhenya Huang, Qi Liu, Yuying Chen, Le Wu, Keli Xiao, Enhong Chen, Haiping Ma, and Guoping Hu. Learning or forgetting? A dynamic approach for tracking the knowledge proficiency of students. ACM Transactions on Information Systems, 38(2):19:1– 19:33, March 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3379507.

He:2025:TTB

[HLH^{+}25] Xi He, Yilin Liu, Weikang He, Xin Xing, Xingyu Lu, and Yanbing Liu. TCKT: Treebased cross-domain knowledge transfer for next POI cold-start recommendation. ACM Transactions on Information Systems, 43(5):119:1– 119:31, September 2025. CO-DEN ATISET. ISSN 1046-8188.

Hou:2017:LAC

[HLL^{+}17] Lei Hou, Juanzi Li, Xiao-Li Li, Jie Tang, and Xiaofei Guo. Learning to align comments to news topics. ACM Transactions on Information Systems, 36(1):9:1–9:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Huang:2025:RAA

[HLL^{+}25] Xu Huang, Jianxun Lian, Yuxuan Lei, Jing Yao, Defu Lian, and Xing Xie. Recommender AI agent: Integrating large language models for interactive recommendations. ACM Transactions on Information Systems, 43(4):96:1– 96:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Hicks:1998:HVC

[HLNS98] David L. Hicks, John J. Leggett, Peter J. Nurnberg, and John L. Schnase. A hypermedia version control framework. ACM Transactions on Information Systems, 16(2):127–160, April 1998. CODEN ATISET. ISSN 1046-8188.

He:2024:MMA

[HLW^{+}24] Liangliang He, Xiao Li, Pancheng Wang, Jintao Tang, and Ting Wang. MAN: Memory-augmented attentive networks for deep learningbased knowledge tracing. ACM Transactions on Information Systems, 42(1):8:1– 8:??, January 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3589340.

Hu:2024:DPD

[HLX^{+}24] Kaixi Hu, Lin Li, Qing Xie, Jianquan Liu, Xiaohui Tao, and Guandong Xu. Decoupled progressive distillation for sequential prediction with interaction dynamics. ACM Transactions on Information Systems, 42(3): 72:1–72:??, May 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3632403.

Hudson:1990:ISF

[HM90] Scott E. Hudson and Shamim P.

Mohamed. Interactive specification of flexible user interface displays. ACM Transactions on Information Systems, 8(3):269–288, July 1990. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Huang:2023:PET

[HML^{+}23] Liwei Huang, Yutao Ma, Yanbo Liu, Bohong Danny Du, Shuliang Wang, and Deyi Li. Position-enhanced and time-aware graph convolutional network for sequential recommendations. ACM Transactions on Information Systems, 41(1):6:1– 6:??, January 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3511700.

Huynh:2025:CUF

[HNN^{+}25] Thanh Trung Huynh, Trong Bang Nguyen, Thanh Toan Nguyen, Phi Le Nguyen, Hongzhi Yin, Quoc Viet Hung Nguyen, and Thanh Tam Nguyen. Certified unlearning for federated recommendation. ACM Transactions on Information Systems, 43(2):48:1–48:??, March 2025. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3706419.

He:2007:SHP

[HO07] Ben He and Iadh Ounis. On setting the hyper-parameters of term frequency normalization for information retrieval. ACM Transactions on Information Systems, 25(3):13:1– 13:??, July 2007. CODEN ATISET. ISSN 1046-8188.

Hofmann:2004:LSM

[Hof04] Thomas Hofmann. Latent semantic models for collaborative filtering. ACM Transactions on Information Systems, 22(1):89–115, January 2004. CODEN ATISET. ISSN 1046-8188.

Huang:2017:ESK

[HQZ17] Minlie Huang, Qiao Qian, and Xiaoyan Zhu. Encoding syntactic knowledge in neural networks for sentiment classification. ACM Transactions on Information Systems, 35(3):26:1–26:??, June 2017. CODEN ATISET. ISSN 1046-8188.

He:2022:GTU

[HRY^{+}22a] Xiangnan He, Zhaochun Ren, Emine Yilmaz, Marc Najork, and Tat-Seng Chua. Graph technologies for user modeling and recommendation: Introduction to the special issue — Part 1. ACM Transactions on Information Systems, 40 (2):21:1–21:5, April 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3477596.

He:2022:ISS

[HRY^{+}22b] Xiangnan He, Zhaochun Ren, Emine Yilmaz, Marc Najork, and Tat-Seng Chua. Introduction to the special section on graph technologies for user modeling and recommendation, Part 2. ACM Transactions on Information Systems, 40(3):42:1–42:5, July 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3490180.

Hartson:1990:UUO

[HSH90] H. Rex Hartson, Antonio C.

Siochi, and Deborah Hix. The UAN: a user-oriented representation for direct manipulation interface designs. ACM Transactions on Information Systems, 8(3):181–203, July 1990. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Hindus:1993:CSR

[HSH93] Debby Hindus, Chris Schmandt, and Chris Horner. Capturing, structuring, and representing ubiquitous audio. ACM Transactions on Information Systems, 11(4):376–400, October 1993. CODEN ATISET. ISSN 1046-8188. URL http:/

                                        /www.acm.org:80. Special Is-

sue on Computer-Supported Cooperative Work (CSCW).

Huang:2009:BCS

[HSS^{+}09] Zi Huang, Heng Tao Shen, Jie Shao, Xiaofang Zhou, and Bin Cui. Bounded coordinate system indexing for real-time video clip search. ACM Transactions on Information Systems, 27(3):17:1–17:??, May 2009. CODEN ATISET. ISSN 1046-8188.

Huang:2022:LLC

[HSYX22] Xiaowen Huang, Jitao Sang, Jian Yu, and Changsheng Xu. Learning to learn a coldstart sequential recommender. ACM Transactions on Information Systems, 40(2):30:1– 30:25, April 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3466753.

Hawking:1999:MIS

[HT99] David Hawking and Paul Thistlewaite. Methods for information server selection. ACM Transactions on Information Systems, 17(1):40– 76, January 1999. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tois/1999-17-1/
                                        p40-hawking/.

Hofmann:2013:FSE

[HWD13] Katja Hofmann, Shimon Whiteson, and Maarten De Rijke. Fidelity, soundness, and efficiency of interleaved comparison methods. ACM Transactions on Information Systems, 31(4):17:1–17:??, November 2013. CODEN ATISET. ISSN 1046-8188.

Hu:2024:SCL

[HWL^{+}24] Yupeng Hu, Kun Wang, Meng Liu, Haoyu Tang, and Liqiang Nie. Semantic collaborative learning for crossmodal moment localization. ACM Transactions on Information Systems, 42(2):50:1– 50:??, March 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3620669.

Huang:2019:QTQ

[HWN^{+}19] Heyan Huang, Xiaochi Wei, Liqiang Nie, Xianling Mao, and Xin-Shun Xu. From question to text: Questionoriented feature attention for answer selection. ACM Transactions on Information Systems, 37(1):6:1–6:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Hu:2025:HTM

[HWT^{+}25] Shirui Hu, Weichang Wu, Zuoli Tang, Zhaoxin Huan, Lin Wang, Xiaolu Zhang, Jun Zhou, Lixin Zou, and Chenliang Li. HORAE: Temporal multi-interest pre-training for sequential recommendation. ACM Transactions on Information Systems, 43(4): 88:1–88:??, July 2025. CO-DEN ATISET. ISSN 1046-8188.

Han:2020:GDR

[HY20] Jungkyu Han and Hayato Yamana. Geographic diversification of recommended POIs in frequently visited areas. ACM Transactions on Information Systems, 38 (1):1:1–1:??, February 2020. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3362505.

Han:2015:USC

[HYH15] Shuguang Han, Zhen Yue, and Daqing He. Understanding and supporting cross-device Web search for exploratory tasks with mobile touch interactions. ACM Transactions on Information Systems, 33(4):16:1–16:??, May 2015. CODEN ATISET. ISSN 1046-8188.

Huang:2024:DLR

[HYLC24] Heyan Huang, Changsen Yuan, Qian Liu, and Yixin Cao. Document-level relation extraction via separate relation representation and logical reasoning. ACM Transactions on Information Systems, 42 (1):22:1–22:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3597610.

Huang:2025:SHL

[HYM^{+}25] Lei Huang, Weijiang Yu, Weitao Ma, Weihong Zhong, Zhangyin Feng, Haotian Wang, Qianglong Chen, Weihua Peng, Xiaocheng Feng, Bing Qin, and Ting Liu. A survey on hallucination in large language models: Principles, taxonomy, challenges, and open questions. ACM Transactions on Information Systems, 43(2):42:1– 42:??, March 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3703155.

Hao:2023:MSB

[HYZ^{+}23] Bowen Hao, Hongzhi Yin, Jing Zhang, Cuiping Li, and Hong Chen. A multi-strategy-based pre-training method for cold-start recommendation. ACM Transactions on Information Systems, 41(2):31:1– 31:??, April 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3544107.

He:2023:ACF

[HZF^{+}23] Xiangnan He, Yang Zhang, Fuli Feng, Chonggang Song, Lingling Yi, Guohui Ling, and Yongdong Zhang. Addressing confounding feature issue for causal recommendation. ACM Transactions on Information Systems, 41(3): 53:1–53:??, July 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3559757.

Huang:2020:CBI

[HZG20] Minlie Huang, Xiaoyan Zhu, and Jianfeng Gao. Challenges in building intelligent open-domain dialog systems. ACM Transactions on Information Systems, 38(3):21:1– 21:32, June 2020. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3383123.

Hu:2025:EER

[HZS^{+}25] Linmei Hu, Xinyu Zhang, Dandan Song, Changzhi Zhou, Hongyu He, and Liqiang Nie. Efficient and effective role player: a compact knowledge-grounded personabased dialogue model enhanced by LLM distillation. ACM Transactions on Information Systems, 43(3):59:1– 59:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Hou:2013:MPH

[HZSL13] Yuexian Hou, Xiaozhao Zhao, Dawei Song, and Wenjie Li. Mining pure high-order word associations via information geometry for information retrieval. ACM Transactions on Information Systems, 31(3): 12:1–12:??, July 2013. CO-DEN ATISET. ISSN 1046-8188.

Heinz:2002:BTF

[HZW02] Steffen Heinz, Justin Zobel, and Hugh E. Williams.

Burst tries: a fast, efficient data structure for string keys. ACM Transactions on Information Systems, 20(2):192– 223, 2002. CODEN ATISET. ISSN 1046-8188.

Ibrahim:2016:CPL

[IC16] Muhammad Ibrahim and Mark Carman. Comparing pointwise and listwise objective functions for randomforest-based learning-to-rank. ACM Transactions on Information Systems, 34(4):20:1– 20:??, September 2016. CO-DEN ATISET. ISSN 1046-8188.

Ipeirotis:2008:CAH

[IG08] Panagiotis G. Ipeirotis and Luis Gravano. Classificationaware hidden-web text database selection. ACM Transactions on Information Systems, 26 (2):6:1–6:??, March 2008. CO-DEN ATISET. ISSN 1046-8188.

Im:2007:DOS

[IH07] Il Im and Alexander Hars.

Does a one-size recommendation system fit all? the effectiveness of collaborative filtering based recommendation systems across different domains and search modes. ACM Transactions on Information Systems, 26(1):4:1– 4:??, November 2007. CO-DEN ATISET. ISSN 1046-

Ishii:1993:IIS

[IKG93] Hiroshi Ishii, Minoru Kobayashi, and Jonathan Grudin. Integration of interpersonal space and shared workspace; Clear-Board design and Experiments. ACM Transactions on Information Systems, 11(4): 349–375, October 1993. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.

org:80. Special Issue on Computer-Supported Cooperative Work (CSCW).

Ivory:2005:EWS

[IM05] Melody Y. Ivory and Rodrick Megraw. Evolution of Web site design patterns. ACM Transactions on Information Systems, 23(4):463–497, October 2005. CODEN ATISET. ISSN 1046-8188.

Isakowitz:1995:TLP

[ISL95] Tomás Isakowitz, Shimon Schocken, and Henry C. Lucas, Jr. Toward a logical/ physical theory of spreadsheet modeling. ACM Transactions on Information Systems, 13 (1):1–37, January 1995. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Ioannidis:1992:CLD

[ISW92] Yannis E. Ioannidis, Tomas Saulys, and Andrew J. Whitsitt. Conceptual learning in database design. ACM Transactions on Information Systems, 10(3):265–293, July 1992. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Imran:2023:RRE

[IYC^{+}23] Mubashir Imran, Hongzhi Yin, Tong Chen, Quoc Viet Hung Nguyen, Alexander Zhou, and Kai Zheng. Re-FRS: Resource-efficient federated recommender system for dynamic and diversified user preferences. ACM Transactions on Information Systems, 41(3):65:1–65:??, July 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3560486.

Jacob:1991:UEM

[Jac91] Robert J. K. Jacob. The use of eye movements in human-computer interaction techniques: What you look at is what you get. ACM Transactions on Information Systems, 9(2):152–169, April 1991. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Computer-Human Interaction.

Jensen:2007:RES

[JBCF07] Eric C. Jensen, Steven M.

Beitzel, Abdur Chowdhury, and Ophir Frieder. Repeatable evaluation of search services in dynamic environments. ACM Transactions on Information Systems, 26 (1):1:1–1:??, November 2007. CODEN ATISET. ISSN 1046-8188.

Jones:2007:TPQ

[JD07] Rosie Jones and Fernando Diaz. Temporal profiles of queries. ACM Transactions on Information Systems, 25 (3):14:1–14:??, July 2007. CODEN ATISET. ISSN 1046-8188.

Jayashree:2022:MWP

[JDA^{+}22] Srivatsa Ramesh Jayashree, Gaël Dias, Judith Jeyafreeda Andrew, Sriparna Saha, Fabrice Maurel, and Stéphane Ferrari. Multimodal Web page segmentation using self-organized multi-objective clustering. ACM Transactions on Information Systems, 40(3):59:1–59:49, July 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3480966.

Jin:2022:GLI

[JDZ^{+}22] Jiarui Jin, Kounianhua Du, Weinan Zhang, Jiarui Qin, Yuchen Fang, Yong Yu, Zheng Zhang, and Alexander J. Smola. GraphHINGE: Learning interaction models of structured neighborhood on heterogeneous information network. ACM Transactions on Information Systems, 40(3):45:1–45:35, July 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3472956.

Jafarzadeh:2025:KGE

[JEAZ25] Parastoo Jafarzadeh, Faezeh Ensan, Mahdiyar Ali Akbar Alavi, and Fattane Zarrinkalam. A knowledge graph embedding model for answering factoid entity questions. ACM Transactions on Information Systems, 43(2): 38:1–38:??, March 2025. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3678003.

Joachims:2007:EAI

[JGP^{+}07] Thorsten Joachims, Laura Granka, Bing Pan, Helene Hembrooke, Filip Radlinski, and Geri Gay. Evaluating the accuracy of implicit feedback from clicks and query reformulations in Web search. ACM Transactions on Information Systems, 25(2):7:1– 7:??, April 2007. CODEN ATISET. ISSN 1046-8188.

Jarvelin:2002:CGB

[JK02] Kalervo Järvelin and Jaana Kekäläinen. Cumulated gainbased evaluation of IR techniques. ACM Transactions on Information Systems, 20 (4):422–446, 2002. CODEN ATISET. ISSN 1046-8188.

Jing:2025:EAP

[JLC^{+}25] Erkang Jing, Yezheng Liu, Yidong Chai, Shuo Yu, Longshun Liu, Yuanchun Jiang, and Yang Wang. Emotion-aware personalized music recommendation with a heterogeneity-aware deep Bayesian network. ACM Transactions on Information Systems, 43(5):124:1–124:43, September 2025. CODEN ATISET. ISSN 1046-8188.

Jagerman:2020:SEO

[JMD20] Rolf Jagerman, Ilya Markov, and Maarten De Rijke. Safe exploration for optimizing contextual bandits. ACM Transactions on Information Systems, 38(3):24:1–24:23, June 2020. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3385670.

Jadidinejad:2022:SPO

[JMO22] Amir H. Jadidinejad, Craig Macdonald, and Iadh Ounis. The Simpson’s Paradox in the offline evaluation of recommendation systems. ACM Transactions on Information Systems, 40(1): 4:1–4:22, January 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3458509.

Jarke:1992:DEE

[JMSV92] M. Jarke, J. Mylopoulos, J. W. Schmidt, and Y. Vassiliou. DAIDA: An environment for evolving information systems. ACM Transactions on Information Systems, 10 (1):1–50, January 1992. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Ji:2022:SAL

[JMZ22] Weiyu Ji, Xiangwu Meng, and Yujie Zhang. STARec:

Adaptive learning with spatiotemporal and activity influence for POI recommendation. ACM Transactions on Information Systems, 40 (4):65:1–65:40, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3485631.

Jarvelin:2024:BIE

[JS24] Kalervo Jarvelin and Eero Sormunen. A blueprint of IR evaluation integrating task and user characteristics. ACM Transactions on Information Systems, 42(6):164:1–164:??, November 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3675162.

Jungclaus:1996:TLO

[JSHS96] Ralf Jungclaus, Gunter Saake, Thorsten Hartmann, and Cristina Sernadas. TROLL — A language for objectoriented specification of information systems. ACM Transactions on Information Systems, 14(2):175– 211, April 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/hartmann.
html.

Jing:2024:SDT

[JSL^{+}24] Liqiang Jing, Xuemeng Song, Xuming Lin, Zhongzhou Zhao, Wei Zhou, and Liqiang Nie. Stylized data-to-text generation: a case study in the e-commerce domain.

ACM Transactions on Information Systems, 42(1):25:1– 25:??, January 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3603374.

Ji:2023:CSD

[JSZL23] Yitong Ji, Aixin Sun, Jie Zhang, and Chenliang Li. A critical study on data leakage in recommender system offline evaluation. ACM Transactions on Information Systems, 41(3):75:1–75:??, July 2023. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3569930.

Jiang:2017:GGS

[JTL^{+}17] Jiawei Jiang, Yunhai Tong, Hua Lu, Bin Cui, Kai Lei, and Lele Yu. GVoS: a general system for near-duplicate video-related applications on storm. ACM Transactions on Information Systems, 36(1): 3:1–3:??, August 2017. CO-DEN ATISET. ISSN 1046-8188.

Jiang:2016:CLT

[JTS16] Di Jiang, Yongxin Tong, and Yuanfeng Song. Cross-lingual topic discovery from multilingual search engine query log. ACM Transactions on Information Systems, 35(2):9:1– 9:??, December 2016. CO-DEN ATISET. ISSN 1046-8188.

Jarvelin:2015:TBI

[JVA^{+}15] Kalervo Järvelin, Pertti Vakkari, Paavo Arvola, Feza Baskaya, Anni Järvelin, Jaana Kekäläinen, Heikki Keskustalo, Sanna Kumpulainen, Miamaria Saastamoinen, Reijo Savolainen, and Eero Sormunen. Taskbased information interaction evaluation: The viewpoint of program theory. ACM Transactions on Information Systems, 33(1):3:1–3:??, March 2015. CODEN ATISET. ISSN 1046-8188.

Jian:2025:GAS

[JWX^{+}25] Meng Jian, Tuo Wang, Zhuoyang Xia, Ge Shi, Richang Hong, and Lifang Wu. Geometric-augmented self-distillation for graphbased recommendation. ACM Transactions on Information Systems, 43(4):93:1–93:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Jiang:2024:TFM

[JXY^{+}24] Yiheng Jiang, Yuanbo Xu, Yongjian Yang, Funing Yang, Pengyang Wang, Chaozhuo Li, Fuzhen Zhuang, and Hui Xiong. TriMLP: a foundational MLP-like architecture for sequential recommendation. ACM Transactions on Information Systems, 42(6): 157:1–157:??, November 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3670995.

Jia:2013:ISD

[JYM13] Lifeng Jia, Clement Yu, and Weiyi Meng. The impacts of structural difference and temporality of tweets on retrieval effectiveness. ACM Transactions on Information Systems, 31(4):21:1–21:??, November 2013. CODEN ATISET. ISSN 1046-8188.

Jiang:2006:ECR

[JZ06] Jing Jiang and Chengxiang Zhai. Extraction of coherent relevant passages using hidden Markov models. ACM Transactions on Information Systems, 24(3):295–319, July 2006. CODEN ATISET. ISSN 1046-8188.

Jing:2024:CSS

[JZZW24] Mengyuan Jing, Yanmin Zhu, Tianzi Zang, and Ke Wang. Contrastive self-supervised learning in recommender systems: a survey. ACM Transactions on Information Systems, 42(2):59:1–59:??, March 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3627158.

Klein:1989:STR

[KBD89] Shmuel T. Klein, Abraham Bookstein, and Scott Deerwester. Storing text retrieval systems on CD-ROM. compression and encryption considerations. ACM Transactions on Information Systems, 7(3):230–245, July 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Kim:1991:DOO

[KBGW91] Won Kim, Nat Ballou, Jorge F. Garza, and Darrell Woelk. A distributed objectoriented database system supporting shared and private databases. ACM Transactions on Information Systems, 9(1): 31–51, January 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Krovetz:1992:LAI

[KC92] Robert Krovetz and W. Bruce Croft. Lexical ambiguity and information retrieval. ACM Transactions on Information Systems, 10(2):115–141, April 1992. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Kong:1995:DDI

[KC95] Q. Kong and G. Chen. On deductive databases with incomplete information. ACM Transactions on Information Systems, 13(3):354–369, July 1995. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Kulkarni:2015:SSE

[KC15] Anagha Kulkarni and Jamie Callan. Selective search: Efficient and effective search of large textual collections. ACM Transactions on Information Systems, 33(4):17:1– 17:??, May 2015. CODEN ATISET. ISSN 1046-8188.

Keller:1995:XAI

[KEL95] Ralf Keller, Wolfgang Effelsberg, and Bernd Lamparter. XMovie: Architecture and implementation of a distributed movie system. ACM Transactions on Information Systems, 13(4):471–499, October 1995. CODEN ATISET. ISSN 1046-8188. URL http:/

/www.acm.org:80. Special Is-

sue on Video Information Retrieval.

Kwok:2001:SQA

[KEW01] Cody Kwok, Oren Etzioni, and Daniel S. Weld. Scaling question answering to the Web. ACM Transactions on Information Systems, 19(3): 242–262, July 2001. CODEN ATISET. ISSN 1046-8188.

Kelly:2010:EPN

[KFS10] Diane Kelly, Xin Fu, and Chirag Shah. Effects of position and number of relevant documents retrieved on users’ evaluations of system performance. ACM Transactions on Information Systems, 28 (2):9:1–9:??, May 2010. CO-DEN ATISET. ISSN 1046-8188.

Kim:2020:ETM

[KJA20] Youngwoo Kim, Myungha Jang, and James Allan. Explaining text matching on neural natural language inference. ACM Transactions on Information Systems, 38 (4):39:1–39:23, October 2020. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3418052.

Krikon:2010:UIP

[KKB10] Eyal Krikon, Oren Kurland, and Michael Bendersky. Utilizing inter-passage and inter-document similarities for reranking search results. ACM Transactions on Information Systems, 29(1): 3:1–3:??, December 2010. CO-DEN ATISET. ISSN 1046-8188.

Kerne:2008:CMI

[KKS^{+}08] Andruid Kerne, Eunyee Koh, Steven M. Smith, Andrew Webb, and Blake Dworaczyk. combinFormation: Mixedinitiative composition of image and text surrogates promotes information discovery. ACM Transactions on Information Systems, 27(1):5:1– 5:??, December 2008. CO-DEN ATISET. ISSN 1046-8188.

Kacmar:1991:PPO

[KL91] Charles J. Kacmar and John J. Leggett. PROXHY: a process-oriented extensible hypertext architecture. ACM Transactions on Information Systems, 9(4):399–419, October 1991. CODEN ATISET.

ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Katzenstein:2000:BSO

[KL00] Gary Katzenstein and F. Javier Lerch. Beneath the surface of organizational processes: a social representation framework for business process redesign. ACM Transactions on Information Systems, 18(4):383–422,

                                        2000. URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tois/2000-18-4/
                                        p383-katzenstein/.

Kazai:2006:ECG

[KL06] Gabriella Kazai and Mounia Lalmas. eXtended cumulated gain measures for the evaluation of content-oriented XML retrieval. ACM Transactions on Information Systems, 24(4):503–542, October 2006. CODEN ATISET. ISSN 1046-8188.

Kurland:2009:CLM

[KL09] Oren Kurland and Lillian Lee. Clusters, language models, and ad hoc information retrieval. ACM Transactions on Information Systems, 27(3): 13:1–13:??, May 2009. CO-DEN ATISET. ISSN 1046-8188.

Kurland:2010:PHS

[KL10] Oren Kurland and Lillian Lee. PageRank without hyperlinks: Structural reranking using links induced by language models. ACM Transactions on Information Systems, 28(4):18:1–18:??, November 2010. CODEN ATISET. ISSN 1046-8188.

Kim:2013:EDE

[KLHK13] Jinhan Kim, Sanghoon Lee, Seung-Won Hwang, and Sunghun Kim. Enriching documents with examples: a corpus mining approach. ACM Transactions on Information Systems, 31(1):1:1–1:??, January 2013. CODEN ATISET. ISSN 1046-8188.

Kling:1994:ISI

[Kli94] R. Kling. Introduction to the special issue on social science perspectives on IS. ACM Transactions on Information Systems, 12(2):117–118, April 1994. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Kimbrough:1997:AMP

[KM97] Steven O. Kimbrough and Scott A. Moore. On automated message processing in electronic commerce and work support systems: Speech act theory and expressive felicity. ACM Transactions on Information Systems, 15(4): 321–367, October 1997. CO-DEN ATISET. ISSN 1046-8188.

Ke:2013:SCP

[KM13] Weimao Ke and Javed Mostafa.

Studying the clustering paradox and scalability of search in highly distributed environments. ACM Transactions on Information Systems, 31(2): 8:1–8:??, May 2013. CODEN ATISET. ISSN 1046-8188.

Kamps:2006:AIN

[KMdRS06] Jaap Kamps, Maarten Marx, Maarten de Rijke, and Börkur Sigurbjörnsson. Articulating information needs in XML query languages. ACM Transactions on Information Systems, 24(4):407–436, October 2006. CODEN ATISET. ISSN 1046-8188.

Kataoka:1992:MIO

[KMKO92] Yutaka Kataoka, Masato Morisaki, Hiroshi Kuribayashi, and Hiroyoshi Ohara. A model for input and output of multilingual text in a windowing environment. ACM Transactions on Information Systems, 10(4):438–451, October 1992. CODEN ATISET. ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Kiesel:2021:MIC

[KMPS21] Johannes Kiesel, Lars Meyer, Martin Potthast, and Benno Stein. Meta-information in conversational search. ACM Transactions on Information Systems, 39(4):50:1–50:44, October 2021. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3468868.

King:1993:DDI

[KN93] Roger King and Michael Novak. Designing database interfaces with DBface. ACM Transactions on Information Systems, 11(2):105–132, April 1993. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Konow:2017:IT

[KNCLO17] Roberto Konow, Gonzalo Navarro, Charles L. A. Clarke, and Alejandro López-Ortíz. Inverted treaps. ACM Transactions on Information Systems, 35(3):22:1–22:??, June 2017. CODEN ATISET. ISSN 1046-8188.

King:2004:DCB

[KNS04] Irwin King, Cheuk Hang Ng, and Ka Cheung Sia. Distributed content-based visual information retrieval system on peer-to-peer networks. ACM Transactions on Information Systems, 22(3):477– 501, July 2004. CODEN ATISET. ISSN 1046-8188.

Kolda:1998:SMD

[KO98] Tamara G. Kolda and Dianne P. O’Leary. A semidiscrete matrix decomposition for latent semantic indexing information retrieval. ACM Transactions on Information Systems, 16(4):322–346, October 1998. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/pubs/citations/journals/

                                        tois/1998-16-4/p322-kolda/

.

Koike:1993:RAS

[Koi93] Hideki Koike. The role of another spatial dimension in software visualization. ACM Transactions on Information Systems, 11(3):266–286, July 1993. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Virtual Worlds.

Koike:1995:FVF

[Koi95] Hideki Koike. Fractal views: a fractal-based method for controlling information display. ACM Transactions on Information Systems, 13(3): 305–323, July 1995. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Konstan:2004:IRS

[Kon04] Joseph A. Konstan. Introduction to recommender systems: Algorithms and evaluation. ACM Transactions on Information Systems, 22(1): 1–4, January 2004. CODEN ATISET. ISSN 1046-8188.

Ko:2010:PMA

[KSNM10] Jeongwoo Ko, Luo Si, Eric Nyberg, and Teruko Mitamura. Probabilistic models for answer-ranking in multilingual question-answering. ACM Transactions on Information Systems, 28(3):16:1– 16:??, June 2010. CODEN ATISET. ISSN 1046-8188.

Kharazmi:2016:EAW

[KSVS16] Sadegh Kharazmi, Falk Scholer, David Vallet, and Mark Sanderson. Examining additivity and weak baselines. ACM Transactions on Information Systems, 34(4):23:1– 23:??, September 2016. CO-DEN ATISET. ISSN 1046-8188.

Keshvari:2024:SDL

[KSYE24] Sanaz Keshvari, Farzan Saeedi, Hadi Sadoghi Yazdi, and Faezeh Ensan. A self-distilled learning to rank model for ad hoc retrieval. ACM Transactions on Information Systems, 42(6):166:1–166:??, November 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3681784.

Korodi:2005:ENM

[KT05] Gergely Korodi and Ioan Tabus. An efficient normalized maximum likelihood algorithm for DNA sequence compression. ACM Transactions on Information Systems, 23(1):3–34, January 2005. CODEN ATISET. ISSN 1046-8188.

Kwok:1990:ECT

[Kwo90] K. L. Kwok. Experiments with a component theory of probabilistic information retrieval based on single terms as document components. ACM Transactions on Information Systems, 8(4):363– 386, October 1990. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80.

Kwok:1995:NAP

[Kwo95] K. L. Kwok. A network approach to probabilistic information retrieval. ACM Transactions on Information Systems, 13(3):324–353, July 1995. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

Krasakis:2024:CEC

[KYK24] Antonios Minas Krasakis, Andrew Yates, and Evangelos Kanoulas. Contextualizing and expanding conversational queries without supervision. ACM Transactions on Information Systems, 42(3):77:1– 77:??, May 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3632622.

Kolbe:2010:ENN

[KZP10] Dashiell Kolbe, Qiang Zhu, and Sakti Pramanik. Efficient k-nearest neighbor searching in nonordered discrete data spaces. ACM Transactions on Information Systems, 28 (2):7:1–7:??, May 2010. CO-DEN ATISET. ISSN 1046-8188.

Kaszkiel:1999:EPR

[KZSD99] Marcin Kaszkiel, Justin Zobel, and Ron Sacks-Davis. Efficient passage ranking for document databases. ACM Transactions on Information Systems, 17(4):406–439, October 1999. CODEN ATISET. ISSN 1046-8188. URL

http://www.acm.org/pubs/
citations/journals/tois/
1999-17-4/p406-kaszkiel/.

Losada:2008:AMB

[LA08] David E. Losada and Leif Azzopardi. Assessing multivariate Bernoulli models for information retrieval. ACM Transactions on Information Systems, 26(3):17:1–17:??, June 2008. CODEN ATISET. ISSN 1046-8188.

Liu:2017:DUP

[LAL^{+}17] Xuanzhe Liu, Wei Ai, Huoran Li, Jian Tang, Gang Huang, Feng Feng, and Qiaozhu Mei. Deriving user preferences of mobile apps from their management activities. ACM Transactions on Information Systems, 35(4):39:1– 39:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Lalor:2024:SFM

[LAO^{+}24] John P. Lalor, Ahmed Abbasi, Kezia Oketch, Yi Yang, and Nicole Forsgren. Should fairness be a metric or a model? A model-based framework for assessing bias in machine learning pipelines. ACM Transactions on Information Systems, 42(4):99:1– 99:??, July 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3641276.

Lukasik:2019:GPR

[LBC^{+}19] Michal Lukasik, Kalina Bontcheva, Trevor Cohn, Arkaitz Zubiaga, Maria Liakata, and Rob Procter. Gaussian processes for rumour stance classification in social media. ACM Transactions on Information Systems, 37(2):20:1– 20:??, March 2019. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3295823.

Li:2024:FBS

[LBM^{+}24] Jun Li, Yi Bin, Yunshan Ma, Yang Yang, Zi Huang, and Tat-Seng Chua. Filter-based stance network for rumor verification. ACM Transactions on Information Systems, 42 (4):108:1–108:??, July 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3649462.

Lau:2008:TBR

[LBS08] Raymond Y. K. Lau, Peter D. Bruza, and Dawei Song. Towards a beliefrevision-based adaptive and context-sensitive information retrieval system. ACM Transactions on Information Systems, 26(2):8:1–8:??, March 2008. CODEN ATISET. ISSN 1046-8188.

Liu:2024:FTI

[LBZ^{+}24] Fei Liu, Chenyang Bu, Haotian Zhang, Le Wu, Kui Yu, and Xuegang Hu. FDKT: Towards an interpretable deep knowledge tracing via fuzzy reasoning. ACM Transactions on Information Systems, 42 (5):139:1–139:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3656167.

Lam:2007:NET

[LCH07] Wai Lam, Shing-Kit Chan, and Ruizhang Huang. Named entity translation matching and learning: With application for mining unseen translations. ACM Transactions on Information Systems, 25(1): 2:1–2:??, February 2007. CO-DEN ATISET. ISSN 1046-8188.

Li:2016:TBI

[LCJ^{+}16] Qing Li, Yuanzhu Chen, Li Ling Jiang, Ping Li, and Hsinchun Chen. A tensorbased information framework for predicting the stock market. ACM Transactions on Information Systems, 34(2): 11:1–11:??, April 2016. CO-DEN ATISET. ISSN 1046-8188.

Lu:2019:DBT

[LCJ^{+}19] Wei Lu, Fu-Lai Chung, Wenhao Jiang, Martin Ester, and Wei Liu. A deep Bayesian tensor-based system for video recommendation. ACM Transactions on Information Systems, 37(1): 7:1–7:??, January 2019. CO-DEN ATISET. ISSN 1046-8188.

Lu:2004:ATM

[LCL04] Wen-Hsiang Lu, Lee-Feng Chien, and Hsi-Jian Lee. Anchor text mining for translation of Web queries: A transitive translation approach. ACM Transactions on Information Systems, 22(2):242– 269, April 2004. CODEN ATISET. ISSN 1046-8188.

Liu:2017:DBE

[LCL^{+}17] Ming Liu, Lei Chen, Bingquan Liu, Guidong Zheng, and Xiaoming Zhang. DBpediabased entity linking via greedy search and adjusted Monte Carlo random walk. ACM Transactions on Information Systems, 36(2):16:1– 16:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Ling:2023:GRI

[LCL^{+}23a] Yanxiang Ling, Fei Cai, Jun Liu, Honghui Chen, and Maarten de Rijke. Generating relevant and informative questions for open-domain conversations. ACM Transactions on Information Systems, 41(1):2:1–2:??, January 2023. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3510612.

Liu:2023:BSI

[LCL^{+}23b] Dugang Liu, Pengxiang Cheng, Zinan Lin, Xiaolian Zhang, Zhenhua Dong, Rui Zhang, Xiuqiang He, Weike Pan, and Zhong Ming. Bounding system-induced biases in recommender systems with a randomized dataset. ACM Transactions on Information Systems, 41(4):108:1–108:??, October 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3582002.

Long:2023:DCL

[LCNY23] Jing Long, Tong Chen, Quoc Viet Hung Nguyen, and Hongzhi Yin. Decentralized collaborative learning framework for next POI recommendation. ACM Transactions on Information Systems, 41(3): 66:1–66:??, July 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3555374.

Li:2019:PRP

[LCPD19] Xinyi Li, Yifan Chen, Benjamin Pettit, and Maarten De Rijke. Personalised reranking of paper recommendations using paper content and user behavior. ACM Transactions on Information Systems, 37(3):31:1–31:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3312528.

Lan:2024:EDR

[LCW^{+}24a] Tian Lan, Deng Cai, Yan Wang, Yixuan Su, Heyan Huang, and Xian-Ling Mao. Exploring dense retrieval for dialogue response selection. ACM Transactions on Information Systems, 42(3):84:1– 84:??, May 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3632750.

Li:2024:MML

[LCW^{+}24b] Shuzhe Li, Wei Chen, Bin Wang, Chao Huang, Yanwei Yu, and Junyu Dong. MCN4Rec: Multi-level collaborative neural network for next location recommendation. ACM Transactions on Information Systems, 42(4): 105:1–105:??, July 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3643669.

Li:2019:SGT

[LCX^{+}19] Chenliang Li, Shiqian Chen, Jian Xing, Aixin Sun, and Zongyang Ma. Seed-guided topic model for document filtering and classification. ACM Transactions on Information Systems, 37(1):9:1–9:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Lipani:2021:HDE

[LCY21] Aldo Lipani, Ben Carterette, and Emine Yilmaz. How am I doing?: Evaluating conversational search systems offline. ACM Transactions on Information Systems, 39(4):51:1– 51:22, October 2021. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3451160.

Liang:2025:LEG

[LCYY25] Xurong Liang, Tong Chen, Wei Yuan, and Hogzhi Yin. Lightweight embeddings with graph rewiring for collaborative filtering. ACM Transactions on Information Systems, 43(4):108:1–108:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Larson:2012:SIS

[LdJKR12] Martha Larson, Franciska de Jong, Wessel Kraaij, and Steve Renals. Special issue on searching speech. ACM Transactions on Information Systems, 30(3):15:1– 15:??, August 2012. CODEN ATISET. ISSN 1046-8188.

Liu:2025:ESP

[LDN^{+}25] Jiongnan Liu, Zhicheng Dou, Jian-Yun Nie, Zhenlin Chen, Guoyu Tang, Sulong Xu, and Ji-Rong Wen. Enhancing sequential personalized product search with external out-of-sequence knowledge. ACM Transactions on Information Systems, 43(4): 84:1–84:??, July 2025. CO-DEN ATISET. ISSN 1046-8188.

Li:2017:ETM

[LDW^{+}17] Chenliang Li, Yu Duan, Haoran Wang, Zhiqian Zhang, Aixin Sun, and Zongyang Ma. Enhancing topic modeling for short texts with auxiliary word embeddings. ACM Transactions on Information Systems, 36(2):11:1– 11:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Lin:2025:HCR

[LDX^{+}25] Jianghao Lin, Xinyi Dai, Yunjia Xi, Weiwen Liu, Bo Chen, Hao Zhang, Yong Liu, Chuhan Wu, Xiangyang Li, Chenxu Zhu, Huifeng Guo, Yong Yu, Ruiming Tang, and Weinan Zhang. How can recommender systems benefit from large language models: a survey. ACM Transactions on Information Systems, 43(2):28:1–28:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3678004.

Lehtonen:2006:PHX

[Leh06] Miro Lehtonen. Preparing heterogeneous XML for fulltext search. ACM Transactions on Information Systems, 24(4):455–474, October 2006. CODEN ATISET. ISSN 1046-8188.

Li:2025:RCD

[LFL^{+}25] Bobo Li, Hao Fei, Fei Li, Shengqiong Wu, Lizi Liao, Yinwei Wei, Tat seng Chua, and Donghong Ji. Revisiting conversation discourse for dialogue disentanglement. ACM Transactions on Information Systems, 43(1):15:1– 15:??, January 2025. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3698191.

Liu:2023:MPB

[LFW23] Zhongzhou Liu, Yuan Fang, and Min Wu. Mitigating popularity bias for users and items with fairness-centric adaptive recommendation. ACM Transactions on Information Systems, 41(3):55:1– 55:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3564286.

Li:2024:BRF

[LGL^{+}24] Hanzhe Li, Jingjing Gu, Xinjiang Lu, Dazhong Shen, Yuting Liu, YaNan Deng, Guoliang Shi, and Hui Xiong. Beyond relevance: Factor-level causal explanation for user travel decisions with counterfactual data augmentation. ACM Transactions on Information Systems, 42(5):128:1– 128:??, September 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3653673.

Liu:2024:AMM

[LGM^{+}24] Qi Liu, Gang Guo, Jiaxin Mao, Zhicheng Dou, Ji-Rong Wen, Hao Jiang, Xinyu Zhang, and Zhao Cao. An analysis on matching mechanisms and token pruning for late-interaction models. ACM Transactions on Information Systems, 42(5):118:1–118:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3639818.

Liu:2019:MVB

[LGQL19] Ming Liu, Gu Gong, Bing Qin, and Ting Liu. A multiview-based collective entity linking method. ACM Transactions on Information Systems, 37(2):23:1–23:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3300197.

Levi:2018:SCP

[LGRK18] Or Levi, Ido Guy, Fiana Raiber, and Oren Kurland. Selective cluster presentation on the search results page. ACM Transactions on Information Systems, 36(3):28:1– 28:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Lu:2025:FSL

[LGY^{+}25] Ziang Lu, Lei Guo, Xu Yu, Zhiyong Cheng, Xiaohui Han, and Lei Zhu. Federated semantic learning for privacypreserving cross-domain recommendation. ACM Transactions on Information Systems, 43(4):103:1–103:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Liu:2024:MPP

[LGZ^{+}24] Hao Liu, Lei Guo, Lei Zhu, Yongqiang Jiang, Min Gao, and Hongzhi Yin. MCRPL: a pretrain, prompt, and fine-tune paradigm for non-overlapping many-to-one cross-domain recommendation. ACM Transactions on Information Systems, 42(4): 97:1–97:??, July 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3641860.

Lewis:1994:GE

[LH94] D. D. Lewis and P. J. Hayes.

Guest editorial. ACM Transactions on Information Systems, 12(3):231–233, July 1994. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Text Categorization.

Liu:2025:GCQ

[LHA^{+}25a] Bulou Liu, Yiran Hu, Qingyao Ai, Yiqun Liu, Yueyue Wu, Chenliang Li, and Weixing Shen. Generating clarifying questions for conversational legal case retrieval without external knowledge. ACM Transactions on Information Systems, 43(4):102:1–102:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Liu:2025:SAC

[LHA^{+}25b] Bulou Liu, Yiran Hu, Qingyao Ai, Yueyue Wu, Yiqun Liu, Chenliang Li, Fan Zhang, Weixing Shen, Chong Chen, and Qi Tian. Structure-aware conversational legal case retrieval. ACM Transactions on Information Systems, 43(3): 65:1–65:??, May 2025. CO-DEN ATISET. ISSN 1046-

Liu:2012:DTT

[LHG^{+}12] Hongyan Liu, Jun He, Yingqin Gu, Hui Xiong, and Xiaoyong Du. Detecting and tracking topics and events from Web search logs. ACM Transactions on Information Systems, 30(4):21:1–21:??, November 2012. CODEN ATISET. ISSN 1046-8188.

Li:2019:NNN

[LHH^{+}19] Xin Li, Dongcheng Han, Jing He, Lejian Liao, and Mingzhong Wang. Next and next new POI recommendation via latent behavior pattern inference. ACM Transactions on Information Systems, 37(4):46:1–46:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3354187.

Li:2023:PAG

[LHL23] Yakun Li, Lei Hou, and Juanzi Li. Preferenceaware graph attention networks for cross-domain recommendations with collaborative knowledge graph. ACM Transactions on Information Systems, 41(3):80:1– 80:??, July 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3576921.

Liu:2011:CBR

[LHSC11] Jiajun Liu, Zi Huang, Heng Tao Shen, and Bin Cui. Correlationbased retrieval for heavily changed near-duplicate videos. ACM Transactions on Information Systems, 29(4): 21:1–21:??, December 2011. CODEN ATISET. ISSN 1046-8188.

Luo:2025:RIT

[LHZ^{+}25] Sichun Luo, Bowei He, Haohan Zhao, Wei Shao, Yanlin Qi, Yinya Huang, Aojun Zhou, Yuxuan Yao, Zongpeng Li, Yuanzhang Xiao, Mingjie Zhan, and Linqi Song. RecRanker: Instruction tuning large language model as ranker for top-k recommendation. ACM Transactions on Information Systems, 43 (5):113:1–113:31, September 2025. CODEN ATISET. ISSN 1046-8188.

Lin:2007:EPU

[Lin07]  Jimmy Lin.  An explo-

ration of the principles underlying redundancy-based factoid question answering. ACM Transactions on Information Systems, 25(2):6:1–6:??, April 2007. CODEN ATISET. ISSN 1046-8188.

Li:2023:NBR

[LJAdR23] Ming Li, Sami Jullien, Mozhdeh Ariannezhad, and Maarten de Rijke. A next basket recommendation reality check. ACM Transactions on Information Systems, 41(4): 116:1–116:??, October 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3587153.

Li:2017:TAP

[LJHL17] Xin Li, Mingming Jiang, Huiting Hong, and Lejian Liao. A time-aware personalized point-of-interest recommendation via high-order tensor factorization. ACM Transactions on Information Systems, 35(4):31:1–31:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Liu:2019:ALL

[LJQY19] Huafeng Liu, Liping Jing, Yuhua Qian, and Jian Yu. Adaptive local low-rank matrix approximation for recommendation. ACM Transactions on Information Systems, 37(4):45:1–45:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3360488.

Li:2025:MGS

[LJZ^{+}25] Xiaoxi Li, Jiajie Jin, Yujia Zhou, Yuyao Zhang, Peitian Zhang, Yutao Zhu, and Zhicheng Dou. From matching to generation: a survey on generative information retrieval. ACM Transactions on Information Systems, 43(3): 83:1–83:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

Li:2020:WSR

[LK20] Dan Li and Evangelos Kanoulas. When to stop reviewing in technology-assisted reviews: Sampling from an adaptive distribution to estimate residual relevant documents. ACM Transactions on Information Systems, 38(4):41:1– 41:36, October 2020. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3411755.

Lee:1989:PSF

[LL89] Dik Lun Lee and Chun-Wu Leng. Partitioned signature files: Design issues and performance evaluation. ACM Transactions on Information Systems, 7(2):158–180, April 1989. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Lim:1999:HDQ

[LL99] Ee-Peng Lim and Ying Lu.

Harp: a distributed query system for legacy public libraries and structured databases. ACM Transactions on Information Systems, 17(3): 294, July 1999. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org/
pubs/citations/journals/
tois/1999-17-3/p294-lim/.

Lin:2023:DRF

[LL23a] Sheng-Chieh Lin and Jimmy Lin. A dense representation framework for lexical and semantic matching. ACM Transactions on Information Systems, 41(4):110:1– 110:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3582426.

Lo:2023:CPR

[LL23b] Pei-Chi Lo and Ee-Peng Lim. Contextual path retrieval: a contextual entity relation embedding-based approach. ACM Transactions on Information Systems, 41(1):1:1– 1:??, January 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3502720.

Liao:2018:JMP

[LLBS18] Yi Liao, Wai Lam, Lidong Bing, and Xin Shen. Joint modeling of participant influence and latent topics for recommendation in event-based social networks. ACM Transactions on Information Systems, 36(3):29:1–29:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Leroy:2003:UDC

[LLC03] Gondy Leroy, Ann M. Lally, and Hsinchun Chen. The use of dynamic contexts to improve casual Internet searching. ACM Transactions on Information Systems, 21(3): 229–253, July 2003. CODEN ATISET. ISSN 1046-8188.

Lin:2025:CMD

[LLC^{+}25a] Xixun Lin, Rui Liu, Yanan Cao, Lixin Zou, Qian Li, Yongxuan Wu, Yang Liu, Dawei Yin, and Guandong Xu. Contrastive modalitydisentangled learning for multimodal recommendation. ACM Transactions on Information Systems, 43(3):70:1– 70:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Liu:2025:UBR

[LLC^{+}25b] Fan Liu, Yaqi Liu, Huilin Chen, Zhiyong Cheng, Liqiang Nie, and Mohan Kankanhalli. Understanding before recommendation: Semantic aspect-aware review exploitation via large language models. ACM Transactions on Information Systems, 43(2): 44:1–44:??, March 2025. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3704999.

Liang:2023:FSA

[LLG^{+}23] Bin Liang, Xiang Li, Lin Gui, Yonghao Fu, Yulan He, Min Yang, and Ruifeng Xu. Few-shot aspect category sentiment analysis via metalearning. ACM Transactions on Information Systems, 41 (1):22:1–22:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3529954.

Liang:2022:PUQ

[LLM22] Shangsong Liang, Yupeng Luo, and Zaiqiao Meng. Profiling users for question answering communities via flow-based constrained coembedding model. ACM Transactions on Information Systems, 40(2):34:1– 34:38, April 2022. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3470565.

Li:2022:DGR

[LLN22] Yongqi Li, Wenjie Li, and Liqiang Nie. Dynamic graph reasoning for conversational open-domain question answering. ACM Transactions on Information Systems, 40 (4):82:1–82:24, October 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3498557.

Lyu:2025:CRC

[LLN^{+}25] Yuanjie Lyu, Zhiyu Li, Simin Niu, Feiyu Xiong, Bo Tang, Wenjin Wang, Hao Wu, Huanyong Liu, Tong Xu, and Enhong Chen. CRUD-RAG: a comprehensive Chinese benchmark for retrievalaugmented generation of large language models. ACM Transactions on Information Systems, 43(2):41:1– 41:??, March 2025. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3701228.

Luo:2024:CSR

[LLP24] Tianze Luo, Yong Liu, and Sinno Jialin Pan. Collaborative sequential recommendations via multi-view GNNtransformers. ACM Transactions on Information Systems, 42(6):141:1–141:??, November 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3649436.

Lauw:2010:SST

[LLPT10] Hady W. Lauw, Ee-Peng Lim, Hweehwa Pang, and Teck-Tim Tan. STEvent: Spatio-temporal event model for social network discovery. ACM Transactions on Information Systems, 28(3):15:1– 15:??, June 2010. CODEN ATISET. ISSN 1046-8188.

Li:2021:DHM

[LLT^{+}21] Juntao Li, Chang Liu, Chongyang Tao, Zhangming Chan, Dongyan Zhao, Min Zhang, and Rui Yan. Dialogue history matters! Personalized response selection in multi-turn retrieval-based chatbots. ACM Transactions on Information Systems, 39 (4):45:1–45:25, October 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3453183.

Li:2021:SUA

[LLW^{+}21] Shijun Li, Wenqiang Lei, Qingyun Wu, Xiangnan He, Peng Jiang, and Tat-Seng Chua. Seamlessly unifying attributes and items: Conversational recommendation for cold-start users. ACM Transactions on Information Systems, 39(4):40:1–40:29, October 2021. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3446427.

Li:2022:SGF

[LLX^{+}22] Hui Li, Lianyun Li, Guipeng Xv, Chen Lin, Ke Li, and Bingchuan Jiang. SPEX: a generic framework for enhancing neural social recommendation. ACM Transactions on Information Systems, 40(2): 37:1–37:33, April 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3473338.

Liu:2024:HEN

[LLY^{+}24] Sannyuya Liu, Shengyingjie Liu, Zongkai Yang, Jianwen Sun, Xiaoxuan Shen, Qing Li, Rui Zou, and Shangheng Du. Heterogeneous evolution network embedding with temporal extension for intelligent tutoring systems. ACM Transactions on Information Systems, 42(2):45:1–45:??, March 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3617828.

Lai:2025:WTE

[LLY^{+}25] Hanyu Lai, Xiao Liu, Hao Yu, Yifan Xu, Iat Long Iong, Shuntian Yao, Aohan Zeng, Zhengxiao Du, Yuxiao Dong, and Jie Tang. WebGLM: Towards an efficient and reliable Web-enhanced questionanswering system. ACM Transactions on Information Systems, 43(5):122:1–122:43, September 2025. CODEN ATISET. ISSN 1046-8188.

Liu:2021:ABD

[LLZ^{+}21] Bulou Liu, Chenliang Li, Wei Zhou, Feng Ji, Yu Duan, and Haiqing Chen. An attentionbased deep relevance model for few-shot document filtering. ACM Transactions on Information Systems, 39(1): 6:1–6:35, January 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3419972.

Li:2022:IAA

[LLZ^{+}22] Siqing Li, Yaliang Li, Wayne Xin Zhao, Bolin Ding, and Ji-Rong Wen. Interpretable aspect-aware capsule network for peer review based citation count prediction. ACM Transactions on Information Systems, 40(1):11:1–11:29, January 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3466640.

Lu:2025:ASB

[LLZ^{+}25] Xiaodong Lu, Mingzhe Liu, Tongyu Zhu, Leilei Sun, Jibin Wang, Weifeng Lv, Yikun Ban, and Deqing Wang. Adaptive sampling-based dynamic graph learning for information diffusion prediction. ACM Transactions on Information Systems, 43 (5):139:1–139:25, September 2025. CODEN ATISET. ISSN 1046-8188.

Lee:1990:PSV

[LM90] Jintae Lee and Thomas W. Malone. Partially shared views: a scheme for communicating among groups that use different type hierarchies. ACM Transactions on Information Systems, 8(1):1– 26, January 1990. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Lempel:2001:SSA

[LM01] R. Lempel and S. Moran. SALSA: the stochastic approach for link-structure analysis. ACM Transactions on Information Systems, 19(2): 131–160, April 2001. CODEN ATISET. ISSN 1046-8188.

Li:2020:MME

[LMDZ20] Chang Li, Ilya Markov, Maarten De Rijke, and Masrour Zoghi. MergeDTS: a method for effective largescale online ranker evaluation. ACM Transactions on Information Systems, 38(4):40:1– 40:28, October 2020. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3411753.

Liu:2023:GNP

[LMMO23] Siwei Liu, Zaiqiao Meng, Craig Macdonald, and Iadh Ounis. Graph neural pretraining for recommendation with side information. ACM Transactions on Information Systems, 41(3):74:1– 74:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3568953.

Leonhardt:2024:ENR

[LMR^{+}24] Jurek Leonhardt, Henrik Müller, Koustav Rudra, Megha Khosla, Abhijit Anand, and Avishek Anand. Efficient neural ranking using forward indexes and lightweight encoders. ACM Transactions on Information Systems, 42 (5):117:1–117:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3631939.

Lu:2014:BSI

[LMW^{+}14] Shiyang Lu, Tao Mei, Jingdong Wang, Jian Zhang, Zhiyong Wang, and Shipeng Li. Browse-to-search: Interactive exploratory search with visual entities. ACM Transactions on Information Systems, 32 (4):18:1–18:??, October 2014. CODEN ATISET. ISSN 1046-8188.

Lan:2021:PNA

[LMW^{+}21] Tian Lan, Xian-Ling Mao, Wei Wei, Xiaoyan Gao, and Heyan Huang. PONE: a novel automatic evaluation metric for open-domain generative dialogue systems. ACM Transactions on Information Systems, 39(1):7:1–7:37, January 2021. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3423168.

Lu:2023:UPR

[LMW^{+}23] Hongyu Lu, Weizhi Ma, Yifan Wang, Min Zhang, Xiang Wang, Yiqun Liu, Tat-Seng Chua, and Shaoping Ma. User perception of recommendation explanation: Are your explanations what users need? ACM Transactions on Information Systems, 41(2): 48:1–48:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3565480.

Lan:2024:TEC

[LMW^{+}24] Tian Lan, Xian-Ling Mao, Wei Wei, Xiaoyan Gao, and Heyan Huang. Towards efficient coarse-grained dialogue response selection. ACM Transactions on Information Systems, 42(2):35:1– 35:??, March 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3597609.

Lv:2020:BAR

[LMZ20] Pengtao Lv, Xiangwu Meng, and Yujie Zhang. BoRe: Adapting to reader consumption behavior instability for news recommendation. ACM Transactions on Information Systems, 38 (1):3:1–3:??, February 2020. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3361217.

Li:2023:PRF

[LMZ^{+}23a] Hang Li, Ahmed Mourad, Shengyao Zhuang, Bevan Koopman, and Guido Zuccon. Pseudo relevance feedback with deep language models and dense retrievers: Successes and pitfalls. ACM Transactions on Information Systems, 41(3):62:1– 62:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3570724.

Liang:2023:ETA

[LMZ23b] Bi Liang, Xiangwu Meng, and Yujie Zhang. Exploring time-aware multi-pattern group venue recommendation in LBSNs. ACM Transactions on Information Systems, 41(3):70:1–70:??, July 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3564280.

Lucchese:2025:EEE

[LNO^{+}25] Claudio Lucchese, Franco Maria Nardini, Salvatore Orlando, Raffaele Perego, and Alberto Veneri. Explainable, effective, and efficient learning-to-rank models using ILMART. ACM Transactions on Information Systems, 43(4):105:1–105:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Lucchese:2013:DTS

[LOP^{+}13] Claudio Lucchese, Salvatore Orlando, Raffaele Perego, Fabrizio Silvestri, and Gabriele Tolomei. Discovering tasks from search engine query logs. ACM Transactions on Information Systems, 31(3):14:1– 14:??, July 2013. CODEN ATISET. ISSN 1046-8188.

Li:2023:PSK

[LPC^{+}23] Minghan Li, Diana Nicoleta Popa, Johan Chagnon, Yagmur Gizem Cinar, and Eric Gaussier. The power of selecting key blocks with local pre-ranking for long document information retrieval. ACM Transactions on Information Systems, 41(3):73:1– 73:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3568394.

Loni:2019:TRM

[LPLH19] Babak Loni, Roberto Pagano, Martha Larson, and Alan Hanjalic. Top-N recommendation with multi-channel positive feedback using factorization machines. ACM Transactions on Information Systems, 37(2):15:1– 15:??, March 2019. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.
acm.org/ft_gateway.cfm?
id=3291756.

Liddy:1994:TCM

[LPY94] Elizabeth D. Liddy, Woojin Paik, and Edmund S. Yu. Text categorization for multiple users based on semantic features from a machinereadable dictionary. ACM Transactions on Information Systems, 12(3):278–295, July 1994. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Text Categorization.

Lin:2023:GFR

[LPYM23] Zhaohao Lin, Weike Pan, Qiang Yang, and Zhong Ming. A generic federated recommendation framework via fake marks and secret sharing. ACM Transactions on Information Systems, 41(2):40:1– 40:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3548456.

Liu:2025:ANI

[LPZ^{+}25] Yuanxing Liu, Jiahuan Pei, Wei-Nan Zhang, Ming Li, Wanxiang Che, and Maarten de Rijke. Augmentation with neighboring information for conversational recommendation. ACM Transactions on Information Systems, 43(3): 62:1–62:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

Li:2024:CCP

[LQJ^{+}24] Xinhang Li, Zhaopeng Qiu, Jiacheng Jiang, Yong Zhang, Chunxiao Xing, and Xian Wu. Conditional cross-platform user engagement prediction. ACM Transactions on Information Systems, 42(1):6:1– 6:??, January 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3589226.

Liu:2025:CPM

[LQZ^{+}25] Qidong Liu, Zhaopeng Qiu, Xiangyu Zhao, Xian Wu, Zijian Zhang, Tong Xu, and Feng Tian. A contrastive pretrain model with prompt tuning for multi-center medication recommendation. ACM Transactions on Information Systems, 43(3):57:1–57:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Lee:1996:DRW

[LR96] Dik Lun Lee and Liming Ren. Document ranking on weight-partitioned signature files. ACM Transactions on Information Systems, 14(2): 109–137, April 1996. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/lee.html.

Leonhardt:2023:EEI

[LRA23] Jurek Leonhardt, Koustav Rudra, and Avishek Anand. Extractive explanations for interpretable text ranking. ACM Transactions on Information Systems, 41(4):88:1– 88:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3576924.

Li:2025:ERF

[LRSD25] Jie Li, Yongli Ren, Mark Sanderson, and Ke Deng. Explaining recommendation fairness from a user/item perspective. ACM Transactions on Information Systems, 43 (1):17:1–17:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3698877.

Liang:2017:IDU

[LRZ^{+}17] Shangsong Liang, Zhaochun Ren, Yukun Zhao, Jun Ma, Emine Yilmaz, and Maarten De Rijke. Inferring dynamic user interests in streams of short texts for user clustering. ACM Transactions on Information Systems, 36(1):10:1– 10:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Lempel:2002:PPA

[LS02] Ronny Lempel and Aya Soffer. PicASHOW: Pictorial authority search by hyperlinks on the Web. ACM Transactions on Information Systems, 20(1):1–24, January 2002. CODEN ATISET. ISSN 1046-8188.

Li:2025:ISD

[LSG^{+}25] Dongyang Li, Jianshan Sun, Chongming Gao, Fuli Feng, and Kun Yuan. Independent or social driven decision? a counterfactual refinement strategy for graphbased social recommendation. ACM Transactions on Information Systems, 43(3):76:1– 76:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Li:2024:DDM

[LSL24] Zihao Li, Aixin Sun, and Chenliang Li. DiffuRec: a diffusion model for sequential recommendation. ACM Transactions on Information Systems, 42(3):66:1– 66:??, May 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3631116.

Lee:2006:UEF

[LSOS06] Hyowon Lee, Alan F. Smeaton, Noel E. O’Connor, and Barry Smyth. User evaluation of Físchlár-News: an automatic broadcast news delivery system. ACM Transactions on Information Systems, 24(2): 145–189, April 2006. CODEN ATISET. ISSN 1046-8188.

Li:2025:CCE

[LSZ^{+}25] Honghao Li, Lei Sang, Yi Zhang, Xuyun Zhang, and Yiwen Zhang. CETN: Contrast-enhanced through network for click-through rate prediction. ACM Transactions on Information Systems, 43(1):4:1–4:??, January 2025. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3688571.

Lin:2025:UHH

[LTC^{+}25] Zhenghong Lin, Yanchao Tan, Jiamin Chen, Hengyu Zhang, Chaochao Chen, Shiping Wang, and Carl Yang. Unified heterogeneous hypergraph construction for incomplete multimedia recommendation. ACM Transactions on Information Systems, 43 (5):132:1–132:31, September 2025. CODEN ATISET. ISSN 1046-8188.

Luz:2012:NSP

[Luz12] Saturnino Luz. The nonverbal structure of patient case discussions in multidisciplinary medical team meetings. ACM Transactions on Information Systems, 30(3): 17:1–17:??, August 2012. CO-DEN ATISET. ISSN 1046-8188.

Liu:2020:FDN

[LW20] Yang Liu and Yi-Fang Brook Wu. FNED: a deep network for fake news early detection on social media. ACM Transactions on Information Systems, 38(3):25:1– 25:33, June 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3386253.

Lin:2008:TAF

[LWA08] Jimmy Lin, Philip Wu, and Eileen Abels. Toward automatic facet analysis and need negotiation: Lessons from mediated search. ACM Transactions on Information Systems, 27(1):6:1–6:??, December 2008. CODEN ATISET. ISSN 1046-8188.

Lv:2025:DCR

[LWC^{+}25] Xiangwei Lv, Guifeng Wang, Jingyuan Chen, Hejian Su, Zhiang Dong, Yumeng Zhu, Beishui Liao, and Fei Wu. Debiased cognition representation learning for knowledge tracing. ACM Transactions on Information Systems, 43 (5):127:1–127:30, September 2025. CODEN ATISET. ISSN 1046-8188.

Liu:2023:FUM

[LWH^{+}23] Qi Liu, Jinze Wu, Zhenya Huang, Hao Wang, Yuting Ning, Ming Chen, Enhong Chen, Jinfeng Yi, and Bowen Zhou. Federated user modeling from hierarchical information. ACM Transactions on Information Systems, 41(2): 46:1–46:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3560485.

Liu:2024:DMF

[LWL^{+}24] Han Liu, Yinwei Wei, Fan Liu, Wenjie Wang, Liqiang Nie, and Tat-Seng Chua. Dynamic multimodal fusion via meta-learning towards micro-video recommendation. ACM Transactions on Information Systems, 42(2):47:1– 47:??, March 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3617827.

Li:2025:BCA

[LWL^{+}25] Xinze Li, Hanbin Wang, Zhenghao Liu, Shi Yu, Shuo Wang, Yukun Yan, Yukai Fu, Yu Gu, and Ge Yu. Building a coding assistant via the retrieval-augmented language model. ACM Transactions on Information Systems, 43(2): 39:1–39:??, March 2025. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3695868.

Liu:2021:TCU

[LWP^{+}21] Hongtao Liu, Wenjun Wang, Qiyao Peng, Nannan Wu, Fangzhao Wu, and Pengfei Jiao. Toward comprehensive user and item representations via three-tier attention network. ACM Transactions on Information Systems, 39(3): 25:1–25:22, July 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3446341.

Li:2010:LCG

[LWSA10] Xiao Li, Ye-Yi Wang, Dou Shen, and Alex Acero. Learning with click graph for query intent classification. ACM Transactions on Information Systems, 28(3):12:1– 12:??, June 2010. CODEN ATISET. ISSN 1046-8188.

Liu:2025:AGE

[LXC^{+}25] Zunlong Liu, Yang Xu, Gao Cong, Lei Zhu, Qinjun Qiu, and Huaxiang Zhang. ARTS: a general and efficient multitask self-prompt framework for explainable sequential recommendation. ACM Transactions on Information Systems, 43(3):73:1–73:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Liu:2023:MME

[LXG^{+}23] Kang Liu, Feng Xue, Dan Guo, Le Wu, Shujie Li, and Richang Hong. MEGCF: Multimodal entity graph collaborative filtering for personalized recommendation. ACM Transactions on Information Systems, 41(2):30:1– 30:??, April 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3544106.

Luo:2016:MSU

[LXLZ16] Xiangfeng Luo, Junyu Xuan, Jie Lu, and Guangquan Zhang. Measuring the semantic uncertainty of news events for evolution potential estimation. ACM Transactions on Information Systems, 34(4): 24:1–24:??, September 2016. CODEN ATISET. ISSN 1046-8188.

Liu:2017:TAC

[LXW^{+}17] Yiqun Liu, Xiaohui Xie, Chao Wang, Jian-Yun Nie, Min Zhang, and Shaoping Ma. Time-aware click model. ACM Transactions on Information Systems, 35(3):16:1– 16:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Li:2022:SCA

[LXZ^{+}22] Dan Li, Tong Xu, Peilun Zhou, Weidong He, Yanbin Hao, Yi Zheng, and Enhong Chen. Social contextaware person search in videos via multi-modal cues. ACM Transactions on Information Systems, 40(3):52:1– 52:25, July 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3480967.

Li:2025:PLR

[LYG^{+}25] Fei Li, Enneng Yang, Guibing Guo, Linying Jiang, Jianzhe Zhao, and Xingwei Wang. Preference logical reasoning with preference operators for explainable recommendations. ACM Transactions on Information Systems, 43(4): 99:1–99:??, July 2025. CO-DEN ATISET. ISSN 1046-8188.

Li:2022:MSE

[LYLS22] Rui Li, Cheng Yang, Tingwei Li, and Sen Su. MiDTD: a simple and effective distillation framework for distantly supervised relation extraction. ACM Transactions on Information Systems, 40 (4):83:1–83:32, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3503917.

Li:2024:PPR

[LYM^{+}24] Canjia Li, Andrew Yates, Sean MacAvaney, Ben He, and Yingfei Sun. PARADE: Passage representation aggregation for document reranking. ACM Transactions on Information Systems, 42(2): 36:1–36:??, March 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3600088.

Lin:2021:MSC

[LYN^{+}21] Sheng-Chieh Lin, Jheng-Hong Yang, Rodrigo Nogueira, Ming-Feng Tsai, Chuan-Ju Wang, and Jimmy Lin. Multistage conversational passage retrieval: an approach to fusing term importance estimation and neural query rewriting. ACM Transactions on Information Systems, 39 (4):48:1–48:29, October 2021. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3446426.

Li:2025:RPB

[LYP^{+}25] Pu Li, Xiaoyan Yu, Hao Peng, Yantuan Xian, Linqin Wang, Li Sun, Jingyun Zhang, and Philip S. Yu. Relational prompt-based pre-trained language models for social event detection. ACM Transactions on Information Systems, 43 (1):12:1–12:??, January 2025. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3695869.

Liang:2017:SRD

[LYS^{+}17] Shangsong Liang, Emine Yilmaz, Hong Shen, Maarten De Rijke, and W. Bruce Croft. Search result diversification in short text streams. ACM Transactions on Information Systems, 36(1):8:1–8:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Li:2024:ISS

[LYY^{+}24] Shujie Li, Guanghu Yuan, Min Yang, Ying Shen, Chengming Li, Ruifeng Xu, and Xiaoyan Zhao. Improving semi-supervised text classification with dual metalearning. ACM Transactions on Information Systems, 42 (4):109:1–109:??, July 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3648612.

Li:2023:HSI

[LYZ^{+}23] Bing Li, Peng Yang, Hanlin Zhao, Penghui Zhang, and Zijian Liu. Hierarchical sliding inference generator for question-driven abstractive answer summarization. ACM Transactions on Information Systems, 41(1): 7:1–7:??, January 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3511891.

Lucarella:1996:VRE

[LZ96] Dario Lucarella and Antonella Zanzi. A visual retrieval environment for hypermedia information systems. ACM Transactions on Information Systems, 14(1):3– 29, January 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/lucarella.
html.

Li:2023:PPL

[LZC23] Lei Li, Yongfeng Zhang, and Li Chen. Personalized prompt learning for explainable recommendation. ACM Transactions on Information Systems, 41(4):103:1–103:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3580488.

Lien:2024:GWS

[LZC24a] Yen-Chieh Lien, Hamed Zamani, and Bruce Croft. Generalized weak supervision for neural information retrieval. ACM Transactions on Information Systems, 42(5):121:1– 121:??, September 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3647639.

Lin:2024:RRC

[LZC^{+}24b] Siyi Lin, Sheng Zhou, Jiawei Chen, Yan Feng, Qihao Shi, Chun Chen, Ying Li, and Can Wang. ReCRec: Reasoning the causes of implicit feedback for debiased recommendation. ACM Transactions on Information Systems, 42(6):158:1– 158:??, November 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3672275.

Liu:2024:CBG

[LZC^{+}24c] Fan Liu, Shuai Zhao, Zhiyong Cheng, Liqiang Nie, and Mohan Kankanhalli. Clusterbased graph collaborative filtering. ACM Transactions on Information Systems, 42(6): 167:1–167:??, November 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3687481.

Lan:2025:CCL

[LZC^{+}25a] Wei Lan, Guoxian Zhou, Qingfeng Chen, Wenguang Wang, Shirui Pan, Yi Pan, and Shichao Zhang. Contrastive clustering learning for multi-behavior recommendation. ACM Transactions on Information Systems, 43 (1):18:1–18:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3698192.

Liu:2025:CTC

[LZC^{+}25b] Zirui Liu, Hailin Zhang, Boxuan Chen, Zihan Jiang, Yikai Zhao, Yangyu Tao, Tong Yang, and Bin Cui. CAFE+: Towards compact, adaptive, and fast embedding for largescale online recommendation models. ACM Transactions on Information Systems, 43 (3):61:1–61:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Lian:2018:GSL

[LZG^{+}18] Defu Lian, Kai Zheng, Yong Ge, Longbing Cao, Enhong Chen, and Xing Xie. GeoMF++: Scalable location recommendation via joint geographical modeling and matrix factorization. ACM Transactions on Information Systems, 36(3):33:1–33:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Liu:2021:MRA

[LZG21] Peng Liu, Lemei Zhang, and Jon Atle Gulla. Multilingual review-aware deep recommender system via aspectbased sentiment analysis. ACM Transactions on Information Systems, 39(2):15:1– 15:33, March 2021. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3432049.

Liao:2025:RNS

[LZL^{+}25] Weibin Liao, Yifan Zhu, Yanyan Li, Qi Zhang, Zhonghong Ou, and Xuesong Li. RevGNN: Negative sampling enhanced contrastive graph learning for academic reviewer recommendation. ACM Transactions on Information Systems, 43(1): 1:1–1:??, January 2025. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3679200.

Lin:2019:EPR

[LZLM19] Xiao Lin, Min Zhang, Yiqun Liu, and Shaoping Ma. Enhancing personalized recommendation by implicit preference communities modeling. ACM Transactions on Information Systems, 37(4): 48:1–48:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3352592.

Liu:2019:SRR

[LZM^{+}19] Yiqun Liu, Junqi Zhang, Jiaxin Mao, Min Zhang, Shaoping Ma, Qi Tian, Yanxiong Lu, and Leyu Lin. Search result reranking with visual and structure information sources. ACM Transactions on Information Systems, 37(3):38:1–38:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3329188.

Lin:2020:EEB

[LZW^{+}20] Hao Lin, Hengshu Zhu, Junjie Wu, Yuan Zuo, Chen Zhu, and Hui Xiong. Enhancing employer brand evaluation with collaborative topic regression models. ACM Transactions on Information Systems, 38(4):32:1–32:33, October 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3392734.

Liu:2021:MEC

[LZW21] Zeyang Liu, Ke Zhou, and Max L. Wilson. Metaevaluation of conversational search evaluation metrics. ACM Transactions on Information Systems, 39(4):52:1– 52:42, October 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3445029.

Li:2024:INR

[LZWZ24] Haoyang Li, Ziwei Zhang, Xin Wang, and Wenwu Zhu. Invariant node representation learning under distribution shifts with multiple latent environments. ACM Transactions on Information Systems, 42(1):26:1– 26:??, January 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3604427.

Li:2023:TAI

[LZZ^{+}23] Qiming Li, Zhao Zhang, Fuzhen Zhuang, Yongjun Xu, and Chao Li. Topicaware intention network for explainable recommendation with knowledge enhancement. ACM Transactions on Information Systems, 41(4):92:1– 92:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3579993.

Meng:2025:QPP

[MAA^{+}25] Chuan Meng, Negar Arabzadeh, Arian Askari, Mohammad Aliannejadi, and Maarten de Rijke. Query performance prediction using relevance judgments generated by large language models. ACM Transactions on Information Systems, 43(4):106:1–106:??, July 2025. CODEN ATISET. ISSN 1046-8188.

McGregor:2024:SRC

[MAH24] Molly McGregor, Leif Azzopardi, and Martin Halvey. A systematic review of cost, effort, and load research in information search and retrieval, 1972–2020. ACM Transactions on Information Systems, 42(1):2:1–2:??, January 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3583069.

Molino:2016:SQA

[MAL16] Piero Molino, Luca Maria Aiello, and Pasquale Lops. Social question answering: Textual, user, and network features for best answer prediction. ACM Transactions on Information Systems, 35(1): 4:1–4:??, October 2016. CO-DEN ATISET. ISSN 1046-8188.

Manber:1997:TCS

[Man97] Udi Manber. A text compression scheme that allows fast searching directly in the compressed file. ACM Transactions on Information Systems, 15(2):124– 136, April 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/manber.
html.

Mansoury:2022:GBA

[MAP^{+}22] Masoud Mansoury, Himan Abdollahpouri, Mykola Pechenizkiy, Bamshad Mobasher, and Robin Burke. A graphbased approach for mitigating multi-sided exposure bias in recommender systems. ACM Transactions on Information Systems, 40(2):32:1– 32:31, April 2022. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3470948.

Markus:1994:FHM

[Mar94] M. L. Markus. Finding a happy medium: Explaining the negative effects of electronic communication on social life at work. ACM Transactions on Information Systems, 12(2):119–149, April 1994. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

Marchionini:2006:TR

[Mar06] Gary Marchionini. TOIS reviewers 2003–2005. ACM Transactions on Information Systems, 24(1):142–143, January 2006. CODEN ATISET. ISSN 1046-8188.

Marchionini:2007:TRJ

[Mar07] Gary Marchionini. TOIS reviewers January 2006 through May 2007. ACM Transactions on Information Systems, 25 (4):15:1–15:??, October 2007. CODEN ATISET. ISSN 1046-8188.

Marchionini:2008:ERM

[Mar08a] Gary Marchionini. Editorial: Reviewer merits and review control in an age of electronic manuscript management systems. ACM Transactions on Information Systems, 26(4): 25:1–25:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Marchionini:2008:TRJ

[Mar08b] Gary Marchionini. TOIS reviewers June 2007 through May 2008. ACM Transactions on Information Systems, 26(4):26:1–26:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Mylopoulos:1990:TRK

[MBJK90] John Mylopoulos, Alex Borgida, Matthias Jarke, and Manolis Koubarakis. Telos: Representing knowledge about information systems. ACM Transactions on Information Systems, 8(4):325–362, October 1990. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Moffat:2017:IUE

[MBST17] Alistair Moffat, Peter Bailey, Falk Scholer, and Paul Thomas. Incorporating user expectations and behavior into the measurement of search effectiveness. ACM Transactions on Information Systems, 35(3):24:1–24:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Marchionini:1994:EHL

[MC94] Gary Marchionini and Gregory Crane. Evaluating hypermedia and learning: Methods and results from the Perseus Project. ACM Transactions on Information Systems, 12(1):5–34, January 1994. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

McDonald:2006:SCS

[MC06] Daniel M. McDonald and Hsinchun Chen. Summary in context: Searching versus browsing. ACM Transactions on Information Systems, 24(1):111–141, January 2006. CODEN ATISET. ISSN 1046-8188.

Maslennikov:2010:CRI

[MC10a] Mstislav Maslennikov and Tat-Seng Chua. Combining relations for information extraction from free text. ACM Transactions on Information Systems, 28(3):14:1– 14:??, June 2010. CODEN ATISET. ISSN 1046-8188.

Minkov:2010:IGW

[MC10b] Einat Minkov and William W. Cohen. Improving graphwalk-based similarity with reranking: Case studies for personal information management. ACM Transactions on Information Systems, 29(1): 4:1–4:??, December 2010. CO-DEN ATISET. ISSN 1046-8188.

Mahdabi:2014:PQF

[MC14a] Parvaz Mahdabi and Fabio Crestani. Patent query formulation by synthesizing multiple sources of relevance evidence. ACM Transactions on Information Systems, 32 (4):16:1–16:??, October 2014. CODEN ATISET. ISSN 1046-8188.

Markov:2014:TQQ

[MC14b] Ilya Markov and Fabio Crestani. Theoretical, qualitative, and quantitative analyses of small-document approaches to resource selection. ACM Transactions on Information Systems, 32(2): 9:1–9:??, April 2014. CODEN ATISET. ISSN 1046-8188.

Miao:2017:CEO

[MCF^{+}17] Zhongchen Miao, Kai Chen, Yi Fang, Jianhua He, Yi Zhou, Wenjun Zhang, and Hongyuan Zha. Cost-effective online trending topic detection and popularity prediction in microblogging. ACM Transactions on Information Systems, 35(3):18:1–18:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Moon:2012:OLF

[MCL^{+}12] Taesup Moon, Wei Chu, Lihong Li, Zhaohui Zheng, and Yi Chang. An online learning framework for refining recency search results with user click feedback. ACM Transactions on Information Systems, 30(4):20:1–20:??, November 2012. CODEN ATISET. ISSN 1046-8188.

Mao:2025:FRB

[MDX^{+}25] Yuren Mao, Xuemei Dong, Wenyi Xu, Yunjun Gao, Bin Wei, and Ying Zhang. FIT-RAG: Black-box RAG with factual information and token reduction. ACM Transactions on Information Systems, 43(2):40:1–40:??, March 2025. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3676957.

Melucci:2008:BIR

[Mel08] Massimo Melucci. A basis for information retrieval in context. ACM Transactions on Information Systems, 26(3): 14:1–14:??, June 2008. CO-DEN ATISET. ISSN 1046-8188.

Metzler:1989:COP

[MH89] Douglas P. Metzler and Stephanie W. Haas. The constituent object parser: Syntactic structure matching for information retrieval. ACM Transactions on Information Systems, 7(4):292–316, October 1989. CODEN ATISET. ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Ma:2023:KGK

[MHLH23] Ting Ma, Longtao Huang, Qianqian Lu, and Songlin Hu. KR-GCN: Knowledgeaware reasoning with graph convolution network for explainable recommendation. ACM Transactions on Information Systems, 41(1):4:1– 4:??, January 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3511019.

Moshfeghi:2022:GTA

[MHR22] Yashar Moshfeghi and Alvaro Francisco Huertas-Rosero. A game theory approach for estimating reliability of crowdsourced relevance assessments. ACM Transactions on Information Systems, 40(3):60:1–60:29, July 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3480965.

Ma:2024:MMV

[MHW^{+}24] Yunshan Ma, Yingzhi He, Xiang Wang, Yinwei Wei, Xiaoyu Du, Yuyangzi Fu, and Tat-Seng Chua. MultiCBR: Multi-view contrastive learning for bundle recommendation. ACM Transactions on Information Systems, 42(4): 100:1–100:??, July 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3640810.

Miao:2016:TPF

[MHZ16] Jun Miao, Jimmy Xiangji Huang, and Jiashu Zhao. Top-PRF: a probabilistic framework for integrating topic space into pseudo relevance feedback. ACM Transactions on Information Systems, 34(4):22:1–22:??, September 2016. CODEN ATISET. ISSN 1046-8188.

Mayer:2015:IOV

[MJH15] Julia M. Mayer, Quentin Jones, and Starr Roxanne Hiltz. Identifying opportunities for valuable encounters: Toward context-aware social matching systems. ACM Transactions on Information Systems, 34(1):1:1–1:??, October 2015. CODEN ATISET. ISSN 1046-8188.

Merz:1994:DQF

[MK94] Ulla Merz and Roger King. DIRECT: a query facility for multiple databases. ACM Transactions on Information Systems, 12(4):339–359, October 1994. CODEN ATISET. ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Middleton:2016:GGS

[MK16a] Stuart E. Middleton and Vadims Krivcovs. Geoparsing and geosemantics for social media: Spatiotemporal grounding of content propagating rumors to support trust and veracity analysis during breaking news. ACM Transactions on Information Systems, 34(3):16:1– 16:??, May 2016. CODEN ATISET. ISSN 1046-8188.

Morsy:2016:ALC

[MK16b] Sara Morsy and George Karypis. Accounting for language changes over time in document similarity search. ACM Transactions on Information Systems, 35(1):1:1– 1:??, October 2016. CODEN ATISET. ISSN 1046-8188.

Mak:1991:EPP

[MKF91] Victor Wing-Kit Mak, Chu Lee Kuo, and Ophir Frieder. Exploiting parallelism in pattern matching: An information retrieval application. ACM Transactions on Information Systems, 9(1):52–74, January 1991. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Miller:2004:PTP

[MKR04] Bradley N. Miller, Joseph A.

Konstan, and John Riedl. PocketLens: Toward a personal recommender system. ACM Transactions on Information Systems, 22(3):437– 476, July 2004. CODEN ATISET. ISSN 1046-8188.

Middleton:2018:LES

[MKZPK18] Stuart E. Middleton, Giorgos Kordopatis-Zilos, Symeon Papadopoulos, and Yiannis Kompatsiaris. Location extraction from social media: Geoparsing, location disambiguation, and geotagging. ACM Transactions on Information Systems, 36(4):40:1– 40:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Mana-Lopez:2004:MSA

[MLDGH04] Manuel J. Maña-López, Manuel De Buenaga, and José M. Gómez-Hidalgo. Multidocument summarization: an added value to clustering in interactive retrieval. ACM Transactions on Information Systems, 22(2):215–241, April 2004. CODEN ATISET. ISSN 1046-8188.

Malone:1995:EOR

[MLF95] Thomas W. Malone, Kum-Yew Lai, and Christopher Fry. Experiments with oval: a radically tailorable tool for cooperative work. ACM Transactions on Information Systems, 13(2):177–205, April 1995. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Mao:2018:HDD

[MLK^{+}18] Jiaxin Mao, Yiqun Liu, Noriko Kando, Min Zhang, and Shaoping Ma. How does domain expertise affect users’ search interaction and outcome in exploratory search? ACM Transactions on Information Systems, 36(4):42:1– 42:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Ma:2024:PDK

[MLL^{+}24] Longxuan Ma, Jiapeng Li, Mingda Li, Wei-Nan Zhang, and Ting Liu. Policydriven knowledge selection and response generation for document-grounded dialogue. ACM Transactions on Information Systems, 42(2):49:1– 49:??, March 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3617829.

Mustar:2022:STQ

[MLP22] Agnès Mustar, Sylvain Lamprier, and Benjamin Piwowarski. On the study of transformers for query suggestion. ACM Transactions on Information Systems, 40 (1):18:1–18:27, January 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3470562.

Meng:2020:JLR

[MLZ^{+}20] Zaiqiao Meng, Shangsong Liang, Xiangliang Zhang, Richard McCreadie, and Iadh Ounis. Jointly learning representations of nodes and attributes for attributed networks. ACM Transactions on Information Systems, 38 (2):16:1–16:32, January 2020. CODEN ATISET. ISSN 1046-

8188. URL https://dl.
acm.org/doi/abs/10.1145/
3377850.

Ma:2022:UTE

[MLZ^{+}22a] Longxuan Ma, Mingda Li, Wei-Nan Zhang, Jiapeng Li, and Ting Liu. Unstructured text enhanced open-domain dialogue system: a systematic survey. ACM Transactions on Information Systems, 40(1):9:1–9:44, January 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3464377.

Mu:2022:KGD

[MLZ^{+}22b] Shanlei Mu, Yaliang Li, Wayne Xin Zhao, Siqing Li, and Ji-Rong Wen. Knowledgeguided disentangled representation learning for recommender systems. ACM Transactions on Information Systems, 40(1):6:1–6:26, January 2022. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3464304.

Mena-Maldonado:2021:PBF

[MMCC^{+}21] Elisa Mena-Maldonado, Rocío Cañamares, Pablo Castells, Yongli Ren, and Mark Sanderson. Popularity bias in falsepositive metrics for recommender systems evaluation. ACM Transactions on Information Systems, 39(3):36:1– 36:43, July 2021. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3452740.

Mei:2024:IFS

[MMH^{+}24] Lang Mei, Jiaxin Mao, Juan Hu, Naiqiang Tan, Hua Chai, and Ji-Rong Wen. Improving first-stage retrieval of pointof-interest search by pretraining models. ACM Transactions on Information Systems, 42(3):74:1–74:??, May 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3631937.

Mostafa:1997:MAI

[MMLP97] J. Mostafa, S. Mukhopadhyay, W. Lam, and M. Palakal. A multilevel approach to intelligent information filtering:

Model, system, and evaluation. ACM Transactions on Information Systems, 15(4): 368–399, October 1997. CO-DEN ATISET. ISSN 1046-8188.

Mcdonald:2021:HAC

[MMO21] Graham Mcdonald, Craig Macdonald, and Iadh Ounis. How the accuracy and confidence of sensitivity classification affects digital sensitivity review. ACM Transactions on Information Systems, 39(1): 4:1–4:34, January 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3417334.

Majumder:2007:YYA

[MMP^{+}07] Prasenjit Majumder, Mandar Mitra, Swapan K. Parui, Gobinda Kole, Pabitra Mitra, and Kalyankumar Datta. YASS: Yet another suffix stripper. ACM Transactions on Information Systems, 25(4):18:1– 18:??, October 2007. CODEN ATISET. ISSN 1046-8188.

Maddalena:2017:CRM

[MMST17] Eddy Maddalena, Stefano Mizzaro, Falk Scholer, and Andrew Turpin. On crowdsourcing relevance magnitudes for information retrieval evaluation. ACM Transactions on Information Systems, 35(3):19:1–19:??, June 2017. CODEN ATISET. ISSN 1046-

Muntean:2021:WPE

[MNP^{+}21a] Cristina Ioana Muntean, Franco Maria Nardini, Raffaele Perego, Nicola Tonellotto, and Ophir Frieder. Weighting passages enhances accuracy. ACM Transactions on Information Systems, 39 (2):11:1–11:11, March 2021. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3428687.

Musto:2021:MDA

[MNP^{+}21b] Cataldo Musto, Fedelucio Narducci, Marco Polignano, Marco De Gemmis, Pasquale Lops, and Giovanni Semeraro. MyrrorBot: a digital assistant based on holistic user models for personalized access to online services. ACM Transactions on Information Systems, 39(4):46:1– 46:34, October 2021. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3447679.

Moffat:1998:ACR

[MNW98] Alistair Moffat, Radford M. Neal, and Ian H. Witten. Arithmetic coding revisited. ACM Transactions on Information Systems, 16(3):256– 294, July 1998. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tois/1998-16-3/p256-moffat/

.

Mic:2019:BSS

[MNZ19] Vladimir Mic, David Novak, and Pavel Zezula. Binary sketches for secondary filtering. ACM Transactions on Information Systems, 37(1):1:1– 1:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Miotto:2012:PMC

[MO12] Riccardo Miotto and Nicola Orio. A probabilistic model to combine tags and acoustic similarity for music retrieval. ACM Transactions on Information Systems, 30 (2):8:1–8:??, May 2012. CO-DEN ATISET. ISSN 1046-8188.

Morrissey:1990:IIU

[Mor90] J. M. Morrissey. Imprecise information and uncertainty in information systems. ACM Transactions on Information Systems, 8(2):159–180, April 1990. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Moss:1990:DMP

[Mos90] J. Eliot B. Moss. Design of the Mneme persistent object store. ACM Transactions on Information Systems, 8(2): 103–139, April 1990. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Macdonald:2011:UBA

[MOT11] Craig Macdonald, Iadh Ounis, and Nicola Tonellotto.

Upper-bound approximations for dynamic pruning. ACM Transactions on Information Systems, 29(4):17:1–17:??, December 2011. CODEN ATISET. ISSN 1046-8188.

Moffat:2020:LAS

[MP20] Alistair Moffat and Matthias Petri. Large-alphabet semistatic entropy coding via asymmetric numeral systems. ACM Transactions on Information Systems, 38(4):33:1– 33:33, October 2020. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3397175.

Moldovan:2003:PIE

[MPHS03] Dan Moldovan, Marius Paşca, Sanda Harabagiu, and Mihai Surdeanu. Performance issues and error analysis in an opendomain question answering system. ACM Transactions on Information Systems, 21 (2):133–154, 2003. CODEN ATISET. ISSN 1046-8188.

Mackenzie:2022:ARD

[MPM22] Joel Mackenzie, Matthias Petri, and Alistair Moffat. Anytime ranking on document-ordered indexes. ACM Transactions on Information Systems, 40(1):13:1– 13:32, January 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3467890.

Ma:2007:IBP

[MPS07] Zhongming Ma, Gautam Pant, and Olivia R. Liu Sheng. Interest-based personalized search. ACM Transactions on Information Systems, 25(1):5:1–5:??, February 2007. CODEN ATISET. ISSN 1046-8188.

Moreo:2023:GFE

[MPS23] Alejandro Moreo, Andrea Pedrotti, and Fabrizio Sebastiani. Generalized funnelling: Ensemble learning and heterogeneous document embeddings for cross-lingual text classification. ACM Transactions on Information Systems, 41(2):36:1–36:??, April 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3544104.

Mousset:2021:EEN

[MPT21] Paul Mousset, Yoann Pitarch, and Lynda Tamine. End-to-end neural matching for semantic location prediction of tweets. ACM Transactions on Information Systems, 39(1): 3:1–3:35, January 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3415149.

Mackinlay:1992:EUI

[MR92] Jock Mackinlay and Jim Rhyne. Editorial: User interface software and technology. ACM Transactions on Information Systems, 10(4):

317–319, October 1992. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Magalhaes:2010:ITF

[MR10] João Magalhães and Stefan Rüger. An informationtheoretic framework for semanticmultimedia retrieval. ACM Transactions on Information Systems, 28(4):19:1–19:??, November 2010. CODEN ATISET. ISSN 1046-8188.

Ma:2023:ITB

[MRC^{+}23] Muyang Ma, Pengjie Ren, Zhumin Chen, Zhaochun Ren, Huasheng Liang, Jun Ma, and Maarten De Rijke. Improving transformer-based sequential recommenders through preference editing. ACM Transactions on Information Systems, 41(3):71:1–71:??, July 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3564282.

Michalkova:2024:UFK

[MRM24] Dominika Michalkova, Mario Parra Rodriguez, and Yashar Moshfeghi. Understanding feelingof-knowing in information search: an EEG study. ACM Transactions on Information Systems, 42(3):79:1– 79:??, May 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3611384.

Melnik:2001:BDF

[MRYGM01] Sergey Melnik, Sriram Raghavan, Beverly Yang, and Hector Garcia-Molina. Building a distributed full-text index for the Web. ACM Transactions on Information Systems, 19 (3):217–241, July 2001. CO-DEN ATISET. ISSN 1046-8188.

Malone:1993:GE

[MS93] Thomas Malone and Norbert Streitz. Guest editorial. ACM Transactions on Information Systems, 11(4):319–320, October 1993. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80. Special

Issue on Computer-Supported Cooperative Work (CSCW).

Meuss:2001:CAA

[MS01] Holger Meuss and Klaus U.

Schulz. Complete answer aggregates for treelike databases: a novel approach to combine querying and navigation. ACM Transactions on Information Systems, 19(2): 161–215, April 2001. CODEN ATISET. ISSN 1046-8188.

McCreadie:2018:EDE

[MSMO18] Richard McCreadie, Rodrygo L. T. Santos, Craig Macdonald, and Iadh Ounis. Explicit diversification of event aspects for temporal summarization. ACM Transactions on Information Systems, 36(3):25:1– 25:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Macdonald:2013:ALM

[MSOH13] Craig Macdonald, Rodrygo L. T. Santos, Iadh Ounis, and Ben He. About learning models with multiple query-dependent features. ACM Transactions on Information Systems, 31(3):11:1– 11:??, July 2013. CODEN ATISET. ISSN 1046-8188.

Middleton:2004:OUP

[MSR04] Stuart E. Middleton, Nigel R. Shadbolt, and David C. De Roure. Ontological user profiling in recommender systems. ACM Transactions on Information Systems, 22(1): 54–88, January 2004. CODEN ATISET. ISSN 1046-8188.

Manioudakis:2021:FSO

[MT21] Kostas Manioudakis and Yannis Tzitzikas. Faceted search with object ranking and answer size constraints. ACM Transactions on Information Systems, 39(1):9:1–9:33, January 2021. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3425603.

Mendoza:2021:BSI

[MTC21] Marcelo Mendoza, Maurizio Tesconi, and Stefano Cresci. Bots in social and interaction networks: Detection and impact estimation. ACM Transactions on Information Systems, 39(1):5:1–5:32, January 2021. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3419369.

Matsuoka:1992:GFB

[MTKY92] Satoshi Matsuoka, Shin Takahashi, Tomihisa Kamada, and Akinori Yonezawa. A general framework for bidirectional translation between abstract and pictorial data. ACM Transactions on Information Systems, 10(4):408–437, October 1992. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Mackenzie:2023:EDT

[MTL23] Joel Mackenzie, Andrew Trotman, and Jimmy Lin. Efficient document-at-a-time and score-at-a-time query evaluation for learned sparse representations. ACM Transactions on Information Systems, 41(4):96:1–96:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3576922.

Ma:2024:ISI

[MWA^{+}24] Yixiao Ma, Yueyue Wu, Qingyao Ai, Yiqun Liu, Yunqiu Shao, Min Zhang, and Shaoping Ma. Incorporating structural information into legal case retrieval. ACM Transactions on Information Systems, 42(2):40:1– 40:??, March 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3609796.

Mao:2025:RPP

[MWC^{+}25] Wenyu Mao, Jiancan Wu, Weijian Chen, Chongming Gao, Xiang Wang, and Xiangnan He. Reinforced prompt personalization for recommendation with large language models. ACM Transactions on Information Systems, 43(3):72:1–72:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Meng:2001:HSE

[MWYL01] Weiyi Meng, Zonghuan Wu, Clement Yu, and Zhuogang Li. A highly scalable and effective method for metasearch. ACM Transactions on Information Systems, 19(3):310–335, July 2001. CODEN ATISET. ISSN 1046-8188.

Ma:2025:HNR

[MWZ^{+}25] Jingwei Ma, Jiahui Wen, Lei Zhu, Mingyang Zhong, Yang Xu, Lei Guo, and Hongzhi Yin. HGDNet: De-noised review-based rating prediction using hierarchical gating and discriminative networks. ACM Transactions on Information Systems, 43(5):140:1– 140:26, September 2025. CO-DEN ATISET. ISSN 1046-8188.

Ma:2024:TSL

[MXM^{+}24] Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Jie Zhou. Triple sequence learning for cross-domain recommendation. ACM Transactions on Information Systems, 42(4): 91:1–91:??, July 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3638351.

Myers:1990:NMH

[Mye90] Brad A. Myers. A new model for handling input. ACM Transactions on Information Systems, 8(3):289–320, July 1990. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Mei:2011:CVR

[MYHL11] Tao Mei, Bo Yang, Xian-Sheng Hua, and Shipeng Li. Contextual video recommendation by multimodal relevance and user feedback. ACM Transactions on Information Systems, 29(2):10:1– 10:??, April 2011. CODEN ATISET. ISSN 1046-8188.

Meng:2023:RNT

[MYL^{+}23] Qing Meng, Hui Yan, Bo Liu, Xiangguo Sun, Mingrui Hu, and Jiuxin Cao. Recognize news transition from collective behavior for news recommendation. ACM Transactions on Information Systems, 41(4):93:1–93:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3578362.

Moffat:1996:SII

[MZ96] Alistair Moffat and Justin Zobel. Self-indexing inverted files for fast text retrieval. ACM Transactions on Information Systems, 14(4):349– 379, October 1996. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/moffat.
                                        html.

Moffat:2008:RBP

[MZ08] Alistair Moffat and Justin Zobel. Rank-biased precision for measurement of retrieval effectiveness. ACM Transactions on Information Systems, 27(1):2:1–2:??, December 2008. CODEN ATISET. ISSN 1046-8188.

Meng:2024:CFK

[MZG^{+}24] Chang Meng, Ziqi Zhao, Wei Guo, Yingxue Zhang, Haolun Wu, Chen Gao, Dong Li, Xiu Li, and Ruiming Tang. Coarse-to-fine knowledge-enhanced multiinterest learning framework for multi-behavior recommendation. ACM Transactions on Information Systems, 42 (1):30:1–30:??, January 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3606369.

Ma:2011:IRS

[MZLK11] Hao Ma, Tom Chao Zhou, Michael R. Lyu, and Irwin King. Improving recommender systems by incorporating social contextual information. ACM Transactions on Information Systems, 29 (2):9:1–9:??, April 2011. CO-DEN ATISET. ISSN 1046-8188.

Na:2015:TSD

[Na15] Seung-Hoon Na. Two-stage document length normalization for information retrieval. ACM Transactions on Information Systems, 33(2):8:1– 8:??, February 2015. CODEN ATISET. ISSN 1046-8188.

Navarro:1997:PNM

[NB97] Gonzalo Navarro and Ricardo Baeza- Yates. Proximal nodes: a model to query document databases by content and structure. ACM Transactions on Information Systems, 15(4):400–435, October 1997. CODEN ATISET. ISSN 1046-8188.

Nong:2015:ISS

[NCHW15] Ge Nong, Wai Hong Chan, Sheng Qing Hu, and Yi Wu. Induced sorting suffixes in external memory. ACM Transactions on Information Systems, 33(3):12:1–12:??, February 2015. CODEN ATISET. ISSN 1046-8188.

Nong:2014:SAC

[NCZG14] Ge Nong, Wai Hong Chan, Sen Zhang, and Xiao Feng Guan. Suffix array construction in external memory using D-critical substrings. ACM Transactions on Information Systems, 32(1):1:1–1:??, January 2014. CODEN ATISET. ISSN 1046-8188.

Niu:2019:UFS

[NFZ19] Xi Niu, Xiangyu Fan, and Tao Zhang. Understanding faceted search from data science and human factor perspectives. ACM Transactions on Information Systems, 37(2):14:1–14:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3284101.

Nguyen:2017:SIS

[NGMD17] Hung T. Nguyen, Preetam Ghosh, Michael L. Mayo, and Thang N. Dinh. Social influence spectrum at scale: Near-optimal solutions for multiple budgets at once. ACM Transactions on Information Systems, 36(2):14:1– 14:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Nie:2020:LSQ

[NLF^{+}20] Liqiang Nie, Yongqi Li, Fuli Feng, Xuemeng Song, Meng Wang, and Yinglong Wang. Large-scale question tagging via joint questiontopic embedding learning. ACM Transactions on Information Systems, 38(2):20:1– 20:23, March 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3380954.

Niu:2024:PAI

[NLL^{+}24] Yanrui Niu, Chao Liang, Ankang Lu, Baojin Huang, Zhongyuan Wang, and Jiahao Guo. Person-action instance search in story videos: an experimental study. ACM Transactions on Information Systems, 42(2):46:1– 46:??, March 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3617892.

Nong:2013:PLT

[Non13] Ge Nong. Practical lineartime O(1)-workspace suffix sorting for constant alphabets. ACM Transactions on Information Systems, 31(3): 15:1–15:??, July 2013. CO-DEN ATISET. ISSN 1046-8188.

Naskar:2020:EDP

[NSK^{+}20] Debashis Naskar, Sanasam Ranbir Singh, Durgesh Kumar, Sukumar Nandi, and Eva Onaindia de la Rivaherrera. Emotion dynamics of public opinions on Twitter. ACM Transactions on Information Systems, 38(2):18:1– 18:24, March 2020. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/abs/10.1145/3379340.

Nelissen:2018:STU

[NSVB18] Klaas Nelissen, Monique Snoeck, Seppe Vanden Broucke, and Bart Baesens. Swipe and tell: Using implicit feedback to predict user engagement on tablets. ACM Transactions on Information Systems, 36 (4):35:1–35:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Nardini:2022:FFS

[NTV22] Franco Maria Nardini, Roberto Trani, and Rossano Venturini. Fast filtering of search results sorted by attribute. ACM Transactions on Information Systems, 40(2):40:1– 40:24, April 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3477982.

Nie:2012:OIS

[NWZC12] Liqiang Nie, Meng Wang, Zheng-Jun Zha, and Tat-Seng Chua. Oracle in image search: a content-based approach to performance prediction. ACM Transactions on Information Systems, 30(2): 13:1–13:??, May 2012. CO-DEN ATISET. ISSN 1046-8188.

Ni:2024:MCD

[NXF^{+}24] Yuxin Ni, Yunwen Xia, Hui Fang, Chong Long, Xinyu Kong, Daqian Li, Dong Yang, and Jie Zhang. Meta-CRS: a dynamic meta-learning approach for effective conversational recommender system. ACM Transactions on Information Systems, 42(1):28:1– 28:??, January 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3604804.

Ni:2024:CPS

[NXP^{+}24] Xuelian Ni, Fei Xiong, Shirui Pan, Jia Wu, Liang Wang, and Hongshu Chen. Community preserving social recommendation with cyclic transfer learning. ACM Transactions on Information Systems, 42(3):67:1–67:??, May 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3631115.

Nie:2014:LRD

[NZW^{+}14] Liqiang Nie, Yi-Liang Zhao, Xiangyu Wang, Jialie Shen, and Tat-Seng Chua. Learning to recommend descriptive tags for questions in social forums. ACM Transactions on Information Systems, 32(1): 5:1–5:??, January 2014. CO-DEN ATISET. ISSN 1046-8188.

Orlikowski:1994:TFM

[OG94] Wanda J. Orlikowski and Debra C. Gash. Technological frames: Making sense of information technology in organizations. ACM Transactions on Information Systems, 12 (2):174–207, April 1994. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Olson:1989:WHC

[Ols89] Margrethe H. Olson. Work at home for computer professionals. current attitudes and future prospects. ACM Transactions on Information Systems, 7(4):317–338, October 1989. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Oosterhuis:2023:DRE

[Oos23] Harrie Oosterhuis. Doubly robust estimation for correcting position bias in click feedback for unbiased learning to rank. ACM Transactions on Information Systems, 41(3): 61:1–61:??, July 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3569453.

Olson:1993:GCC

[OOSC93] Judith S. Olson, Gary M. Olson, Marianne Storrosten, and Mark Carter. Groupwork close up: a comparison of the group design process with and without a simple group editor. ACM Transactions on Information Systems, 11(4):321–348, October 1993. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Computer-Supported Cooperative Work (CSCW).

Oberweis:1996:ISB

[OS96] Andreas Oberweis and Peter Sander. Information system behavior specification by high-level Petri nets. ACM Transactions on Information Systems, 14(4):380– 420, October 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/oberweis.
html.

Oard:2019:JME

[OSV19] Douglas W. Oard, Fabrizio Sebastiani, and Jyothi K. Vinjumur. Jointly minimizing the expected costs of review for responsiveness and privilege in e-discovery. ACM Transactions on Information Systems, 37(1):11:1–11:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Owei:2002:IAH

[Owe02]  Vesper Owei.  An intel-

ligent approach to handling imperfect information in concept-based natural language queries. ACM Transactions on Information Systems, 20(3):291–328, 2002. CODEN ATISET. ISSN 1046-8188.

Pal:2015:MAR

[Pal15] Aditya Pal. Metrics and algorithms for routing questions to user communities. ACM Transactions on Information Systems, 33(3):14:1– 14:??, March 2015. CODEN ATISET. ISSN 1046-8188.

Paik:2022:TMP

[PARS22] Jiaul H. Paik, Yash Agrawal, Sahil Rishi, and Vaishal Shah. Truncated models for probabilistic weighted retrieval. ACM Transactions on Information Systems, 40(3):56:1– 56:24, July 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3476837.

Pernici:1989:CTA

[PBF^{+}89] B. Pernici, F. Barbic, M. G. Fugini, R. Maiocchi, J. R. Rames, and C. Rolland. C-TODOS: An automatic tool for office system conceptual design. ACM Transactions on Information Systems, 7(4): 378–419, October 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Papadopoulos:2016:OSI

[PBJ^{+}16] Symeon Papadopoulos, Kalina Bontcheva, Eva Jaho, Mihai Lupu, and Carlos Castillo. Overview of the special issue on trust and veracity of information in social media. ACM Transactions on Information Systems, 34(3):14:1– 14:??, May 2016. CODEN ATISET. ISSN 1046-8188.

Patil:2022:GTA

[PBS22] Shubham Patil, Debopriyo Banerjee, and Shamik Sural. A graph theoretic approach for multi-objective budget constrained capsule wardrobe recommendation. ACM Transactions on Information Systems, 40(1):1:1–1:33, January 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3457182.

Pan:2022:CGL

[PCC^{+}22] Zhiqiang Pan, Fei Cai, Wanyu Chen, Chonghao Chen, and Honghui Chen. Collaborative graph learning for session-based recommendation. ACM Transactions on Information Systems, 40 (4):72:1–72:26, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3490479.

Pan:2022:GCA

[PCCC22] Zhiqiang Pan, Fei Cai, Wanyu Chen, and Honghui Chen. Graph co-attentive sessionbased recommendation. ACM Transactions on Information Systems, 40(4):67:1–67:31, October 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3486711.

Pausch:1992:LLS

[PCD92] Randy Pausch, Matthew Conway, and Robert DeLine. Lessons learned from SUIT, the Simple User Interface Toolkit. ACM Transactions on Information Systems, 10 (4):320–344, October 1992. CODEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Powell:2003:CPC

[PF03] Allison L. Powell and James C.

French. Comparing the performance of collection selection algorithms. ACM Transactions on Information Systems, 21(4):412–456, October 2003. CODEN ATISET. ISSN 1046-8188.

Pan:2025:LDG

[PGC^{+}25] Zhiqiang Pan, Chen Gao, Fei Cai, Honghui Chen, and Yong Li. Light dynamic graph learning on temporal networks. ACM Transactions on Information Systems, 43 (5):138:1–138:27, September 2025. CODEN ATISET. ISSN 1046-8188.

Piwowarski:2007:PRU

[PGD07] B. Piwowarski, P. Gallinari, and G. Dupret. Precision recall with user modeling (PRUM): Application to structured information retrieval. ACM Transactions on Information Systems, 25(1): 1:1–1:??, February 2007. CO-DEN ATISET. ISSN 1046-8188.

Peng:2017:PMT

[PGW^{+}17] Min Peng, Wang Gao, Hua Wang, Yanchun Zhang, Jiajia Huang, Qianqian Xie, Gang Hu, and Gang Tian. Parallelization of massive textstream compression based on compressed sensing. ACM Transactions on Information Systems, 36(2):17:1–17:??, September 2017. CODEN ATISET. ISSN 1046-8188.

Peng:2025:DAL

[PGZ^{+}25] Yingtao Peng, Chen Gao, Yu Zhang, Tangpeng Dan, Xiaoyi Du, Hengliang Luo, Yong Li, and Xiaofeng Meng. Denoising alignment with large language model for recommendation. ACM Transactions on Information Systems, 43(2):32:1–32:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3696662.

Pinto:2007:NXM

[PH07] Alberto Pinto and Goffredo Haus. A novel XML music information retrieval method using graph invariants. ACM Transactions on Information Systems, 25(4):19:1– 19:??, October 2007. CODEN ATISET. ISSN 1046-8188.

Pal:2012:EQS

[PHK12] Aditya Pal, F. Maxwell Harper, and Joseph A. Konstan. Exploring question selection bias to identify experts and potential experts in community question answering. ACM Transactions on Information Systems, 30(2):10:1– 10:??, May 2012. CODEN ATISET. ISSN 1046-8188.

Poulovassilis:1994:NGM

[PL94] Alexandra Poulovassilis and Mark Levene. A nested-graph model for the representation and manipulation of complex objects. ACM Transactions on Information Systems, 12 (1):35–68, January 1994. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Pan:2022:PSA

[PLR^{+}22] Yaoxin Pan, Shangsong Liang, Jiaxin Ren, Zaiqiao Meng, and Qiang Zhang. Personalized, sequential, attentive, metric-aware product search. ACM Transactions on Information Systems, 40(2):36:1– 36:29, April 2022. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3473337.

Pu:2024:EEL

[PLS^{+}24] Yanjun Pu, Fang Liu, Rongye Shi, Haitao Yuan, Ruibo Chen, Tianhao Peng, and Wenjun Wu. ELAKT: Enhancing locality for attentive knowledge tracing. ACM Transactions on Information Systems, 42(4):112:1– 112:??, July 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3652601.

Papadias:2001:AST

[PMD01] Dimitris Papadias, Nikos Mamoulis, and Vasilis Delis. Approximate spatio-temporal retrieval. ACM Transactions on Information Systems, 19 (1):53–96, 2001. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org/
                                        pubs/articles/journals/
                                        tois/2001-19-1/p53-papadias/
                                        p53-papadias.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tois/2001-19-1/
                                        p53-papadias/.

Paik:2011:GEE

[PMPJ11] Jiaul H. Paik, Mandar Mitra, Swapan K. Parui, and Kalervo Järvelin. GRAS: an effective and efficient stemming algorithm for information retrieval. ACM Transactions on Information Systems, 29(4): 19:1–19:??, December 2011. CODEN ATISET. ISSN 1046-8188.

Park:2004:ALS

[PPGK04] Seung-Taek Park, David M.

Pennock, C. Lee Giles, and Robert Krovetz. Analysis of lexical signatures for improving information persistence on the World Wide Web. ACM Transactions on Information Systems, 22(4):540–572, October 2004. CODEN ATISET. ISSN 1046-8188.

Paik:2013:ERQ

[PPPR13] Jiaul H. Paik, Swapan K.

Parui, Dipasree Pal, and Stephen E. Robertson. Effective and robust query-based stemming. ACM Transactions on Information Systems, 31(4):18:1–18:??, November 2013. CODEN ATISET. ISSN 1046-8188.

Park:2004:ISI

[PR04] Jinsoo Park and Sudha Ram.

Information systems interoperability: What lies beneath? ACM Transactions on Information Systems, 22(4):595– 632, October 2004. CODEN ATISET. ISSN 1046-8188.

Park:2009:ALS

[PR09] Laurence A. F. Park and Kotagiri Ramamohanarao. An analysis of latent semantic term self-correlation. ACM Transactions on Information Systems, 27(2):8:1–8:??, February 2009. CODEN ATISET. ISSN 1046-8188.

Piao:2016:SFA

[PR16] Minghao Piao and Keun Ho Ryu. Subspace frequency analysis-based field indices extraction for electricity customer classification. ACM Transactions on Information Systems, 34(2):12:1–12:??, April 2016. CODEN ATISET. ISSN 1046-8188.

Park:2005:NDR

[PRP05] Laurence A. F. Park, Kotagiri Ramamohanarao, and Marimuthu Palaniswami. A novel document retrieval method using the discrete wavelet transform. ACM Transactions on Information Systems, 23(3):267–298, July 2005. CODEN ATISET. ISSN 1046-8188.

Pant:2005:LCC

[PS05] Gautam Pant and Padmini Srinivasan. Learning to crawl: Comparing classification schemes. ACM Transactions on Information Systems, 23(4):430–462, October 2005. CODEN ATISET. ISSN 1046-8188.

Plaisant:1999:IDA

[PSDB99] Catherine Plaisant, Ben Shneiderman, Khoa Doan, and Tom Bruns. Interface and data architecture for query preview in networked information systems. ACM Transactions on Information Systems, 17(3):320, July 1999. CODEN ATISET. ISSN 1046-

8188. URL http://www.
acm.org/pubs/citations/
journals/tois/1999-17-3/
p320-plaisant/.

Petersen:2016:PLD

[PSL16] Casper Petersen, Jakob Grue Simonsen, and Christina Lioma. Power law distributions in information retrieval. ACM Transactions on Information Systems, 34(2):8:1–8:??, April 2016. CODEN ATISET. ISSN 1046-8188.

Peng:2024:LMR

[PSM24] Shaowen Peng, Kazunari Sugiyama, and Tsunenori Mine. Less is more: Removing redundancy of graph convolutional networks for recommendation. ACM Transactions on Information Systems, 42(3):85:1–85:??, May 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3632751.

Puppin:2010:TCS

[PSPBY10] Diego Puppin, Fabrizio Silvestri, Raffaele Perego, and Ricardo Baeza-Yates. Tuning the capacity of search engines:

Load-driven routing and incremental caching to reduce and balance the load. ACM Transactions on Information Systems, 28(2):5:1–5:??, May 2010. CODEN ATISET. ISSN 1046-8188.

Pirkola:2007:FBI

[PTKJ07] Ari Pirkola, Jarmo Toivonen, Heikki Keskustalo, and Kalervo Järvelin. Frequencybased identification of correct translation equivalents (FITE) obtained through transformation rules. ACM Transactions on Information Systems, 26(1):2:1–2:??, November 2007. CODEN ATISET. ISSN 1046-8188.

Piwowarski:2008:SCR

[PTL08] Benjamin Piwowarski, Andrew Trotman, and Mounia Lalmas. Sound and complete relevance assessment for XML retrieval. ACM Transactions on Information Systems, 27 (1):1:1–1:??, December 2008. CODEN ATISET. ISSN 1046-8188.

Pan:2013:TJE

[PTS13] Sinno Jialin Pan, Zhiqiang Toh, and Jian Su. Transfer joint embedding for crossdomain named entity recognition. ACM Transactions on Information Systems, 31(2): 7:1–7:??, May 2013. CODEN ATISET. ISSN 1046-8188.

Pibiri:2017:CEF

[PV17] Giulio Ermanno Pibiri and Rossano Venturini. Clustered Elias–Fano indexes. ACM Transactions on Information Systems, 36(1):2:1–2:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Pibiri:2019:HMG

[PV19] Giulio Ermanno Pibiri and Rossano Venturini. Handling massive N-gram datasets efficiently. ACM Transactions on Information Systems, 37 (2):25:1–25:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3302913.

Parameswaran:2011:RSC

[PVGM11] Aditya Parameswaran, Petros Venetis, and Hector Garcia-Molina. Recommendation systems with complex constraints: a course recommendation perspective. ACM Transactions on Information Systems, 29(4):20:1–20:??, December 2011. CODEN ATISET. ISSN 1046-8188.

Peng:2025:PEF

[PXW^{+}25] Qiyao Peng, Hongyan Xu, Yinghui Wang, Hongtao Liu, Cuiying Huo, and Wenjun Wang. PEPT: Expert finding meets personalized pretraining. ACM Transactions on Information Systems, 43 (1):8:1–8:??, January 2025. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3690380.

Pan:2019:TRH

[PYC^{+}19] Weike Pan, Qiang Yang, Wanling Cai, Yaofeng Chen, Qing Zhang, Xiaogang Peng, and Zhong Ming. Transfer to rank for heterogeneous oneclass collaborative filtering. ACM Transactions on Information Systems, 37(1):10:1– 10:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Pei:2024:MLR

[PYDR24] Jiahuan Pei, Guojun Yan, Maarten De Rijke, and Pengjie Ren. Mixture-of-languages routing for multilingual dialogues. ACM Transactions on Information Systems, 42(6):165:1–165:??, November 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3676956.

Palaniappan:1992:EFO

[PYF^{+}92] Murugappan Palaniappan, Nicole Yankelovich, George Fitzmaurice, Anne Loomis, Bernard Haan, James Coombs, and Norman Meyrowitz. The envoy framework: An open architecture for agents. ACM Transactions on Information Systems, 10(3):233–264, July 1992. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

Peng:2022:RNS

[PZD^{+}22] Hao Peng, Ruitong Zhang, Yingtong Dou, Renyu Yang, Jingyi Zhang, and Philip S. Yu. Reinforced neighborhood selection guided multirelational graph neural networks. ACM Transactions on Information Systems, 40 (4):69:1–69:46, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3490181.

Peng:2024:USB

[PZH^{+}24] Hao Peng, Jingyun Zhang, Xiang Huang, Zhifeng Hao, Angsheng Li, Zhengtao Yu, and Philip S. Yu. Unsupervised social bot detection via structural information theory. ACM Transactions on Information Systems, 42(6):148:1– 148:??, November 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3660522.

Possas:2005:SBV

[PZMRN05] Bruno Pôssas, Nivio Ziviani, Wagner Meira, Jr., and Berthier Ribeiro-Neto. Setbased vector model: an efficient approach for correlationbased ranking. ACM Transactions on Information Systems, 23(4):397–429, October 2005. CODEN ATISET. ISSN 1046-8188.

Quan:2024:AVL

[QDG^{+}24] Yuhan Quan, Jingtao Ding, Chen Gao, Nian Li, Lingling Yi, Depeng Jin, and Yong Li. Alleviating video-length effect for micro-video recommendation. ACM Transactions on Information Systems, 42(2): 44:1–44:??, March 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3617826.

Qin:2023:GGD

[QDZW23] Xubo Qin, Zhicheng Dou, Yutao Zhu, and Ji-Rong Wen. GDESA: Greedy diversity encoder with self-attention for search results diversification. ACM Transactions on Information Systems, 41(2):34:1– 34:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3544103.

Qu:2019:PBN

[QFZ^{+}19] Yanru Qu, Bohui Fang, Weinan Zhang, Ruiming Tang, Minzhe Niu, Huifeng Guo, Yong Yu, and Xiuqiang He. Product-based neural networks for user response prediction over multi-field categorical data. ACM Transactions on Information Systems, 37(1):5:1–5:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Qin:2024:LHS

[QGW^{+}24] Yingrong Qin, Chen Gao, Shuangqing Wei, Yue Wang, Depeng Jin, Jian Yuan, Lin Zhang, Dong Li, Jianye Hao, and Yong Li. Learning from hierarchical structure of knowledge graph for recommendation. ACM Transactions on Information Systems, 42(1):18:1–18:??, January 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3595632.

Qiu:2022:EPI

[QHCY22] Ruihong Qiu, Zi Huang, Tong Chen, and Hongzhi Yin. Exploiting positional information for session-based recommendation. ACM Transactions on Information Systems, 40(2):35:1–35:24, April 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3473339.

Qiu:2020:ECS

[QHLY20] Ruihong Qiu, Zi Huang, Jingjing Li, and Hongzhi Yin. Exploiting cross-session information for session-based recommendation with graph neural networks. ACM Transactions on Information Systems, 38(3):22:1–22:23, June 2020. CODEN ATISET. ISSN 1046-8188. URL https://dl.

acm.org/doi/abs/10.1145/
3382764.

Qin:2025:PTO

[QJG^{+}25] Yifang Qin, Wei Ju, Yiyang Gu, Ziyue Qiao, Zhiping Xiao, and Ming Zhang. PolyCF: Towards optimal spectral graph filters for collaborative filtering. ACM Transactions on Information Systems, 43(4):

94:1–94:??, July 2025. CO-DEN ATISET. ISSN 1046-8188.

Qian:2019:SRL

[QLNY19] Tieyun Qian, Bei Liu, Quoc Viet Hung Nguyen, and Hongzhi Yin. Spatiotemporal representation learning for translation-based POI recommendation. ACM Transactions on Information Systems, 37(2):18:1–18:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3295499.

Qin:2024:DMP

[QWJ^{+}24] Yifang Qin, Hongjun Wu, Wei Ju, Xiao Luo, and Ming Zhang. A diffusion model for POI recommendation. ACM Transactions on Information Systems, 42(2):54:1– 54:??, March 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3624475.

Quan:2015:LDM

[QWZ^{+}15] Xiaojun Quan, Qifan Wang, Ying Zhang, Luo Si, and Liu Wenyin. Latent discriminative models for social emotion detection with emotional dependency. ACM Transactions on Information Systems, 34(1):2:1–2:??, October 2015. CODEN ATISET. ISSN 1046-

Qu:2025:TMM

[QXX^{+}25] Zekai Qu, Ruobing Xie, Chaojun Xiao, Yuan Yao, Zhiyuan Liu, Fengzong Lian, Zhanhui Kang, and Jie Zhou. Thoroughly modeling multidomain pre-trained recommendation as language. ACM Transactions on Information Systems, 43(2):54:1– 54:??, March 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3708883.

Qin:2023:LRU

[QZS^{+}23] Jiarui Qin, Weinan Zhang, Rong Su, Zhirong Liu, Weiwen Liu, Guangpeng Zhao, Hao Li, Ruiming Tang, Xiuqiang He, and Yong Yu. Learning to retrieve user behaviors for click-through rate estimation. ACM Transactions on Information Systems, 41(4):98:1–98:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3579354.

Qin:2024:ASO

[QZS^{+}24] Chuan Qin, Hengshu Zhu, Dazhong Shen, Ying Sun, Kaichun Yao, Peng Wang, and Hui Xiong. Automatic skill-oriented question generation and recommendation for intelligent job interviews. ACM Transactions on Information Systems, 42(1):27:1– 27:??, January 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3604552.

Qiao:2025:MVI

[QZW^{+}25] Shutong Qiao, Wei Zhou, Junhao Wen, Chen Gao, Qun Luo, Peixuan Chen, and Yong Li. Multi-view intent learning and alignment with large language models for session-based recommendation. ACM Transactions on Information Systems, 43(4): 91:1–91:??, July 2025. CO-DEN ATISET. ISSN 1046-8188.

Qin:2020:ENN

[QZX^{+}20] Chuan Qin, Hengshu Zhu, Tong Xu, Chen Zhu, Chao Ma, Enhong Chen, and Hui Xiong. An enhanced neural network approach to personjob fit in talent recruitment. ACM Transactions on Information Systems, 38(2):15:1– 15:33, March 2020. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3376927.

Qian:2006:SPB

[QZXP06] Gang Qian, Qiang Zhu, Qiang Xue, and Sakti Pramanik. A space-partitioning-based indexing method for multidimensional non-ordered discrete data spaces. ACM Transactions on Information Systems, 24(1):79–110, January 2006. CODEN ATISET. ISSN 1046-8188.

Rahmani:2022:SAI

[RABC22] Hossein A. Rahmani, Mohammad Aliannejadi, Mitra Baratchi, and Fabio Crestani. A systematic analysis on the impact of contextual information on pointof-interest recommendation. ACM Transactions on Information Systems, 40(4):88:1– 88:35, October 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3508478.

Rada:1992:CTH

[Rad92] Roy Rada. Converting a textbook to hypertext. ACM Transactions on Information Systems, 10(3):294–315, July 1992. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Raman:2014:UID

[RBCT14] Karthik Raman, Paul N. Bennett, and Kevyn Collins-Thompson. Understanding intrinsic diversity in Web search: Improving wholesession relevance. ACM Transactions on Information Systems, 32(4):20:1– 20:??, October 2014. CODEN ATISET. ISSN 1046-8188.

Rizzo:2023:RMT

[RBM23] Stefano Giovanni Rizzo, Matteo Brucato, and Danilo Montesi. Ranking models for the temporal dimension of text. ACM Transactions on Information Systems, 41(2):

49:1–49:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3565481.

Rodriguez:2009:AMG

[RBV09] Marko A. Rodriguez, Johan Bollen, and Herbert Van De Sompel. Automatic metadata generation using associative networks. ACM Transactions on Information Systems, 27(2):7:1–7:??, February 2009. CODEN ATISET. ISSN 1046-8188.

Ren:2018:SRE

[RCR^{+}18] Pengjie Ren, Zhumin Chen, Zhaochun Ren, Furu Wei, Liqiang Nie, Jun Ma, and Maarten de Rijke. Sentence relations for extractive summarization with deep neural networks. ACM Transactions on Information Systems, 36 (4):39:1–39:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Ren:2021:CSE

[RCR^{+}21] Pengjie Ren, Zhumin Chen, Zhaochun Ren, Evangelos Kanoulas, Christof Monz, and Maarten De Rijke. Conversations with search engines: SERP-based conversational response generation. ACM Transactions on Information Systems, 39(4):47:1– 47:29, October 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3432726.

Rau:2024:RBW

[RDK24] David Rau, Mostafa Dehghani, and Jaap Kamps. Revisiting bag of words document representations for efficient ranking with transformers. ACM Transactions on Information Systems, 42 (5):114:1–114:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3640460.

Resnick:1993:PBC

[Res93] Paul Resnick. Phone-based CSCW: Tools and trials. ACM Transactions on Information Systems, 11(4):401– 424, October 1993. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Computer-Supported Cooperative Work (CSCW).

Rad:2024:VNA

[RFB^{+}24] Radin Hamidi Rad, Hossein Fani, Ebrahim Bagheri, Mehdi Kargar, Divesh Srivastava, and Jaroslaw Szlichta. A variational neural architecture for skill-based team formation. ACM Transactions on Information Systems, 42 (1):7:1–7:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3589762.

Raghavan:1989:CIR

[RJB89] Vijay V. Raghavan, Gwang S.

Jung, and Peter Bollmann.

A critical investigation of recall and precision as measures of retrieval system performance. ACM Transactions on Information Systems, 7(3): 205–229, July 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Raiber:2019:RFW

[RK19] Fiana Raiber and Oren Kurland. Relevance feedback: The whole is inferior to the sum of its parts. ACM Transactions on Information Systems, 37(4):44:1– 44:??, December 2019. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.
acm.org/ft_gateway.cfm?
id=3360487.

Riloff:1994:IEB

[RL94] Ellen Riloff and Wendy Lehnert. Information extraction as a basis for high-precision text classification. ACM Transactions on Information Systems, 12(3):296–333, July 1994. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Text Categorization.

Ruan:2025:CCP

[RLC^{+}25] Shulan Ruan, Huijie Liu, Zhao Chen, Bin Feng, Kun Zhang, Caleb Chen Cao, Enhong Chen, and Lei Chen. CPWS: Confident programmatic weak supervision for high-quality data labeling. ACM Transactions on Information Systems, 43(4):87:1– 87:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Roitero:2024:HMC

[RLS^{+}24] Kevin Roitero, David La Barbera, Michael Soprano, Gianluca Demartini, Stefano Mizzaro, and Tetsuya Sakai. How many crowd workers do I need? On statistical power when crowdsourcing relevance judgments. ACM Transactions on Information Systems, 42(1):21:1– 21:??, January 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3597201.

Robey:1996:SPI

[RN96] Daniel Robey and Michael Newman. Sequential patterns in information systems development: An application of a social process model. ACM Transactions on Information Systems, 14(1):30– 63, January 1996. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/robey.
html.

Rowe:1996:ULO

[Row96] Neil C. Rowe. Using local optimality criteria for efficient information retrieval with redundant information filters. ACM Transactions on Information Systems, 14(2):138– 174, April 1996. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/rowe.
                                        html.

Romm:1998:EMC

[RP98] Celia T. Romm and Nava Pliskin. Electronic mail as a coalition-building information technology. ACM Transactions on Information Systems, 16(1):82–100, January 1998. CODEN ATISET. ISSN 1046-8188.

Ruotsalo:2018:IIM

[RPE^{+}18] Tuukka Ruotsalo, Jaakko Peltonen, Manuel J. A. Eugster, Dorota Głowacka, Patrik Floréen, Petri Myllymäki, Giulio Jacucci, and Samuel Kaski. Interactive intent modeling for exploratory search. ACM Transactions on Information Systems, 36(4):44:1– 44:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Ram:1998:CCS

[RR98] Sudha Ram and V. Ramesh. Collaborative conceptual schema design: a process model and prototype system. ACM Transactions on Information Systems, 16(4):347–371, October 1998. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tois/1998-16-4/p347-ram/.

Rangan:1995:FTC

[RRS95] P. Venkat Rangan, Srinivas Ramanathan, and Srihari Sampathkumar. Feedback techniques for continuity and synchronization in multimedia information retrieval. ACM Transactions on Information Systems, 13(2):145– 176, April 1995. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80.

Rama:1993:ICR

[RS93] D. V. Rama and Padmini Srinivasan. An investigation of content representation using text grammars. ACM Transactions on Information Systems, 11(1):51–75, January 1993. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Rus:1997:CIC

[RS97] Daniela Rus and Devika Subramanian. Customizing information capture and access. ACM Transactions on Information Systems, 15(1):67– 101, January 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/rus.html.

Radinsky:2013:BDW

[RSD^{+}13] Kira Radinsky, Krysta M.

Svore, Susan T. Dumais, Milad Shokouhi, Jaime Teevan, Alex Bocharov, and Eric Horvitz. Behavioral dynamics on the Web: Learning, modeling, and prediction. ACM Transactions on Information Systems, 31(3):16:1– 16:??, July 2013. CODEN ATISET. ISSN 1046-8188.

Rosaci:2009:MDR

[RSG09] Domenico Rosaci, Giuseppe M. L. Sarné, and Salvatore Garruzzo. MUADDIB: a distributed recommender system supporting device adaptivity. ACM Transactions on Information Systems, 27(4):24:1– 24:??, November 2009. CO-DEN ATISET. ISSN 1046-8188.

Ruhleder:1994:RLR

[Ruh94] Karen Ruhleder. Rich and lean representations of information for knowledge work: The role of computing packages in the work of classical scholars. ACM Transactions on Information Systems, 12 (2):208–230, April 1994. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Razgallah:2024:UNG

[RVA^{+}24] Hédi Razgallah, Michalis Vlachos, Ahmad Ajalloeian, Ninghao Liu, Johannes Schneider, and Alexis Steinmann. Using neural and graph neural recommender systems to overcome choice overload: Evidence from a music education platform. ACM Transactions on Information Systems, 42(4):92:1–92:??, July 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3637873.

Ren:2020:CCR

[RYC^{+}20] Xuhui Ren, Hongzhi Yin, Tong Chen, Hao Wang, Nguyen Quoc Viet Hung, Zi Huang, and Xiangliang Zhang. CRSAL: Conversational recommender systems with adversarial learning. ACM Transactions on Information Systems, 38(4): 34:1–34:40, October 2020. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3394592.

Rosen-Zvi:2010:LAT

[RZCG^{+}10] Michal Rosen-Zvi, Chaitanya Chemudugunta, Thomas Griffiths, Padhraic Smyth, and Mark Steyvers. Learning author-topic models from text corpora. ACM Transactions on Information Systems, 28 (1):4:1–4:??, January 2010. CODEN ATISET. ISSN 1046-8188.

Rashidi:2024:IJV

[RZM24] Lida Rashidi, Justin Zobel, and Alistair Moffat. The impact of judgment variability on the consistency of offline effectiveness measures. ACM Transactions on Information Systems, 42(1):19:1– 19:??, January 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3596511.

Siro:2024:UPU

[SAD24] Clemencia Siro, Mohammad Aliannejadi, and Maarten De Rijke. Understanding and predicting user satisfaction with conversational recommender systems. ACM Transactions on Information Systems, 42(2):55:1–55:??, March 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3624989.

Somanchi:2023:EUH

[SAK^{+}23] Sriram Somanchi, Ahmed Abbasi, Ken Kelley, David Dobolyi, and Ted Tao Yuan. Examining user heterogeneity in digital experiments. ACM Transactions on Information Systems, 41(4):100:1– 100:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3578931.

Savoy:2012:AAB

[Sav12] Jacques Savoy. Authorship attribution based on specific vocabulary. ACM Transactions on Information Systems, 30(2):12:1–12:??, May 2012. CODEN ATISET. ISSN 1046-8188.

Su:2025:PTL

[SAW^{+}25] Weihang Su, Qingyao Ai, Yueyue Wu, Anzhe Xie, Changyue Wang, Yixiao Ma, Haitao Li, Zhijing Wu, Yiqun Liu, and Min Zhang. Pretraining for legal case retrieval based on inter-case distinctions. ACM Transactions on Information Systems, 43 (5):129:1–129:27, September 2025. CODEN ATISET. ISSN 1046-8188.

Stevenson:2024:SMT

[SBH24] Mark Stevenson and Reem Bin-Hezam. Stopping methods for technology-assisted reviews based on point processes. ACM Transactions on Information Systems, 42(3): 73:1–73:??, May 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3631990.

Sadeghi:2017:RFB

[SBM^{+}17] Seyedeh Sargol Sadeghi, Roi Blanco, Peter Mika, Mark Sanderson, Falk Scholer, and David Vallet. Re-finding behaviour in vertical domains. ACM Transactions on Information Systems, 35(3):21:1– 21:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Si:2003:SLM

[SC03] Luo Si and Jamie Callan.

A semisupervised learning method to merge search engine results. ACM Transactions on Information Systems, 21(4):457–491, October 2003. CODEN ATISET. ISSN 1046-8188.

Schumaker:2009:TAS

[SC09] Robert P. Schumaker and Hsinchun Chen. Textual analysis of stock market prediction using breaking financial news: The AZFin text system. ACM Transactions on Information Systems, 27(2):

12:1–12:??, February 2009. CODEN ATISET. ISSN 1046-8188.

Safran:2019:ELB

[SC19] Mejdl Safran and Dunren Che. Efficient learning-based recommendation algorithms for top- n tasks and top- n workers in large-scale crowdsourcing systems. ACM Transactions on Information Systems, 37(1):2:1–2:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Sousa:2019:RSL

[SCG^{+}19] Daniel Xavier Sousa, Sérgio Canuto, Marcos André Gonçalves, Thierson Couto Rosa, and Wellington Santos Martins. Risk-sensitive learning to rank with evolutionary multiobjective feature selection. ACM Transactions on Information Systems, 37(2):24:1– 24:??, March 2019. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3300196.

Sciore:1989:OS

[Sci89] Edward Sciore. Object specialization. ACM Transactions on Information Systems, 7(2):103–122, April 1989. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Shen:2022:UPB

[SCL^{+}22] Jiaxing Shen, Jiannong Cao, Oren Lederman, Shaojie Tang, and Alex “Sandy” Pentland. User profiling based on nonlinguistic audio data. ACM Transactions on Information Systems, 40(1):17:1– 17:23, January 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3474826.

Su:2024:PAS

[SDZW24] Zhan Su, Zhicheng Dou, Yutao Zhu, and Ji-Rong Wen. Passage-aware search result diversification. ACM Transactions on Information Systems, 42(5):136:1–136:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3653672.

Siochi:1991:CAU

[SE91] Antonio C. Siochi and Roger W.

Ehrich. Computer analysis of user interfaces based on repetition in transcripts of user sessions. ACM Transactions on Information Systems, 9(4): 309–335, October 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Stotts:1989:PNB

[SF89] P. David Stotts and Richard Furuta. Petri net based hypertext: Document structure with browsing semantics. ACM Transactions on Information Systems, 7(1):3– 29, January 1989. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Stotts:1998:HAV

[SFR98] P. David Stotts, Richard Furuta, and Cyrano Ruiz Cabarrus. Hyperdocuments as automata: Verification of tracebased browsing properties by model checking. ACM Transactions on Information Systems, 16(1):1–30, January 1998. CODEN ATISET. ISSN 1046-8188.

Skaggs:2014:TMW

[SG14] Bradley Skaggs and Lise Getoor. Topic modeling for Wikipedia link disambiguation. ACM Transactions on Information Systems, 32(3): 10:1–10:??, June 2014. CO-DEN ATISET. ISSN 1046-8188.

Sen:2022:KWY

[SGJ22] Procheta Sen, Debasis Ganguly, and Gareth J. F. Jones. I know what you need: Investigating document retrieval effectiveness with partial session contexts. ACM Transactions on Information Systems, 40(3):53:1–53:30, July 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3488667.

Shaw:1993:DSV

[SGLS93] Chris Shaw, Mark Green, Jiandong Liang, and Yunqi Sun. Decoupled simulation in virtual reality with the MR toolkit. ACM Transactions on Information Systems, 11 (3):287–317, July 1993. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Virtual Worlds.

Sun:2024:MUS

[SGZ^{+}24] Weiwei Sun, Shuyu Guo, Shuo Zhang, Pengjie Ren, Zhumin Chen, Maarten de Rijke, and Zhaochun Ren. Metaphorical user simulators for evaluating task-oriented dialogue systems. ACM Transactions on Information Systems, 42 (1):17:1–17:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3596510.

Shirakawa:2017:IWG

[SHN17] Masumi Shirakawa, Takahiro Hara, and Shojiro Nishio. IDF for word N-grams. ACM Transactions on Information Systems, 36(1):5:1–5:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Shang:2023:LRT

[SHS^{+}23] Yu-Ming Shang, Heyan Huang, Xin Sun, Wei Wei, and Xian-Ling Mao. Learning relation ties with a forcedirected graph in distant supervised relation extraction. ACM Transactions on Information Systems, 41(1):10:1– 10:??, January 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3520082.

Su:2025:IEF

[SJL^{+}25] Yixin Su, Wei Jiang, Fangquan Lin, Cheng Yang, Sarah M. Erfani, Junhao Gan, Yunxiang Zhao, Ruixuan Li, and Rui Zhang. Intrinsic and extrinsic factor disentanglement for recommendation in various context scenarios. ACM Transactions on Information Systems, 43(3):82:1– 82:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Sun:2025:MAL

[SJZ^{+}25] Ying Sun, Yang Ji, Hengshu Zhu, Fuzhen Zhuang, Qing He, and Hui Xiong. Marketaware long-term job skill recommendation with explainable deep reinforcement learning. ACM Transactions on Information Systems, 43(2): 46:1–46:??, March 2025. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3704998.

Shtok:2012:PQP

[SKC^{+}12] Anna Shtok, Oren Kurland, David Carmel, Fiana Raiber, and Gad Markovits. Predicting query performance by query-drift estimation. ACM Transactions on Information Systems, 30(2):11:1– 11:??, May 2012. CODEN ATISET. ISSN 1046-8188.

Shtok:2016:QPP

[SKC16] Anna Shtok, Oren Kurland, and David Carmel. Query performance prediction using reference lists. ACM Transactions on Information Systems, 34(4):19:1–19:??, September 2016. CODEN ATISET. ISSN 1046-8188.

Sakai:2024:VFE

[SKK24] Tetsuya Sakai, Jin Young Kim, and Inho Kang. A versatile framework for evaluating ranked lists in terms of group fairness and relevance. ACM Transactions on Information Systems, 42(1):11:1– 11:??, January 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3589763.

Stevens:1995:ISI

[SL95] Scott Stevens and Thomas Little. Introduction to the special issue on video information retrieval. ACM Transactions on Information Systems, 13(4):371–372, October 1995. CODEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Su:2024:CDR

[SLD^{+}24] Hongzu Su, Jingjing Li, Zhekai Du, Lei Zhu, Ke Lu, and Heng Tao Shen. Crossdomain recommendation via dual adversarial adaptation. ACM Transactions on Information Systems, 42(3):83:1– 83:??, May 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3632524.

Schnase:1993:SDM

[SLHS93] John L. Schnase, John J. Leggett, David L. Hicks, and Ron L. Szabo. Semantic data modeling of hypermedia associations. ACM Transactions on Information Systems, 11 (1):27–50, January 1993. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Shi:2024:BTN

[SLM^{+}24] Haitao Shi, Meng Liu, Xiaoxuan Mu, Xuemeng Song, Yupeng Hu, and Liqiang Nie. Breaking through the noisy correspondence: a robust model for image-text matching. ACM Transactions on Information Systems, 42(6): 149:1–149:??, November 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3662732.

Sun:2024:CMD

[SLQ24] KE Sun, Chenliang Li, and Tieyun Qian. City matters! A dual-target cross-city sequential POI recommendation model. ACM Transactions on Information Systems, 42(6): 153:1–153:??, November 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3664284.

Sun:2020:ECG

[SLW^{+}20] Xiao Sun, Jia Li, Xing Wei, Changliang Li, and Jianhua Tao. Emotional conversation generation based on a Bayesian deep neural network. ACM Transactions on Information Systems, 38(1): 8:1–8:24, February 2020. CO-DEN ATISET. ISSN 1046-8188.

Shao:2024:ITL

[SLW^{+}24] Yunqiu Shao, Haitao Li, Yueyue Wu, Yiqun Liu, Qingyao Ai, Jiaxin Mao, Yixiao Ma, and Shaoping Ma. An intent taxonomy of legal case retrieval. ACM Transactions on Information Systems, 42(2):62:1–62:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3626093.

Shi:2024:RPB

[SLX^{+}24] Xiaoyu Shi, Quanliang Liu, Hong Xie, Di Wu, Bo Peng, MingSheng Shang, and Defu Lian. Relieving popularity bias in interactive recommendation: a diversity-novelty-aware reinforcement learning approach. ACM Transactions on Information Systems, 42 (2):52:1–52:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3618107.

Shao:2019:AMI

[SLZ^{+}19] Yunqiu Shao, Yiqun Liu, Fan Zhang, Min Zhang, and Shaoping Ma. On annotation methodologies for image search evaluation. ACM Transactions on Information Systems, 37(3):29:1–29:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3309994.

Sun:2024:MCN

[SLZ^{+}24] Zhu Sun, Yu Lei, Lu Zhang, Chen Li, Yew-Soon Ong, and Jie Zhang. A multichannel next POI recommendation framework with multigranularity check-in signals. ACM Transactions on Information Systems, 42(1):15:1– 15:??, January 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3592789.

Sang:2025:AAG

[SLZ^{+}25] Lei Sang, Honghao Li, Yiwen Zhang, Yi Zhang, and Yun Yang. AdaGIN: Adaptive graph interaction network for click-through rate prediction. ACM Transactions on Information Systems, 43(1): 3:1–3:??, January 2025. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3681785.

Strong:1995:EEH

[SM95] Diane M. Strong and Steven M. Miller. Exceptions and exception handling in computerized information processes. ACM Transactions on Information Systems, 13(2):206–233, April 1995. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

Shipman:1999:IFH

[SM99] Frank M. Shipman and Raymond J. McCall. Incremental formalization with the hyperobject substrate. ACM Transactions on Information Systems, 17(2):199–227, April 1999. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org/pubs/citations/
journals/tois/1999-17-2/
p199-shipman/.

Sun:2011:IIR

[SMGM11] Bingjun Sun, Prasenjit Mitra, C. Lee Giles, and Karl T. Mueller. Identifying, indexing, and ranking chemical formulae and chemical names in digital documents. ACM Transactions on Information Systems, 29(2):12:1– 12:??, April 2011. CODEN ATISET. ISSN 1046-8188.

Song:2016:VTP

[SMN^{+}16] Xuemeng Song, Zhao-Yan Ming, Liqiang Nie, Yi-Liang Zhao, and Tat-Seng Chua. Volunteerism tendency prediction via harvesting multiple social networks. ACM Transactions on Information Systems, 34(2):10:1–10:??, April 2016. CODEN ATISET. ISSN 1046-8188.

Sritrakool:2025:QDL

[SMT25] Nakarin Sritrakool, Saranya Maneeroj, and Atsuhiro Takasu. QUADEN: Discovering latent neighbors for sparse users and items across interaction quadrants in recommender system. ACM Transactions on Information Systems, 43(4):92:1–92:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Sander:2005:MAS

[SNS^{+}05] Jörg Sander, Raymond T. Ng, Monica C. Sleumer, Man Saint Yuen, and Steven J. Jones. A methodology for analyzing SAGE libraries for cancer profiling. ACM Transactions on Information Systems, 23(1):35–60, January 2005. CODEN ATISET. ISSN 1046-8188. SilvadeMoura:2000:FFW

[SNZBY00] Edleno Silva de Moura, Gonzalo Navarro, Nivio Ziviani, and Ricardo Baeza-Yates. Fast and flexible word searching on compressed text. ACM Transactions on Information Systems, 18(2):113–139, April 2000. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org/pubs/citations/
journals/tois/2000-18-2/
p113-silva_de_moura/.

Straube:1990:QQP

[SÖ90] Dave D. Straube and M. Tamer Özsu. Queries and query processing in object-oriented database systems. ACM Transactions on Information Systems, 8(4):387–430, October 1990. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Schedl:2011:EMS

[SPKW11] Markus Schedl, Tim Pohle, Peter Knees, and Gerhard Widmer. Exploring the music similarity space on the Web. ACM Transactions on Information Systems, 29(3):14:1– 14:??, July 2011. CODEN ATISET. ISSN 1046-8188.

Shen:2006:QEW

[SPS^{+}06] Dou Shen, Rong Pan, Jian-Tao Sun, Jeffrey Junfeng Pan, Kangheng Wu, Jie Yin, and Qiang Yang. Query enrichment for web-query classification. ACM Transactions on Information Systems, 24(3): 320–352, July 2006. CODEN ATISET. ISSN 1046-8188.

Shen:2022:JRL

[SQZ^{+}22] Dazhong Shen, Chuan Qin, Hengshu Zhu, Tong Xu, Enhong Chen, and Hui Xiong. Joint representation learning with relation-enhanced topic models for intelligent job interview assessment. ACM Transactions on Information Systems, 40(1):15:1– 15:36, January 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3469654.

Shi:2024:DSR

[SRF^{+}24] Chaoyu Shi, Pengjie Ren, Dongjie Fu, Xin Xin, Shansong Yang, Fei Cai, Zhaochun Ren, and Zhumin Chen. Diversifying sequential recommendation with retrospective and prospective transformers. ACM Transactions on Information Systems, 42(5):132:1– 132:??, September 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3653016.

Szczur:1993:TPT

[SS93] Martha R. Szczur and Sylvia B. Sheppard. TAE plus: Transportable applications environment plus: a user interface development environment. ACM Transactions on Information Systems, 11(1):76–101, January 1993. CODEN ATISET. ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Siersdorfer:2008:MMM

[SS08] Stefan Siersdorfer and Sergej Sizov. Meta methods for model sharing in personal information systems. ACM Transactions on Information Systems, 26(4):22:1–22:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Shen:2009:NFE

[SSCT09] Jialie Shen, John Shepherd, Bin Cui, and Kian-Lee Tan. A novel framework for efficient automated singer identification in large music databases. ACM Transactions on Information Systems, 27(3):18:1– 18:??, May 2009. CODEN ATISET. ISSN 1046-8188.

Smith:1989:KBS

[SSGC89] Philip J. Smith, Steven J. Shute, Deb Galdes, and Mark H. Chignell. Knowledgebased search tactics for an intelligent intermediary system. ACM Transactions on Information Systems, 7(3): 246–270, July 1989. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:

80. Special Issue on Research and Development in Information Retrieval.

SanPedro:2011:CRY

[SSS11] Jose San Pedro, Stefan Siersdorfer, and Mark Sanderson. Content redundancy in YouTube and its application to video tagging. ACM Transactions on Information Systems, 29(3):13:1–13:??, July 2011. CODEN ATISET. ISSN 1046-8188.

Song:2017:DLD

[SSY^{+}17] Xuan Song, Ryosuke Shibasaki, Nicholos Jing Yuan, Xing Xie, Tao Li, and Ryutaro Adachi. DeepMob: Learning deep knowledge of human emergency behavior and mobility from big and heterogeneous data. ACM Transactions on Information Systems, 35(4):41:1–41:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Si:2023:ERS

[SSZ^{+}23] Zihua Si, Zhongxiang Sun, Xiao Zhang, Jun Xu, Yang Song, Xiaoxue Zang, and Ji-Rong Wen. Enhancing recommendation with search data in a causal learning manner.

ACM Transactions on Information Systems, 41(4):111:1– 111:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3582425.

Sakai:2024:OPW

[STC^{+}24] Tetsuya Sakai, Sijie Tao, Nuo Chen, Yujing Li, Maria Maistro, Zhumin Chu, and Nicola Ferro. On the ordering of pooled Web pages, gold assessments, and bronze assessments. ACM Transactions on Information Systems, 42 (1):23:1–23:??, January 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3600227.

Salminen:1995:THI

[STSM95] Airi Salminen, Jean Tague-Sutcliffe, and Charles McClellan. From text to hypertext by indexing. ACM Transactions on Information Systems, 13 (1):69–99, January 1995. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Sakai:2022:RAW

[STZ22] Tetsuya Sakai, Sijie Tao, and Zhaohao Zeng. Relevance assessments for Web search evaluation: Should we randomise or prioritise the pooled documents? ACM Transactions on Information Systems, 40 (4):76:1–76:35, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3494833.

Sanderson:1999:IRE

[SV99] Mark Sanderson and C. J.

Van Rijsbergen. The impact on retrieval effectiveness of skewed frequency distributions. ACM Transactions on Information Systems, 17(4): 440–465, October 1999. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.
acm.org/pubs/citations/
journals/tois/1999-17-4/
p440-sanderson/.

Shasha:1990:NTB

[SW90] Dennis Shasha and Tsong-Li Wang. New techniques for best-match retrieval. ACM Transactions on Information Systems, 8(2):140–158, April 1990. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Schauble:1994:EPQ

[SW94] Peter Schauble and Beat Wuthrich. On the expressive power of query languages. ACM Transactions on Information Systems, 12(1):69– 91, January 1994. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80.

Shao:2023:URJ

[SWL^{+}23] Yunqiu Shao, Yueyue Wu, Yiqun Liu, Jiaxin Mao, and Shaoping Ma. Understanding relevance judgments in legal case retrieval. ACM Transactions on Information Systems, 41(3):76:1–76:??, July 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3569929.

Sun:2022:UAA

[SWZ^{+}22] Peijie Sun, Le Wu, Kun Zhang, Yu Su, and Meng Wang. An unsupervised aspect-aware recommendation model with explanation text generation. ACM Transactions on Information Systems, 40(3):63:1–63:29, July 2022. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3483611.

Shao:2024:AUS

[SWZ^{+}24] Pengyang Shao, Le Wu, Kun Zhang, Defu Lian, Richang Hong, Yong Li, and Meng Wang. Average user-side counterfactual fairness for collaborative filtering. ACM Transactions on Information Systems, 42(5):140:1–140:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3656639.

Sang:2025:DHG

[SWZW25] Lei Sang, Yu Wang, Yiwen Zhang, and Xindong Wu. Denoising heterogeneous graph pre-training framework for recommendation. ACM Transactions on Information Systems, 43(5):121:1–121:31, September 2025. CODEN ATISET. ISSN 1046-8188.

Sun:2024:BGS

[SYSW24] Yatong Sun, Xiaochun Yang, Zhu Sun, and Bin Wang. BERD+: a generic sequential recommendation framework by eliminating unreliable data with item- and attribute-level signals. ACM Transactions on Information Systems, 42(2): 41:1–41:??, March 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3611008.

Sun:2017:CIP

[SYX^{+}17] Yu Sun, Nicholas Jing Yuan, Xing Xie, Kieran McDonald, and Rui Zhang. Collaborative intent prediction with real-time contextual data. ACM Transactions on Information Systems, 35(4):30:1– 30:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Sturman:1993:DMW

[SZ93] David J. Sturman and David Zeltzer. A design method for “whole hand” humancomputer interaction. ACM Transactions on Information Systems, 11(3):219–238, July 1993. CODEN ATISET. ISSN 1046-8188. URL http://www. acm.org:80. Special Issue on Virtual Worlds.

Shokouhi:2009:RRM

[SZ09] Milad Shokouhi and Justin Zobel. Robust result merging using sample-based score estimates. ACM Transactions on Information Systems, 27(3): 14:1–14:??, May 2009. CO-DEN ATISET. ISSN 1046-8188.

Sakai:2021:REM

[SZ21] Tetsuya Sakai and Zhaohao Zeng. Retrieval evaluation measures that agree with users’ SERP preferences: Traditional, preference-based, and diversity measures. ACM Transactions on Information Systems, 39(2):14:1– 14:35, March 2021. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3431813.

Shen:2023:RWC

[SZC^{+}23] Yanyan Shen, Lifan Zhao, Weiyu Cheng, Zibin Zhang, Wenwen Zhou, and Lin Kangyi. RESUS: Warm-up cold users via meta-learning residual user preferences in CTR prediction. ACM Transactions on Information Systems, 41(3):69:1–69:??, July 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3564283.

Sun:2023:LIE

[SZM^{+}23] Kai Sun, Richong Zhang, Samuel Mensah, Yongyi Mao, and Xudong Liu. Learning implicit and explicit multi-task interactions for information extraction. ACM Transactions on Information Systems, 41(2):27:1–27:??, April 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3533020.

Shi:2017:LRB

[SZS17] Lei Shi, Wayne Xin Zhao, and Yi-Dong Shen. Local representative-based matrix factorization for coldstart recommendation. ACM Transactions on Information Systems, 36(2):22:1–22:??, September 2017. CODEN ATISET. ISSN 1046-8188.

Shi:2024:UGN

[SZY^{+}24] Chuan Shi, Meiqi Zhu, Yue Yu, Xiao Wang, and Junping Du. Unifying graph neural networks with a generalized optimization framework. ACM Transactions on Information Systems, 42(6):147:1– 147:??, November 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3660852.

Song:2023:SPF

[SZZL23] Haoyu Song, Wei-Nan Zhang, Kaiyan Zhang, and Ting Liu. A stack-propagation framework for low-resource personalized dialogue generation. ACM Transactions on Information Systems, 41(3):68:1– 68:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3563389.

Taghva:1996:EMB

[TBC96] Kazem Taghva, Julie Borsack, and Allen Condit. Evaluation of model-based retrieval effectiveness with OCR text. ACM Transactions on Information Systems, 14(1):64– 93, January 1996. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/taghva.
                                        html.

Thomas:2020:ISM

[TBCW20] Paul Thomas, Bodo Billerbeck, Nick Craswell, and Ryen W. White. Investigating searchers’ mental models to inform search explanations. ACM Transactions on Information Systems, 38(1):10:1– 10:25, February 2020. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3371390.

Turtle:1991:EIN

[TC91] Howard Turtle and W. Bruce Croft. Evaluation of an inference network=based retrieval model. ACM Transactions on Information Systems, 9(3): 187–222, July 1991. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.

org:80. Special Issue on Research and Development in Information Retrieval.

Thomas:2021:TCC

[TCMC21] Paul Thomas, Mary Czerwinksi, Daniel Mcduff, and Nick Craswell. Theories of conversation for conversational IR. ACM Transactions on Information Systems, 39 (4):39:1–39:23, October 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3439869.

Tran:2025:TPB

[TCQ^{+}25] Hung Vinh Tran, Tong Chen, Nguyen Quoc Viet Hung, Zi Huang, Lizhen Cui, and Hongzhi Yin. A thorough performance benchmarking on lightweight embeddingbased recommender systems. ACM Transactions on Information Systems, 43(3):63:1– 63:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Tang:2024:DRD

[TCSL24] Xing Tang, Ling Chen, Hongyu Shi, and Dandan Lyu. DHyper: a recurrent dual hypergraph neural network for event prediction in temporal knowledge graphs. ACM Transactions on Information Systems, 42(5):129:1–129:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3653015.

Teevan:2008:HPR

[Tee08] Jaime Teevan. How people recall, recognize, and reuse search results. ACM Transactions on Information Systems, 26(4):19:1–19:??, September 2008. CODEN ATISET. ISSN 1046-8188.

Tejedor:2012:CML

[TFS^{+}12] Javier Tejedor, Michal Fapso, Igor Szöke, Jan ‘Honza’ Cernocký, and Frantisek Grézl. Comparison of methods for language-dependent and language-independent query-by-example spoken term detection. ACM Transactions on Information Systems, 30 (3):18:1–18:??, August 2012. CODEN ATISET. ISSN 1046-8188.

Tagarelli:2010:SCX

[TG10] Andrea Tagarelli and Sergio Greco. Semantic clustering of XML documents. ACM Transactions on Information Systems, 28(1):3:1–3:??, January 2010. CODEN ATISET. ISSN 1046-8188.

Tomasic:1997:DSE

[TGL^{+}97] Anthony Tomasic, Luis Gravano, Calvin Lue, Peter Schwarz, and Laura Haas. Data structures for efficient broker implementation. ACM Transactions on Information Systems, 15(3):223– 253, July 1997. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80/tois/abstracts/tomasic.
                                        html.

Thomason:2016:CTA

[TGS16] Alasdair Thomason, Nathan Griffiths, and Victor Sanchez. Context trees: Augmenting geospatial trajectories with context. ACM Transactions on Information Systems, 35(2):14:1–14:??, December 2016. CODEN ATISET. ISSN 1046-8188.

Tang:2025:OMA

[THL^{+}25] Zuoli Tang, Zhaoxin Huan, Zihao Li, Xiaolu Zhang, Jun Hu, Chilin Fu, Jun Zhou, Lixin Zou, and Chenliang Li. One model for all: Large language models are domainagnostic recommendation systems. ACM Transactions on Information Systems, 43 (5):118:1–118:27, September 2025. CODEN ATISET. ISSN 1046-8188.

Tigelaar:2012:PPI

[THT12] Almer S. Tigelaar, Djoerd Hiemstra, and Dolf Trieschnigg. Peer-to-peer information retrieval: An overview. ACM Transactions on Information Systems, 30 (2):9:1–9:??, May 2012. CO-DEN ATISET. ISSN 1046-8188.

Tryfonopoulos:2009:IFQ

[TKD09] Christos Tryfonopoulos, Manolis Koubarakis, and Yannis Drougas. Information filtering and query indexing for an information retrieval model. ACM Transactions on Information Systems, 27(2):10:1– 10:??, February 2009. CO-DEN ATISET. ISSN 1046-

Tan:1999:EIG

[TkWW99] Bernard C. Y. Tan, Kwok kee Wei, and Richard T. Watson. The equalizing impact of a group support system on status differentials. ACM Transactions on Information Systems, 17(1):77– 100, January 1999. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tois/1999-17-1/p77-tan/.

Turney:2003:MPC

[TL03] Peter D. Turney and Michael L. Littman. Measuring praise and criticism: Inference of semantic orientation from association. ACM Transactions on Information Systems, 21(4): 315–346, October 2003. CO-DEN ATISET. ISSN 1046-8188.

Ture:2014:ERS

[TL14] Ferhan Ture and Jimmy Lin. Exploiting representations from statistical machine translation for cross-language information retrieval. ACM Transactions on Information Systems, 32(4):19:1– 19:??, October 2014. CODEN ATISET. ISSN 1046-8188.

Talvensaari:2007:CEC

[TLJ^{+}07] Tuomas Talvensaari, Jorma Laurikkala, Kalervo Järvelin, Martti Juhola, and Heikki Keskustalo. Creating and exploiting a comparable corpus in cross-language information retrieval. ACM Transactions on Information Systems, 25 (1):4:1–4:??, February 2007. CODEN ATISET. ISSN 1046-8188.

Tang:2016:TLI

[TLKW16] Jie Tang, Tiancheng Lou, Jon Kleinberg, and Sen Wu. Transfer learning to infer social ties across heterogeneous networks. ACM Transactions on Information Systems, 34 (2):7:1–7:??, April 2016. CO-DEN ATISET. ISSN 1046-8188.

Tian:2022:EGI

[TLS^{+}22] Zhiqiang Tian, Yezheng Liu, Jianshan Sun, Yuanchun Jiang, and Mingyue Zhu. Exploiting group information for personalized recommendation with graph neural networks. ACM Transactions on Information Systems, 40(2):27:1– 27:23, April 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3464764.

Tang:1991:VVI

[TM91] John C. Tang and Scott L.

Minneman. VideoDraw: a video interface for collaborative drawing. ACM Transactions on Information Systems, 9(2):170–184, April 1991. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.

acm.org:80. Special Issue on Computer-Human Interaction.

Tan:2010:CBI

[TM10] Qingzhao Tan and Prasenjit Mitra. Clustering-based incremental Web crawling. ACM Transactions on Information Systems, 28(4):17:1– 17:??, November 2010. CO-DEN ATISET. ISSN 1046-8188.

Tymoshenko:2019:SDS

[TM19] Kateryna Tymoshenko and Alessandro Moschitti. Shallow and deep syntactic/ semantic structures for passage reranking in questionanswering systems. ACM Transactions on Information Systems, 37(1):8:1–8:??, January 2019. CODEN ATISET. ISSN 1046-8188.

Tonellotto:2020:UII

[TM20] Nicola Tonellotto and Craig Macdonald. Using an inverted index synopsis for query latency and performance prediction. ACM Transactions on Information Systems, 38 (3):29:1–29:33, June 2020. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3389795.

Tsai:2006:CMS

[TMT06] Chih-Fong Tsai, Ken Mc-Garry, and John Tait. CLAIRE: a modular support vector image indexing and classification system. ACM Transactions on Information Systems, 24 (3):353–379, July 2006. CO-DEN ATISET. ISSN 1046-8188.

Tompa:1989:DMF

[Tom89] Frank Wm. Tompa. A data model for flexible hypertext database systems. ACM Transactions on Information Systems, 7(1):85–100, January 1989. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Tao:2005:HST

[TP05] Yufei Tao and Dimitris Papadias. Historical spatiotemporal aggregation. ACM Transactions on Information Systems, 23(1):61–102, January 2005. CODEN ATISET. ISSN 1046-8188.

Tian:2015:TTA

[TQH15] Yonghong Tian, Mengren Qian, and Tiejun Huang. TASC: a transformationaware soft cascading approach for multimodal video copy detection. ACM Transactions on Information Systems, 33 (2):7:1–7:??, February 2015. CODEN ATISET. ISSN 1046-8188.

Thanh:2023:PGB

[TQN^{+}23] Toan Nguyen Thanh, Nguyen Duc Khang Quach, Thanh Tam Nguyen, Thanh Trung Huynh, Viet Hung Vu, Phi Le Nguyen, Jun Jo, and Quoc Viet Hung Nguyen. Poisoning GNN-based recommender systems with generative surrogate-based attacks.

ACM Transactions on Information Systems, 41(3):58:1– 58:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3567420.

Transier:2010:EBA

[TS10] Frederik Transier and Peter Sanders. Engineering basic algorithms of an inmemory text search engine. ACM Transactions on Information Systems, 29(1):2:1– 2:??, December 2010. CO-DEN ATISET. ISSN 1046-8188.

Tamine:2019:OVO

[TSNS19] Lynda Tamine, Laure Soulier, Gia-Hung Nguyen, and Nathalie Souf. Offline versus online representation learning of documents using external knowledge. ACM Transactions on Information Systems, 37(4):42:1–42:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3349527.

Tavakoli:2025:OOE

[TTZ^{+}25] Leila Tavakoli, Johanne R. Trippas, Hamed Zamani, Falk Scholer, and Mark Sanderson. Online and offline evaluation in search clarification. ACM Transactions on Information Systems, 43(1):2:1– 2:??, January 2025. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3681786.

Tuzhilin:1995:TKB

[Tuz95] Alexander Tuzhilin. Templar:

a knowledge-based language for software specifications using temporal logic. ACM Transactions on Information Systems, 13(3):269–304, July 1995. CODEN ATISET. ISSN 1046-8188. URL http://www.

acm.org:80.

Tan:2018:QTQ

[TWLX18] Jiwei Tan, Xiaojun Wan, Hui Liu, and Jianguo Xiao. QuoteRec: Toward quote recommendation for writing. ACM Transactions on Information Systems, 36(3):34:1– 34:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Tang:2025:TTC

[TWS^{+}25] Haoran Tang, Shiqing Wu, Xueyao Sun, Jun Zeng, Guandong Xu, and Qing Li. TCGC: Temporal collaboration-aware graph coevolution learning for dynamic recommendation. ACM Transactions on Information Systems, 43(1):5:1–5:??, January 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3687470.

Tian:2024:PPC

[TXC^{+}24] Changxin Tian, Yuexiang Xie, Xu Chen, Yaliang Li, and Xin Zhao. Privacy-preserving cross-domain recommendation with federated graph learning. ACM Transactions on Information Systems, 42 (5):135:1–135:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3653448.

Tang:2022:RCI

[TY22] Zhiwen Tang and Grace Hui Yang. A re-classification of information seeking tasks and their computational solutions. ACM Transactions on Information Systems, 40(4):80:1– 80:32, October 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3497875.

Tang:2024:LGR

[TZG^{+}24] Yubao Tang, Ruqing Zhang, Jiafeng Guo, Maarten de Rijke, Wei Chen, and Xueqi Cheng. Listwise generative retrieval models via a sequential learning process. ACM Transactions on Information Systems, 42(5):133:1–133:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3653712.

Tian:2022:WHL

[TZP22] Yuan Tian, Ke Zhou, and Dan Pelleg. What and how long: Prediction of mobile app engagement. ACM Transactions on Information Systems, 40(1):8:1–8:38, January 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3464301.

Tian:2023:CPM

[TZP23] Yuan Tian, Ke Zhou, and Dan Pelleg. Characterization and prediction of mobile tasks. ACM Transactions on Information Systems, 41 (1):13:1–13:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3522711.

Urgo:2022:UPT

[UA22] Kelsey Urgo and Jaime Arguello. Understanding the “Pathway” towards a searcher’s learning objective. ACM Transactions on Information Systems, 40(4):77:1– 77:42, October 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3495222.

Upstill:2003:QIE

[UCH03] Trystan Upstill, Nick Craswell, and David Hawking. Queryindependent evidence in home page finding. ACM Transactions on Information Systems, 21(3):286–313, July 2003. CODEN ATISET. ISSN 1046-8188.

Umemoto:2017:SSU

[USN^{+}17] Kazutoshi Umemoto, Ruihua Song, Jian-Yun Nie, Xing Xie, Katsumi Tanaka, and Yong Rui. Search by screenshots for universal article clipping in mobile apps. ACM Transactions on Information Systems, 35(4):34:1–34:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Utting:1989:COH

[UY89] Kenneth Utting and Nicole Yankelovich. Context and orientation in hypermedia networks. ACM Transactions on Information Systems, 7(1): 58–84, January 1989. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Vuong:2021:SCC

[VAJR21] Tung Vuong, Salvatore Andolina, Giulio Jacucci, and Tuukka Ruotsalo. Spoken conversational context improves query auto-completion in Web search. ACM Transactions on Information Systems, 39(3):31:1–31:32, July 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3447875.

Vuong:2022:DMC

[VAJR22] Tung Vuong, Salvatore Andolina, Giulio Jacucci, and Tuukka Ruotsalo. Does more context help? Effects of context window and application source on retrieval performance. ACM Transactions on Information Systems, 40(2): 39:1–39:40, April 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3474055.

Vujovic:1998:EAF

[VB98] N. Vujovic and D. Brzakovic. Evaluation of an algorithm for finding a match of a distorted texture pattern in a large image database. ACM Transactions on Information Systems, 16(1):31–60, January 1998. CODEN ATISET. ISSN 1046-8188.

Vuong:2025:ICA

[VDR25] Tung Vuong, Pritom Kumar Das, and Tuukka Ruotsalo. Incorporating cognitive abilities into Web search reranking. ACM Transactions on Information Systems, 43 (5):126:1–126:35, September 2025. CODEN ATISET. ISSN 1046-8188.

VanGysel:2018:NVS

[VdRK18] Christophe Van Gysel, Maarten de Rijke, and Evangelos Kanoulas. Neural vector spaces for unsupervised information retrieval. ACM Transactions on Information Systems, 36(4):38:1–38:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Vardasbi:2017:SSW

[VFA17] Ali Vardasbi, Heshaam Faili, and Masoud Asadpour. SWIM: Stepped weighted shell decomposition influence maximization for large-scale networks. ACM Transactions on Information Systems, 36(1): 6:1–6:??, August 2017. CO-DEN ATISET. ISSN 1046-

Vallet:2011:EUB

[VHJC11] David Vallet, Frank Hopfgartner, Joemon M. Jose, and Pablo Castells. Effects of usage-based feedback on video retrieval: A simulation-based study. ACM Transactions on Information Systems, 29(2): 11:1–11:??, April 2011. CO-DEN ATISET. ISSN 1046-8188.

Vakulenko:2021:LSA

[VKD21] Svitlana Vakulenko, Evangelos Kanoulas, and Maarten De Rijke. A large-scale analysis of mixed initiative in information-seeking dialogues for conversational search. ACM Transactions on Information Systems, 39(4):49:1– 49:32, October 2021. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3466796.

Vlissides:1990:UFB

[VL90] John M. Vlissides and Mark A. Linton. Unidraw: a framework for building domainspecific graphical editors. ACM Transactions on Information Systems, 8(3):237– 268, July 1990. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org:
                                        80.

Vuong:2024:PRI

[VR24] Tung Vuong and Tuukka Ruotsalo. Predicting representations of information needs from digital activity context. ACM Transactions on Information Systems, 42(4):95:1–95:??, July 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3639819.

Voorhees:2017:URI

[VSS17] Ellen M. Voorhees, Daniel Samarov, and Ian Soboroff. Using replicates in information retrieval evaluation. ACM Transactions on Information Systems, 36(2):12:1– 12:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Laere:2014:GWD

[VST^{+}14] Olivier Van Laere, Steven Schockaert, Vlad Tanasescu, Bart Dhoedt, and Christopher B. Jones. Georeferencing Wikipedia documents using data from social media sources. ACM Transactions on Information Systems, 32(3):12:1–12:??, June 2014. CODEN ATISET. ISSN 1046-8188.

Wang:2022:SMR

[WA22] Zhenduo Wang and Qingyao Ai. Simulating and modeling the risk of conversational search. ACM Transactions on Information Systems, 40 (4):85:1–85:33, October 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3507357.

Wilkinson:2021:WWE

[WAL^{+}21] Daricia Wilkinson, Öznur Alkan, Q. Vera Liao, Massimiliano Mattetti, Inge Vejsbjerg, Bart P. Knijnenburg, and Elizabeth Daly. Why or why not? The effect of justification styles on chatbot recommendations. ACM Transactions on Information Systems, 39(4):42:1– 42:21, October 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3441715.

Wiecha:1990:TRD

[WBB^{+}90] Charles Wiecha, William Bennett, Stephen Boies, John Gould, and Sharon Greene. ITS: a tool for rapidly developing interactive applications. ACM Transactions on Information Systems, 8(3): 204–236, July 1990. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Webb:2016:DWP

[WBP^{+}16] Helena Webb, Pete Burnap, Rob Procter, Omer Rana, Bernd Carsten Stahl, Matthew Williams, William Housley, Adam Edwards, and Marina Jirotka. Digital wildfires: Propagation, verification, regulation, and responsible innovation. ACM Transactions on Information Systems, 34(3):15:1–15:??, May 2016. CODEN ATISET. ISSN 1046-8188.

Wang:2021:ISB

[WC21a] Wei Wang and Longbing Cao. Interactive sequential basket recommendation by learning basket couplings and positive/negative feedback. ACM Transactions on Information Systems, 39(3): 24:1–24:26, July 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3444368.

Wang:2021:VNV

[WC21b] Wei Wang and Longbing Cao. VM-NSP: Vertical negative sequential pattern mining with loose negative element constraints. ACM Transactions on Information Systems, 39(2):22:1–22:27, March 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3440874.

Wang:2025:ADS

[WCP^{+}25] Xin Wang, Hong Chen, Zirui Pan, Yuwei Zhou, Chaoyu Guan, Lifeng Sun, and Wenwu Zhu. Automated disentangled sequential recommendation with large language models. ACM Transactions on Information Systems, 43(2):29:1–29:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3675164.

Wen:2024:PAE

[WCS^{+}24] Zhiyuan Wen, Jiannong Cao, Jiaxing Shen, Ruosong Yang, Shuaiqi Liu, and Maosong Sun. Personality-affected emotion generation in dialog systems. ACM Transactions on Information Systems, 42 (5):134:1–134:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3655616.

Wu:2024:FDF

[WCX24] Yao Wu, Jian Cao, and Guandong Xu. FASTER: a dynamic fairness-assurance strategy for session-based recommender systems. ACM Transactions on Information Systems, 42(1):5:1–5:??, January 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3586993.

Wang:2025:NPE

[WCZ^{+}25] Tairan Wang, Xiuying Chen, Qingqing Zhu, Taicheng Guo, Shen Gao, Zhiyong Lu, Xin Gao, and Xiangliang Zhang. New paradigm for evaluating scholar summaries: a facetaware metric and a metaevaluation benchmark. ACM Transactions on Information Systems, 43(4):100:1–100:??, July 2025. CODEN ATISET. ISSN 1046-8188.

White:2017:SRP

[WDG17] Ryen W. White, Fernando Diaz, and Qi Guo. Search result prefetching on desktop and mobile. ACM Transactions on Information Systems, 35(3):23:1–23:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Wang:2024:PDR

[WDL^{+}24] Shuting Wang, Zhicheng Dou, Jiongnan Liu, Qiannan Zhu, and Ji-Rong Wen. Personalized and diversified: Ranking search results in an integrated way. ACM Transactions on Information Systems, 42(3): 81:1–81:??, May 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3631989.

Wang:2008:URM

[WdVR08] Jun Wang, Arjen P. de Vries, and Marcel J. T. Reinders. Unified relevance models for rating prediction in collaborative filtering. ACM Transactions on Information Systems, 26(3):16:1–16:??, June 2008. CODEN ATISET. ISSN 1046-8188.

Webber:2013:ARC

[Web13] William Webber. Approximate recall confidence intervals. ACM Transactions on Information Systems, 31(1): 2:1–2:??, January 2013. CO-DEN ATISET. ISSN 1046-8188.

Wang:2025:MAA

[WFW^{+}25] Shijie Wang, Wenqi Fan, Xiao-Yong Wei, Xiaowei Mei, Shanru Lin, and Qing Li. Multi-agent attacks for blackbox social recommendations.

ACM Transactions on Information Systems, 43(1):24:1– 24:??, January 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3696105.

Wang:2024:MRC

[WGL^{+}24] Yanan Wang, Yong Ge, Zhepeng Li, Li Li, and Rui Chen. M^{3}Rec: a context-aware offline meta-level model-based reinforcement learning approach for cold-start recommendation. ACM Transactions on Information Systems, 42(6):146:1–146:??, November 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3659947.

Wang:2025:EHH

[WGZ^{+}25] Hao Wang, Bin Guo, Yating Zeng, Mengqi Chen, Yasan Ding, Ying Zhang, Lina Yao, and Zhiwen Yu. Enabling harmonious human-machine interaction with visual-context augmented dialogue system: a review. ACM Transactions on Information Systems, 43(3): 68:1–68:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

White:2009:CSE

[WH09] Ryen W. White and Eric Horvitz. Cyberchondria: Studies of the escalation of medical concerns in Web search. ACM Transactions on Information Systems, 27(4): 23:1–23:??, November 2009.

CODEN ATISET. ISSN 1046-8188.

White:2015:BDB

[WH15] Ryen W. White and Eric Horvitz. Belief dynamics and biases in Web search. ACM Transactions on Information Systems, 33(4):18:1– 18:??, May 2015. CODEN ATISET. ISSN 1046-8188.

Want:1992:ABL

[WHFG92] Roy Want, Andy Hopper, Veronica Falcao, and Jonathan Gibbons. The active badge location system. ACM Transactions on Information Systems, 10(1):91–102, January 1992. CODEN ATISET. ISSN 1046-8188. URL http:

//www.acm.org:80.

Wang:2016:ROC

[WHL^{+}16] Shuaiqiang Wang, Shanshan Huang, Tie-Yan Liu, Jun Ma, Zhumin Chen, and Jari Veijalainen. Ranking-oriented collaborative filtering: a listwise approach. ACM Transactions on Information Systems, 35(2):10:1–10:??, December 2016. CODEN ATISET. ISSN 1046-8188.

Wang:2022:DSR

[WHL^{+}22] Lili Wang, Chenghan Huang, Ying Lu, Weicheng Ma, Ruibo Liu, and Soroush Vosoughi. Dynamic structural role node embedding for user modeling in evolving networks. ACM Transactions on Information Systems, 40(3):46:1– 46:21, July 2022. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3472955.

Wiecha:1992:UIC

[Wie92] Charles Wiecha. ITS and user interface consistency: a response to Grudin. ACM Transactions on Information Systems, 10(1):112–114, January 1992. CODEN ATISET. ISSN 1046-8188. URL http:

                                        //www.acm.org:80.

Wilbur:1998:KMH

[Wil98] W. John Wilbur. The knowledge in multiple human relevance judgments. ACM Transactions on Information Systems, 16(2):101–126, April 1998. CODEN ATISET. ISSN 1046-8188.

Williams:2005:IGI

[Wil05] Hugh E. Williams. Introduction to genomic information retrieval. ACM Transactions on Information Systems, 23 (1):1–2, January 2005. CO-DEN ATISET. ISSN 1046-8188.

Wang:2012:DPC

[WKF^{+}12] Dong Wang, Simon King, Joe Frankel, Ravichander Vipperla, Nicholas Evans, and Raphaël Troncy. Direct posterior confidence for out-of-vocabulary spoken term detection. ACM Transactions on Information Systems, 30 (3):16:1–16:??, August 2012.

CODEN ATISET. ISSN 1046-8188.

Wiil:1997:HHS

[WL97] Uffe K. Wiil and John J.

Leggett. Hyperform: a hypermedia system development environment. ACM Transactions on Information Systems, 15(1):1–31, January 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/wiil.
html.

Wang:2022:CGC

[WL22] Hongwei Wang and Jure Leskovec. Combining graph convolutional neural networks and label propagation. ACM Transactions on Information Systems, 40(4):73:1–73:27, October 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3490478.

Wang:2025:PTMa

[WLF^{+}25a] Wenjie Wang, Zheng Liu, Fuli Feng, Zhicheng Dou, Qingyao Ai, Grace Hui Yang, Defu Lian, Lu Hou, Aixin Sun, Hamed Zamani, Donald Metzler, and Maarten de Rijke. Pre-trained models for search and recommendation: Introduction to the special issue — Part 1. ACM Transactions on Information Systems, 43(2):27:1–27:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3709134.

Wang:2025:PTMb

[WLF^{+}25b] Wenjie Wang, Zheng Liu, Fuli Feng, Zhicheng Dou, Qingyao Ai, Grace Hui Yang, Defu Lian, Lu Hou, Aixin Sun, Hamed Zamani, Donald Metzler, and Maarten de Rijke. Pre-trained models for search and recommendation: Introduction to the special issue — Part 2. ACM Transactions on Information Systems, 43(5):111:1–111:5, September 2025. CODEN ATISET. ISSN 1046-8188.

Wang:2018:LAI

[WLFT18] Yanhao Wang, Yuchen Li, Ju Fan, and Kian-Lee Tan. Location-aware influence maximization over dynamic social streams. ACM Transactions on Information Systems, 36 (4):43:1–43:??, October 2018. CODEN ATISET. ISSN 1046-8188.

Wang:2017:UPP

[WLG^{+}17] Haoyu Wang, Yuanchun Li, Yao Guo, Yuvraj Agarwal, and Jason I. Hong. Understanding the purpose of permission use in mobile apps. ACM Transactions on Information Systems, 35(4):43:1– 43:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Wang:2023:AGP

[WLL^{+}23a] Yiqi Wang, Chaozhuo Li, Zheng Liu, Mingzheng Li, Jiliang Tang, Xing Xie, Lei Chen, and Philip S. Yu. An adaptive graph pre-training framework for localized collaborative filtering. ACM Transactions on Information Systems, 41(2):43:1– 43:??, April 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3555372.

Wu:2023:AAE

[WLL^{+}23b] Hanrui Wu, Jinyi Long, Nuosi Li, Dahai Yu, and Michael K. Ng. Adversarial auto-encoder domain adaptation for coldstart recommendation with positive and negative hypergraphs. ACM Transactions on Information Systems, 41(2): 33:1–33:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3544105.

Wang:2024:TCB

[WLL24] Jian Wang, Dongding Lin, and Wenjie Li. Targetconstrained bidirectional planning for generation of targetoriented proactive dialogue. ACM Transactions on Information Systems, 42(5):124:1– 124:??, September 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3652598.

Wei:2021:TGE

[WLM^{+}21] Wei Wei, Jiayi Liu, Xianling Mao, Guibing Guo, Feida Zhu, Pan Zhou, Yuchong Hu, and Shanshan Feng. Targetguided emotion-aware chat machine. ACM Transactions on Information Systems, 39 (4):43:1–43:24, October 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3456414.

Wang:2025:PPS

[WLR^{+}25] Wei Wang, Yujie Lin, Pengjie Ren, Zhumin Chen, Tsunenori Mine, Jianli Zhao, Qiang Zhao, Moyan Zhang, Xianye Ben, and Yujun Li. Privacypreserving sequential recommendation with collaborative confusion. ACM Transactions on Information Systems, 43 (2):50:1–50:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3707204.

Wang:2017:SMT

[WLS^{+}17] Hongning Wang, Rui Li, Milad Shokouhi, Hang Li, and Yi Chang. Search, mining, and their applications on mobile devices: Introduction to the special issue. ACM Transactions on Information Systems, 35(4):29:1–29:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Wang:2022:HEH

[WLT^{+}22] Hao Wang, Defu Lian, Hanghang Tong, Qi Liu, Zhenya Huang, and Enhong Chen. HyperSoRec: Exploiting hyperbolic user and item representations with multiple aspects for social-aware recommendation. ACM Transactions on Information Systems, 40(2):24:1–24:28, April 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3463913.

Wan:2024:STC

[WLW^{+}24a] Zhongwei Wan, Xin Liu, Benyou Wang, Jiezhong Qiu, Boyu Li, Ting Guo, Guangyong Chen, and Yang Wang. Spatio-temporal contrastive learning-enhanced GNNs for session-based recommendation. ACM Transactions on Information Systems, 42(2): 58:1–58:??, March 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3626091.

Wang:2024:CDR

[WLW^{+}24b] Wenjie Wang, Xinyu Lin, Liuhui Wang, Fuli Feng, Yunshan Ma, and Tat-Seng Chua. Causal disentangled recommendation against user preference shifts. ACM Transactions on Information Systems, 42(1):12:1–12:??, January 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3593022.

Wu:2008:ITI

[WLWK08] Ho Chung Wu, Robert Wing Pong Luk, Kam Fai Wong, and Kui Lam Kwok. Interpreting TF-IDF term weights as making relevance decisions. ACM Transactions on Information Systems, 26(3):13:1–13:??, June 2008.

CODEN ATISET. ISSN 1046-8188.

Wang:2014:CDS

[WLY^{+}14] Jianguo Wang, Eric Lo, Man Lung Yiu, Jiancong Tong, Gang Wang, and Xiaoguang Liu. Cache design of SSD-based search engine architectures: an experimental study. ACM Transactions on Information Systems, 32 (4):21:1–21:??, October 2014. CODEN ATISET. ISSN 1046-8188.

Wang:2024:CEF

[WLY^{+}24] Xiangmeng Wang, Qian Li, Dianer Yu, Qing Li, and Guandong Xu. Counterfactual explanation for fairness in recommendation. ACM Transactions on Information Systems, 42(4):106:1– 106:??, July 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3643670.

Wang:2024:IBB

[WLZ^{+}24] Tianshi Wang, Fengling Li, Lei Zhu, Jingjing Li, Zheng Zhang, and Heng Tao Shen. Invisible black-box backdoor attack against deep crossmodal hashing retrieval. ACM Transactions on Information Systems, 42(4):111:1– 111:??, July 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3650205.

Wang:2021:DPR

[WLZW21] Jingyuan Wang, Xin Lin, Yuan Zuo, and Junjie Wu. DGeye: Probabilistic risk perception and prediction for urban dangerous goods management. ACM Transactions on Information Systems, 39(3): 28:1–28:30, July 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3448256.

Wang:2023:SRM

[WMC^{+}23] Chenyang Wang, Weizhi Ma, Chong Chen, Min Zhang, Yiqun Liu, and Shaoping Ma. Sequential recommendation with multiple contrast signals. ACM Transactions on Information Systems, 41 (1):11:1–11:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3522673.

Wang:2024:DLR

[WMGZ24] Quan Wang, Zhendong Mao, Jie Gao, and Yongdong Zhang. Document-level relation extraction with progressive self-distillation. ACM Transactions on Information Systems, 42(6):143:1–143:??, November 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3656168.

Wei:2005:MBA

[WMJ05] Yan Zheng Wei, Luc Moreau, and Nicholas R. Jennings. A market-based approach to recommender systems. ACM Transactions on Information Systems, 23(3):227–266, July 2005. CODEN ATISET. ISSN 1046-8188.

Wu:2023:MOO

[WMM^{+}23] Haolun Wu, Chen Ma, Bhaskar Mitra, Fernando Diaz, and Xue Liu. A multiobjective optimization framework for multi-stakeholder fairness-aware recommendation. ACM Transactions on Information Systems, 41(2): 47:1–47:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3564285.

Webber:2010:SMI

[WMZ10] William Webber, Alistair Moffat, and Justin Zobel. A similarity measure for indefinite rankings. ACM Transactions on Information Systems, 28(4):20:1–20:??, November 2010. CODEN ATISET. ISSN 1046-8188.

Wang:2021:TDU

[WMZ^{+}21] Chenyang Wang, Weizhi Ma, Min Zhang, Chong Chen, Yiqun Liu, and Shaoping Ma. Toward dynamic user intention: Temporal evolutionary effects of item relations in sequential recommendation. ACM Transactions on Information Systems, 39(2): 16:1–16:33, March 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3432244.

Wang:2023:SFR

[WMZ^{+}23] Yifan Wang, Weizhi Ma, Min Zhang, Yiqun Liu, and Shaoping Ma. A survey on the fairness of recommender systems. ACM Transactions on Information Systems, 41(3): 52:1–52:??, July 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3547333.

Wu:2024:LTI

[WNCH24] Jiaxin Wu, Chong-Wah Ngo, Wing-Kwong Chan, and Zhijian Hou. (Un)likelihood training for interpretable embedding. ACM Transactions on Information Systems, 42(3):75:1–75:??, May 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3632752.

Wang:2017:UVP

[WNS^{+}17] Xiang Wang, Liqiang Nie, Xuemeng Song, Dongxiang Zhang, and Tat-Seng Chua. Unifying virtual and physical worlds: Learning toward local and global consistency. ACM Transactions on Information Systems, 36(1):4:1–4:??, August 2017. CODEN ATISET. ISSN 1046-8188.

White:2021:MD

[WNWL^{+}21] Ryen W. White, Elnaz Nouri, James Woffinden-Luey, Mark EncarnacióN, and Sujay Kumar Jauhar. Microtask detection. ACM Transactions on Information Systems, 39(2): 17:1–17:29, March 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3432290.

Wen:2002:QCU

[WNZ02] Ji-Rong Wen, Jian-Yun Nie, and Hong-Jiang Zhang. Query clustering using user logs. ACM Transactions on Information Systems, 20(1):59– 81, January 2002. CODEN ATISET. ISSN 1046-8188.

Wong:1994:PBD

[Won94] Stephen T. C. Wong. Preferencebased decision making for cooperative knowledge-based systems. ACM Transactions on Information Systems, 12 (4):407–435, October 1994. CODEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Wu:2025:MLR

[WPS^{+}25] Xuyang Wu, Ajit Puthenputhussery, Hongwei Shang, Changsung Kang, and Yi Fang. Meta-learning to rank for sparsely supervised queries. ACM Transactions on Information Systems, 43(1):14:1– 14:??, January 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3698876.

Wang:2025:EEE

[WQH^{+}25] Jinguang Wang, Shengsheng Qian, Jun Hu, Wenxiang Dong, Xudong Huang, and Richang Hong. End-to-end explainable fake news detection via evidence-claim variational causal inference. ACM Transactions on Information Systems, 43(4):89:1– 89:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Wu:2019:CAU

[WQL^{+}19] Libing Wu, Cong Quan, Chenliang Li, Qian Wang, Bolong Zheng, and Xiangyang Luo. A context-aware useritem representation learning for item recommendation. ACM Transactions on Information Systems, 37 (2):22:1–22:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3298988.

Wang:1998:SHD

[WR98] Weigang Wang and Roy Rada. Structured hypertext with domain semantics. ACM Transactions on Information Systems, 16(4):372– 412, October 1998. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/pubs/citations/journals/
tois/1998-16-4/p372-wang/

.

White:2005:EIF

[WRJV05] Ryen W. White, Ian Ruthven, Joemon M. Jose, and C. J. Van Rijsbergen. Evaluating implicit feedback models using searcher simulations. ACM Transactions on Information Systems, 23(3):325– 361, July 2005. CODEN ATISET. ISSN 1046-8188.

Wang:2022:FFM

[WRLD22] Jinze Wang, Yongli Ren, Jie Li, and Ke Deng. The footprint of factorization models and their applications in collaborative filtering. ACM Transactions on Information Systems, 40(4):71:1–71:32, October 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3490475.

Watters:1990:THB

[WS90] Carolyn Watters and Michael A. Shepherd. A transient hypergraph-based model for data access. ACM Transactions on Information Systems, 8(2):77–102, April 1990. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Wong:2001:AAF

[WSBC01] Kam-Fai Wong, Dawei Song, Peter Bruza, and Chun-Hung Cheng. Application of aboutness to functional benchmarking in information retrieval. ACM Transactions on Information Systems, 19(4):337– 370, October 2001. CODEN ATISET. ISSN 1046-8188.

Wang:2008:DSZ

[WTS^{+}08] Xuanhui Wang, Tao Tao, Jian-Tao Sun, Azadeh Shakery, and Chengxiang Zhai. DirichletRank: Solving the zero-one gap problem of PageRank. ACM Transactions on Information Systems, 26(2):10:1–10:??, March 2008. CODEN ATISET. ISSN 1046-8188.

Walsham:1994:ISS

[WW94] G. Walsham and T. Waema.

Information systems strategy and implementation: a case study of a building society. ACM Transactions on Information Systems, 12(2):150– 173, April 1994. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80.

Wu:2010:AVG

[WW10] Gang Wu and Yimin Wei.

Arnoldi versus GMRES for computing PageRank: a theoretical contribution to Google’s PageRank problem. ACM Transactions on Information Systems, 28(3):11:1– 11:28, June 2010. CODEN ATISET. ISSN 1046-8188.

Wu:2024:ESS

[WWG^{+}24] Jiancan Wu, Xiang Wang, Xingyu Gao, Jiawei Chen, Hongcheng Fu, and Tianyu Qiu. On the effectiveness of sampled softmax loss for item recommendation. ACM Transactions on Information Systems, 42(4):98:1– 98:??, July 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3637061.

Wang:2025:LHF

[WWHQ25] Huili Wang, Chuhan Wu, Yongfeng Huang, and Tao Qi. Learning human feedback from large language models for content quality-aware recommendation. ACM Transactions on Information Systems, 43(4):86:1–86:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Wu:2023:PNR

[WWHX23] Chuhan Wu, Fangzhao Wu, Yongfeng Huang, and Xing Xie. Personalized news recommendation: Methods and challenges. ACM Transactions on Information Systems, 41(1):24:1–24:??, January 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3530257.

Wu:2025:CSU

[WWL^{+}25] Hanrui Wu, Yanxin Wu, Nuosi Li, Jia Zhang, Yonghui Xu, Michael K. Ng, and Jinyi Long. Cold-start user recommendation via heterogeneous domain adaptation. ACM Transactions on Information Systems, 43(5):141:1–141:26, September 2025. CODEN ATISET. ISSN 1046-8188.

Wu:2022:KPD

[WWN22] Hanrui Wu, Qingyao Wu, and Michael K. Ng. Knowledge preserving and distribution alignment for heterogeneous domain adaptation. ACM Transactions on Information Systems, 40(1):16:1– 16:29, January 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3469856.

Wu:2025:TTF

[WWN^{+}25] Lianwei Wu, Kang Wang, Kunlin Nie, Sensen Guo, Chao Gao, Zhen Wang, and Shudong Li. TFGIN: Tightfitting graph inference network for table-based fact verification. ACM Transactions on Information Systems, 43 (5):130:1–130:26, September 2025. CODEN ATISET. ISSN 1046-8188.

Wang:2025:GEP

[WWR^{+}25] Zihan Wang, Hanbing Wang, Pengjie Ren, Zhumin Chen, Maarten de Rijke, and Zhaochun Ren. Graphenhanced prompt learning for cross-domain contract element extraction. ACM Transactions on Information Systems, 43(3):66:1–66:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Wang:2025:UVM

[WWT^{+}25] Guolong Wang, Xun Wu, Xun Tu, Zhaoyuan Liu, and Junchi Yan. Unsupervised video moment retrieval with knowledge-based pseudo-supervision construction. ACM Transactions on Information Systems, 43 (1):23:1–23:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3701229.

Wan:2023:MCH

[WWX^{+}23] Qizhi Wan, Changxuan Wan, Keli Xiao, Rong Hu, and Dexi Liu. A multi-channel hierarchical graph attention network for open event extraction. ACM Transactions on Information Systems, 41 (1):20:1–20:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3528668.

Wan:2024:TER

[WWX^{+}24] Qizhi Wan, Changxuan Wan, Keli Xiao, Hui Xiong, Dexi Liu, Xiping Liu, and Rong Hu. Token-event-role structurebased multi-channel documentlevel event extraction. ACM Transactions on Information Systems, 42(4):104:1– 104:??, July 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3643885.

Wan:2025:MGB

[WWX^{+}25] Qizhi Wan, Changxuan Wan, Keli Xiao, Rong Hu, Dexi Liu, Guoqiong Liao, Xiping Liu, and Yuxin Shuai. A multifocal graph-based neural network scheme for topic event extraction. ACM Transactions on Information Systems, 43 (1):13:1–13:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3696353.

Wan:2010:ENK

[WX10] Xiaojun Wan and Jianguo Xiao. Exploiting neighborhood knowledge for single document summarization and keyphrase extraction. ACM Transactions on Information Systems, 28(2):8:1–8:??, May 2010. CODEN ATISET. ISSN 1046-8188.

Wang:2013:RLS

[WXLC13] Quan Wang, Jun Xu, Hang Li, and Nick Craswell. Regularized latent semantic indexing: a new approach to largescale topic modeling. ACM Transactions on Information Systems, 31(1):5:1–5:??, January 2013. CODEN ATISET. ISSN 1046-8188.

Wu:2025:LRS

[WXR^{+}25] Shiguang Wu, Xin Xin, Pengjie Ren, Zhumin Chen, Jun Ma, Maarten de Rijke, and Zhaochun Ren. Learning robust sequential recommenders through confident soft labels. ACM Transactions on Information Systems, 43(1):21:1–21:??, January 2025. CODEN ATISET. ISSN 1046-8188. URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3700876.

Wu:2025:ICP

[WXZ^{+}25] Yiqing Wu, Ruobing Xie, Zhao Zhang, Xu Zhang, Fuzhen Zhuang, Leyu Lin, Zhanhui Kang, Zhulin An, and Yongjun Xu. ID-centric pre-training for recommendation. ACM Transactions on Information Systems, 43 (5):125:1–125:29, September 2025. CODEN ATISET. ISSN 1046-8188.

Wong:1995:MIR

[WY95] S. K. M. Wong and Y. Y. Yao. On modeling information retrieval with probabilistic inference. ACM Transactions on Information Systems, 13 (1):38–68, January 1995. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Wang:2025:MGM

[WYZ^{+}25] Hao Wang, Mingjia Yin, Luankang Zhang, Sirui Zhao, and Enhong Chen. MF-GSLAE: a multi-factor user representation pre-training framework for dual-target cross-domain recommendation. ACM Transactions on Information Systems, 43(2): 30:1–30:??, March 2025. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3690382.

Williams:2004:FPQ

[WZB04] Hugh E. Williams, Justin Zobel, and Dirk Bahle. Fast phrase querying with combined indexes. ACM Transactions on Information Systems, 22(4):573–594, October 2004. CODEN ATISET. ISSN 1046-8188.

Wang:2017:CUT

[WZC^{+}17] Senzhang Wang, Xiaoming Zhang, Jianping Cao, Lifang He, Leon Stenneth, Philip S. Yu, Zhoujun Li, and Zhiqiu Huang. Computing urban traffic congestions by incorporating sparse GPS probe data and social media data. ACM Transactions on Information Systems, 35(4):40:1– 40:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Wang:2025:ULL

[WZC^{+}25] Junmei Wang, Fengjing Zhang, Xiadan Chen, Puyu He, Ellen Anne Huang, and Jimmy Xiangji Huang. Utilizing large language model for conversational information seeking via dual-query generation and joint-encoding. ACM Transactions on Information Systems, 43(5):135:1–135:30, September 2025. CODEN ATISET. ISSN 1046-8188.

Wu:2023:PPB

[WZG^{+}23a] Chen Wu, Ruqing Zhang, Jiafeng Guo, Maarten De Rijke, Yixing Fan, and Xueqi Cheng. PRADA: Practical black-box adversarial attacks against neural ranking models. ACM Transactions on Information Systems, 41 (4):89:1–89:??, October 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3576923.

Wu:2023:NRM

[WZG^{+}23b] Chen Wu, Ruqing Zhang, Jiafeng Guo, Yixing Fan, and Xueqi Cheng. Are neural ranking models robust? ACM Transactions on Information Systems, 41(2):29:1– 29:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3534928.

Wei:2024:NAS

[WZHY24] Lanning Wei, Huan Zhao, Zhiqiang He, and Quanming Yao. Neural architecture search for GNN-based graph classification. ACM Transactions on Information Systems, 42(1):1:1–1:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3584945.

Wu:2019:DDA

[WZL^{+}19] Zhijing Wu, Ke Zhou, Yiqun Liu, Min Zhang, and Shaoping Ma. Does diversity affect user satisfaction in image search. ACM Transactions on Information Systems, 37(3):35:1–35:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3320118.

Wang:2024:TBA

[WZL^{+}24] Zhidan Wang, Lixin Zou, Chenliang Li, Shuaiqiang Wang, Xu Chen, Dawei Yin, and Weidong Liu. Toward bias-agnostic recommender systems: a universal generative framework. ACM Transactions on Information Systems, 42(6):142:1–142:??, November 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3655617.

Wang:2019:EHO

[WZW^{+}19] Hongwei Wang, Fuzheng Zhang, Jialin Wang, Miao Zhao, Wenjie Li, Xing Xie, and Minyi Guo. Exploring high-order user preference on the knowledge graph for recommender systems. ACM Transactions on Information Systems, 37(3):32:1–32:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3312738.

Wang:2022:PEE

[WZW^{+}22] Chao Wang, Hengshu Zhu, Peng Wang, Chen Zhu, Xi Zhang, Enhong Chen, and Hui Xiong. Personalized and explainable employee training course recommendations: a Bayesian variational approach. ACM Transactions on Information Systems, 40 (4):70:1–70:32, October 2022.

CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3490476.

Wang:2023:QRM

[WZW23] Lingzhi Wang, Xingshan Zeng, and Kam-Fai Wong. Quotation recommendation for multi-party online conversations based on semantic and topic fusion. ACM Transactions on Information Systems, 41(4):94:1– 94:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3594633.

Wang:2025:HGN

[WZW^{+}25] Shuliang Wang, Jiabao Zhu, Yi Wang, Chen Ma, Xin Zhao, Yansen Zhang, Ziqiang Yuan, and Sijie Ruan. Hierarchical gating network for crossdomain sequential recommendation. ACM Transactions on Information Systems, 43(4): 90:1–90:??, July 2025. CO-DEN ATISET. ISSN 1046-8188.

Wang:2024:MVE

[WZY^{+}24] Dongjing Wang, Xin Zhang, Yuyu Yin, Dongjin Yu, Guandong Xu, and Shuiguang Deng. Multi-view enhanced graph attention network for session-based music recommendation. ACM Transactions on Information Systems, 42(1):16:1–16:??, January 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3592853.

Wang:2025:UBS

[WZY^{+}25] Lei Wang, Jingsen Zhang, Hao Yang, Zhi-Yuan Chen, Jiakai Tang, Zeyu Zhang, Xu Chen, Yankai Lin, Hao Sun, Ruihua Song, Xin Zhao, Jun Xu, Zhicheng Dou, Jun Wang, and Ji-Rong Wen. User behavior simulation with large language model-based agents. ACM Transactions on Information Systems, 43(2): 55:1–55:??, March 2025. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3708985.

Wu:2004:EMB

[WZZ04] Xindong Wu, Chengqi Zhang, and Shichao Zhang. Efficient mining of both positive and negative association rules. ACM Transactions on Information Systems, 22(3): 381–405, July 2004. CODEN ATISET. ISSN 1046-8188.

Wang:2023:CPT

[WZZ^{+}23] Hui Wang, Kun Zhou, Xin Zhao, Jingyuan Wang, and Ji-Rong Wen. Curriculum pre-training heterogeneous subgraph transformer for top-N recommendation. ACM Transactions on Information Systems, 41(1):19:1– 19:??, January 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3528667.

Wang:2024:SSB

[WZZ^{+}24a] Chao Wang, Hengshu Zhu, Chen Zhu, Chuan Qin, Enhong Chen, and Hui Xiong. SetRank: a setwise Bayesian approach for collaborative ranking in recommender system. ACM Transactions on Information Systems, 42(2): 56:1–56:??, March 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3626194.

Wang:2024:MAG

[WZZ^{+}24b] Ke Wang, Yanmin Zhu, Tianzi Zang, Chunyang Wang, Kuan Liu, and Peibo Ma. Multi-aspect graph contrastive learning for review-enhanced recommendation. ACM Transactions on Information Systems, 42(2):51:1– 51:??, March 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3618106.

Xu:1998:CBS

[XC98] Jinxi Xu and W. Bruce Croft. Corpus-based stemming using cooccurrence of word variants. ACM Transactions on Information Systems, 16(1): 61–81, January 1998. CODEN ATISET. ISSN 1046-8188.

Xu:2000:IEI

[XC00] Jinxi Xu and W. Bruce Croft. Improving the effectiveness of information retrieval with local context analysis. ACM Transactions on Information Systems, 18(1):79–112, January 2000. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org/
pubs/citations/journals/
tois/2000-18-1/p79-xu/.

Xu:2005:CEF

[XC05] Jennifer J. Xu and Hsinchun Chen. CrimeNet explorer: a framework for criminal network knowledge discovery. ACM Transactions on Information Systems, 23(2):201– 226, April 2005. CODEN ATISET. ISSN 1046-8188.

Xue:2013:MRU

[XC13] Xiaobing Xue and W. Bruce Croft. Modeling reformulation using query distributions. ACM Transactions on Information Systems, 31(2): 6:1–6:??, May 2013. CODEN ATISET. ISSN 1046-8188.

Xie:2024:HTS

[XC24] Jiayi Xie and Zhenzhong Chen. Hierarchical transformer with spatio-temporal context aggregation for next point-of-interest recommendation. ACM Transactions on Information Systems, 42(2): 37:1–37:??, March 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3597930.

Xu:2023:TTG

[XCG^{+}23] Ronghui Xu, Meng Chen, Yongshun Gong, Yang Liu, Xiaohui Yu, and Liqiang Nie. TME: Tree-guided multi-task embedding learning towards semantic venue annotation. ACM Transactions on Information Systems, 41(4):112:1– 112:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3582553.

Xu:2024:ISE

[XHSD24] Jiechen Xu, Lei Han, Shazia Sadiq, and Gianluca Demartini. On the impact of showing evidence from peers in crowdsourced truthfulness assessments. ACM Transactions on Information Systems, 42(3):87:1–87:??, May 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3637872.

Xue:2019:DIB

[XHW^{+}19] Feng Xue, Xiangnan He, Xiang Wang, Jiandong Xu, Kai Liu, and Richang Hong. Deep item-based collaborative filtering for top-N recommendation. ACM Transactions on Information Systems, 37(3):33:1–33:??, July 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3314578.

Xia:2022:CRA

[XHX^{+}22] Lianghao Xia, Chao Huang, Yong Xu, Huance Xu, Xiang Li, and Weiguo Zhang. Collaborative reflection-augmented autoencoder network for recommender systems. ACM Transactions on Information Systems, 40(1):10:1–10:22, January 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3467023.

Xue:2009:ULM

[XHYY09] Gui-Rong Xue, Jie Han, Yong Yu, and Qiang Yang. User language model for collaborative personalized search. ACM Transactions on Information Systems, 27(2):11:1– 11:??, February 2009. CO-DEN ATISET. ISSN 1046-8188.

Xu:2023:VGL

[XLL^{+}23] Senrong Xu, Liangyue Li, Zenan Li, Yuan Yao, Feng Xu, Zulong Chen, Quan Lu, and Hanghang Tong. On the vulnerability of graph learningbased collaborative filtering. ACM Transactions on Information Systems, 41(4):87:1– 87:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3572834.

Xie:2022:LTI

[XLW^{+}22] Zhongwei Xie, Ling Liu, Yanzhao Wu, Luo Zhong, and Lin Li. Learning textimage joint embedding for efficient cross-modal retrieval with deep feature engineering. ACM Transactions on Information Systems, 40(4):74:1– 74:27, October 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3490519.

Xu:2025:GAE

[XLX^{+}25] Lixiang Xu, Yusheng Liu, Tong Xu, Enhong Chen, and Yuanyan Tang. Graph augmentation empowered contrastive learning for recommendation. ACM Transactions on Information Systems, 43(2):34:1–34:??, March 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3677377.

Xu:2025:ATO

[XMS^{+}25] Heng-Da Xu, Xian-Ling Mao, Fanshu Sun, Tian-Yi Che, Chun Xu, and Heyan Huang. AgentTOD: a task-oriented dialogue agent with a flexible and adaptive API calling paradigm. ACM Transactions on Information Systems, 43 (5):136:1–136:32, September 2025. CODEN ATISET. ISSN 1046-8188.

Xu:2024:GBP

[XPL^{+}24] Hongyan Xu, Qiyao Peng, Hongtao Liu, Yueheng Sun, and Wenjun Wang. Groupbased personalized news recommendation with longand short-term fine-grained matching. ACM Transactions on Information Systems, 42 (1):3:1–3:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3584946.

Xu:2024:NPM

[XPSC24] Shicheng Xu, Liang Pang, Huawei Shen, and Xueqi Cheng. NIR-Prompt: a multitask generalized neural information retrieval training framework. ACM Transactions on Information Systems, 42(2):57:1–57:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3626092.

Xin:2025:LEC

[XSWX25] Haoran Xin, Ying Sun, Chao Wang, and Hui Xiong. LLMCDSR: Enhancing crossdomain sequential recommendation with large language models. ACM Transactions on Information Systems, 43 (5):120:1–120:33, September 2025. CODEN ATISET. ISSN 1046-8188.

Xu:2021:RRM

[XTF^{+}21] Ruijian Xu, Chongyang Tao, Jiazhan Feng, Wei Wu, Rui Yan, and Dongyan Zhao. Response ranking with multitypes of deep interactive representations in retrieval-based dialogues. ACM Transactions on Information Systems, 39 (4):44:1–44:28, October 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3462207.

Xu:2025:HVJ

[XWL^{+}25] Nuo Xu, Pinghui Wang, Zi Liang, Junzhou Zhao, and Xiaohong Guan. How vital is the jurisprudential relevance: Law article-intervened legal case retrieval and matching. ACM Transactions on Information Systems, 43(4):85:1– 85:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Xu:2025:DCL

[XWZ^{+}25] Nuo Xu, Pinghui Wang, Junzhou Zhao, Feiyang Sun, Lin Lan, Jing Tao, Li Pan, and Xiaohong Guan. Distinguish confusion in legal judgment prediction via revised relation knowledge. ACM Transactions on Information Systems, 43(1):6:1–6:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3689628.

Xu:2024:SIG

[XXDW24] Chen Xu, Jun Xu, Zhenhua Dong, and Ji-Rong Wen. Syntactic-informed graph networks for sentence matching. ACM Transactions on Information Systems, 42(2):38:1– 38:??, March 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3609795.

Xu:2024:DCL

[XYCC24] Jingyun Xu, Junnan Yu, Yi Cai, and Tat-Seng Chua. Dual contrastive learning for cross-domain named entity recognition. ACM Transactions on Information Systems, 42(6):163:1–163:??, November 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3678879.

Xia:2023:EDS

[XYW^{+}23a] Xin Xia, Junliang Yu, Qinyong Wang, Chaoqun Yang, Nguyen Quoc Viet Hung, and Hongzhi Yin. Efficient ondevice session-based recommendation. ACM Transactions on Information Systems, 41(4):102:1–102:??, October 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3580364.

Xin:2023:UBL

[XYW^{+}23b] Xin Xin, Jiyuan Yang, Hanbing Wang, Jun Ma, Pengjie Ren, Hengliang Luo, Xinlei Shi, Zhumin Chen, and Zhaochun Ren. On the user behavior leakage from recommender system exposure. ACM Transactions on Information Systems, 41(3):57:1– 57:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3568954.

Xu:2025:LMT

[XYX^{+}25] Chen Xu, Xiaopeng Ye, Jun Xu, Xiao Zhang, Weiran Shen, and Ji-Rong Wen. LTP-MMF: Toward long-term provider max-min fairness under recommendation feedback loops. ACM Transactions on Information Systems, 43 (1):11:1–11:??, January 2025.

CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3695867.

Xue:2023:LDV

[XYZ^{+}23] Lyuxin Xue, Deqing Yang, Shuoyao Zhai, Yuxin Li, and Yanghua Xiao. Learning dual-view user representations for enhanced sequential recommendation. ACM Transactions on Information Systems, 41(4):86:1– 86:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3572028.

Xiao:2025:ICR

[XZCZ25] Likang Xiao, Richong Zhang, Junfan Chen, and Lei Zhang. Including co-relation via concatenate operator for static and temporal knowledge graph embedding. ACM Transactions on Information Systems, 43(5):123:1–123:26, September 2025. CODEN ATISET. ISSN 1046-8188.

Xiang:2025:CGP

[XZD^{+}25] Zongyi Xiang, Yan Zhang, Lixin Duan, Hongzhi Yin, and Ivor W. Tsang. Coherenceguided preference disentanglement for cross-domain recommendations. ACM Transactions on Information Systems, 43(4):109:1–109:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Xie:2022:GNC

[XZH^{+}22] Qianqian Xie, Yutao Zhu, Jimin Huang, Pan Du, and Jian-Yun Nie. Graph neural collaborative topic model for citation recommendation. ACM Transactions on Information Systems, 40(3):48:1– 48:30, July 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3473973.

Xu:2024:RMA

[XZT^{+}24] Zhichao Xu, Hansi Zeng, Juntao Tan, Zuohui Fu, Yongfeng Zhang, and Qingyao Ai. A reusable modelagnostic framework for faithfully explainable recommendation and system scrutability. ACM Transactions on Information Systems, 42(1): 29:1–29:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3605357.

Xu:2021:MMC

[XZY21] Yanan Xu, Yanmin Zhu, and Jiadi Yu. Modeling multiple coexisting category-level intentions for next item recommendation. ACM Transactions on Information Systems, 39(3):23:1–23:24, July 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3441642.

Xie:2022:DGG

[XZZ^{+}22] Zhiwen Xie, Runjie Zhu, Kunsong Zhao, Jin Liu, Guangyou Zhou, and Jimmy Xiangji Huang. Dual gated graph attention networks with dynamic iterative training for cross-lingual entity alignment. ACM Transactions on Information Systems, 40(3):44:1– 44:30, July 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3471165.

Yang:2015:BHC

[Yan15] Hui Yang. Browsing hierarchy construction by minimum evolution. ACM Transactions on Information Systems, 33(3):13:1–13:??, March 2015. CODEN ATISET. ISSN 1046-8188.

Yang:1994:EBM

[YC94] Yiming Yang and Christopher G. Chute. An examplebased mapping method for text categorization and retrieval. ACM Transactions on Information Systems, 12 (3):252–277, July 1994. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.

org:80. Special Issue on Text Categorization.

Yan:2024:CRG

[YCG^{+}24] Mingshi Yan, Zhiyong Cheng, Chen Gao, Jing Sun, Fan Liu, Fuming Sun, and Haojie Li. Cascading residual graph convolutional network for multi-behavior recommendation. ACM Transactions on Information Systems, 42 (1):10:1–10:??, January 2024.

CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3587693.

Yang:2023:DRU

[YCL^{+}23] Mengyue Yang, Guohao Cai, Furui Liu, Jiarui Jin, Zhenhua Dong, Xiuqiang He, Jianye Hao, Weiqi Shao, Jun Wang, and Xu Chen. Debiased recommendation with user feature balancing. ACM Transactions on Information Systems, 41(4):114:1–114:??, October 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3580594.

Yan:2025:UIP

[YCL^{+}25] Mingshi Yan, Zhiyong Cheng, Fan Liu, Yingda Lyu, and Yahong Han. User invariant preference learning for multi-behavior recommendation. ACM Transactions on Information Systems, 43(4): 110:1–110:??, July 2025. CO-DEN ATISET. ISSN 1046-8188.

Yin:2014:LSI

[YCS^{+}14] Hongzhi Yin, Bin Cui, Yizhou Sun, Zhiting Hu, and Ling Chen. LCARS: a spatial item recommender system. ACM Transactions on Information Systems, 32(3):11:1– 11:??, June 2014. CODEN ATISET. ISSN 1046-8188.

Yuan:2015:WWW

[YCZ^{+}15] Quan Yuan, Gao Cong, Kaiqi Zhao, Zongyang Ma, and Aixin Sun. Who, where, when, and what: a nonparametric Bayesian approach to context-aware recommendation and search for Twitter users. ACM Transactions on Information Systems, 33(1): 2:1–2:??, March 2015. CO-DEN ATISET. ISSN 1046-8188.

Yin:2016:JMU

[YCZ^{+}16] Hongzhi Yin, Bin Cui, Xiaofang Zhou, Weiqing Wang, Zi Huang, and Shazia Sadiq. Joint modeling of user checkin behaviors for real-time point-of-interest recommendation. ACM Transactions on Information Systems, 35(2): 11:1–11:??, December 2016. CODEN ATISET. ISSN 1046-8188.

Yao:2022:CAK

[YDW22] Jing Yao, Zhicheng Dou, and Ji-Rong Wen. Clarifying ambiguous keywords with personal word embeddings for personalized search. ACM Transactions on Information Systems, 40(3):43:1– 43:29, July 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3470564.

Yao:2021:RRL

[YDXW21] Jing Yao, Zhicheng Dou, Jun Xu, and Ji-Rong Wen. RLPS: a reinforcement learningbased framework for personalized search. ACM Transactions on Information Systems, 39(3):27:1–27:29, July 2021. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3446617.

Yang:2018:ULP

[YE18] Jing Yang and Carsten Eickhoff. Unsupervised learning of parsimonious general-purpose embeddings for user and location modeling. ACM Transactions on Information Systems, 36(3):32:1–32:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Yang:2025:FPL

[YFO^{+}25] Qi Yang, Aleksandr Farseev, Marlo Ongpin, Alfred Huang, Yu-Yi Chu-Farseeva, Da-Min You, Kirill Lepikhin, and Sergey Nikolenko. Fusing predictive and large language models for actionable recommendations in creative marketing. ACM Transactions on Information Systems, 43 (5):116:1–116:31, September 2025. CODEN ATISET. ISSN 1046-8188.

Yin:2024:UDS

[YFSO24] Qing Yin, Hui Fang, Zhu Sun, and Yew-Soon Ong. Understanding diversity in session-based recommendation. ACM Transactions on Information Systems, 42 (1):24:1–24:??, January 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3600226.

Yang:2015:QCM

[YGZ15] Hui Yang, Dongyi Guan, and Sicong Zhang. The query change model: Modeling session search as a Markov decision process. ACM Transactions on Information Systems, 33(4):20:1–20:??, May 2015. CODEN ATISET. ISSN 1046-8188.

Yu:2025:MMV

[YHDN25] Li Yu, Jianyong Hu, Qihan Du, and Xi Niu. MVideoRec: Micro video recommendations through modality decomposition and contrastive learning. ACM Transactions on Information Systems, 43(3):60:1– 60:??, May 2025. CODEN ATISET. ISSN 1046-8188.

You:2012:EET

[YHS^{+}12] Gae-Won You, Seung-Won Hwang, Young-In Song, Long Jiang, and Zaiqing Nie. Efficient entity translation mining: a parallelized graph alignment approach. ACM Transactions on Information Systems, 30(4):25:1–25:??, November 2012. CODEN ATISET. ISSN 1046-8188.

Yang:2021:HHG

[YHS^{+}21] Tianchi Yang, Linmei Hu, Chuan Shi, Houye Ji, Xiaoli Li, and Liqiang Nie. HGAT: Heterogeneous graph attention networks for semisupervised short text classification. ACM Transactions on Information Systems, 39(3):

32:1–32:29, July 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3450352.

Yin:2024:HHH

[YHWZ24] Zhizhuo Yin, Kai Han, Pengzi Wang, and Xi Zhu. H3GNN: Hybrid hierarchical Hyper-Graph neural network for personalized session-based recommendation. ACM Transactions on Information Systems, 42(3):63:1–63:??, May 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3630002.

Yang:2017:YMP

[YHY^{+}17] Longqi Yang, Cheng-Kang Hsieh, Hongjian Yang, John P. Pollak, Nicola Dell, Serge Belongie, Curtis Cole, and Deborah Estrin. Yum-Me: a personalized nutrient-based meal recommender system. ACM Transactions on Information Systems, 36(1):7:1–7:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Yoshioka:2001:GTK

[YHYO01] Takeshi Yoshioka, George Herman, JoAnne Yates, and Wanda Orlikowski. Genre taxonomy: a knowledge repository of communicative actions. ACM Transactions on Information Systems, 19 (4):431–456, October 2001. CODEN ATISET. ISSN 1046-8188.

Yu:2006:LSC

[YKHW06] Hong Yu, Won Kim, Vasileios Hatzivassiloglou, and John Wilbur. A large scale, corpusbased approach for automatically disambiguating biomedical abbreviations. ACM Transactions on Information Systems, 24(3):380–404, July 2006. CODEN ATISET. ISSN 1046-8188.

Yan:2011:TSG

[YLS^{+}11] Xin Yan, Raymond Y. K. Lau, Dawei Song, Xue Li, and Jian Ma. Toward a semantic granularity model for domainspecific information retrieval. ACM Transactions on Information Systems, 29(3):15:1– 15:??, July 2011. CODEN ATISET. ISSN 1046-8188.

Yao:2023:OAP

[YLW^{+}23] Zijun Yao, Bin Liu, Fei Wang, Daby Sow, and Ying Li. Ontology-aware prescription recommendation in treatment pathways using multievidence healthcare data. ACM Transactions on Information Systems, 41(4):99:1– 99:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3579994.

Ye:2023:TRN

[YLYT23] Haibo Ye, Xinjie Li, Yuan Yao, and Hanghang Tong. Towards robust neural graph collaborative filtering via structure denoising and embedding perturbation. ACM Transactions on Information Systems, 41(3):59:1–59:??, July 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3568396.

Yan:2021:MRA

[YLZW21] Rui Yan, Weiheng Liao, Dongyan Zhao, and Ji-Rong Wen. Multi-response awareness for retrieval-based conversations: Respond with diversity via dynamic representation learning. ACM Transactions on Information Systems, 39(4):54:1–54:29, October 2021. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3470450.

Yan:2022:LTL

[YLZW22] Surong Yan, Kwei-Jay Lin, Xiaolin Zheng, and Haosen Wang. LkeRec: Toward lightweight end-to-end joint representation learning for building accurate and effective recommendation. ACM Transactions on Information Systems, 40(3):54:1– 54:28, July 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3486673.

Yang:2022:LLI

[YMZ^{+}22] Jun Yang, Weizhi Ma, Min Zhang, Xin Zhou, Yiqun Liu, and Shaoping Ma. LegalGNN: Legal information enhanced graph neural network for recommendation. ACM Transactions on Information Systems, 40(2):33:1–33:29, April 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3469887.

Yu:2022:SHQ

[YMZF22] Weiren Yu, Julie McCann, Chengyuan Zhang, and Hakan Ferhatosmanoglu. Scaling high-quality pairwise linkbased similarity retrieval on billion-edge graphs. ACM Transactions on Information Systems, 40(4):78:1–78:45, October 2022. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3495209.

Yi:2024:CGP

[YOM24] Zixuan Yi, Iadh Ounis, and Craig MacDonald. Contrastive graph prompt-tuning for cross-domain recommendation. ACM Transactions on Information Systems, 42(2): 60:1–60:??, March 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3618298.

Yang:2025:OVF

[YPF^{+}25] Enyue Yang, Weike Pan, Lixin Fan, Hanlin Gu, Zhitao Li, Qiang Yang, and Zhong Ming. Ownership verification for federated recommendation. ACM Transactions on Information Systems, 43(3): 69:1–69:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

Yang:2024:DFM

[YPYM24] Enyue Yang, Weike Pan, Qiang Yang, and Zhong Ming. Discrete federated multi-behavior recommendation for privacy-preserving heterogeneous one-class collaborative filtering. ACM Transactions on Information Systems, 42(5):125:1–125:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3652853.

Yi:2023:MAA

[YRC23] Jing Yi, Xubin Ren, and Zhenzhong Chen. Multiauxiliary augmented collaborative variational autoencoder for tag recommendation. ACM Transactions on Information Systems, 41(4): 106:1–106:??, October 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3578932.

Yu:2025:DCD

[YSJ25] Yi Yu, Kazunari Sugiyama, and Adam Jatowt. Domain counterfactual data augmentation for explainable recommendation. ACM Transactions on Information Systems, 43(3):58:1–58:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Yan:2024:TEM

[YSW^{+}24] Surong Yan, Chenglong Shi, Haosen Wang, Lei Chen, Ling Jiang, Ruilin Guo, and Kwei-Jay Lin. Teach and explore: a multiplex information-guided effective and efficient reinforcement learning for sequential recommendation. ACM Transactions on Information Systems, 42(5):120:1–120:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3630003.

Yin:2023:TRSa

[YSXK23a] Hongzhi Yin, Yizhou Sun, Guandong Xu, and Evangelos Kanoulas. Trustworthy recommendation and search: Introduction to the special issue — Part 1. ACM Transactions on Information Systems, 41(3):51:1–51:??, July 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3579995.

Yin:2023:TRSb

[YSXK23b] Hongzhi Yin, Yizhou Sun, Guandong Xu, and Evangelos Kanoulas. Trustworthy recommendation and search: Introduction to the special section — Part 2. ACM Transactions on Information Systems, 41(4):82:1– 82:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3604776.

Yang:2017:NNA

[YSZ^{+}17] Cheng Yang, Maosong Sun, Wayne Xin Zhao, Zhiyuan Liu, and Edward Y. Chang. A neural network approach to jointly modeling social networks and mobile trajectories. ACM Transactions on Information Systems, 35(4):36:1– 36:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Yom-Tov:2013:ESP

[YTD13] Elad Yom-Tov and Fernando Diaz. The effect of social and physical detachment on information need. ACM Transactions on Information Systems, 31(1):4:1–4:??, January 2013. CODEN ATISET. ISSN 1046-8188.

Yom-Tov:2025:SEO

[YTL25] Elad Yom-Tov and Liat Levontin. The in-situ effect of offensive ads on search engine users. ACM Transactions on Information Systems, 43(3): 56:1–56:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

Yan:2015:DDS

[YW15] Su Yan and Xiaojun Wan.

Deep dependency substructurebased learning for multidocument summarization. ACM Transactions on Information Systems, 34(1):3:1–3:??, October 2015. CODEN ATISET. ISSN 1046-8188.

Yao:2025:NRR

[YWL^{+}25] Jing Yao, Xiting Wang, Jianxun Lian, Xiaoyuan Yi, and Xing Xie. Neural recommendation reasoning with logic rules. ACM Transactions on Information Systems, 43 (5):133:1–133:28, September 2025. CODEN ATISET. ISSN 1046-8188.

Yang:2024:DFA

[YWQ^{+}24] Hao Yang, Xian Wu, Zhaopeng Qiu, Yefeng Zheng, and Xu Chen. Distributional fairness-aware recommendation. ACM Transactions on Information Systems, 42 (5):131:1–131:??, September 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3652854.

Yeh:2008:EPH

[YWYL08] Jui-Feng Yeh, Chung-Hsien Wu, Liang-Chih Yu, and Yu-Sheng Lai. Extended probabilistic HAL with close temporal association for psychiatric query document retrieval. ACM Transactions on Information Systems, 27 (1):4:1–4:??, December 2008. CODEN ATISET. ISSN 1046-8188.

Ye:2024:RFB

[YXA^{+}24] Ziyi Ye, Xiaohui Xie, Qingyao Ai, Yiqun Liu, Zhihong Wang, Weihang Su, and Min Zhang. Relevance feedback with brain signals. ACM Transactions on Information Systems, 42(4): 93:1–93:??, July 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3637874.

Yan:2023:AHP

[YXL^{+}23] Ming Yan, Haiyang Xu, Chenliang Li, Junfeng Tian, Bin Bi, Wei Wang, Xianzhe Xu, Ji Zhang, Songfang Huang, Fei Huang, Luo Si, and Rong Jin. Achieving human parity on visual question answering. ACM Transactions on Information Systems, 41(3): 79:1–79:??, July 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3572833.

Yuan:2025:PFP

[YYQ^{+}25] Wei Yuan, Chaoqun Yang, Liang Qu, Quoc Viet Hung Nguyen, Guanhua Ye, and Hongzhi Yin. PTF-FSR: a parameter transmission-free federated sequential recommender system. ACM Transactions on Information Systems, 43(2):52:1–52:??, March 2025. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3708344.

Yuan:2024:MVA

[YYY^{+}24] Wei Yuan, Shilong Yuan, Chaoqun Yang, Nguyen Quoc Viet hung, and Hongzhi Yin. Manipulating visually aware federated recommender systems and its countermeasures. ACM Transactions on Information Systems, 42(3): 64:1–64:??, May 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3630005.

Yuan:2025:HCH

[YZC^{+}25] Meng Yuan, Zhao Zhang, Wei Chen, Chu Zhao, Tong Cai, Deqing Wang, Rui Liu, and Fuzhen Zhuang. HEK-CL: Hierarchical enhanced knowledge-aware contrastive learning for recommendation. ACM Transactions on Information Systems, 43(4):95:1– 95:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Yi:2024:DCN

[YZH^{+}24] Kun Yi, Qi Zhang, Hui He, Kaize Shi, Liang Hu, Ning An, and Zhendong Niu. Deep coupling network for multivariate time series forecasting. ACM Transactions on Information Systems, 42(5):127:1– 127:??, September 2024. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3653447.

Yu:2025:UGP

[YZL^{+}25] Qing Yu, Lixin Zou, Xiangyang Luo, Xiangyu Zhao, and Chenliang Li. Uniform graph pre-training and prompting for transferable recommendation. ACM Transactions on Information Systems, 43(4):101:1–101:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Yang:2024:CKG

[YZS^{+}24] Yang Yang, Chubing Zhang, Xin Song, Zheng Dong, Hengshu Zhu, and Wenjie Li. Contextualized knowledge graph embedding for explainable talent training course recommendation. ACM Transactions on Information Systems, 42(2):33:1–33:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3597022.

Yao:2017:VAR

[YZW^{+}17] Yuan Yao, Wayne Xin Zhao, Yaojing Wang, Hanghang Tong, Feng Xu, and Jian Lu. Version-aware rating prediction for mobile app recommendation. ACM Transactions on Information Systems, 35(4):38:1–38:??, August 2017. CODEN ATISET. ISSN 1046-8188.

Yao:2024:DBL

[YZZS24] Yitong Yao, Jing Zhang, Peng Zhang, and Yueheng Sun. A dual-branch learning model with gradient-balanced loss for long-tailed multi-label text classification. ACM Transactions on Information Systems, 42(2):34:1–34:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3597416.

Zou:2023:UMC

[ZAK^{+}23] Jie Zou, Mohammad Aliannejadi, Evangelos Kanoulas, Maria Soledad Pera, and Yiqun Liu. Users meet clarifying questions: Toward a better understanding of user interactions for search clarification. ACM Transactions on Information Systems, 41 (1):16:1–16:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3524110.

Zhang:2016:MOS

[ZAL^{+}16] Huiling Zhang, Md Abdul Alim, Xiang Li, My T. Thai, and Hien T. Nguyen. Misinformation in online social networks: Detect them all with a limited budget. ACM Transactions on Information Systems, 34(3):18:1–18:??, May 2016. CODEN ATISET. ISSN 1046-8188.

Zhang:2023:PPP

[ZBYH23] Peng-Fei Zhang, Guangdong Bai, Hongzhi Yin, and Zi Huang. Proactive privacypreserving learning for crossmodal retrieval. ACM Transactions on Information Systems, 41(2):35:1–35:??, April 2023. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3545799.

Zhou:2006:ERF

[ZCD06] Zhi-Hua Zhou, Ke-Jia Chen, and Hong-Bin Dai. Enhancing relevance feedback in image retrieval using unlabeled data. ACM Transactions on Information Systems, 24(2):

219–244, April 2006. CODEN ATISET. ISSN 1046-8188.

Zheng:2023:SAK

[ZCLC23] Jianming Zheng, Fei Cai, Yanxiang Ling, and Honghui Chen. Sequence-aware knowledge distillation for a lightweight event representation. ACM Transactions on Information Systems, 41(2):37:1– 37:??, April 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3545798.

Zhu:2022:LHI

[ZCLX22] Nengjun Zhu, Jian Cao, Xinjiang Lu, and Hui Xiong. Learning a hierarchical intent model for next-item recommendation. ACM Transactions on Information Systems, 40(2):38:1–38:28, April 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3473972.

Zhou:2023:EHT

[ZCS^{+}23] Yuchen Zhou, Yanan Cao, Yanmin Shang, Chuan Zhou, Shirui Pan, Zheng Lin, and Qian Li. Explainable hyperbolic temporal point process for user-item interaction sequence generation. ACM Transactions on Information Systems, 41(4):83:1– 83:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3570501.

Zhang:2023:SDA

[ZCZ^{+}23] Weinan Zhang, Yiming Cui, Kaiyan Zhang, Yifa Wang, Qingfu Zhu, Lingzhi Li, and Ting Liu. A static and dynamic attention framework for multi turn dialogue generation. ACM Transactions on Information Systems, 41 (1):15:1–15:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3522763.

Zhang:2022:LSC

[ZCZW22] Wei Zhang, Zeyuan Chen, Hongyuan Zha, and Jianyong Wang. Learning from substitutable and complementary relations for graph-based sequential product recommendation. ACM Transactions on Information Systems, 40(2): 26:1–26:28, April 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3464302.

Zhao:2022:BFG

[ZDW^{+}22] Minghao Zhao, Qilin Deng, Kai Wang, Runze Wu, Jianrong Tao, Changjie Fan, Liang Chen, and Peng Cui. Bilateral filtering graph convolutional network for multirelational social recommendation in the power-law networks. ACM Transactions on Information Systems, 40(2): 31:1–31:24, April 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3469799.

Zhao:2023:AAT

[ZDZ23] Yan Zhao, Liwei Deng, and Kai Zheng. AdaTaskRec: an adaptive task recommendation framework in spatial crowdsourcing. ACM Transactions on Information Systems, 41(4):95:1– 95:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3593582.

Zhang:2023:CLL

[ZDZW23] Han Zhang, Zhicheng Dou, Yutao Zhu, and Ji-Rong Wen. Contrastive learning for legal judgment prediction. ACM Transactions on Information Systems, 41(4):113:1–113:??, October 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3580489.

Zhou:2007:DPM

[ZFL^{+}07] Changqing Zhou, Dan Frankowski, Pamela Ludford, Shashi Shekhar, and Loren Terveen. Discovering personally meaningful places: an interactive clustering approach. ACM Transactions on Information Systems, 25(3):12:1– 12:??, July 2007. CODEN ATISET. ISSN 1046-8188.

Zhang:2022:RQG

[ZGC^{+}22] Ruqing Zhang, Jiafeng Guo, Lu Chen, Yixing Fan, and Xueqi Cheng. A review on question generation from natural language text. ACM Transactions on Information Systems, 40(1):14:1–14:43, January 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3468889.

Zhao:2018:TEF

[ZGCC18] Jingwen Zhao, Yunjun Gao, Gang Chen, and Rui Chen. Towards efficient framework for time-aware spatial keyword queries on road networks. ACM Transactions on Information Systems, 36(3): 24:1–24:??, April 2018. CO-DEN ATISET. ISSN 1046-8188.

Zhang:2020:DFG

[ZGF^{+}20] Ruqing Zhang, Jiafeng Guo, Yixing Fan, Yanyan Lan, and Xueqi Cheng. Dual-factor generation model for conversation. ACM Transactions on Information Systems, 38(3): 31:1–31:31, June 2020. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.
acm.org/doi/abs/10.1145/
3394052.

Zhang:2025:VPP

[ZGL^{+}25] Gangyi Zhang, Chongming Gao, Wenqiang Lei, Xiaojie Guo, Shijun Li, Hongshen Chen, Zhuozhi Ding, Sulong Xu, and Lingfei Wu. Vague preference policy learning for conversational recommendation. ACM Transactions on Information Systems, 43(3):78:1–78:??, May 2025.

CODEN ATISET. ISSN 1046-8188.

Zhang:2017:TAP

[ZGN^{+}17] Dongxiang Zhang, Long Guo, Liqiang Nie, Jie Shao, Sai Wu, and Heng Tao Shen. Targeted advertising in public transportation systems with quantitative evaluation. ACM Transactions on Information Systems, 35(3):20:1– 20:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Zhou:2017:MMD

[ZH17] Guang-You Zhou and Jimmy Xiangji Huang. Modeling and mining domain shared knowledge for sentiment analysis. ACM Transactions on Information Systems, 36(2):18:1– 18:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Zhang:2021:EMN

[Zha21] Min Zhang. Editorial message from the new Editor-in-Chief. ACM Transactions on Information Systems, 39(3): 23e:1–23e:2, July 2021. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3447945.

Zhu:2013:SHF

[ZHC^{+}13] Xiaofeng Zhu, Zi Huang, Hong Cheng, Jiangtao Cui, and Heng Tao Shen. Sparse hashing for fast multimedia search. ACM Transactions on Information Systems, 31(2):

9:1–9:??, May 2013. CODEN ATISET. ISSN 1046-8188.

Zhao:2020:LSR

[ZHC^{+}20] Wayne Xin Zhao, Yupeng Hou, Junhua Chen, Jonathan J. H. Zhu, Eddy Jing Yin, Hanting Su, and Ji-Rong Wen. Learning semantic representations from directed social links to tag microblog users at scale. ACM Transactions on Information Systems, 38(2):17:1–17:30, March 2020. CODEN ATISET. ISSN 1046-8188. URL https://dl.

acm.org/doi/abs/10.1145/
3377550.

Zhao:2022:TIL

[ZHD^{+}22] Jiashu Zhao, Jimmy Xiangji Huang, Hongbo Deng, Yi Chang, and Long Xia. Are topics interesting or not? An LDA-based topic-graph probabilistic model for Web search personalization. ACM Transactions on Information Systems, 40(3):51:1–51:24, July 2022. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3476106.

Zang:2022:CCH

[ZHL^{+}22] Hongyu Zang, Dongcheng Han, Xin Li, Zhifeng Wan, and Mingzhong Wang. CHA: Categorical hierarchy-based attention for next POI recommendation. ACM Transactions on Information Systems, 40(1):7:1–7:22, January 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3464300.

Zou:2023:LAC

[ZHRK23] Jie Zou, Jimmy Huang, Zhaochun Ren, and Evangelos Kanoulas. Learning to ask: Conversational product search via representation learning. ACM Transactions on Information Systems, 41 (2):45:1–45:??, April 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3555371.

Zhang:2022:CVN

[ZHW^{+}22] Peng Zhang, Wenjie Hui, Benyou Wang, Donghao Zhao, Dawei Song, Christina Lioma, and Jakob Grue Simonsen. Complex-valued neural network-based quantum language models. ACM Transactions on Information Systems, 40(4):84:1–84:31, October 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3505138.

Zhao:2014:MTA

[ZHY14] Jiashu Zhao, Jimmy Xiangji Huang, and Zheng Ye. Modeling term associations for probabilistic information retrieval. ACM Transactions on Information Systems, 32 (2):7:1–7:??, April 2014. CO-DEN ATISET. ISSN 1046-

Zhang:2019:MAD

[ZHZ^{+}19] Zheng Zhang, Minlie Huang, Zhongzhou Zhao, Feng Ji, Haiqing Chen, and Xiaoyan Zhu. Memory-augmented dialogue management for taskoriented dialogue systems. ACM Transactions on Information Systems, 37(3): 34:1–34:??, July 2019. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.
acm.org/ft_gateway.cfm?
id=3317612.

Zhang:2024:SSL

[ZJK^{+}24] Shengyu Zhang, Tan Jiang, Kun Kuang, Fuli Feng, Jin Yu, Jianxin Ma, Zhou Zhao, Jianke Zhu, Hongxia Yang, Tat-Seng Chua, and Fei Wu. SLED: Structure learning based denoising for recommendation. ACM Transactions on Information Systems, 42(2):43:1–43:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3611385.

Zeng:2024:XLS

[ZJL^{+}24] Kaisheng Zeng, Hailong Jin, Xin Lv, Fangwei Zhu, Lei Hou, Yi Zhang, Fan Pang, Yu Qi, Dingxiao Liu, Juanzi Li, and Ling Feng. XLORE 3: a large-scale multilingual knowledge graph from heterogeneous Wiki knowledge resources. ACM Transactions on Information Systems, 42 (6):145:1–145:??, November 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3660521.

Zhang:2016:CRD

[ZJT16] Yating Zhang, Adam Jatowt, and Katsumi Tanaka. Causal relationship detection in archival collections of product reviews for understanding technology evolution. ACM Transactions on Information Systems, 35(1):3:1–3:??, October 2016. CODEN ATISET. ISSN 1046-8188.

Zhang:2024:MMM

[ZJW^{+}24] Yazhou Zhang, Ao Jia, Bo Wang, Peng Zhang, Dongming Zhao, Pu Li, Yuexian Hou, Xiaojia Jin, Dawei Song, and Jing Qin. M3GAT: a multi-modal, multi-task interactive graph attention network for conversational sentiment analysis and emotion recognition. ACM Transactions on Information Systems, 42(1):13:1–13:??, January 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3593583.

Zou:2020:TQB

[ZK20] Jie Zou and Evangelos Kanoulas.

Towards question-based highrecall information retrieval: Locating the last few relevant documents for technologyassisted reviews. ACM Transactions on Information Systems, 38(3):27:1–27:35, June 2020. CODEN ATISET. ISSN 1046-8188. URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3388640.

Zheleva:2008:TSR

[ZKG08] Elena Zheleva, Aleksander Kolcz, and Lise Getoor. Trusting spam reporters: a reporter-based reputation system for email filtering. ACM Transactions on Information Systems, 27(1):3:1–3:??, December 2008. CODEN ATISET. ISSN 1046-8188.

Zhang:2022:GTP

[ZKJW22] Chuxu Zhang, Julia Kiseleva, Sujay Kumar Jauhar, and Ryen W. White. Grounded task prioritization with contextaware sequential ranking. ACM Transactions on Information Systems, 40(4):68:1– 68:28, October 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3486861.

Zhang:2023:KBE

[ZKMM23] Richong Zhang, Jaein Kim, Jiajie Mei, and Yongyi Mao. Knowledge base embedding for sampling-based prediction. ACM Transactions on Information Systems, 41(2):28:1– 28:??, April 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3533769.

Zhai:2004:SSM

[ZL04] Chengxiang Zhai and John Lafferty. A study of smoothing methods for language models applied to information retrieval. ACM Transactions on Information Systems, 22 (2):179–214, April 2004. CO-DEN ATISET. ISSN 1046-8188.

Zhang:2017:PLQ

[ZLF^{+}17] Dongxiang Zhang, Yuchen Li, Ju Fan, Lianli Gao, Fumin Shen, and Heng Tao Shen. Processing long queries against short text: Topk advertisement matching in news stream applications. ACM Transactions on Information Systems, 35(3):28:1– 28:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Zhao:2023:RSA

[ZLF^{+}23] Wayne Xin Zhao, Zihan Lin, Zhichao Feng, Pengfei Wang, and Ji-Rong Wen. A revisiting study of appropriate offline evaluation for top-N recommendation algorithms. ACM Transactions on Information Systems, 41(2):32:1– 32:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3545796.

Zhang:2022:ECF

[ZLH^{+}22] Ge Zhang, Zhao Li, Jiaming Huang, Jia Wu, Chuan Zhou, Jian Yang, and Jianliang Gao. eFraudCom: an e-commerce fraud detection system via competitive graph neural networks. ACM Transactions on Information Systems, 40(3):

47:1–47:29, July 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3474379.

Zhang:2025:SMS

[ZLH^{+}25] Xu Zhang, Zexu Lin, Xiaoyu Hu, Jianlei Wang, Wenpeng Lu, and De-Yu Zhou. SECON: Maintaining semantic consistency in data augmentation for code search. ACM Transactions on Information Systems, 43(2):36:1– 36:??, March 2025. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3686151.

Zeng:2020:NIR

[ZLL^{+}20] Zijie Zeng, Jing Lin, Lin Li, Weike Pan, and Zhong Ming. Next-item recommendation via collaborative filtering with bidirectional item similarity. ACM Transactions on Information Systems, 38(1):7:1– 7:22, February 2020. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3366172.

Zhang:2022:JPF

[ZLL^{+}22] Peng Zhang, Baoxi Liu, Tun Lu, Xianghua Ding, Hansu Gu, and Ning Gu. Jointly predicting future content in multiple social media sites based on multi-task learning. ACM Transactions on Information Systems, 40(4):79:1– 79:28, October 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3495530.

Zhang:2023:UBS

[ZLM^{+}23] Junqi Zhang, Yiqun Liu, Jiaxin Mao, Weizhi Ma, Jiazheng Xu, Shaoping Ma, and Qi Tian. User behavior simulation for search result re-ranking. ACM Transactions on Information Systems, 41(1):5:1–5:??, January 2023. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3511469.

Zhao:2024:DTR

[ZLRW24] Wayne Xin Zhao, Jing Liu, Ruiyang Ren, and Ji-Rong Wen. Dense text retrieval based on pretrained language models: a survey. ACM Transactions on Information Systems, 42(4):89:1– 89:??, July 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3637870.

Zhang:2023:MLA

[ZLS^{+}23] Xinyue Zhang, Jingjing Li, Hongzu Su, Lei Zhu, and Heng Tao Shen. Multilevel attention-based domain disentanglement for BCDR. ACM Transactions on Information Systems, 41(4):97:1– 97:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3576925.

Zhang:2016:TRM

[ZLW16] Chenyi Zhang, Hongwei Liang, and Ke Wang. Trip recommendation meets realworld constraints: POI availability, diversity, and traveling time uncertainty. ACM Transactions on Information Systems, 35(1):5:1–5:??, October 2016. CODEN ATISET. ISSN 1046-8188.

Zhang:2023:LLF

[ZLW^{+}23] Honglei Zhang, Fangyuan Luo, Jun Wu, Xiangnan He, and Yidong Li. LightFR: Lightweight federated recommendation with privacypreserving matrix factorization. ACM Transactions on Information Systems, 41 (4):90:1–90:??, October 2023. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3578361.

Zhang:2024:RSM

[ZLW24] Jin Zhang, Xinrui Li, and Liye Wang. A review selection method based on consumer decision phases in ecommerce. ACM Transactions on Information Systems, 42(1):4:1–4:??, January 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3587265.

Zeng:2022:MGL

[ZLWW22] Xingshan Zeng, Jing Li, Lingzhi Wang, and Kam-Fai Wong. Modeling global and local interactions for online conversation recommendation. ACM Transactions on Information Systems, 40(3): 49:1–49:33, July 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3473970.

Zhang:2022:QTG

[ZLY^{+}22] Xiao Zhang, Meng Liu, Jianhua Yin, Zhaochun Ren, and Liqiang Nie. Question tagging via graph-guided ranking. ACM Transactions on Information Systems, 40(1): 12:1–12:23, January 2022. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3468270.

Zhang:2025:LYE

[ZLZ25a] Jinyu Zhang, Chao Li, and Zhongying Zhao. Lightweight yet efficient: an external attentive graph convolutional network with positional prompts for sequential recommendation. ACM Transactions on Information Systems, 43(3):79:1–79:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Zhu:2025:MBR

[ZLZ^{+}25b] Xi Zhu, Fake Lin, Ziwei Zhao, Tong Xu, Xiangyu Zhao, Zikai Yin, Xueying Li, and Enhong Chen. Multi-behavior recommendation with personalized directed acyclic behavior graphs. ACM Transactions on Information Systems, 43 (1):20:1–20:??, January 2025. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3696417.

Zhang:2014:TPB

[ZM14] Richong Zhang and Yongyi Mao. Trust prediction via belief propagation. ACM Transactions on Information Systems, 32(3):15:1–15:??, June 2014. CODEN ATISET. ISSN 1046-8188.

Zhang:2020:PLP

[ZMK^{+}20] Richong Zhang, Samuel Mensah, Fanshuang Kong, Zhiyuan Hu, Yongyi Mao, and Xudong Liu. Pairwise link prediction model for out of vocabulary knowledge base entities. ACM Transactions on Information Systems, 38(4):36:1– 36:28, October 2020. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3406116.

Zheng:2019:CCM

[ZML^{+}19] Yukun Zheng, Jiaxin Mao, Yiqun Liu, Cheng Luo, Min Zhang, and Shaoping Ma. Constructing click model for mobile search with viewport time. ACM Transactions on Information Systems, 37(4): 43:1–43:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3360486.

Zhang:2024:TCM

[ZMN^{+}24] Shengyu Zhang, Qiaowei Miao, Ping Nie, Mengze Li, Zhengyu Chen, Fuli Feng, Kun Kuang, and Fei Wu. Transferring causal mechanism over meta-representations for target-unknown crossdomain recommendation. ACM Transactions on Information Systems, 42(4):103:1– 103:??, July 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3643807.

Zhang:2024:RCF

[ZMZ^{+}24] An Zhang, Wenchang Ma, Jingnan Zheng, Xiang Wang, and Tat-Seng Chua. Robust collaborative filtering to popularity distribution shift. ACM Transactions on Information Systems, 42(3):80:1– 80:??, May 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3627159.

Zhang:2025:VTG

[ZMZ25] Shuo Zhang, Xiangwu Meng, and Yujie Zhang. Variational type graph autoencoder for denoising on event recommendation. ACM Transactions on Information Systems, 43 (1):26:1–26:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3703156.

Zhang:2017:CIJ

[ZNL^{+}17] Dongxiang Zhang, Liqiang Nie, Huanbo Luan, Kian-Lee Tan, Tat-Seng Chua, and Heng Tao Shen. Compact indexing and judicious searching for billion-scale microblog retrieval. ACM Transactions on Information Systems, 35(3):27:1–27:??, June 2017. CODEN ATISET. ISSN 1046-8188.

Zobel:2006:DVS

[Zob06] J. Zobel. Detection of video sequences using compact signatures. ACM Transactions on Information Systems, 24 (1):1–50, January 2006. CO-DEN ATISET. ISSN 1046-8188.

Zhang:2024:TBP

[ZOML24] Xinyu Zhang, Kelechi Ogueji, Xueguang Ma, and Jimmy Lin. Toward best practices for training multilingual dense retrieval models. ACM Transactions on Information Systems, 42(2):39:1–39:??, March 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3613447.

Zhu:2024:CMC

[ZPW^{+}24] Yunchang Zhu, Liang Pang, Kangxi Wu, Yanyan Lan, Huawei Shen, and Xueqi Cheng. Cross-model comparative loss for enhancing neuronal utility in language understanding. ACM Transactions on Information Systems, 42(5):123:1–123:??, September 2024. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3652599.

Zheng:2023:ADR

[ZQC^{+}23] Ruiqi Zheng, Liang Qu, Bin Cui, Yuhui Shi, and Hongzhi Yin. AutoML for deep recommender systems: a survey. ACM Transactions on Information Systems, 41(4):101:1– 101:??, October 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3579355.

Zhu:2024:UMR

[ZQH^{+}24a] Zhengbang Zhu, Rongjun Qin, Junjie Huang, Xinyi Dai, Yang Yu, Yong Yu, and Weinan Zhang. Understanding or manipulation: Rethinking online performance gains of modern recommender systems. ACM Transactions on Information Systems, 42(4): 90:1–90:??, July 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3637869.

Zhuo:2024:MHM

[ZQH^{+}24b] Xingrui Zhuo, Shengsheng Qian, Jun Hu, Fuxin Dai, Kangyi Lin, and Gongqing Wu. Multi-hop multiview memory transformer for session-based recommendation. ACM Transactions on Information Systems, 42(6): 144:1–144:??, November 2024. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3663760.

Zezula:1991:DPS

[ZRT91] P. Zezula, F. Rabitti, and P. Tiberio. Dynamic partitioning of signature files. ACM Transactions on Information Systems, 9(4):336– 369, October 1991. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Zhu:2002:TKB

[ZRZ02] Lei Zhu, Aibing Rao, and Aidong Zhang. Theory of keyblock-based image retrieval. ACM Transactions on Information Systems, 20 (2):224–257, 2002. CODEN ATISET. ISSN 1046-8188.

Zhang:2025:REU

[ZSH^{+}25] Junjie Zhang, Wenqi Sun, Yupeng Hou, Xin Zhao, and Ji-Rong Wen. Review-enhanced universal sequence representation learning for recommender systems. ACM Transactions on Information Systems, 43(3):74:1–74:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Zhang:2025:FEN

[ZSL^{+}25] Xiaoyu Zhang, Shaoyun Shi, Yishan Li, Weizhi Ma, Peijie Sun, and Min Zhang. Featureenhanced neural collaborative reasoning for explainable recommendation. ACM Transactions on Information Systems, 43(1):7:1–7:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3690381.

Zou:2024:KEC

[ZSLK24] Jie Zou, Aixin Sun, Cheng Long, and Evangelos Kanoulas. Knowledge-enhanced conversational recommendation via transformer-based sequential modeling. ACM Transactions on Information Systems, 42 (6):162:1–162:??, November 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3677376.

Zhu:2022:LNG

[ZSN^{+}22] Yutao Zhu, Ruihua Song, Jian-Yun Nie, Pan Du, Zhicheng Dou, and Jin Zhou. Leveraging narrative to generate movie script. ACM Transactions on Information Systems, 40(4):86:1–86:32, October 2022. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3507356.

Zha:2024:TUR

[ZSQ^{+}24] Rui Zha, Ying Sun, Chuan Qin, Le Zhang, Tong Xu, Hengshu Zhu, and Enhong Chen. Towards unified representation learning for career mobility analysis with trajectory hypergraph. ACM Transactions on Information Systems, 42(4):110:1–110:??, July 2024. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3651158.

Zhang:2021:GBR

[ZSY^{+}21] Yuan Zhang, Fei Sun, Xiaoyong Yang, Chen Xu, Wenwu Ou, and Yan Zhang. Graphbased regularization on embedding layers for recommendation. ACM Transactions on Information Systems, 39(1): 2:1–2:27, January 2021. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3414067.

Zhang:2024:TDL

[ZSZ^{+}24] Yuting Zhang, Ying Sun, Fuzhen Zhuang, Yongchun Zhu, Zhulin An, and Yongjun Xu. Triple dual learning for opinion-based explainable recommendation. ACM Transactions on Information Systems, 42(3):70:1–70:??, May 2024. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3631521.

Zhang:2017:FFT

[ZTM^{+}17] Jing Zhang, Jie Tang, Cong Ma, Hanghang Tong, Yu Jing, Juanzi Li, Walter Luyten, and Marie-Francine Moens. Fast and flexible top-k similarity search on large networks. ACM Transactions on Information Systems, 36(2):13:1– 13:??, September 2017. CO-DEN ATISET. ISSN 1046-8188.

Zhao:2025:BSP

[ZTZL25] Chuang Zhao, Hui Tang, Hongke Zhao, and Xiaomeng Li. Beyond sequential patterns: Rethinking healthcare predictions with contextual insights. ACM Transactions on Information Systems, 43 (4):107:1–107:??, July 2025. CODEN ATISET. ISSN 1046-8188.

Zhao:2023:TAP

[ZWC^{+}23] Yuyue Zhao, Xiang Wang, Jiawei Chen, Yashen Wang, Wei Tang, Xiangnan He, and Haiyong Xie. Time-aware path reasoning on knowledge graph for recommendation. ACM Transactions on Information Systems, 41(2):26:1– 26:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3531267.

Zhang:2025:DDG

[ZWC^{+}25] Zeyang Zhang, Xin Wang, Haibo Chen, Haoyang Li, and Wenwu Zhu. Disentangled dynamic graph attention network for out-of-distribution sequential recommendation. ACM Transactions on Information Systems, 43(1):19:1– 19:??, January 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3701988.

Zhou:2022:DAU

[ZWE^{+}22] Sheng Zhou, Xin Wang, Martin Ester, Bolang Li, Chen Ye, Zhen Zhang, Can Wang, and Jiajun Bu. Directionaware user recommendation based on asymmetric network embedding. ACM Transactions on Information Systems, 40(2):29:1–29:23, April 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3466754.

Zhang:2023:KEA

[ZWF^{+}23] Yingying Zhang, Xian Wu, Quan Fang, Shengsheng Qian, and Changsheng Xu. Knowledge-enhanced attributed multi-task learning for medicine recommendation. ACM Transactions on Information Systems, 41(1):17:1– 17:??, January 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3527662.

Zhu:2022:BHD

[ZWL^{+}22] Yadong Zhu, Xiliang Wang, Qing Li, Tianjun Yao, and Shangsong Liang. BotSpot++: a hierarchical deep ensemble model for bots install fraud detection in mobile advertising. ACM Transactions on Information Systems, 40(3): 50:1–50:28, July 2022. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3476107.

Zhu:2023:EQB

[ZWL^{+}23] Lei Zhu, Tianshi Wang, Jingjing Li, Zheng Zhang, Jialie Shen, and Xinhua Wang. Efficient query-based black-box attack against cross-modal hashing retrieval.

ACM Transactions on Information Systems, 41(3):54:1– 54:??, July 2023. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3559758.

Zhang:2019:QAK

[ZWMH19] Richong Zhang, Yue Wang, Yongyi Mao, and Jinpeng Huai. Question answering in knowledge bases: a verification assisted model with iterative training. ACM Transactions on Information Systems, 37(4):40:1–40:??, December 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3345557.

Zhan:2019:LMA

[ZWRL19] Xueying Zhan, Yaowei Wang, Yanghui Rao, and Qing Li. Learning from multiannotator data: a noiseaware classification framework. ACM Transactions on Information Systems, 37 (2):26:1–26:??, March 2019. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3309543.

Zhou:2023:EMV

[ZWrWZ23] Kun Zhou, Hui Wang, Ji rong Wen, and Wayne Xin Zhao. Enhancing multiview smoothness for sequential recommendation models. ACM Transactions on Information Systems, 41(4):107:1– 107:??, October 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3582495.

Zhang:2024:ELM

[ZWS24] Chen Zhang, Benyou Wang, and Dawei Song. On elastic language models. ACM Transactions on Information Systems, 42(6):161:1–161:??, November 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3677375.

Zhang:2025:KDK

[ZWS^{+}25] Haonan Zhang, Dongxia Wang, Zhu Sun, Yanhui Li, Youcheng Sun, Huizhi Liang, and Wenhai Wang. KG4RecEval: Does knowledge graph really matter for recommender systems? ACM Transactions on Information Systems, 43(3):64:1– 64:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Zhao:2023:PGS

[ZWT^{+}23] Shiwei Zhao, Runze Wu, Jianrong Tao, Manhu Qu, Minghao Zhao, Changjie Fan, and Hongke Zhao. perCLTV: a general system for personalized customer lifetime value prediction in online games. ACM Transactions on Information Systems, 41(1):23:1– 23:??, January 2023. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3530012.

Zheng:2023:IAD

[ZWX^{+}23] Zhi Zheng, Chao Wang, Tong Xu, Dazhong Shen, Penggang Qin, Xiangyu Zhao, Baoxing Huai, Xian Wu, and Enhong Chen. Interactionaware drug package recommendation via policy gradient. ACM Transactions on Information Systems, 41(1): 3:1–3:??, January 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3511020.

Zhang:2024:RAS

[ZWX^{+}24] Yitao Zhang, Changxuan Wan, Keli Xiao, Qizhi Wan, Dexi Liu, and Xiping Liu. rHDP: an aspect sharingenhanced hierarchical topic model for multi-domain corpus. ACM Transactions on Information Systems, 42(3): 71:1–71:??, May 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3631352.

Zhang:2024:SCS

[ZWY^{+}24] Yabin Zhang, Zhenlei Wang, Wenhui Yu, Lantao Hu, Peng Jiang, Kun Gai, and Xu Chen. Soft contrastive sequential recommendation. ACM Transactions on Information Systems, 42(6):154:1– 154:??, November 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3665325.

Zhang:2022:MGH

[ZWZ^{+}22] Chengyuan Zhang, Yang Wang, Lei Zhu, Jiayu Song, and Hongzhi Yin. Multigraph heterogeneous interaction fusion for social recommendation. ACM Transactions on Information Systems, 40(2):28:1–28:26, April 2022. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3466641.

Zhang:2024:SSN

[ZWZ^{+}24] Yuxiang Zhang, Junjie Wang, Xinyu Zhu, Tetsuya Sakai, and Hayato Yamana. SSR: Solving named entity recognition problems via a singlestream reasoner. ACM Transactions on Information Systems, 42(5):138:1–138:??, September 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3655619.

Zhu:2025:CDC

[ZWZS25] Jiajie Zhu, Yan Wang, Feng Zhu, and Zhu Sun. Causal deconfounding via confounder disentanglement for dualtarget cross-domain recommendation. ACM Transactions on Information Systems, 43(5):128:1–128:33, September 2025. CODEN ATISET. ISSN 1046-8188.

Zhang:2025:RIF

[ZXH^{+}25] Junjie Zhang, Ruobing Xie, Yupeng Hou, Xin Zhao, Leyu Lin, and Ji-Rong Wen. Recommendation as instruction following: a large language model empowered recommendation approach. ACM Transactions on Information Systems, 43(5):114:1–114:37, September 2025. CODEN ATISET. ISSN 1046-8188.

Zhang:2024:BPL

[ZXM^{+}24] Xiaokun Zhang, Bo Xu, Fenglong Ma, Chenliang Li, Yuan Lin, and Hongfei Lin. Bipreference learning heterogeneous hypergraph networks for session-based recommendation. ACM Transactions on Information Systems, 42(3): 68:1–68:??, May 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3631940.

Zhang:2020:UAC

[ZXZL20] Xiaoying Zhang, Hong Xie, Junzhou Zhao, and John C. S. Lui. Understanding assimilation-contrast effects in online rating systems: Modelling, debiasing, and applications. ACM Transactions on Information Systems, 38(1):2:1–2:??, February 2020. CODEN ATISET. ISSN 1046-8188. URL https:

//dl.acm.org/ft_gateway.
cfm?id=3362651.

Zhao:2023:MUR

[ZXZL23] Yu Zhao, Qiang Xu, Ying Zou, and Wei Li. Modeling user reviews through Bayesian graph attention networks for recommendation. ACM Transactions on Information Systems, 41(3):78:1– 78:??, July 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3570500.

Zhao:2021:PES

[ZY21] Guangzhen Zhao and Peng Yang. Pretrained embeddings for stance detection with hierarchical capsule network on social media. ACM Transactions on Information Systems, 39(1):1:1–1:32, January 2021. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3412362.

Zhang:2024:CDN

[ZYC^{+}24] Xinghua Zhang, Bowen Yu, Xin Cong, Taoyu Su, Quangang Li, Tingwen Liu, and Hongbo Xu. Cross-domain NER under a divide-and-transfer paradigm. ACM Transactions on Information Systems, 42(5):137:1–137:??, September 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3655618.

Zhou:2024:RRO

[ZYD^{+}24] Yujia Zhou, Jing Yao, Zhicheng Dou, Yiteng Tu, Ledell Wu, Tat-Seng Chua, and Ji-Rong Wen. ROGER: Rankingoriented generative retrieval. ACM Transactions on Information Systems, 42(6):155:1– 155:??, November 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3603167.

Zhao:2025:SGC

[ZYL^{+}25] Chu Zhao, Enneng Yang, Yuliang Liang, Jianzhe Zhao, Guibing Guo, and Xingwei Wang. Symmetric graph contrastive learning against noisy views for recommendation. ACM Transactions on Information Systems, 43(3): 80:1–80:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

Zhang:2023:MML

[ZYW^{+}23] Hang Zhang, Yajun Yang, Xin Wang, Hong Gao, and Qinghua Hu. MLI: a multilevel inference mechanism for user attributes in social networks. ACM Transactions on Information Systems, 41(2): 39:1–39:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3545797.

Zhang:2021:ICR

[ZYY^{+}21] Chuxu Zhang, Huaxiu Yao, Lu Yu, Chao Huang, Dongjin Song, Haifeng Chen, Meng Jiang, and Nitesh V. Chawla. Inductive contextual relation learning for personalization. ACM Transactions on Information Systems, 39(3):35:1– 35:22, July 2021. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3450353.

Zou:2024:ATL

[ZYY^{+}24] Tao Zou, Le Yu, Junchen Ye, Leilei Sun, Bowen Du, and Deqing Wang. Adaptive taxonomy learning and historical patterns modeling for patent classification. ACM Transactions on Information Systems, 42(6):160:1–160:??, November 2024. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3674834.

Zhang:2025:SDA

[ZZC^{+}25] Qin Zhang, Mengqi Zheng, Shangsi Chen, Han Liu, and Meng Fang. Self data augmentation for open domain question answering. ACM Transactions on Information Systems, 43(2):51:1– 51:??, March 2025. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3707449.

Zhang:2020:MLC

[ZZD^{+}20] Dong Zhang, Shu Zhao, Zhen Duan, Jie Chen, Yanping Zhang, and Jie Tang. A multi-label classification method using a hierarchical and transparent representation for paper-reviewer recommendation. ACM Transactions on Information Systems, 38(1):5:1–5:20, February 2020. CODEN ATISET. ISSN 1046-8188. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3361719.

Zhang:2025:UPA

[ZZD25] Haobo Zhang, Qiannan Zhu, and Zhicheng Dou. A unified prompt-aware framework for personalized search and explanation generation. ACM Transactions on Information Systems, 43(3):71:1– 71:??, May 2025. CODEN ATISET. ISSN 1046-8188.

Zhou:2025:ECS

[ZZG^{+}25] Qi Zhou, Peng Zhang, Hansu Gu, Tun Lu, and Ning Gu. Exploring cross-site user modeling without cross-site user identity linkage: a case study of content preference prediction. ACM Transactions on Information Systems, 43 (1):16:1–16:??, January 2025. CODEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3697832.

Zhao:2018:ALT

[ZZH^{+}18] Wayne Xin Zhao, Wenhui Zhang, Yulan He, Xing Xie, and Ji-Rong Wen. Automatically learning topics and difficulty levels of problems in online judge systems. ACM Transactions on Information Systems, 36(3):27:1– 27:??, April 2018. CODEN ATISET. ISSN 1046-8188.

Zhang:2025:MMR

[ZZH^{+}25] Chengde Zhang, Zihan Zhang, Jiaying Huang, Yan Liu, Dawei Jin, Xia Xiao, and Zuwu Shen. MKCRec: Metarelation guided knowledge coupling for paper recommendation. ACM Transactions on Information Systems, 43(3): 77:1–77:??, May 2025. CO-DEN ATISET. ISSN 1046-8188.

Zhao:2015:GSB

[ZZL^{+}15] Wayne Xin Zhao, Xudong Zhang, Daniel Lemire, Dongdong Shan, Jian-Yun Nie, Hongfei Yan, and Ji-Rong Wen. A general SIMDbased approach to accelerating compression algorithms. ACM Transactions on Information Systems, 33(3):15:1– 15:??, March 2015. CODEN ATISET. ISSN 1046-8188.

Zhang:2021:ERT

[ZZL^{+}21] Qi Zhang, Hengshu Zhu, Qi Liu, Enhong Chen, and Hui Xiong. Exploiting realtime search engine queries for earthquake detection: a summary of results. ACM Transactions on Information Systems, 39(3):37:1–37:32, July 2021. CODEN ATISET. ISSN 1046-8188. URL https:/

/dl.acm.org/doi/10.1145/
3453842.

Zhu:2022:EMM

[ZZL^{+}22] Lei Zhu, Chaoqun Zheng, Xu Lu, Zhiyong Cheng, Liqiang Nie, and Huaxiang Zhang. Efficient multi-modal hashing with online query adaption for multimedia retrieval. ACM Transactions on Information Systems, 40(2): 41:1–41:36, April 2022. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3477180.

Zang:2023:SCD

[ZZL^{+}23] Tianzi Zang, Yanmin Zhu, Haobing Liu, Ruohan Zhang, and Jiadi Yu. A survey on cross-domain recommendation: Taxonomies, methods, and future directions. ACM Transactions on Information Systems, 41(2):42:1– 42:??, April 2023. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3548455.

Zhang:2025:SLE

[ZZSS25] Yi Zhang, Yiwen Zhang, Lei Sang, and Victor S. Sheng. Simplify to the limit! Embedding-less graph collaborative filtering for recommender systems. ACM Transactions on Information Systems, 43(1):22:1– 22:??, January 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3701230.

Zeng:2021:RLB

[ZZT^{+}21] Weixin Zeng, Xiang Zhao, Jiuyang Tang, Xuemin Lin, and Paul Groth. Reinforcement learning-based collective entity alignment with adaptive features. ACM Transactions on Information Systems, 39(3):26:1–26:31, July 2021. CODEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3446428.

Zhang:2025:ACM

[ZZW^{+}25] Leping Zhang, Xiao Zhang, Yichao Wang, Xuan Li, Zhenhua Dong, and Jun Xu. Adapting constrained Markov decision process for OCPC bidding with delayed conversions. ACM Transactions on Information Systems, 43(2): 47:1–47:??, March 2025. CO-DEN ATISET. ISSN 1046-

                                        8188. URL https://dl.acm.
                                        org/doi/10.1145/3706420.

Zhang:2023:RGB

[ZZY^{+}23] Yi Zhang, Yiwen Zhang, Dengcheng Yan, Shuiguang Deng, and Yun Yang. Revisiting graph-based recommender systems from the perspective of variational auto-encoder. ACM Transactions on Information Systems, 41(3):81:1– 81:??, July 2023. CODEN ATISET. ISSN 1046-8188.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3573385.

Zhao:2016:PLB

[ZZZ^{+}16] Wayne Xin Zhao, Ningnan Zhou, Wenhui Zhang, Ji-Rong Wen, Shan Wang, and Edward Y. Chang. A probabilistic lifestyle-based trajectory model for social strength inference from human trajectory data. ACM Transactions on Information Systems, 35(1):8:1–8:??, October 2016.

CODEN ATISET. ISSN 1046-8188.

Zang:2024:CMV

[ZZZ^{+}24] Tianzi Zang, Yanmin Zhu, Ruohan Zhang, Chunyang Wang, Ke Wang, and Jiadi Yu. Contrastive multi-view interest learning for crossdomain sequential recommendation. ACM Transactions on Information Systems, 42(3): 76:1–76:??, May 2024. CO-DEN ATISET. ISSN 1046-

8188. URL https://dl.acm.
org/doi/10.1145/3632402.

Zhang:2025:MLP

[ZZZ^{+}25] Dan Zhang, Shaojie Zheng, Yifan Zhu, Huihui Yuan, Jibing Gong, and Jie Tang. MCAP: Low-pass GNNs with matrix completion for academic recommendations. ACM Transactions on Information Systems, 43(2):33:1– 33:??, March 2025. CODEN ATISET. ISSN 1046-8188.

URL https://dl.acm.org/
doi/10.1145/3698193.